### v2

# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

Abstract
The application of machine learning in wireless communications has been extensively explored, with deep unfolding emerging as a powerful model-based technique. Deep unfolding enhances interpretability by transforming complex iterative algorithms into structured layers of deep neural networks (DNNs). This approach seamlessly integrates domain knowledge with deep learning (DL), leveraging the strengths of both methods to simplify complex signal processing tasks in communication systems. To provide a solid foundation, we first present a brief overview of DL and deep unfolding. We then explore the applications of deep unfolding in key areas, including signal detection, channel estimation, beamforming design, decoding for error-correcting codes, sensing and communication, power allocation, and security. Each section focuses on a specific task, highlighting its significance in emerging 6G technologies and reviewing recent advancements in deep unfolding-based solutions. Finally, we discuss the challenges associated with developing deep unfolding techniques and propose potential improvements to enhance their applicability across diverse wireless communication scenarios.
Index Terms:
Deep learning (DL), deep unfolding, model-driven machine learning, wireless communication signal processing, ISAC, IRS, THz, mmWave.
I Introduction
The future mobile generations are expected to experience a surge in resource demand due to the exponential growth of connected devices. With the advent of 5G, billions of devices require seamless connectivity, intensifying competition for efficient spectrum allocation [1, 2, 3]. Each network has unique requirements and constraints, making resource allocation a complex challenge. Given the inherent limitations of available resources, problem formulation must prioritize flexibility, adaptability, and efficient utilization. Significant advancements in physical layer technologies have enhanced wireless communication systems, including massive multiple-input multiple-output (MIMO), intelligent reflecting surfaces (IRS), millimeter-wave (mmWave) communication, physical layer security, and non-orthogonal multiple access (NOMA). Additionally, orthogonal time-frequency space (OTFS) modulation is emerging as a promising alternative to orthogonal frequency-division multiplexing (OFDM), further improving spectral efficiency and resilience in next-generation networks.
In the communication system, traditional model-based analytical algorithms often employ iterative optimization techniques to achieve accurate solutions. Increasing the number of iterations improves performance as the algorithm converges closer to the optimal solution. Deriving an optimal solution in the form of a closed-form expression can be highly challenging in specific arbitrary communication scenarios. Hence, we resort to a sub-optimal solution in the traditional model-based method. Moreover, practical implementation on hardware imposes significant constraints on computational resources, latency, and power consumption, which limit the feasibility of deploying highly iterative methods. The classical iterative approach often relies on mathematically tractable models and algorithms for specific scenarios. Hence, adapting to dynamic and heterogeneous conditions, such as rapidly changing channels, interference, and diverse network configurations, can be challenging. Machine learning techniques offer a transformative approach by leveraging data-driven learning to extract patterns, adapt to new conditions, and approximate solutions efficiently. These methods can generalize across varying scenarios, learn optimal strategies from experience, and enable the design of intelligent communication systems that operate effectively in complex, real-world environments. These methods can potentially learn task-specific patterns and approximate solutions with significantly reduced computational complexity. Machine learning models can provide efficient and scalable alternatives to traditional iterative algorithms, enabling better performance under hardware limitations. This paradigm shift has opened new avenues for addressing complex problems in communication systems, bridging the gap between theoretical advancements and practical deployment.
Deep Learning (DL), a machine learning subclass maps the relationship between input and output without explicitly modeling it in the architecture [4]. It involves the use of deep neural networks (DNN) with multiple hidden layers (hence “deep”) to analyze and interpret data. Numerous architectures have been proposed such as multilayer perceptron (MLP), convolutional NN (CNN), deep reinforcement learning (DRL), generative adversarial networks (GANs), etc. DNNs have been applied to various components of wireless communication, including detection, channel estimation, beamforming designs, and power allocation [5, 6, 7, 8, 9]. DL architectures exhibit several inherent limitations. These include complex training procedures and a black-box nature. The training process for these models requires extensive datasets, considerable time, and intricate fine-tuning. The complexity of implementing DNN architectures often constrains their applicability in real-world systems. Therefore, DL models are highly sensitive to the specific configurations of wireless systems, such as the number of antennas, users, and operating frequencies, requiring retraining or reimplementation whenever these parameters change. These challenges are not typically encountered in conventional optimization techniques, underscoring the potential advantages of using traditional algorithms [10, 11].
Communication problems are inherently influenced by dynamic physical constraints that evolve in real-time. These constraints stem from factors such as the physical environment, hardware limitations, and the specific operational requirements of communication systems. Different communication scenarios impose varying demands on bandwidth, power consumption, and latency, which fluctuate with user density, data rates, and service types. Interpreting and addressing these constraints using conventional DNNs presents significant challenges. Traditional data-driven neural network architectures were developed primarily to process language and vision data, rendering them suboptimal for wireless communication tasks that involve channels, signals, and outputs governed by complex constraints. Although certain constraints can be handled through the application of specific activation functions, for example, employing the tanh function to enforce a constant modulus constraint in analog beamforming, standard activation functions generally lack the versatility to manage the intricate constraints characteristic of wireless communication systems. To mitigate these challenges, model-driven DL approaches facilitate the development of problem-specific network architectures and customized activation functions, optimized for the unique signal processing requirements of wireless communication applications [12].
Compared to the conventional fully-connected DL architectures, model-driven DL have fewer trainable parameters, implying less training time and computational complexity. Deep unfolding is an approach that falls into the category of model-driven DL [13]. The main motive behind developing a hybrid model (model-driven DL) for decision mapping, is to utilize the benefits of both optimization and data-driven DL. Model-driven DL is categorized into three frameworks as given in [13] : (1) learned optimizers [14, 15] (2) deep unfolding [16] (3) DNN-aided optimizers [17]. Here, we will focus on the second category i.e., deep unfolding. A deep unfolded network can be trained in an easier and faster fashion with smaller datasets than conventional DNNs. This is because deep unfolding incorporates domain knowledge from iterative algorithms into layers of DNNs, offering improved performance and interpretability [18]. The main idea in solving communication signal processing tasks lies in optimizing key hyperparameters of conventional algorithms based on data training. This includes elements such as step sizes in gradient-based methods, initial solutions, searching directions, criteria for early termination, and other relevant parameters. Deep unfolding uses suitable loss functions to train learnable parameters that are interpretable. Each iteration of the iterative algorithm is incorporated as a hidden layer of the DNN. These layers are then stacked to form a network, mimicking the optimization flow of algorithms. Furthermore, in many cases, the trainable parameters are independent of the system configurations, implying that the trained unfolding model can work for various scenarios.
I-A Related Works
The growing application of deep unfolding in communication systems highlights the need for a comprehensive survey in this area. Wang et al. in [19] presented a review of DL and deep unfolding approaches for addressing physical layer signal processing challenges, including MIMO detection, channel estimation, and channel decoding. The authors of [20] provided an overview of deep unfolding applications in transceiver design, MIMO detection, and CSI estimation and feedback. Similarly, in [21], the authors highlighted the role of DL in communication, with a focus on DNNs, GANs, RNNs, and end-to-end learning. This work also explored deep unfolding for Bayesian optimal estimation, using MIMO detection and channel estimation as case studies. The authors in [22] focused primarily on MIMO detection while exploring its use in multiuser MIMO precoding and belief propagation-based channel decoding. A survey on model-based, DL, and hybrid approaches in the context of wireless communication networks is provided in [23]. In [16], Monga et al. conducted a survey of deep unfolding algorithms specifically implemented for applications in signal and image processing. The authors in [24] reviewed the implementation of deep unfolding in signal processing tasks. A brief review on deep unfolding in MIMO detection is presented in [17]. The authors in [13] have presented a general framework for optimization that leverages the benefit of deep unfolding with signal detection as one of the case studies. In [12], a chapter is dedicated to model-based machine learning for communication. The deep unfolding networks have been discussed for signal detection, channel estimation, and channel coding. Machine learning-based optimization framework is discussed in [25] with an emphasis on deep unfolding in specific design paradigms such as signal detection. A comparative study between purely data-driven and deep unfolding is conducted with complete emphasis on hybrid beamforming optimization in [26]. An overview of existing surveys on deep unfolding is shown in TABLE I.
I-B Contributions
Existing surveys lack coverage of recent studies addressing critical 6G technologies, such as mmwave spectrum, new air interface that includes massive MIMO, IRS, next generation multiple access, and are often limited to specific application areas within the physical layer. Consequently, a comprehensive survey is needed to encompass the broad applications of deep unfolding techniques across various aspects of the physical layer in wireless communication. This paper provides an in-depth review of deep unfolding methods applied to the physical layer of wireless communication. The contributions are summarized below.
  1. 1.
    We provide a comprehensive overview of deep unfolding algorithms and their applications in physical layer wireless communication. Part A of the paper introduces the fundamentals of DL and deep unfolding to establish a foundation, outlining their general structures. We describe the DL architectures, including MLP, CNN, RNN, DRL, GNN, and GAN. Additionally, we detail a systematic, step-by-step approach for developing deep unfolding algorithms. Finally, we highlight the advantages and motivations for adopting hybrid approaches to tackle optimization tasks effectively.
  2. 2.
    We survey the application of deep unfolding in the areas of signal detection, channel estimation, precoder design, sensing and communication, decoding, power allocation, and physical layer security in Part B. Within each section, we present these areas with the key 6G technologies (like IRS, mmWave, etc.) used to aid readers to design frameworks that adapt to future technologies. Additionally, this survey explores the application of deep unfolding in integrated sensing and communication (ISAC).Also, we discuss the joint techniques in the physical layer in Part C. The different areas are joint detection and channel estimation, joint activity detection and channel estimation, joint channel estimation and precoder design, and joint detection and decoding.
  3. 3.
    We outline a set of challenges in developing deep unfolding models and propose future directions that could guide advancements in deep unfolding techniques to enhance their performance.
TABLE I: Existing Survey Papers on Deep Unfolding for Communication.
SNo. Reference Year Detection Channel Estimation Precoder Design Sensing and Communication Decoding in Error Correcting Codes Power Allocation Security
1 Deep Learning for Wireless Physical Layer: Opportunities and Challenges [19] 2017 $\checkmark$ $\checkmark$ $\checkmark$
2 Model-Driven Deep Learning for Physical Layer Communications [20] 2019 $\checkmark$ $\checkmark$
3 Deep Learning in Physical Layer Communications [21] 2019 $\checkmark$ $\checkmark$
4 Deep Unfolding for Communications Systems: A Survey and Some New Directions [22] 2019 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
5 Wireless Networks Design in the Era of Deep Learning: Model-Based, AI-Based, or Both? [23] 2019 $\checkmark$ $\checkmark$
6 Algorithm Unrolling: Interpretable, Efficient Deep Learning for Signal and Image Processing [16] 2021 $\checkmark$ $\checkmark$
7 Intelligent Radio Signal Processing: A Survey [27] 2021 $\checkmark$ $\checkmark$ $\checkmark$
8 Redefining Wireless Communication for 6G: Signal Processing Meets Deep Learning With Deep Unfolding [24] 2021 $\checkmark$ $\checkmark$ $\checkmark$
9 Model-Based Deep Learning: Key Approaches and Design Guidelines [17] 2021 $\checkmark$
10 Model-Based Deep Learning: On the Intersection of Deep Learning and Optimization [13] 2022 $\checkmark$
11 Machine Learning and Wireless Communications [12] 2022 $\checkmark$ $\checkmark$ $\checkmark$
12 Machine Learning for Large-Scale Optimization in 6G Wireless Networks [25] 2023 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
13 Artificial Intelligence-Empowered Hybrid Multiple-Input/Multiple-Output Beamforming: Learning to Optimize for High-Throughput Scalable MIMO [26] 2024 $\checkmark$
14 This Survey $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
I-C Paper Outline
The skeleton of this paper is depicted in Fig.  1.
Figure 1: Outline of the paper.
This survey paper is organized as follows:
  • Part A: We discuss the general structure of DNN and deep unfolding models. Section II covers the architecture of DNN and outlines the steps in designing a DNN. Section III presents the deep unfolding framework, which combines classical iterative algorithms with DNNs.
  • Part B: The applications of deep unfolding in different physical layer signal processing tasks are thoroughly reviewed. Specifically, Section IV explains signal detection with perfect knowledge of channel gains in the receiver, followed by channel estimation in Section V. Section VI covers the precoder design. Section VII reviews different deep unfolding approaches for ISAC. The decoding for error correction codes is discussed in Section VIII, followed by power allocation in Section IX and security aspects in Section X.
  • Part C : Section XI surveys the joint techniques in the physical layer. In this section, we discuss the application of deep unfolding in different joint techniques such as joint detection and channel estimation, joint activity detection and channel estimation, joint channel estimation and precoder design, and joint detection and decoding.
  • Part D : Section XII discusses the challenges in adopting the deep unfolding approach in wireless communication and the future scopes of research. Section XIII concludes the paper.
TABLE II: Acronyms and Corresponding Full Names.
Abbreviation Description Abbreviation Description
AMP Approximate message passing ADMM Alternating direction method of multipliers
BP Belief propagation BS Base station
BER Bit error rate CIR Channel impulse response
CNN Convolutional neural network CSI Channel state information
DL Deep learning DNN Deep neural network
DRL Deep reinforcement learning EM Expectation maximization
EP Expectation propagation FDD Frequency division duplexing
FFT Fast Fourier transform FDMA Frequency division multiple access
GAN Generative adversarial network GNN Graph neural network
HetNets Heterogeneous networks IRS Intelligent reflecting surface
IoT Internet of things ICI Inter-carrier interference
ISI Inter-symbol interference ISAC Integrated sensing and communication
ISTA Iterative shrinkage-thresholding algorithm JCEHB Joint channel estimation hybrid beamforming
JCDD Joint channel estimation detection and decoding LLR Log-likelihood ratio
LDPC Low-density parity check LPBN Long term passive beamforming
LMMSE Linear minimum mean square error LR Low resolution
LS Least squares LMS Least mean squares
LSTM Long short term memory MAP Maximum a posteriori
MDP Markov decision process MF Matched filtering
MO Manifold optimization MM Majorization minimization
MS Mobile station MMSE Minimum mean square error
MIMO Multiple input multiple output ML Maximum likelihood
MPD Message passing detector MSE Mean square error
MSER Minimum symbol error rate mMTC Massive machine type communication
NN Neural network NOMA Non-orthogonal multiple access
OAMP Orthogonal approximate message passing OFDM Orthogonal frequency division multiplexing
OMP Orthogonal matching pursuit OTFS Orthogonal time frequency space
PGA Projected gradient ascent PGD Projected gradient descent
RF Radio frequency RNN Recurrent neural network
RL Reinforcement learning RSMA Rate splitting multiple access
SSCA Stochastic successive convex approximation SABN Short term active beamforming network
SBRNN Sliding bidirectional recurrent neural network SCDMA Sparse code division multiple access
SE Spectral efficiency SGD Stochastic gradient descent
SIC Successive interfernce cancellation SISO Single-input single-output
SIMO Single-input multiple-output SINR Signal-to-interference and noise ratio
SNR Signal-to-noise ratio TAS Transmit antenna selection
TDD Time division duplexing THz Tera Hertz
UAV Unmanned aerial vehicle URLLC Ultra reliable low latency communication
WMMSE Weighted minimum mean square error ZF Zero forcing
I-D Notation and Acronyms
[…]
II Deep Neural Network: General Structure
Artificial intelligence originated with the famous Turing test in the 1950s, marking a breakthrough in machine thinking capability [28]. Over the decades, around the 1990s, machine learning formed a subclass under artificial intelligence. Machine learning requires manual feature extraction of structured data. Most machine learning algorithms can be categorized into supervised, unsupervised, and reinforcement learning [4].
Figure 2: Feedforward NN (MLP) architecture.
Supervised learning aims to learn how to map inputs to outputs to predict labels for new data. In unsupervised learning, the dataset includes many features that are not labeled. The algorithm tries to identify hidden patterns within the data without the guidance of labeled examples. The third category, reinforcement learning, consists of an agent that learns to make decisions based on the environment. Since there is no predefined dataset, the agent interacts with the environment and learns from the feedback it receives for every action [4],[29]. The output of a machine learning algorithm can be of two types: regression and classification.
In the 2010s, DL emerged as a crucial technology in designing intelligent networks with unstructured data [30]. Researchers have adopted DL techniques to solve complex problems. It has been widely used in computer vision, natural language processing, and image processing. This success has inspired researchers to use DL in the physical layer of wireless communication, especially in various signal-processing tasks.
An NN consists of interconnected layers of neurons that work together to learn patterns from the given data. Different types of artificial NN architectures have been developed for various learning tasks, such as feed-forward NNs, CNN, recurrent NN (RNN), graph NN (GNN), etc. Fig. 2 shows the feedforward NN architecture with $K$ ($K=2$ here) hidden layers and $N_{k}$ neurons in layer $k$, for all $k=0,\dots,K$. The architecture has an input layer, hidden layers, and an output layer.
We now discuss the steps in a feedforward NN:
  1. 1.
    The network begins with an input vector
    ${\bf{s}}=[s_{1},s_{2},\dots,s_{n}]$, at the input layer where $n$ is the number of input features.
  2. 2.
    For each neuron in a layer, the weighted sum of inputs is computed:
    $z^{(k)}_{j}=\sum_{i=1}^{n}w_{ji}^{(k)}s_{j}+b_{j}^{(k)},$
    where $w_{ji}^{(k)}$ is the weight connecting neuron $i$ to neuron $j$ in the $k^{\text{th}}$ layer and $b_{j}^{(k)}$ is the bias for neuron $j$ at the $k^{\text{th}}$ layer.
  3. 3.
    A non-linear activation function $\sigma(\cdot)$ is applied:
    $a_{j}^{(k)}=\sigma(z^{(k)}_{j}),$
    where $a_{j}^{(k)}$ is the activation output of the neuron $j$ in the layer $k$. Common activation functions include ReLU, sigmoid, tanh, softmax etc. [4].
  4. 4.
    At the output layer, after applying the weighted sum and activation process, the output is given by:
    $\hat{\bf{y}}=\sigma(\mathbf{W}^{K}\mathbf{s}^{K}+\mathbf{b}^{K})$
    where ${\bf{W}}$ is the weight matrix at the output layer and $\bf{b}$ is the bias vector at the output layer. The activation function, ($\sigma(\cdot)$) for the output often depends on the task, whether it is a classification or regression task.
  5. 5.
    The loss is computed to measure the network’s performance:
    $\mathcal{L}(\hat{\bf{y}},\bf{y})$
    where $\hat{\bf{y}}$ is the predicted output, and $\bf{y}$ is the true output. The common loss functions are mean squared error (MSE), binary cross entropy, categorical cross entropy, etc.
To train the NN, backpropagation is used by adjusting the parameters of the network, i.e., the weight and biases to minimize the loss. The parameters are updated using an optimization algorithm like stochastic gradient descent (SGD).
On the other hand, CNN is designed to process grid-like data by leveraging spatial structure. It uses layers of convolutional filters to extract features like edges followed by pooling layers to reduce spatial dimensions and retain key information [31]. The extracted features are passed through fully connected layers. One of the early applications of CNN was modulation classification based on signal constellation diagrams; the authors have used the CNN-based DL method [32]. They have developed and analyzed methods to represent modulated signals in data formats with grid-like topologies to conform with the CNN architecture.
RNN is another type of NN designed explicitly for processing sequential data. In RNN, the output from the previous step is fed as an input to the current step, incorporating sequential dependencies between the nodes. A recurrent neuron performs this function by maintaining a hidden state. Long short-term memory (LSTM) is one version of RNN [4]. The LSTM network was adopted into the NOMA framework in [33] to learn the channel characteristics of the NOMA system.
DRL integrates RL with DL to develop intelligent agents capable of making sequential decisions in complex environments. In DRL, an agent interacts with its environment to maximize cumulative rewards, learning optimal policies through exploration and exploitation. DNNs are employed to approximate value functions or policies, allowing DRL to efficiently handle high-dimensional state and action spaces. Key techniques such as Deep Q-Networks (DQN), policy gradient methods, and actor-critic frameworks enable agents to improve their performance iteratively through trial and error, guided by reward signals [4, 34]. In the domain of communications and networking, DRL has emerged as a powerful tool for addressing challenges such as resource allocation, spectrum management, and network optimization. Modern network architectures, including the Internet of Things (IoT), heterogeneous networks (HetNets), and unmanned aerial vehicle (UAV) networks, exhibit decentralized, dynamic, and autonomous characteristics, making them well-suited for reinforcement learning-based solutions [35].
The concept of using autoencoders to design end-to-end communication systems has been widely studied [36]. End-to-end learning trains a single model to map raw input directly to the desired output. It avoids intermediate steps or manual feature engineering. Autoencoders are neural networks that optimize both the transmitter and receiver together in this framework. In communication systems, the goal is to design the transmitter and receiver jointly for a given channel model. This helps achieve better performance, such as faster convergence and lower bit error rates [37].
GNNs are specialized networks designed for graph-structured data, where nodes are connected by edges [38]. Unlike traditional neural networks, GNNs work on non-Euclidean data and capture relationships between graph nodes. They learn rich representations of nodes, edges, or the whole graph by aggregating and transforming features from neighboring nodes. This ability to use graph structure helps GNNs perform better in tasks that rely on relational data.
In 2020, Goodfellow et al. proposed a new AI algorithm for generative modeling [39]. GANs use two neural networks: a generator and a discriminator. The generator creates realistic outputs, while the discriminator identifies if they are real or fake. By competing and learning together, GANs can produce high-quality outputs that are often indistinguishable from real data.
For further reading on application of deep learning in physical layer of communication, we invite readers to refer [40, 41, 42, 27].
III Deep Unfolding: General Structure
Deep unfolding, also known as algorithm unrolling [16], is a typical model-driven DL method [10],[20],[43, 44, 45, 46, 47, 48]. It has been frequently considered for various tasks such as signal detection, design of beamforming matrices, and many more optimization tasks in wireless communication. Owing to reduced training time and dataset, deep unfolding is one promising alternate approach to purely model-driven or data-driven.
Figure 3: Schematics of data-driven, classical and deep-unfolding approaches.
Deep unfolding integrates model-based approaches and domain knowledge into the design of DNN architectures. The training is carried out in an end-to-end manner. In deep unfolding, an iterative algorithm is mapped to DNN layers. Then, DL techniques such as SGD are used to learn the parameters. Fig. 3 shows a schematic diagram comparing the classical approach, data-driven approach, and the deep unfolding approach. Furthermore, Fig. 4 depicts the deep unfolding approach in greater details.
Figure 4: General framework of deep unfolding.
For any iterative algorithm, the update equation for each iteration is given as:
${\bf{s}}^{(t+1)}=h({\bf{s}}^{t};{\bf{\Psi}}^{t})$
where ${\bf{s}}^{t}\in{\mathbb{R}}^{n}$ denotes the variable vector to be updated in each iteration in the range $t=1,\dots,T$. The iterative function for the iterative algorithm is represented by $h(\cdot;\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. The trainable parameters, ${\bf{\Psi}}^{t}\in\mathbb{R}^{n}$ include the model parameters and the step size for the regularization coefficients of the algorithm [25]. Each iteration function $h(\cdot;\cdot)$ is mapped to a single network layer, and a finite number of layers are stacked together as shown in Fig. 4. To learn the trainable parameters ${\bf{\Psi}}^{t}$ in an end-to-end approach, the following optimization is carried out:
$\text{minimize}_{{\bf{\Psi}}_{T}}\mathcal{L}({\bf{s}}^{T}({\bf{\Psi}}_{T}))$
where $\mathcal{L}(\cdot)$ denotes the loss function for training, ${\bf{\Psi}}_{T}=\{\Psi^{t}\}_{t=0}^{T}$ is the set of trainable parameters of the total $T$-layer network and ${\bf{s}}^{T}(\cdot)$ is the output of the unfolded network. Feeding the data forward through the unfolded network with trainable parameters in the testing stage is similar to executing the parameter-optimized iterative algorithm for a finite number of iterations. Thus, the number of layers in deep unfolding models is fixed, resulting in a predefined computational complexity. Within these fixed layers, the model is trained to optimize the trainable parameters, enabling the problem to be addressed using a structured algorithm with parameters tailored for optimal performance.
The procedure to solve an optimization problem using a deep unfolding approach is outlined below.
  1. 1.
    Identify a principle algorithm that serves as the foundation for deep unfolding.
  2. 2.
    Select trainable parameters or specific blocks in the principle algorithm. Replace these with data-trained operations optimized through deep unfolding.
  3. 3.
    Define the loss function and represent the algorithm as a DNN. Each layer corresponds to an iteration of the original algorithm. Train the network using supervised or unsupervised methods. Since the number of iterations is predetermined, the network’s layers match these iterations. The number of layers can also be fine-tuned to balance performance and computational complexity.
TABLE III: Key technologies envisioned for 6G using deep unfolding.
Key technology Task Reference
Massive MIMO Detection [49],[50], [51], [52], [53], [44], [54], [55], [56], [57], [58], [59], [60], [61]
Channel estimation [62], [63], [64], [65], [66], [67]
Precoder Design [68], [69], [70], [71], [72], [73], [74], [75], [76], [77], [78], [79], [80], [81]
ISAC [82], [83], [84]
mmWave Channel Estimation [85]
Precoder design [86], [87], [88]
ISAC [89]
Modern Multicarrier Modulation OFDM Detection [90],[91], [92]
OTFS Detection [46]
OFDM Channel Estimation [93], [94], [95], [45], [96], [97], [98], [99], [47] [100]
OTFS Channel Estimation [101]
OFDM-ISAC [102], [103]
OTFS-ISAC [104]
Next Generation Multiple Access Detection [105], [106]
Precoder design [107]
Orbital Angular Momentum Detection [108]
IRS Channel Estimation [109], [110], [111], [112]
Precoder design [113], [114], [115], [116]
ISAC [117]
THz Precoder design [118], [119], [120]
Principle algorithms are typically iterative and based on optimization techniques. However, they are often complex, computationally demanding, and slow to converge. Deep unfolding improves these algorithms by leveraging data training. The key contributions of deep unfolding are:
  1. 1.
    Optimizing hyperparameters to enhance convergence, leading to better objective values or faster processing [26, 76, 84, 11].
  2. 2.
    Simplifying complex tasks, such as matrix inversion or decomposition, with lower computational complexity [121, 70, 122].
The first application is common in unfolded projected gradient ascent/descent (PGA/PGD) methods, while the second is widely used in unfolded weighted minimum mean square error (WMMSE) methods. Another approach involves learning the output at each layer, which corresponds to the iterative solution of the principal algorithm. This technique is widely applied in detection networks [123, 118, 49], following the PGA/PGD framework.
The main advantage of deep unfolding is its ability to incorporate domain knowledge into the network architecture. This improves performance while reducing complexity by using fewer parameters and a fixed number of layers. As a result, deep unfolding requires a smaller training dataset compared to generic DNNs. Next, we provide a comprehensive survey of deep unfolding algorithms designed for various signal processing tasks in wireless communication. Additionally, TABLE III lists key 6G technologies relevant to deep unfolding-based signal processing.
IV Signal Detection
This section discusses the application of deep unfolding techniques for signal detection. We segregate these signal detection techniques into various key technologies, as described in the following.
IV-A Massive MIMO
MIMO transmission uses multiple antennas at both the transmitter and receiver. It improves spectral and energy efficiency by utilizing a larger spatial domain. Adding more antennas (tens or hundreds) introduces the concept of massive MIMO [3]. However, detecting multiple symbols becomes more complex due to large-scale inter-symbol interference. Designing a low-complexity, high-performance detector for massive MIMO remains a challenge. For a MIMO communication system with $n_{\text{t}}$ transmit and $n_{\text{r}}$ receive antennas, the received signal can be given as:
$\mathbf{y}=\mathbf{Hx}+\mathbf{n}$ (1)
where $\mathbf{x}\in\mathbb{C}^{n_{\text{t}}}$ is the transmitted baseband signal vector, $\mathbf{H}\in\mathbb{C}^{n_{\text{r}}\times n_{\text{t}}}$ is the channel matrix and $\mathbf{y}\in\mathbb{C}^{n_{\text{r}}}$ is the received baseband signal vector. The detector extracts the desired transmitted information $\mathbf{x}$ from the received signal vector $\mathbf{y}$ with given $\mathbf{H}$ and the Gaussian noise vector $\mathbf{n}$.
The maximum a posteriori (MAP) and maximum likelihood (ML) estimations are Bayesian-based detection methods. These are the most commonly used classical detection techniques. The MAP rule applies when the transmitted signals have unequal a priori probabilities. The ML rule is used when these probabilities are equal [124]. ML minimizes the average error probability, providing an optimal solution. The ML detection rule is given as:
$\mathbf{\hat{x}_{\text{ML}}}=\text{arg}\min_{\mathbf{x}\in\mathcal{A}^{n_{t}}}||\mathbf{y}-\mathbf{Hx}||^{2}.$ (2)
This rule minimizes the Euclidean distance between $\bf{y}$ and $\bf{Hx}$. The minimization is performed over all possible combinations of the transmitted symbol vector $\bf{x}$ from the modulation alphabet $\mathcal{A}$. Since ML performs an exhaustive search and examines all possible signals, the computational complexity increases exponentially as $n_{\text{t}}$ increases.
MIMO detection can be linear or non-linear. Common linear detectors include matched filter (MF), zero-forcing (ZF), and minimum mean squared error (MMSE) [125]. These detectors require matrix inversion. Lattice reduction techniques can improve the performance of ZF and MMSE. Non-linear detectors, such as those using successive interference cancellation (SIC), perform interference estimation and removal in multiple stages [3], [126]. SIC and parallel interference cancellation are well-known techniques for this. In SIC, a detector like MF, ZF, or MMSE is used to detect the transmitted symbol in a data stream, leading to MF-SIC, ZF-SIC, or MMSE-SIC schemes. The data streams are arranged based on the received SNR. After detecting the symbol and estimating the channel coefficients, the interference caused by the detected symbol is estimated. This interference is then removed from the received signal, which is used to detect the next data stream in terms of received SNR. This process repeats until all symbols are detected. Sphere decoding is another non-linear detector [127]. It is an approximate ML detection method, where detection happens in a subspace of the modulation alphabet. The method searches within a bounded hypersphere around the received lattice point, using a radius for iterative symbol searching. The logarithmic likelihood ratio (LLR) is computed based on the symbol candidate found. This search region is smaller than the exhaustive ML search [128, 129].
The channel matrix for a MIMO fading channel can be modeled as a graph. The MIMO detection problem is framed as an inference task to calculate the posterior probability of system variables. Low-complexity algorithms are designed to compute these probabilities efficiently. The graph model combines probability and graph theory, showing the relationships between different nodes. Common graphical models include Bayesian belief networks, Markov random fields, and factor graphs [3]. Message-passing algorithms, like belief propagation (BP), are crucial for representing the MIMO system’s graph. Probability messages are exchanged between nodes representing transmitted symbols and received signals along the graph edges. The system is often modeled as a factor graph, where performance depends on the graph structure and algorithm convergence. Iterative algorithms, like message-passing detectors (MPD) with BP, approximate message passing (AMP), and channel hardening-exploiting message passing, provide solutions with lower complexity. Expectation propagation (EP) extends BP to handle larger constellations and higher-dimensional MIMO systems. EP approximates the posterior distribution of transmitted symbols iteratively, enabling efficient detection in polynomial time [130, 131].
The conventional detectors (linear and non-linear) involve computationally intensive matrix inversions or factorizations (like singular value decomposition or QR decomposition). Although these detectors yield nearly optimal solutions, they come at the cost of high complexity and high latency. Hence, with the aid of DL techniques, a deep unfolding approach has been adopted in literature to achieve low complexity and low latency.
PGD is an iterative optimization algorithm. Each step includes a gradient descent update to minimize a cost function, followed by a projection step to enforce constraints like sparsity or feasibility. In MIMO detection, PGD minimizes the error between the predicted and actual received signals while projecting the solution onto a feasible set, such as valid constellation points [132]. To achieve ML optimization, the PGD algorithm is unfolded into the DetNet architecture for MIMO detection [49, 133]. The gradient step size is learned across layers. The authors claimed that this approach reduces computational complexity compared to the sphere decoder and is robust to ill-conditioned channels. By leveraging the benefits of graphical models, He et al. proposed model-driven DL networks, OAMP-Net [10] and OAMP-Net2 [11], for MIMO detection with imperfect channel state information (CSI). OAMP-Net2 outperforms OAMP-Net by adapting to varying channel conditions. The orthogonal AMP (OAMP) algorithm is unfolded, and the LMMSE channel estimator is used to design the receiver. Learnable parameters, such as the step size for updating the prior mean and variance in the MMSE estimator, are trained using SGD. These parameters in OAMP-Net2 help compensate for channel estimation errors, improving performance compared to DetNet [49]. In [134], the authors introduced MMNet, a deep unfolded detector designed for a wide range of realistic channel matrices, including ill-conditioned and non-i.i.d. scenarios. MMNet is based on an iterative soft-thresholding algorithm, alternating between a linear detector and a non-linear denoising step. It balances architectural flexibility and computational complexity, outperforming DetNet [49] and OAMP-Net2 [11]. Overloaded massive MIMO systems, where the number of transmit antennas exceeds the number of receive antennas, require detectors with low complexity and good performance. In [61], the trainable projected gradient (TPG) algorithm was introduced for these systems. TPG uses a DL-aided iterative PGD method with trainable parameters, such as step size, optimized via backpropagation and SGD. TPG achieves lower computational cost and fewer trainable parameters compared to DetNet [49], OAMP-Net2 [11], and others. Additionally, a trainable detector based on the Hubbard–Stratonovich transformation is discussed in [135]. Takabe et al. avoided matrix inversion in this approach. Compared to the TPG detector [61] and OAMP-Net2 [10], the proposed network achieves faster training and lower execution cost.
In [136], the concrete MAP method is introduced, based on a continuous relaxation of the probability mass function for discrete random variables, as derived in [137]. This approach solves the MAP detection problem using continuous optimization instead of exhaustive search, approximating the MAP solution for any probabilistic nonlinear model. By unfolding the gradient descent algorithm, the detector provides soft outputs (detection probabilities), enabling better decoding compared to DetNet [49], OAMP-Net2 [10], and MMNet [134]. Yun et al. in [138] enhanced the unfolded PGD design by adding momentum components to the iterative hidden layers. Momentum helps guide parameter updates in the right direction and reduces oscillations during PGD optimization [139]. This modification improves convergence speed and overall performance, outperforming DetNet [49] and OAMP-Net2 [10, 11], while maintaining generality.
In wireless communication, physical layer signal processing involves continuous-time analog waveforms, which must be quantized for hardware implementation. Low-resolution one-bit analog-to-digital converters are commonly used in massive MIMO systems due to their high sampling rates, low cost, and low energy consumption [140]. To recover high-dimensional signals from noisy, low-resolution measurements, Khobahi et al. [141] developed a low-resolution detection network. This network extracts signals from noisy one-bit measurements without requiring prior knowledge of the channel matrix. It is based on unfolding the PGD method for symbol detection using the ML estimator, eliminating the need for CSI. The network outperforms traditional ML estimators [142] and DL networks like DeepSIC [143], even with small datasets of about 500 samples. In [59], a generalized PGD algorithm was proposed, eliminating the need for separate loss functions for projection and gradient steps. Instead of alternating after each step, projection is applied after several gradient steps, achieving a better balance. By unfolding this approach, the authors created a self-correcting auto detector that enhances performance [59].
Homotopy optimization transforms the original problem into a different one using a homotopy map. This method solves a series of simpler problems step by step, gradually progressing to the harder ones [144]. In MIMO detection, non-convex algorithms often struggle with the ML detection problem due to bad local minima [50]. By using a homotopy map to transition from easier to harder problems, detection performance is significantly improved. The study in [50] focuses on ML MIMO detection with one-bit quantized observations and binary symbol constellations. Using homotopy optimization, the authors apply a non-convex continuous reformulation of binary optimization [145] to address the detection problem. Their algorithm is designed to run efficiently and adapt to different channels. Building on this work, the authors introduced a proximal gradient-based homotopy method in [53]. This approach connects the structure of DetNet [49] with homotopy optimization for improved MIMO detection.
GNNs, as structured DNNs, excel in adaptability and flexibility. In MIMO detection, the graph structure of wireless networks makes GNNs particularly effective. Tan et al. [55] unfold MPD iterations to update messages at symbol nodes using prior and posterior LLR. Learnable parameters, such as prior probabilities and correction factors (damping, rescaling, offset), are optimized via mini-batch SGD to minimize the loss function. A greedy search determines the number of layers to balance performance and complexity. In [60], a low-complexity AMP-GNN model is proposed for massive MIMO detection. It outperforms OAMP-Net2 [10] by avoiding matrix inversion. For single-carrier FDMA systems, common in 5G to reduce peak-to-average power ratios, message-passing algorithms are computationally demanding. To address this, the unfolded network in [146] simplifies generalized AMP iterations [147], maintaining high performance even with channel mismatches. To enhance EP-based detection, [51] optimizes parameter selection using a modified moment-matching scheme, showing superior performance in high modulation cases compared to OAMP-Net2. In [52], approximate EP from [148] is unfolded, achieving better detection performance and robust convergence.
Liao et al. [44] incorporated interference cancellation into DNN layers using iterative embedding. Inspired by [149], their sequential detector trains residual error coefficients with MSE loss, reducing complexity by avoiding matrix inversion and relying on matrix addition and multiplication. Compared to DetNet [49, 133], which requires 90 layers for convergence, their model achieves convergence with just 8 layers, offering significant computational simplicity. Massive MIMO systems improve spectral efficiency (SE), coverage, and range but demand higher computational power. LcgNet, a model-driven DL architecture for massive MIMO detection, unfolds conjugate gradient descent iterations and learns step sizes through offline training [54]. Its quantized version, QLcgNet, minimizes memory usage with negligible performance degradation. LcgNet outperforms DetNet in terms of symbol error rate, computational complexity, and memory cost, making it a robust solution for massive MIMO systems.
The quantized version, QLcgNetV, reduces memory requirements while maintaining high detection performance. To enhance LcgNet, Olutayo et al. [150] incorporated a preconditioner during training, designing a preconditioned LcgNet. Preconditioning improves the condition number of a matrix [151], optimizing the receiver filter’s spectrum and minimizing eigenvalue spread. This enhancement significantly boosts the convergence speed compared to the original LcgNet [54]. Additionally, the dynamic conjugate gradient network utilizes self-supervised learning with forward error correction codes, adapting effectively to non-stationary, time-varying channels.
Sphere decoding achieves ML optimality but suffers from high computational complexity in massive MIMO systems. To address this, Nguyen et al. [123] proposed a fast DL-aided sphere decoding algorithm inspired by [152]. The DNN generates reliable initial candidates for sphere decoding, prioritizing candidates closer to the DNN’s output for testing. This approach accelerates sphere shrinkage, reducing computational cost. The sparse network architecture eliminates matrix inversion, relying only on element-wise matrix multiplications. The proposed method performs better in highly correlated channels than DetNet [49, 133], OAMP-Net2 [10, 11], MMNet [134], and LcgNet [54]. A MIMO detector based on annealed Langevin dynamics is introduced in [57]. It replaces the traditional two-step iterative framework (continuous-space optimization followed by projection) with a single-step approximation of the MAP estimator by sampling from Langevin dynamics. This approach shows superior performance in correlated channels compared to DetNet [49], OAMP-Net2 [11], and MMNet [134]. To mitigate the computational costs of matrix inversion in classical detectors like ZF and MMSE, Berra et al. [58] proposed unfolding accelerated Chebyshev successive over-relaxation and over-relaxation algorithms. This network requires fewer trainable parameters and features a fast, stable training scheme. The unfolded detector achieves better performance compared to DetNet [133], OAMP-Net2 [10], TPG-detector [61], MEPD [51], EP-Net [46], and other methods like [53, 44].
To avoid the computational burden incurred due to matrix inversion, the authors in [56] have proposed a sparse refinement architecture for symbol detection in uplink massive MIMO. The authors have used deep unfolding for various crucial components in matrix inversion, significantly reducing the detection complexity. For detection, the alternating direction method of multipliers (ADMM) algorithm is utilized in [153]. ADMM decomposes a complex optimization problem into simpler subproblems, which are solved iteratively. The algorithm combines the principles of dual ascent and the method of multipliers. Each iteration involves solving the subproblems for the primal variables in an alternating manner, followed by updating the dual variables (Lagrange multipliers) to enforce consistency with the original problem’s constraints [154]. This iterative process ensures that the subproblem solutions converge to the global solution of the main optimization problem. The authors in [155] developed an unfolded ADMM-based network for MIMO detection. They employed an inexact ADMM update [154, 156] to reduce complexity while achieving near-optimal performance. Compared to DetNet [49, 133], the proposed network requires fewer trainable parameters, resulting in improved performance in terms of symbol error.
IV-B Modern Multicarrier Modulation
An OFDM-based MIMO detector faces many challenges. The increased number of antennas at the transceiver exhibits enhanced spatial multiplexing but at the cost of increased inter-symbol interference (ISI). In standard OFDM multicarrier modulation, the cyclic prefix is inserted between OFDM blocks to mitigate ISI. Zhou et al. in [90], have unfolded a conjugate gradient method based OAMP detector for the OFDM-MIMO system. To address the issues faced in detecting coarsely quantized or low-resolution signals in OFDM-based MIMO systems, Shao et al. [91] have unfolded the expectation maximization (EM) algorithm. By unfolding the likelihood ascent search algorithm, Ullah et al. [92] are able to jointly detect and estimate the soft output in MIMO OFDM systems.
Li et al. [46] proposed a deep unfolding network for OTFS detection by converting the iterative EP algorithm into structured network layers. The core idea is to approximate the conditional probability distributions $p(\bf{y}|\bf{x},\bf{H})$ and $p(\bf{x})$ where $p(\cdot)$ represents the probability, and $\bf{y},\bf{x},\bf{H}$ denote the received signal vector, transmitted signal vector, and channel matrix, respectively. The trainable parameters in this network include the step size, damping factor, and offset values for the mean ($m$) and variance ($v$) of the probability distribution. During offline training, the number of layers in the unfolding network is selected to balance performance and computational complexity. To enhance convergence and prevent the network from being trapped in local optima, the Adam optimizer with warm restarts is employed. This technique adaptively adjusts the step sizes, improving both training stability and performance. In [101], an unfolded OAMP-based detector for OTFS is introduced to handle channel estimation errors. This detector is trained using imperfect CSI and optimizes three trainable parameters: step sizes for updating the linear estimate, nonlinear estimate, and noise variance.
IV-C Next Generation Multiple Access
A cellular network must simultaneously support a growing number of active subscribers within a limited time-frequency spectrum [157]. The current trend in wireless communication is to increase the signal bandwidth to achieve higher transmission rates. However, increasing bandwidth results in shorter symbol periods due to the inverse relationship between bandwidth and symbol period. Consequently, in a multi-path fading channel, the reduced symbol period exacerbates ISI. Multi-carrier modulation techniques like OFDM provide an effective solution to mitigate ISI by dividing a wideband signal into a set of orthogonal narrowband subcarriers [158]. This approach not only minimizes interference but also enhances robustness in multipath environments.
Multiple access techniques allow multiple users to share the same communication channel and resources, such as time, frequency, or code, to transmit data efficiently. Traditional multiple access schemes achieve this by dividing signaling dimensions orthogonally across time, frequency, or code [159]. This orthogonality enables low-complexity multi-user detection at the receiver. However, orthogonal multiple access cannot achieve the desired sum-rate capacity for a multiuser wireless system. To address this, NOMA has been introduced, allowing more users to be served than the number of orthogonal resource units by leveraging non-orthogonal resource sharing [157]. NOMA is categorized into power-domain NOMA and code-domain NOMA [160]. In power-domain NOMA, the transmitter superimposes information symbols into a single waveform using superposition coding. At the receiver, SIC is employed to iteratively decode signals until the desired signal is recovered.
A deep unfolded detector, named the learned preconditioned conjugate gradient descent network with SIC, is proposed for downlink MIMO-NOMA systems in [105]. To achieve a robust model with low dataset requirements, the algorithm incorporates a data transfer learning strategy. This approach adapts the model efficiently to variations in environmental factors, such as channel models, modulation schemes, or power allocation, ensuring consistent performance across diverse scenarios.
In sparse spread CDMA, each user is assigned spreading codes with only a few non-zero elements. This sparsity minimizes overlap and enables multiple users to transmit simultaneously on the same resources with minimal interference. In sparse CDMA, the spreading matrix contains sparse non-zero elements, as initially studied in [161]. Each user’s transmitted symbols are encoded using unique sparse signature sequences, which facilitate decoding at the BS [162]. Takabe et al. proposed a complex sparse trainable projected gradient detector (C-STPG) for SCDMA networks in [106]. The detector builds upon the PGD framework, delivering significantly improved detection performance compared to conventional LMMSE detectors. Simulation results demonstrate reduced computational complexity during execution and minimal training costs due to the limited number of trainable parameters. The authors also introduced gradual sparsification—a joint learning technique that designs sparse signature sequences using deep unfolding. This technique optimizes the weights of the signature matrix while reducing error during training through MSE loss. The trainable parameters include the step size for gradient descent ($\gamma$) and a softness parameter ($\alpha$) to adjust the strength of the soft projection.
IV-D Orbital Angular Momentum
Orbital angular momentum (OAM) MIMO multiplexing leverages electromagnetic waves with different orthogonal OAM modes [163]. These modes, theoretically infinite in number, form an orthogonal basis that can be used to multiplex distinct signals. A comprehensive review of OAM wave generation and detection is presented in [153]. In the context of OAM MIMO systems, Kamiya et al. [108] developed a signal detection algorithm specifically designed for systems employing uniform circular array antennas. These systems often face inter-mode interference, particularly when the transmitter and receiver antennas are misaligned, which degrades both MMSE capacity and detection performance. To address this issue, the ADMM algorithm is unfolded with parallel interference cancellation. This approach significantly enhances detection performance, even in scenarios where alignment imperfections introduce severe inter-mode interference.
In addition to detection, managing the peak-to-average power ratio (PAPR) presents a significant challenge in OAM multiplexing systems. To address this, Kamiya et al. [164] proposed a deliberate amplitude clipping method for OAM systems employing uniform circular array antennas. This approach involves applying amplitude clipping to digital baseband signals before OAM beamforming and pulse-shaping filtering. By doing so, it effectively reduces peak power, minimizes out-of-band radiation, and preserves the orthogonality of OAM modes. To counteract the bit error rate (BER) degradation introduced by clipping, the authors developed an iterative distortion recovery algorithm. Drawing inspiration from clipping noise cancellation techniques used in OFDM systems, the algorithm is implemented using deep unfolding. This implementation allows for trainable parameters that adaptively mitigate clipping-induced distortions, leading to enhanced detection performance and improved robustness in practical OAM systems.
V Channel Estimation
Estimating the channel in massive MIMO systems presents significant challenges, primarily because of the large number of antennas at the transceiver. The complexity of estimation grows with the scale of the system, and noise further exacerbates the issue, resulting in imperfect channel estimates and rendering estimation errors unavoidable [3], [165], [166], [167]. Pilot-based channel estimation techniques are commonly used to address these challenges. These techniques involve transmitting known pilot symbols, followed by processing at the receiver to estimate the channel’s characteristics and compensate for its effects. A block diagram illustrating the sequence of operations in a pilot-based channel estimation technique is provided in Fig. 5. It typically includes the generation and transmission of pilot symbols, reception at the receiver, estimation of the channel matrix using algorithms such as least squares (LS) or MMSE, and application of the estimated channel to process the received signals.
Figure 5: Block diagram for pilot-based channel estimation.
In the following, we discuss deep unfolding aided channel estimation methods.
V-A Massive MIMO
MMSE-based channel estimation is commonly used in MIMO systems to estimate the channel response in noisy environments. It minimizes the mean squared error between the actual channel and its estimate using the second-order statistics of the channel and noise. In massive MIMO systems with FDD, reducing CSI feedback overhead is challenging. A method in [62] combines deep learning and superposition coding to address this. Downlink CSI is encoded, superimposed on uplink user data, and sent to the base station. A multi-task neural network unfolds the MMSE criterion to reduce interference and reconstruct the CSI efficiently.
The iterative shrinkage-thresholding algorithm (ISTA) is commonly used for sparse channel estimation, enabling efficient CSI feedback. Each ISTA iteration includes a linear operation followed by a nonlinear soft-thresholding step, similar to the ReLU activation function [16]. In [63], Chen et al. proposed a trainable deep learning architecture for sparse recovery problems using an adaptive depth approach. Instead of a fixed number of layers, the method dynamically adjusts the depth based on the problem’s complexity. Highly sparse channels use fewer layers, while more complex channels require additional layers. This adaptive approach improves accuracy and lowers computational costs compared to traditional methods and fixed-depth networks. Results in [63] demonstrate that the method reduces reconstruction error and speeds up channel estimation, making it more efficient.
A scenario-adaptive CSI compression framework is proposed in [168], requiring minimal channel measurements for training. The authors introduced a CSI feedback architecture designed for efficient encoder updates in dynamic wireless environments, which is essential for downlink precoding in massive MIMO systems. The encoding network focuses on improving the robustness and efficiency of CSI compression. It employs a deep unfolding-based feedback network to iteratively refine the feedback process and enhance feature extraction. Similarly, the decoding network uses a deep unfolding structure inspired by the ISTA algorithm. This iterative approach improves CSI recovery accuracy, making the process more efficient and adaptable to changing conditions.
In massive MIMO systems, where the channel matrix is often sparse, the majorization-minimization (MM) approach is effective for addressing sparse channel estimation [169]. In [64], the authors address the critical challenge of downlink channel estimation in massive MIMO, essential for efficient communication. Traditional methods often rely on angular-domain channel sparsity to reduce pilot overhead but struggle in dynamic environments. To overcome this, the authors propose an online two-timescale joint uplink/downlink dictionary learning and channel estimation algorithm. This method leverages channel sparsity and angle reciprocity between uplink and downlink transmissions to adapt to dynamic scenarios and significantly reduce pilot overhead. The two-timescale strategy includes learning a long-term dictionary to capture spatial channel characteristics and frequently estimating short-term sparse channel vectors. The short-term channel estimation subproblems are solved using an unfolded MM algorithm, while the long-term dictionary optimization uses standard convex methods. Deep unfolding is employed to jointly optimize the dictionary learning and channel estimation process for both uplink and downlink. Simulations in [64] show that this approach reduces pilot overhead and outperforms existing methods in static and dynamic environments, especially in scenarios with mobile users.
The matching pursuit algorithm is a greedy iterative method often applied in sparse channel scenarios like massive MIMO. It identifies channel components (basis functions) by matching them to the residual, which is the difference between the observed signal and the signal generated by the current channel estimate. Yassine et al. [65] introduced a deep unfolded algorithm, called mpNet, based on the classical matching pursuit algorithm. This network estimates the multiuser massive MIMO channel in an unsupervised manner, initializing with a dictionary of imperfect steering vectors. The system operates in time division duplex (TDD) mode to leverage channel reciprocity. The network dynamically adjusts the number of layers based on incoming data, ensuring efficient performance in realistic environments.
The basis pursuit algorithm is a widely used method for sparse signal recovery, particularly in compressed sensing. It identifies the sparsest solution to an underdetermined system by minimizing the $\ell_{1}$-norm of the signal. This approach is crucial in scenarios where data is acquired under constraints, and the signal is assumed to be sparse in certain domains. In [66], the authors observed that existing sparse channel estimation schemes for downlink CSI often deliver suboptimal reconstruction due to the use of random measurement matrices. To address this, they proposed deep unfolding-based autoencoders that optimize the measurement matrix by unfolding the basis pursuit algorithm. This customized design enhances reconstruction performance in sparse channel estimation tasks.
Sparse Bayesian learning is another effective approach for channel recovery [170]. It estimates the non-zero elements of the channel matrix while utilizing prior knowledge of sparsity. Hu et al. [67] introduced a deep unfolding framework called deep deterministic policy gradient (DDPG) for multi-user MIMO systems. DDPG, a variant of deep Q-networks, is designed to handle problems in continuous action spaces [171]. Unlike traditional deep unfolding neural networks with fixed depth, the authors leveraged DDPG to dynamically adjust the network depth based on the complexity of the problem. This adaptive design enhances the efficiency and performance of sparse channel recovery in varying scenarios. The channel estimation problem is formulated using an off-grid sparse Bayesian learning model. The sparse Bayesian learning algorithm is unfolded layer by layer, introducing trainable parameters such as grid points and off-grid variables. Simulation results show that the proposed method significantly outperforms fixed-depth models, achieving a 30% reduction in the number of layers while improving normalized mean square error performance.
V-B mmWave Massive MIMO
mmWave massive MIMO operates in the high-frequency range of 30 GHz to 300 GHz, allowing higher dimensionality and the arrangement of many antennas in a compact space due to compressed wavelengths. The BS, equipped with numerous antennas, must acquire downlink CSI. However, the large amount of feedback required creates significant overhead, posing a challenge in mmWave massive MIMO systems. Compressed sensing techniques are widely used to reduce feedback overhead by exploiting channel sparsity [172]. The limited scattering properties of mmWave propagation and the power-focusing capability of lens antenna arrays transform the spatial domain channel into a sparse beamspace representation. Consequently, the channel estimation problem becomes a sparse signal recovery task, often addressed through compressed sensing methods. However, performing beamspace channel estimation in massive MIMO systems with limited RF chains remains a challenging task.
Compressive sensing aided deep unfolding is explored in [173], where the authors apply Onsager correction for AMP [174] and construct an unfolded model similar to [18]. The network in [175] uses deep unfolding to learn beamspace channel estimates. The authors demonstrate that this algorithm performs excellently with large antenna arrays and limited RF chains. Yi et al. in [176] introduced an unfolded network called LampResNet (Learned AMP with deep residual learning). In this model, the AMP algorithm is unfolded to estimate the channel, and deep residual learning is employed to mitigate channel noise. Normalized MSE analysis shows that the proposed networks outperform LAMP [173] and achieve performance similar to LDAMP [175]. However, exploiting channel sparsity does not always yield optimal estimation performance. To address this, the authors in [85] investigate mmWave channel sparsity in the beamspace. They formulate the estimation problem as a sparse signal recovery task and improve convergence speed by unfolding the iterative PGD algorithm. In this approach, the gradient descent step size is treated as a trainable parameter.
V-C Modern Multicarrier Modulation
For frequency-selective channels, OFDM is a preferred multicarrier modulation technique. It spreads the data across multiple orthogonal subcarriers in the frequency domain, helping to mitigate the effects of frequency-selective fading. The channel is typically modeled on a per-subcarrier basis.
To enable accurate CSI feedback by transforming the channel matrix into a sparse representation, Yiyun et al. [93] proposed a neural network consisting of compression and reconstruction modules for CSI compression and feedback. They designed the reconstruction module as an unfolded ISTA network and found that higher SNR improves reconstruction performance compared to the traditional BP method. Hu et al. [94] introduced an unfolded algorithm called the learnable optimization and regularization algorithm (LORA) to enhance CSI feedback speed and accuracy. In this approach, the conventional ISTA algorithm is unfolded, and the regularization term is learned through end-to-end training. They modeled the regularization term as a learnable, differentiable transform, further refined in [94]. By incorporating this learnable regularization module, the model better captures CSI characteristics and adapts to different levels of channel estimation error. To improve CSI feedback accuracy and efficiency in OFDM FDD-based massive MIMO systems, while reducing complexity at the user equipment, Yangyang et al. [95] proposed an unfolded ISTA network. Their study showed that the real and imaginary components of the channel differential term share similar correlations, simplifying network design. Unfolding ISTA significantly enhances the CSI encoder’s reconstruction performance.
In [45], Guo et al. proposed a two-stage low-rank CSI feedback scheme for a low-rank OFDM mmWave channel. In massive MIMO systems operating in FDD mode, CSI is estimated by unfolding the fast ISTA algorithm. Fast ISTA is an improved version of the conventional ISTA algorithm. A set of thresholds is derived from the input features of the network, and to reduce the uncertainty in CSI reconstruction, the iterative steps are mapped to layers of a neural network, with $L$ stages corresponding to $L$ iterations of fast ISTA. In [96], the authors explore quantized compressed CSI at the user equipment before feedback. They propose a deep unfolding-based bit-level CSI feedback method to mitigate the effects of accumulated quantization errors. Building upon ISTA, the authors use tied learned ISTA to reduce the number of learnable parameters, offering a faster and more accurate solution to the signal reduction problem. In tied learned ISTA, the trainable parameters are shared across all layers [177].
Jianqiao et al. [97] developed an unfolded inverse-free variational Bayesian learning method to reduce the high computational complexity of matrix inversion. The sparse channel recovery performance is enhanced by learning the a-priori parameters of the sparsely represented dictionary for channel estimation. In [98], based on approximate Bayesian inference and an extension of expectation consistent approximation, the authors proposed a generalized expectation consistent algorithm, an iterative message-passing approach for sparse channel estimation. In a lens-based beamspace massive MIMO system, this approach compensates for large channel attenuation in the mmWave band. The authors introduced a deep unfolded unsupervised learning network called the learned denoising-based generalized expectation consistent signal recovery network. This network unfolds the generalized expectation consistent signal recovery algorithm and formulates the wideband beamspace channel estimation problem as a compressed image recovery task. Stein’s unbiased risk estimator is used as the loss function. In [99], the authors leveraged angle-domain channel sparsity to reduce uplink pilot overhead during channel estimation. They designed an unfolded multiple-measurement-vectors learned AMP network by jointly training the phase shift network and channel estimator, integrating the channel sparsity for improved performance.
Underwater acoustic communication is widely used for long-range transmissions, reaching up to tens of kilometers [178, 179, 180]. The main challenges in underwater communication include frequency attenuation, time-varying multipath, and large Doppler effects. Zhao et al. [47] explored integrating the MMSE structure within the layers of a DNN to create an interpretable model. This approach aims to build an equalizer framework for improving underwater communication performance.
Figure 6: Analog, digital and hybrid beamformer [181].
Cognitive radio is a wireless technology that continuously monitors the spectrum to identify unused channels, or spectrum holes, allowing for more efficient spectrum usage. In [100], a wideband signal compressed spectrum reconstruction algorithm is proposed to detect spectrum holes across a large band using a sub-Nyquist sampling rate. The authors integrate ISTA into layers of a CNN, improving the performance in both the probability of detection and the probability of false alarm for spectrum hole detection.
V-D IRS
To mitigate the blockage effect and extend the connectivity range for mmWave communication, IRS is a key technology proposed to enhance both energy and SE. IRS consists of many reflecting elements that steer the propagation of electromagnetic waves by adjusting their reflection coefficients. However, to unlock the full potential of IRS, accurate channel estimation is required for both individual and cascaded links. This is a challenging task because IRS-based channel estimation involves estimating multiple channels simultaneously: the direct channels between the BS and each user, the channel between the BS and IRS elements, and the channel between IRS elements and users.
In [109], Jiguang et al. proposed a method for estimating the rank-deficient cascaded channel in an IRS-assisted SIMO system. They developed a deep unfolding network based on a gradient descent algorithm. The model structure is inspired by DetNet [49]. In massive MIMO, increasing the number of IRS elements and BS antennas enhances the system’s capacity. However, this leads to increased training overhead to obtain accurate channel estimates within the channel coherence time. The proposed network addresses this challenge by compensating for the computational complexity.
Parallel factor decomposition is an efficient technique for estimating multiple large channel matrices in massive MIMO communication systems. It allows a high-dimensional matrix to be decomposed into a linear combination of several rank-one matrices, making the estimation of low-rank matrices more computationally efficient [182]. In [110], the parallel factor decomposition is unfolded to estimate multiple large channel matrices in the uplink of an IRS-empowered multi-user MISO communication system.
Awais et al. [111] developed two networks based on the MAP estimation criterion. The first network is a deep denoising network designed for noise-dominated SNR regions, providing support at various noise levels. The second network is a deep unfolded half-quadratic splitting network that alternates between optimizing extrapolation and denoising for interference-dominated regions. As the SNR increases, the proposed network shows significant improvement in normalized MSE. Under severe channel variations, the network delivers a performance gain of 20% in the high SNR range.
Tsai et al. [112] proposed a two-stage learned AMP network with row compression to address the joint estimation problem of direct and cascaded channels in IRS-aided mmWave systems. By utilizing the deep unfolding technique, the network combines the advantages of compressive sensing and deep learning. This approach enables reliable channel estimation with reduced training, which is crucial for efficiently reconfiguring wireless environments under IRS reflection control.
VI Precoder Design
This section reviews the applications of deep unfolding for precoder design used along with key 6G technologies.
VI-A Massive MIMO
In MIMO systems, beamforming enhances signal transmission and reception by directing signals toward specific directions using multiple antennas. This improves the received SNR [181, 183]. There are three types of beamformers: analog, digital, and hybrid. Fig.  6 illustrates the architectures for analog, digital, and hybrid beamforming [181]. The hybrid beamformer system model is described in [184]. A digital beamformer assigns one RF chain per antenna. However, in large-scale arrays, it has high power consumption and complex architecture. Analog beamformers use time delay units or phase shifters for antenna weighting but struggle with inter-user interference in multi-user scenarios. To overcome these issues, hybrid beamforming was introduced [185, 186].
Beamforming design is an optimization problem where the objective is to maximize a network utility function, such as sum rate, mean rate, or harmonic rate. This optimization is subject to resource and hardware constraints, including power budget and constant modulus analog beamforming coefficients.
Consider a base station (BS) with $n_{\text{t}}$ antennas serving $M$ single-antenna users [27]. The received signal at user $m$ is given by:
$y_{m}=\boldsymbol{{h}}_{m}^{H}\sum_{m=1}^{M}\boldsymbol{{f}}_{m}x_{m}+n_{m}$ (3)
where, $x_{m}$ is the transmitted signal with an average power of $\mathbb{E}\{|x_{m}|^{2}\}=1$, $\mathbf{h}_{m}\in\mathbb{C}^{n_{t}\times 1}$ is the channel vector, $n_{m}\sim\mathcal{CN}(0,\sigma^{2})$ represents the AWGN noise at user $m$, and $\boldsymbol{f}_{m}\in\mathbb{C}^{n_{t}\times 1}$ is the beamforming vector for user $m$. The signal-to-interference-plus-noise ratio (SINR) at user $m$ is given by:
$\text{SINR}_{m}=\frac{\left|{\boldsymbol{{h}}_{m}^{H}\boldsymbol{{f}}_{m}}\right|^{2}}{\sum_{i=1,i\neq m}^{M}\left|{\boldsymbol{{h}}_{i}^{H}\boldsymbol{{f}}_{i}}\right|^{2}+\sigma^{2}}.$ (4)
This expression quantifies the received signal quality by measuring the desired signal power relative to interference and noise. The hybrid beamformer design problem can be formulated as the maximization of the sum rate, given by:
$(\mathcal{P}_{\text{HBF}})\quad\max_{\boldsymbol{F}^{\text{DB}},\boldsymbol{F}^{\text{AB}}}$ $\sum_{m=1}^{M}\log_{2}\left(1+\text{SINR}_{m}\right)$ (5)
s.t. $\left\lVert\boldsymbol{F}^{\text{AB}}\boldsymbol{F}^{\text{DB}}\right\rVert_{F}^{2}\leq P,\quad\boldsymbol{F}^{\text{AB}}\in\mathcal{F},$
$|\boldsymbol{F}^{\text{AB}}|^{2}=1.$
where $\boldsymbol{f}_{m}=\mathbf{F}^{\text{AB}}\boldsymbol{f}_{m}^{\text{DB}}$, $\mathbf{F}^{\text{AB}}\in\mathbb{C}^{n_{t}\times N_{\text{RF}}}$ is the analog precoding matrix with $N_{\text{RF}}$ RF chains, belonging to a codebook $\mathcal{F}$. The digital precoding vector for the $m^{\text{th}}$ user is $\boldsymbol{f}_{m}^{\text{DB}}\in\mathbb{C}^{N_{\text{RF}}\times 1}$, and the digital precoding matrix $\boldsymbol{F}^{\text{DB}}$ consists of column vectors $\boldsymbol{f}_{m}^{\text{DB}}$. These problems are nonconvex and NP-hard, meaning they lack closed-form solutions. The unit-modulus constraints on the analog precoding matrix make precoder design even more challenging. To address this, the algorithm unrolling technique combines model-driven and data-driven approaches. It optimizes beamforming matrices by integrating the interpretability of model-based methods with the adaptability of data-driven techniques. Shlezinger et al. [26] discuss the architectural constraints and challenges of hybrid beamforming optimization using both conventional and data-driven methods.
The classic iterative WMMSE algorithm is widely used for precoder design to maximize the system’s weighted sum rate. It assigns different weights to the MSE of each user based on priority or quality of service. By adjusting these weights iteratively, WMMSE balances user performance and improves overall efficiency. The algorithm provides closed-form, iterative update rules to solve a surrogate optimization problem with the same local optima as the sum-rate maximization problem [122]. Maximizing the weighted sum rate while satisfying a power constraint is a nonconvex problem [187]. The approach proposed in [187] exploits the separable structure of power constraints. To achieve a locally optimal beamforming strategy, the authors introduce a simple and efficient distributed inexact cyclic coordinate descent method.
A deep unfolded version of the iterative WMMSE algorithm is proposed for precoder design in multi-user MIMO systems [69]. The authors call this network architecture the iterative algorithm induced deep unfolding network (IAIDNN). To improve efficiency, WMMSE is reformulated to approximate matrix inversions. A generalized chain rule in matrix form helps visualize the gradient recurrence between adjacent layers during backpropagation. Using this rule, the gradients of trainable parameters are computed effectively. The WMMSE algorithm [122] involves matrix inversions, eigendecompositions, and bisection searches in its iterations. Since any MMSE solution requires matrix inversion, WMMSE also relies on it. To improve convergence while avoiding matrix inversion, a downlink beamforming method using an unfolded PGD-WMMSE algorithm for multi-user MISO systems is proposed in [68]. The authors replace matrix inversions with PGD and compute gradients in parallel. Unlike IAIDNN [69], this approach eliminates matrix inversion, making hardware implementation more practical. Additionally, it achieves a high weighted sum rate in fully loaded scenarios, where the number of users matches the number of transmit antennas at the BS.
In [69], learnable parameters are used to approximate matrix multiplication and inversion in WMMSE update rules. However, the network fails to exploit the inherent graph structure of wireless networks. This limitation is addressed in [70]. An unfolded WMMSE algorithm for transceiver design in multicell MU-MIMO interference channels with local CSI is proposed in [70]. The authors use graph convolutional networks (GCNs) to unfold WMMSE, enabling coordination in multicell MU-MIMO systems. Graph filters and convolutional networks approximate certain WMMSE variables using graph learners. However, the approach in [70] lacks flexibility in learning the exact transformation of WMMSE variables. To overcome this, [71] introduces a hybrid unfolded WMMSE algorithm for transmit beamforming in MU-MIMO ad-hoc networks. This method leverages GNNs to learn functional transformations of WMMSE parameters. Compared to the classical WMMSE algorithm, the proposed network reduces the number of layers, leading to faster convergence of the optimization problem. In [188], the joint optimization of precoders for multicell MU-MIMO systems is studied. The WMMSE algorithm is unfolded to enable distributed and synchronized precoding across cooperative BSs. Numerical analysis compares this approach with the networks in [69, 68]. The methods in [69, 68] rely on centralized optimization, leading to high complexity and backhaul overhead. In contrast, the proposed network in [188] reduces computational cost while maintaining efficiency.
The gradient descent algorithm is a widely used method for finding a stationary point. Shi et al. [72] designed a hybrid analog-digital transceiver using iterative gradient descent to obtain minimum symbol error rate (MSER) in massive MIMO systems. They unfolded the iterative gradient descent algorithm into a DNN, where learnable parameters serve as step sizes for the gradients. By using a fixed number of DNN layers with optimized step sizes, the approach reduces the number of iterations and speeds up convergence compared to conventional methods. The input to this MSER-based algorithm includes a user’s transmit beamforming matrix, an analog transmit beamforming matrix, an analog receive beamforming matrix, and a digital receive beamforming matrix. The step sizes guide the gradient descent update for each matrix. The gradients for these parameters are updated using SGD. The authors claim that their proposed network matches the performance of the MSER-based iterative gradient descent algorithm while significantly reducing complexity. Zhu et al. [73] proposed a beamforming learning architecture by unfolding a parallel gradient projection algorithm for MISO heterogeneous systems. Their design optimizes scenarios where network parameters, such as the number of antennas, BSs, or overall network size, differ from those used during training. Gradient-based methods have been successfully applied to beamforming optimization with reduced complexity. However, these approaches can be further improved using advanced optimization frameworks like PGA.
A fast and robust precoding network for multi-user MIMO systems with imperfect CSI and per-antenna power constraints is proposed in [74]. The objective function is formulated as a geometric program that minimizes the MSE between the transmitted and recovered symbol vectors at the user. This optimization leverages the duality between downlink and uplink MSE feasible regions. The authors map the iterative optimization process into a layer-wise deep unfolding framework, enabling fast and robust downlink precoding prediction.
PGA is a variant of the gradient ascent algorithm used to maximize an objective function under constraints. It updates the solution in the gradient direction to increase the objective function and then projects it back onto the feasible set. Lavi et al. [75] unfolded the PGA algorithm to optimize hybrid precoding for multiuser downlink MIMO, assuming perfect CSI availability. To handle noisy CSI, they implemented the projected conceptual mirror prox algorithm [189, 190]. The step sizes for updating the analog and digital precoder matrices are tuned using mini-batch SGD, a data-driven approach. The solution is computed within a fixed number of iterations, defined as network layers, ensuring a balance between interpretability and efficiency. In [76], Bilbao et al. proposed an efficient analog beamforming design to maximize the achievable rate of massive MIMO in-band full-duplex systems. The beamforming problem is non-convex due to constant modulus constraints on analog beamforming coefficients. To address this, they introduced an unfolded algorithm combining PGA with ZF projection onto convex sets, achieving faster convergence.
Wang et al. [191] proposed a universal algorithm for downlink massive MU-MIMO systems with finite-alphabet precoding to minimize inter-user interference. The algorithm is an improved version of the original ADMM framework. The authors observed that the standard ADMM struggles with nonlinear discrete optimization due to biased alternating updates. To address this, they introduced iterative discrete estimation [191], ensuring unbiased updates and better convergence. To reduce hardware costs and power consumption in multi-user massive MIMO, He et al. [77] developed an unfolded network using the IDE-based algorithm with finite-alphabet precoding. Additionally, constructive interference-based symbol-level precoding for multi-user MISO downlink transmission is discussed in [78]. In [79], the conjugate gradient algorithm is unfolded for constant envelope precoding to reduce computational overhead. The Riemannian manifold optimization (MO)-based conjugate gradient descent is implemented across network layers. Simulations show that this approach effectively suppresses multi-user interference, improving overall performance.
In [192], the authors use deep unfolding to optimize user scheduling, BS clustering, and beamforming in a dense C-RAN system. Reducing the number of BSs per user helps ease backhaul congestion. The goal is to maximize the weighted sum rate while minimizing BS clustering, all within power constraints. The PGD algorithm is unfolded with step sizes learned in an unsupervised manner. This approach avoids matrix inversion and bisection search, reducing iterations and computational complexity.
In [80], the interior point method is unfolded in an unsupervised manner for precoding design. It leverages interference in multi-user MISO systems to minimize power while meeting SINR constraints. In [81], the authors address the challenges of mutual coupling in beamforming. They propose mitigating its effects through covariance matrix reconstruction by unfolding an augmented coprime array.
VI-B mmWave
mmWave wireless communication, operating in the 30–300 GHz band, offers abundant spectral resources to alleviate the bandwidth crunch at sub-6 GHz frequencies [157]. However, mmWave signals suffer from severe propagation and penetration losses. To compensate for these losses, large-scale antenna arrays with tens to hundreds of elements are required to achieve high power gains. The short wavelength (1–10 mm) of mmWave signals allows compact integration of these arrays into small transceivers. These features introduce unique challenges in designing physical-layer transmission algorithms [193]. mmWave MIMO systems rely on beamforming to mitigate multi-user interference while enhancing diversity and spectrum efficiency.
Typically, the beam selection and digital precoding matrices are designed separately, leading to suboptimal solutions. To address this, Hu et al. [86] proposed a joint design approach combining DL and model-driven techniques. DRL optimizes beam selection, while a deep unfolding neural network, based on iterative WMMSE, optimizes the digital precoding matrix. The goal is to maximize the system sum rate while satisfying beam selection constraints. The DRL and deep unfolding networks are trained sequentially, with one network’s output serving as the input to the other. Backpropagation and SGD are used for training and parameter updates.
Lin et al. [87] studied a linear MMSE transceiver design for uplink massive MU-MIMO with one-bit ADCs. To minimize the system MSE by optimizing the precoder at each user equipment, the PGD algorithm is unfolded. The authors focused on training only the step size in each PGD iteration, significantly reducing the training overhead. In [88], the authors combined deep generative models with unfolding techniques to achieve near-optimal hybrid beamforming while minimizing feedback and computational complexity. They represented the channel as a low-dimensional manifold using a GAN, enabling efficient optimization of digital and analog precoders. The hybrid beamforming problem is formulated as a bilevel optimization, where an iterative rule updates both precoders. The upper-level optimization refines the digital precoder, leveraging full knowledge of the lower-level task constraints. Simultaneously, the lower-level optimization adjusts the analog precoder based on the digital precoder’s decisions. This structured optimization approach ensures efficient beamforming design with reduced computational costs.
VI-C Next Generation Multiple Access
Rate-Splitting Multiple Access (RSMA) is a flexible multiple access technology that merges the strengths of space division multiple access (SDMA) and NOMA. It uses linearly precoded rate-splitting at the transmitter and SIC at the receivers. RSMA balances interference management by partially decoding some interference while treating the rest as noise, bridging the gap between SDMA (which treats all interference as noise) and NOMA (which fully decodes interference) [194]. In [107], the authors applied fractional programming with hyperplane fixed-point iteration to design the RSMA beamformer. The algorithm is unfolded into a neural network, replacing its inner loop with a small-scale DNN to predict Lagrangian dual variables. Simulations show that the proposed approach achieves high performance while reducing computational complexity.
VI-D IRS
An IRS consists of a programmable array of passive elements that dynamically adjust the phase shifts of incoming signals. This enables signals to align in phase at the receiver, enhancing reception. IRS can mitigate interference, create alternative propagation paths to bypass obstacles, and extend signal coverage to previously unreachable areas. The phase shifts are optimized to reflect signals coherently toward the intended user while minimizing interference in other directions. Therefore, jointly optimizing IRS phase control and transmit beamforming enhances the performance of IRS-assisted systems [195].
With the aid of IRS, Chen et al. [113] addressed the joint design of the hybrid beamforming matrix and the phase-shifting matrix of the IRS elements. They proposed the deep unfolded WMMSE MO algorithm (DU-WMMSE-MO) using a deep unfolding methodology. The algorithm showed improved performance in IRS phase-shift design, utilizing two loops to jointly update the step size of the beamforming and IRS matrices. Compared to the Armijo backtracking line search method used in conventional WMMSE-MO [196], DU-WMMSE-MO provides more flexibility but with a higher runtime. Training is performed in an unsupervised manner, with a loss function that aligns with the objective function. The trainable parameters include the weights for the precoder, phase configuration of the IRS elements and combiner, and biases for the precoder and phase-shift matrix of the IRS.
An IRS-assisted multiuser MIMO full-duplex system is explored in [114]. The active beamforming matrices at the BS and uplink users, along with the passive beamforming matrix at the IRS, are jointly optimized to maximize the weighted sum-rate of the system. To address the high computational complexity of matrix inversion in the conventional stochastic successive convex approximation (SSCA)-based optimization algorithm, a deep-unfolding neural network is proposed. Fig. 7 shows two such netwroks from [114] . The network consists of a long-term passive beamforming network (LPBN) and a short-term active beamforming network (SABN). The learnable parameters of the LPBN and SABN are the IRS phase shift or passive beamforming vector $\boldsymbol{\theta}$ and the active beamforming vector $\boldsymbol{\Psi}$, respectively.
Figure 7: Two networks LPBN and SABN [114].
In [115], Ming-Min et al. introduced a two-timescale transmission approach to reduce the signal processing complexity and channel training overhead for the IRS-aided multiuser MISO system. To minimize the average transmit power at the AP while ensuring the individual quality of service constraints for the users, a primal-dual decomposition-based two-timescale hybrid beamforming method is proposed and unfolded. This approach allows for efficient beamforming optimization by separating the optimization into two timescales, improving both computational efficiency and performance.
In [197], a low-complexity IRS-assisted downlink MISO system is designed using an unfolded block coordinate descent algorithm to address the joint optimization problem of transmit beamforming and IRS phase shifts. The multi-ratio fractional programming structure of the objective function is handled by decoupling the problem into two subproblems: one for optimizing the beamforming at the base station and another for optimizing the phase shifts of the IRS. This is achieved through the use of Lagrange dual transformation, effectively simplifying the optimization process. The numerical analysis validated the achieved weighted sum rate is comparable to [198] but with reduced computational complexity. Furthermore, the authors in [199], have developed an unfolded network to address the joint beamforming task for an IRS-assisted MISO system. The WMMSE and power iteration algorithm are used as an alternating optimization strategy to derive the local optimal solution for the active and passive beamforming. The proposed network reduce the computational runtime to approximately 3% of the block coordinate descent algorithm in [198].
In [116], an IRS-assisted MISO covert symbiotic radio communication system is explored, and a deep unfolding algorithm based on gradient descent is proposed for beamforming design. The goal is to optimize the achievable transmission rate of the primary transmitter by refining both active and passive beamforming strategies while satisfying two key constraints: (1) ensuring covertness and (2) maintaining the SNR of the secondary transmission. The complex optimization problem, along with its constraints, is transformed into the Lagrangian dual domain. Each layer in the deep unfolded network consists of four modules that update: (1) the active beamforming, (2) the passive beamforming, and (3) two Lagrange multipliers corresponding to the two constraints.
Gangyong et al. in [200] have designed the IRS phase shift matrix by maximizing the sum path gain of the channel. This maximization is achieved by unfolding the Riemannian gradient descent algorithm. The network takes channel matrices and the initial reflection coefficients as inputs and outputs the optimized IRS coefficients. By configuring the IRS element phase shifts, the system effectively directs signal propagation paths to maximize the sum-path-gain, which is a key metric for evaluating beamforming quality and learning optimal IRS phase shifts to enhance the overall beamforming pattern.
VI-E THz
6G wireless networks are expected to achieve terabit-per-second (Tbps) data rates to support ultra-reliable low-latency communication. To enhance spectral and energy efficiency, THz-enabled massive MIMO has emerged as a promising solution, leveraging the abundant bandwidth at THz frequencies to significantly boost data rates. Additionally, hybrid beamforming offers a cost-effective approach by reducing the number of RF chains while maintaining high performance, making large-scale deployment of high-frequency systems more practical [201].
Nguyen et al. [118] proposed a deep unfolding framework by unrolling the iterations of Riemannian manifold optimization (Riemannian MO), specifically the MO-AltMin algorithm [202]. Instead of directly minimizing WMMSE and maximizing SE, the problem is transformed into an approximate matrix factorization problem, which is then solved using LS. This approach is computationally more efficient, as it requires fewer layers to compute the weights of the precoding matrix, significantly reducing the training time. Compared to the approach in [75], the proposed network achieves better performance while maintaining much lower time and computational complexity.
Minghui et al. [119] proposed a DL-based RSMA scheme for IRS-aided THz multi-user MIMO systems, focusing on enhancing robustness against CSI imperfections to improve system performance. The approximate WMMSE algorithm for RSMA is unfolded in this approach to boost precoding performance while reducing computational complexity. The deep unfolding algorithm processes the estimated equivalent channel through a series of layers to optimize the digital active precoding. This strategy aims to mitigate performance loss caused by model mismatches in real-world systems. By deep unfolding the approximate WMMSE algorithm, the method achieves improved precoding performance and lower computational complexity compared to the original WMMSE approach. This iterative deep unfolding process refines the precoding strategy based on the estimated equivalent channel, optimizing performance in IRS-aided THz massive MIMO-OFDM systems.
In [120], Adam et al. introduced a method to jointly optimize active beamforming, passive beamforming, and UAV trajectory to minimize power consumption while accounting for UAV jitters and imperfect hardware constraints. They designed a hybrid semi-unfolding DNN to address these constraints during the real-time CSI gain estimation for IRS and UAV links. The approach is termed “semi-unfolded” because it combines two types of networks: one based on deep unfolding and the other based on a multi-kernel convolutional LSTM (MK-CLSTM) NN. For active beamforming, the authors employed semidefinite programming to obtain a closed-form solution, which was then transformed into NN layers to replicate the iterative beamforming solution in real time. The Lagrangian functions with well-defined Karush-Kuhn-Tucker (KKT) conditions were derived to design the layers of the deep unfolded architecture. The UAV trajectory and beamforming vectors are the parameters updated within these layers. For passive beamforming, a MK-CLSTM-based NN was used. Training was done using mini-batch SGD. The authors claim that their model-driven approach achieved $99.24\%$ accuracy and is $26.21\%$ faster than the classical iterative methods, offering significant improvements in both efficiency and speed for UAV-based communication systems.
VII Sensing and Communication
The independent operation of sensing and communication as separate systems has been widely studied [203]. However, with the deployment of wireless communication services in higher frequency bands, such as mmWave and THz, spectrum congestion has become a major issue. In the 6G era, independent sensing systems experience interference from communication signals, leading researchers to recognize the shared system architecture and signal processing techniques between sensing and communication [204], [205], [206]. This realization has driven the development of ISAC, where both functionalities are merged into a unified system.
ISAC can be categorized based on sensing methods into active sensing and passive sensing [207].
  • Active sensing (Mono-static sensing): The transmitter and receiver are co-located, allowing the system to extract target parameters using both the transmitted signal and received echoes. However, this approach suffers from significant self-interference at the ISAC receiver due to the strong reflections from the transmitted signals.
  • Passive sensing (Bi-static sensing): The sensing echoes are received at a different location than the transmitter, eliminating the need for in-band full-duplex operation at the receiver. This approach is more suitable for applications where reducing self-interference is a priority.
ISAC provides a promising solution for optimizing spectral efficiency and enhancing system performance in future wireless networks. However, understanding the fundamental performance limits of ISAC systems is crucial for their practical implementation. The estimation accuracy of key sensing parameters, such as direction of arrival, Doppler frequency, velocity, and target position, was thoroughly investigated in [208]. This study provided a detailed analysis of both theoretical and practical limits of ISAC systems. The authors emphasized that the performance limits of ISAC channels cannot be derived by simply merging the existing bounds from independent sensing and communication systems. Instead, a unified framework is necessary to characterize these limits effectively. To address this, the authors proposed deriving unified Cramér-Rao bounds for ISAC, which establish the fundamental lower bounds on sensing accuracy given a specific allocation of orthogonal resources between communication and sensing. These unified expressions provide insights into how ISAC systems can optimize their resource allocation strategies to achieve superior performance in joint sensing and communication tasks.
VII-A MIMO ISAC
The integration of deep unfolding in ISAC systems has shown promising advancements in optimizing waveform design, beamforming, and communication-sensing tradeoffs. Ziwei et al. [82] applied deep unfolding techniques to optimize waveforms in MIMO radar systems, specifically focusing on minimizing the range-integrated sidelobe level. This is crucial for reducing interference and improving target detection performance. Their deep unfolding network efficiently optimizes radar waveforms, enhancing overall radar performance. In [83], the challenge of designing constant-modulus waveforms for multiuser MIMO ISAC systems was explored. Since constant-modulus waveforms help avoid signal distortion in nonlinear power amplifiers, their design is crucial. However, the non-convex nature of the constant-modulus constraint makes optimization challenging. To address this, the authors employed a deep unfolded network with a sparsely connected structure, trained in an unsupervised manner, ensuring efficient waveform design under these constraints. Hybrid beamforming in ISAC systems poses additional complexity due to the need to balance communication rates and sensing accuracy. The authors in [84] tackled this challenge by unfolding the iterative PGA algorithm. Their deep unfolding-based method accelerated convergence and enhanced overall system performance, achieving an optimal balance between communication and sensing objectives. Further, joint optimization of radar-communication beamforming was addressed in [209]. The authors incorporated symbol-level precoding with constructive interference constraints to improve energy efficiency by exploiting multi-user interference. They developed iterative algorithms to leverage problem separability, enabling parallel and recursive computation. Deep unfolding was utilized to reduce computational complexity, improve scalability, and support real-time implementation.
VII-B mmWave MIMO ISAC
mmWave MIMO is a key technology in 6G wireless communication, offering high transmission capacity and sub-meter user localization. Its large-scale antenna arrays enable precise angular resolution. Traditional localization methods estimate position-related parameters before reconstructing the user’s location. To improve this, Fan et al. proposed a fast direct localization method in [89]. They use an unfolded ADMM solver with low complexity for direct localization. The trainable parameters include $\rho$ (penalty parameter), $w$ (weight coefficient for channel gain between the base station and user), and $\tau_{1}$, $\tau_{2}$ (two positive proximal parameters).
VII-C Modern Multicarrier Modulation for ISAC
5G and beyond 5G wireless communication systems operate at carrier frequencies in the microwave regime, where radar applications are dominant. This has led to the development of joint waveforms for sensing and communication to address limited spectral resources [210]. OFDM-based radars in LTE and 5G NR mobile networks’ BS are studied for radar and sensing in [211], [212], [213]. OTFS modulation for joint radar parameter estimation and communication is explored in [214], [215], [216].
An ISAC signal processing problem was proposed in [102] to process pilot and data signals together. The optimization problem requires the simultaneous recovery of CSI, transmitted data symbols, and target information, making it non-convex. To solve this, the authors used an iterative optimization algorithm that alternates between passive sensing and communication detection. Jingcai et al. proposed a deep unfolded network for target parameter estimation in MIMO-OFDM-based ISAC systems [103]. Their method leverages the sparsity of target parameters, using a compressed sensing approach. The problem is reformulated as a least absolute shrinkage and selection operator (LASSO) problem. The fast ISTA algorithm is then unfolded, with learnable parameters enabling high-precision solutions with low computational complexity.
The problem of estimating parameters of moving targets in high-speed scenarios using OTFS modulation for ISAC systems is explored in [104]. The target parameter estimation is reformulated as a compressed sensing problem. For the deep unfolded network, ADMM iterations are unfolded into the network’s layers.
VII-D IRS-Assisted ISAC
Liu et al. explored beamforming design for an IRS-assisted ISAC system with bounded angle uncertainty of sensing targets and statistical CSI uncertainty related to user equipment in [117]. They formulated a bi-level optimization problem, unfolding the projection method and Riemannian MO approach. These methods handle the maximum transmit power and unit-modulus constraints. The bi-level problem consists of two parts: the lower-level problems focus on identifying the worst-case CSI and target sensing angles, while the upper-level problem addresses beamforming design and optimization of Lagrange multipliers based on the given CSI and target angles. A double-loop deep unfolding NN is applied, unfolding the iterative gradient descent into a multi-layer structure, with each layer containing an upper-level and a lower-level loop. Trainable step sizes and offset parameters are introduced as network parameters.
VIII Decoding in Error Correcting Codes
Error control codes are essential for reliable and efficient communication. During data transmission, bits may flip due to channel noise, causing errors. Error correction coding helps detect and correct these errors at the receiver. Claude Shannon revolutionized channel coding with his landmark paper [217].
Mobile communication has evolved through several generations with significant technological advancements. In the first generation (1G), communication was primarily analog, leaving no need for channel coding. With the second generation (2G), convolutional and block codes were introduced to compress voice signals and convert them into digital form. Turbo codes were adopted as standard coding schemes in the third (3G) and fourth (4G) generations, offering near-Shannon limit performance with moderate decoding complexity [218]. Low-density parity-check (LDPC) codes, first introduced in the 1960s, were later rediscovered and extensively studied [219], [220], [221]. They are widely used in 4G, 5G, and beyond. LDPC codes are represented using a bipartite graph, placing them in the “codes on graphs” category, designed for capacity-approaching codes with iterative decoding algorithms. Since LDPC codes use sparse graphs, the number of edges grows linearly with block length, making them low-density and less complex [222]. Polar codes, introduced by Erdal Arıkan in 2009, are a modern class of error-correcting codes designed to achieve the capacity of symmetric binary-input discrete memoryless channels [223]. They are based on channel polarization and are known for their capacity-achieving property, meaning they can approach the Shannon limit as block length increases.
BP is widely used to decode modern error correction codes [224]. This decoder is constructed from the Tanner graph, introduced by Michael Tanner in his landmark paper [225]. He developed the min-sum and max-sum algorithms, which are derived from the generic sum-product decoding [226]. Polar codes use successive cancellation decoding, where received bits are decoded sequentially by leveraging frozen bits and previously decoded bits. More advanced techniques, such as successive list decoding assisted by cyclic redundancy check (CRC), further improve performance [227].
In practical wireless communication, channels often deviate from mathematical models or are not precisely known at the decoder. In such cases, a DL-based decoder offers a robust and adaptive solution. Additionally, model-driven DL is well-suited for providing an interpretable analysis of iterative decoding algorithms. Wei et al. [228] proposed an unfolded ADMM-based penalized algorithm for binary linear codes. Their approach introduces two networks that transform the ADMM penalty parameters $\{\mu,\alpha\}$ into iteration-dependent parameters, along with an adjustable penalty function. Each stage of the learnable ADMM decoding network (LADN) corresponds to one ADMM iteration, where iteration-dependent parameters enhance convergence and reduce the number of iterations. The key idea is that increasing the number of learnable parameters improves generalization capability. For the second network, a piecewise linear penalty function with adjustable slopes is designed to suppress pseudo codewords effectively, improving decoding performance beyond conventional $\ell_{1}$ and $\ell_{2}$ norm functions. Parameter updates are trained using backpropagation via the Adam optimizer. A customized loss function based on MSE ensures effectiveness at every iteration by considering outputs from all layers. The authors claim that their decoder outperforms the conventional ADMM decoder. Additionally, decoding irregular binary LDPC codes via unfolded ADMM is explored in [229].
In [48], the authors propose the TurboNet architecture, which combines DL with the max-log MAP algorithm. While most existing model-driven decoding algorithms rely on BP and are iterative, the authors enhance the traditional MAP algorithm by unfolding it, with each iteration represented as a DNN decoding unit. As a result, the soft-input soft-output decoder, which traditionally uses the max-log MAP algorithm and extrinsic LLRs, is replaced by two subnets based on a neural log-MAP algorithm for each iteration.
In [230], the authors tackle the decoding challenges of marker codes in insertion and deletion channels, with a focus on overcoming CSI uncertainty. To improve detection performance, they propose DL techniques, implementing both RNNs and a deep unfolded network. The forward-backward algorithm is unfolded into multiple layers within the network to enhance the overall decoding process.
Gao et al. [231] have unfolded the factor graph of the BP decoder using a one-bit quantizer for polar codes. In [232], a double-loop iterative decoding algorithm for LDPC codes is unfolded within a penalty dual decomposition framework. Additionally, by unfolding the iterative decoding process between check nodes and variable nodes into a feed-forward propagation network, a deep unfolded normalized min-sum LDPC decoder is derived in [233].
Tanner-graph-based transmission schemes for massive multiple access, incorporating NOMA and grant-based random access, are discussed in [234]. To enhance decoding efficiency, a message-passing decoder is unfolded for both transmission schemes, allowing the decoder to improve as more information about the transmitted data is learned during the training epochs. In [235], theoretical results for the unfolded BP decoder are presented. The authors define the generalization gap of a decoder as the difference between the empirical and expected bit-error rates. They provide a set of theoretical results that bound this gap and examine its dependence on decoder complexity, which is influenced by code parameters (such as block length, message length, and variable/check node degrees), the number of decoding iterations, and the size of the training dataset.
IX Power allocation
In a multi-antenna wireless communication system with multiple users and limited resources, ensuring reliable transmission of massive amounts of data presents significant challenges. One of the primary constraints is the power budget, which must be allocated efficiently among data streams to maintain a high quality of service. The water-filling algorithm is a widely recognized technique for optimal power allocation across subchannels, aiming to maximize the sum capacity or sum rate. Since power allocation directly impacts channel capacity, it is formulated as an optimization problem, where maximizing the capacity rate leads to improved system performance. The power allocation problem can thus be formulated as an optimization problem as shown in (6) below:
$\max_{\left\{P_{i}\right\}}$ $\frac{\sum_{i=1}^{K}w_{i}R_{i}(P_{i})}{\sum_{i=1}^{K}P_{i}}$ (6)
s.t. $R_{i}(P_{i})\geq R_{\min,i},\quad\forall i,$
$\sum_{i=1}^{K}P_{i}\leq P_{\text{total}},$
$P_{i}\geq 0,\quad\forall i.$
where $K$ is the number of users, $w_{i}$ is the weight associated to user $i$ decided by user priority or fairness, $R_{i}(P_{i})$ is the rate achieved by user $i$ given the power $P_{i}$ allocated to them, $R_{\text{mini},i}$ is the minimum required rate for fairness and $P_{\text{total}}$ is the sum of all users’ power values.
Several classical approaches, such as the Dinkelbach algorithm, Lagrange dual decomposition, interference pricing, successive convex approximation (SCA), WMMSE, and matching theory, have been employed for power allocation [122, 236, 237, 238, 239, 240]. However, a major drawback of these methods is their reliance on iterative algorithms, which are computationally intensive and result in long runtimes, ultimately degrading system energy efficiency. The deep unfolding approach mitigates this issue by reducing the number of iterations required to achieve a near-optimal solution, thereby enhancing computational efficiency and energy savings.
Wireless networks expand in terms of spatial coverage, but traditional DNN architectures, such as multi-layer perceptrons and CNNs, struggle to capture the inherent topology of these networks. To address this limitation, Chowdhury et al. [241] propose leveraging GNNs to parameterize the power allocation function. In their approach, the WMMSE algorithm is unfolded into layers, resulting in a distributed, efficient, and scalable solution. The unfolded WMMSE method integrates domain-specific elements with trainable components, which are parameterized by GNNs to exploit the natural graph structure of wireless communication networks. The objective is to map the channel matrix and approximate optimal power allocation by training the GNN using a block coordinate descent algorithm.
In [242], Li et al. unroll the SCA algorithm into NN layers for energy-efficient power allocation in wireless interference networks. The proposed unfolded algorithm is highly adaptable, as it generalizes across varying network sizes and channel distributions while also enabling transfer learning between different path loss models. Given that the system operates in a multi-user wireless interference network, it can be naturally modeled as a graph neural network GNN. The power allocation policy is formulated based on the CSI matrix ${\bf{H}}$ and the total power constraint $P_{\text{total}}$, within a layered architecture incorporating trainable weights.
Hu et al. [243] propose an unsupervised power allocation optimization approach for ad-hoc networks by integrating attentive graph representation with deep unfolded WMMSE. The key advantage of attentive graph representation is its ability to allow network nodes to selectively focus on the most relevant connections during learning, rather than treating all links equally. The power allocation framework is built upon this attentive graph representation, as described in [244, 245]. Meanwhile, the deep unfolded WMMSE maps the iterative steps of the WMMSE algorithm into neural network layers within an unsupervised learning paradigm. This approach effectively combines deep unfolding with graph representation learning, significantly reducing the search space for the objective function. As a result, the model achieves improved interpretability and accelerated training while eliminating the need for supervised data.
X Physical Layer Security
Physical layer security is a critical research area aimed at ensuring secure and reliable transmission between the transmitter and receiver while preventing eavesdropping and impersonation attacks. Due to the broadcast nature of wireless communication, adversarial users pose a significant threat. These adversaries generally fall into two categories: eavesdroppers (or middleman attackers), who attempt to extract information from an ongoing transmission without detection, and jammers, who deliberately degrade the signal quality to prevent the receiver from successfully decoding the intended message [246]. Developing robust security mechanisms at the physical layer is essential to mitigate these threats and enhance communication reliability.
The fundamental network model for analyzing security and confidentiality is a three-terminal system, as illustrated in Fig. 8. In this setup, the legitimate transmitter (Alice), the intended receiver (Bob), and the unauthorized eavesdropper (Eve) are equipped with $N_{\text{t}}$, $N_{\text{r}}$ and $N_{\text{e}}$ antennas, respectively. As introduced by Wyner in his seminal work on wiretap channels [247], this model captures the imperfections introduced by the communication channel. Alice transmits an information signal, $\bf{x}$, a random vector intended for Bob, over the main channel, which is modeled as a discrete memoryless channel. However, due to the broadcast nature of wireless communication, the signal received by Bob, $\bf{y}$, is also partially intercepted by Eve through an additional wiretap channel, resulting in her received signal, $\bf{z}$, as shown in Fig.  8.
Figure 8: General wiretap channel [246].
Traditionally, secret key-based encryption has been the standard approach to ensuring message confidentiality, where the encryption key is shared exclusively between the transmitter and the intended receiver. As defined in [248], a secrecy system can be abstractly described as a set of transformations that map a set of possible messages to a set of possible cryptograms. Each transformation corresponds to an encryption process that utilizes a unique, secure key to encipher the message, ensuring that unauthorized entities cannot decipher the transmitted information.
The primary objective of a secure communication system is to maximize the secrecy rate, ensuring that the transmission rate to the intended user remains high while minimizing information leakage to the eavesdropper. The secrecy capacity $C_{s}$ is defined as the maximum rate $R_{s}$ at which secure communication can be achieved, guaranteeing that the eavesdropper receives negligible information about the transmitted message. This capacity represents the fundamental limit of secure data transmission, balancing efficiency and confidentiality in the presence of adversarial interception. The secrecy capacity can be expressed as:
$C_{s}=\max_{P_{X}}\left\{I\left({\bf{X};\bf{Y}}\right)-I\left({\bf{X};\bf{Z}}\right)\right\}$
where, $I(\bf{X};\bf{Y})$ is the mutual information between the transmitter and the intended user, and $I(\bf{X};\bf{Z})$ represents the information leakage to the eavesdropper. The transmitter can maximize the secrecy capacity by carefully choosing the probability distribution $P_{X}$ for the transmitted signal $\bf{x}$ to maximize the difference $I(\bf{X};\bf{Y})$-$I(\bf{X};\bf{Z})$ i.e, the rate $R_{s}$.
Jamming attacks pose a serious threat to the reliable operation of wireless communication systems [249, 250]. A jamming attacker deliberately transmits interference or noise on the same frequencies as legitimate users. This disrupts communication channels, reduces signal quality, and can block data transmission entirely. The interference makes it harder for users to decode messages, leading to connection losses or forced retransmissions. As a result, network efficiency decreases, and latency increases. Marti et al. in [251] explored solutions to combat smart jamming in massive multi-user MIMO systems. In these systems, jamming attacks can selectively interfere at specific times to avoid detection. The authors proposed an optimization-based approach for joint jammer estimation, channel estimation, and data detection, using the BS’s multi-antenna configuration. This method does not require prior knowledge of the jammer’s behavior, making communication more resilient even during sporadic or bursty jamming. Using a deep unfolding approach, the authors developed two iterative algorithms. The first is mitigation estimation and detection (MAED), which uses forward-backward splitting for estimation. The second is soft-output MAED (SO-MAED), an enhanced version that applies deep unfolding to optimize parameters. SO-MAED improves detection accuracy, especially in high-interference environments, and does not require knowledge of the attack type. Error precision and gradient step sizes are key parameters for improving learning stability.
XI Joint Techniques
This section reviews different joint techniques that are aided by deep unfolding method.
XI-A Joint Detection and Channel Estimation
An iterative method for estimating joint channel parameters and selecting symbols using the EM algorithm is presented in [252]. This method performs symbol detection in the presence of Gaussian noise and intersymbol interference in a dispersive channel. The authors model the missing data using a Markov chain, and channel parameter estimation follows the ML criterion. The Markovian properties of the channel state sequence allow for calculating the required likelihoods using the forward-backward algorithm. In [253], the authors explore a V-BLAST-type combination of OFDM with MIMO to improve SE and multi-user downlink throughput. They propose a joint data detection and channel estimation algorithm for MIMO-OFDM that integrates the QR decomposition M-algorithm and Kalman filter. The M-algorithm, combined with QR decomposition, reduces the computational complexity of tree-based search methods [254].
As mentioned in [6], the data-driven approach uses a fully connected network, where each sub-block of the OFDM transceiver is designed. Gao et al. [255] proposed a deep unfolding approach to address the issue of DNN performance being dependent on large training datasets. In contrast to the approach in [6], where a DL architecture completely replaces the conventional modules, the ComNet architecture shown in Fig. 9 incorporates both channel estimation and detection modules. Both modules use conventional communication solutions as inputs to the network. The DL network then refines these coarse inputs. For example, the LS channel estimate, $\bf{\hat{h}_{\text{LS}}}$, is refined using pilot symbols ${\bf{x}}_{p}$ and corresponding received symbols ${\bf{y}}_{p}$ as inputs to the LS sub-network within the channel estimation module. The channel estimate, $\bf{\hat{h}_{\text{LS}}}$, is passed from the LS sub-network to the LS Refine Net sub-network to generate an accurate channel estimate, $\bf{\hat{h}}$. A single-layer DNN is used to design the LS Refine Net. The refined channel estimate, $\bf{\hat{h}}$, from the LS Refine Net is then passed as input to the ZF detector sub-network in the signal detection module. The ZF detector sub-network is also fed with the received data vector, $\bf{y_{\text{D}}}$, and the LS channel estimate, $\bf{\hat{h}_{\text{LS}}}$. This is followed by the ZF Refine Net. The refined channel estimate, $\bf{\hat{h}}$, the received data vector, $\bf{y_{\text{D}}}$, and the ZF estimate, $\bf{\hat{x}_{\text{ZF}}}$, from the ZF detector sub-network are passed to the ZF Refine Net to estimate the transmit binary data, $\hat{b}$. The short dotted line from the LS sub-network to the ZF detector sub-network in Fig. 9 provides a robust candidate for the recovered binary symbols, $\hat{b}^{{}^{\prime}}$, at deQAM. Both the channel estimation and detection modules are trained using an $\ell_{2}$ loss between the estimated binary data, $\hat{b}$, and the raw transmit binary symbol, $b$. Simulations have shown that this design outperforms both LMMSE and FC-DNN [6].
Figure 9: ComNet architecture [255].
An artificial intelligence-aided receiver for a cyclic prefix-free OFDM system with joint detection and channel estimation is explored in [256]. Zhang et al. [257] unfolded the OAMP algorithm for generalized frequency division multiplexing (GFDM) systems with low-resolution analog-to-digital converters. In [258], a general EM algorithm is derived and unfolded to address the impact of channel estimation errors on signal detection. The expectation step estimates the CSI, while the maximization step detects the transmitted signal using the estimated CSI. A modified, trainable projected gradient detector is used in the maximization step. In [259], an ADMM-based unfolded algorithm is proposed for robust MIMO detection, incorporating channel estimation errors. To reduce error propagation due to channel aging, a robust data-aided Kalman filter-based channel tracking method is developed. The pre-estimated data symbols are used to improve channel estimation accuracy, enhancing robust detection performance.
Recently, the energy harvesting technique has gained significant attention as a promising feature for green communication networks. Ambient backscatter communication is an energy-efficient technology that allows devices to transmit data by reflecting existing ambient signals, such as those from Wi-Fi routers or cellular networks, instead of generating their own radio waves. A backscatter device modulates the signal over the ambient radio frequency without an active radio frequency component. However, the backscatter link is weaker than the direct link, which limits its coverage and data rate [260]. In [261], Wang et al. explored an IRS-based ambient backscatter communication system. The proposed network consists of two parts. The first subnetwork calculates the IRS phase shift matrix for binary symbols. The phase shift matrix weights, or phase angle vectors, are updated using a DNN training tool. These updated phase angle vectors are then used to compute the received signal, which is passed to the second subnetwork, the detection module. The EM algorithm is unfolded for symbol detection.
XI-B Joint Activity Detection and Channel Estimation
In [262], Ke et al. addressed massive access in massive MIMO systems by introducing an adaptive scheme for active user detection and channel estimation using compressive sensing techniques. This scheme takes advantage of the sparse traffic patterns of many user devices and the sparse nature of massive MIMO channels in the virtual angular domain. The access latency is reduced due to the use of non-orthogonal pseudo-random pilots for uplink communication. The problem of detecting active users and estimating their channels is formulated as a generalized multiple measurement vector compressive sensing problem. A generalized multiple measurement vector AMP (GMMV-AMP) algorithm is designed to leverage the structured sparsity in the uplink channel matrix. Additionally, a Turbo-GMMV-AMP algorithm is developed to exploit channel sparsity in both the spatial and angular domains. In [263], an EP-based joint active user detection and channel estimation technique for mMTC networks is proposed. The EP algorithm iteratively minimizes the Kullback-Leibler divergence between the target and approximate distributions via moment matching, efficiently finding a tractable close approximation of the target distribution.
Grant-free random access is used for active device detection at the BS to address the latency issue in massive access. The active device detection is formulated as a sparse signal recovery problem. To improve the performance of massive device detection, Qiang et al. [43] proposed a deep unfolded network called AMP-Net, as shown in Fig. 10. The basic principle of AMP-Net is based on the AMP algorithm. The input variables passed through the $T$ layers include the state evolution parameter, $\tau^{2}$, the device state matrix $\bf{X}$, the received signal $\bf{Y}$ at the BS, and the residual matrix $\bf{R}$. The pilot vectors are denoted by $\bf{a_{\text{n}}}$. Each layer contains an MMSE estimator, $\eta_{t}(\cdot)$. The set of trainable parameters consists of four scalar values. The parameters $\lambda_{n}^{t}$ and $\gamma_{n}^{t}$ estimate the activity probability and channel variance, respectively. The parameter $\theta_{n}^{t}$ regulates the denoiser, while $\mu_{n}^{t}$ controls the state evolution parameter, helping facilitate an accurate prediction of the detected target.
Figure 10: The structure of AMP-Net [43].
Traditional orthogonal multiple access fails to achieve massive connectivity for mMTC systems. Random and sporadic data traffic means that only a small fraction of users access the BS during uplink transmission. The joint channel estimation and active user detection problem for uplink mMTC is investigated in [264]. The authors developed a pilot transmission scheme for nonorthogonal identification. Due to the sporadic activity of users, where only a small fraction are active at any time, traditional orthogonal pilots are inefficient. Instead, nonorthogonal pilots are used, and the joint estimation and detection problem is formulated as a compressive sensing problem. The work considers single and multiple antenna configurations at the base station, modeling them as single measurement vector and multiple measurement vector problems, respectively. The authors designed a deep unfolding model based on the ADMM algorithm. For ADMM, the trainable parameters include: (1) an auxiliary variable to compute matrix inversions, (2) a regularization parameter that controls the penalty term in the ADMM optimization, influencing the balance between sparsity and accuracy in the recovered signal, and (3) a threshold parameter in the shrinkage function that determines the sparsity of the output in each iteration. In [265], two deep unfolded neural network architectures are proposed for sporadic user detection and channel estimation in mMTC. The first architecture unfolds a linearized ADMM, applied to a constrained convex problem. Linearized ADMM addresses computationally intractable block coordinate sub-problems by executing a single iteration of the proximal gradient method for each subproblem instead of solving them explicitly [156]. The second architecture unfolds a vector AMP with a denoiser based on ISTA. This unfolded network reduces complexity and improves accuracy compared to the original algorithm.
In [266], a deep unfolded algorithm is proposed for the joint design of compression modules, quantization modules at the APs, and the decompression and detection modules at the CPU for cell-free massive MIMO systems with capacity-limited fronthauls. The objective is to address activity detection in an end-to-end learning framework. Ma et al. [267] investigated joint device activity and data detection in mMTC. They proposed a one-phase non-coherent scheme, where data bits are embedded in the pilot sequences. The BS simultaneously detects active devices and their embedded data bits without explicit channel estimation. The AMP algorithm is unfolded to enhance detection performance by effectively exploiting the pilot activity correlation. In [268], a contention-based asynchronous random access scheme with data length diversity is considered for mMTC. The paper unfolds AMP with backpropagation to address the composite problem of activity detection, channel estimation, and data recovery. Zou et al. [269] formulated the joint activity detection and channel estimation problem as a group-row-sparse matrix recovery problem. A proximal gradient-based unfolding NN is developed to parameterize the algorithmic iterations and solve this problem. In [270], compressive sensing recovery algorithms are used to address the joint active device detection and channel estimation problem for uplink transmission in grant-free based vehicular networks. The ADMM algorithm is unfolded to improve convergence rate and recovery accuracy. Shi et al. [271] considered grant-free massive access in an IoT network. A large number of single-antenna IoT devices sporadically transmit non-orthogonal sequences to a multi-antenna BS. An unfolded DNN framework is developed using the multidimensional shrinkage thresholding operator inherited from the ISTA structure. This network simultaneously achieves low computational complexity and high robustness for solving the joint activity detection and channel estimation problem.
Cell-free technology for massive MIMO enhances mmWave signal coverage [272]. Using deep unfolding techniques, the authors in [273] developed a grant-free access protocol based on mmWave cell-free massive MIMO. The proposed hybrid structure leverages learned vector AMP to reduce the need for prior channel information and achieve faster convergence.
XI-C Joint Channel Estimation and Precoder Design
A codebook design for channel estimation in mmWave communications with a hybrid precoding structure is studied in [274]. To evaluate the performance of a codeword, the authors propose a metric called generalized detection probability (GDP), which considers both the beam pattern’s flatness and the maximum transmission power. Yuan et al. [275] introduced a two-stage wideband hybrid beamforming scheme for THz wireless systems. The model features a multi-antenna BS communicating with a multi-antenna user over frequency-selective fading. Additionally, a probabilistic robust hybrid beamforming scheme is designed to mitigate channel estimation errors.
(a) Long-term
(b) Short-term
Figure 11: JCEHB architecture [276].
Instead of modeling and optimizing hybrid beamforming and channel estimation separately, Kang et al. [276] proposed the joint channel estimation hybrid beamforming (JCEHB) algorithm. The sum rate is maximized for TDD massive MIMO system in a mixed timescale framework. The algorithm’s architecture is shown in Fig. 11. Observe from Fig. LABEL:sub@fig:fig1, the deep-unfolding NN architecture consists of two parts : channel estimation and hybrid beamforming. For channel estimation, the pilot training module is designed by unfolding the recursive least squares (RLS) algorithm. The SSCA algorithm is unfolded to replace the high complexity computation operations in hybrid beamforming. The long-term analog beamforming matrices, are optimized offline with CSI statistics that is constant over long time slots or a superframe. After fixing the analog beamformers, the deep-unfolding NN for channel estimation is trained at each time slot to optimize the digital beamformers, as shown in Fig. LABEL:sub@fig:fig2. In contrast to long-term analog beamforming, short-term digital beamforming matrices are optimized every time slot using real-time low-dimensional effective CSI. Analog beamforming matrices are optimized offline using CSI statistics, while digital beamforming matrices are updated within each time slot based on estimated low-dimensional CSI. The authors claimed that the hybrid approach in a mixed (long-term and short term) timescale framework reduce signaling overhead as compared to separate design of NN in a single timescale.
An unfolded WMMSE-based NN for massive multi-user MIMO under imperfect CSI is studied in [277]. The proposed network enhances performance with low complexity by leveraging CSI statistics for near-optimal precoder solutions, accelerating convergence. In [278], an unfolded sparse Bayesian learning algorithm is derived to capture channel sparsity in a mmWave massive MIMO system with hybrid beamforming. Similarly, [279] proposes a learning-based channel semantic acquisition and beamforming solution for cell-free massive MIMO. A multi-layer perceptron estimates the channel, while a successive over-relaxation-based linear beamforming scheme reduces estimation errors in hybrid beamforming. Wideband precoding under imperfect CSI is discussed in [280]. A robust frequency-selective precoding algorithm maximizes the expected weighted sum rate while considering per-antenna power constraints. This algorithm is unfolded to reduce computational complexity. To improve beamspace channel estimation, [281] introduces a deep unfolding algorithm using the top $\ell_{K,2}$ norm. This norm, a variation of the traditional $\ell_{2}$ norm, focuses on the largest elements in the sparse beamspace channel vector [282]. It helps determine sparsity levels, reducing noise and computational complexity. The iterative trimmed-ridge regression algorithm is unfolded to allow parallel computation. These algorithms are applicable to both narrowband and wideband massive MIMO systems, ensuring versatility in channel estimation.
XI-D Joint Detection and Decoding
An adaptive Bayesian receiver using sequential Monte Carlo for signal detection and decoding in fading channels with known channel statistics is proposed in [283]. The transmitted signal is treated as missing data, and multiple signal samples are sequentially estimated based on observed signals. These estimated sequences approximate the Bayesian estimate of both transmitted signals and channel states. Adaptive receiver algorithms are developed for both uncoded and convolutionally coded systems. El Gamal et al. [284] present an iterative receiver for joint detection and decoding of CDMA signals. The multiuser detector follows the MMSE criterion. A modified symbol-aided demodulation technique extracts soft information from previous decoding iterations to improve channel estimates, enhancing joint detection and decoding efficiency.
Traditionally, channel estimation, signal detection, and channel decoding in a MIMO receiver are performed sequentially and independently. This disjoint approach leads to suboptimal performance. To address this, Sun et al. [285] propose receivers that integrate these functions, reducing error propagation and processing delay. Their design targets MIMO ultra-reliable low-latency communication (URLLC) systems. The model uses short channel codes and a limited number of pilot symbols for channel estimation. The authors employ the ADMM algorithm to solve large-scale non-convex optimization problems, deriving closed-form solutions for each iteration. Two NNs, JCDDNet-G for Gaussian channels and JCDDNet-S for sparse mmWave channels, are built by unfolding the ADMM algorithms as shown in Fig. 12. Simulations show that the proposed trainable JCDD receivers significantly outperform traditional turbo receivers.
Figure 12: JCDDNet-S/G architecture [285].
In [286], Zhang et al. propose a receiver framework based on a deep unfolded NN structure. The joint signal detection and channel decoding modules are modeled as multi-layer deep feed-forward networks. Signal detection is achieved by unfolding the EP algorithm, while channel decoding is based on unfolded turbo decoding. Building on this work, Zhang et al. develop an unfolding-based framework for MIMO detectors in [287]. This framework adaptively determines trainable parameters, allowing the detector to adjust effectively to varying channel conditions.
XII Challenges in Developing Deep Unfolding Models and Future Research Scopes
The extensive review of recently emerging techniques and applications of deep unfolding have demonstrated its potential in addressing complex signal processing problems with reduced complexity and runtime. Despite these advantages, the adoption of deep unfolding in practical scenarios remains limited. In this section, we outline the main challenges in developing deep unfolding models, explore the reasons behind these challenges, and propose potential improvements to make deep unfolding a more powerful and widely applicable technique in wireless communications.
XII-A Limitations of Deep Unfolding Techniques
Although deep unfolding has found applications in various design problems, it has limitations. Most of these limitations arise from the fact that deep unfolding is not an independent design methodology; it must be built upon existing conventional optimization frameworks. Consequently, developing deep unfolding solutions faces the following challenges:
(C1) Contrasts between conventional optimization and deep unfolding: There is a notable contradiction between the development methodologies of conventional optimizers and deep unfolding models. Most conventional solutions for transceiver design and signal processing in wireless communications are based on convex optimization techniques. However, many wireless design problems are inherently non-convex and highly coupled due to factors such as fractional SINR, constant-modulus constraints (e.g., for analog beamforming), and mixed-integer variables (e.g., in symbol detection or antenna selection). These challenges are typically addressed using a two-layer approach. In the first layer, design variables are decoupled through alternating optimization (AO) or block coordinate ascent/descent, breaking the original problem into smaller sub-problems, each focusing on a subset of variables that are solved alternately. In the second layer, techniques such as SCA or the ADMM are applied to solve each sub-problem, often introducing additional variables. Although this two-layer approach transforms the original non-convex problem into a series of convex ones that can be solved using convex optimization tools like CVX, it significantly increases the number of variables and sub-problems and leads to nested loops in the solution process. These loops, corresponding to the two layers, must iterate until convergence, adding considerable complexity. In contrast, deep unfolding models typically approach the problem as a unified whole, solving all variables in parallel, facilitating the development of unfolding architectures. This fundamental difference makes it challenging to develop suitable algorithms for unfolding. Additionally, the increased complexity in variables and sub-problems inherent in conventional methods further complicates the design of deep unfolding frameworks, directly contributing to their overall complexity.
(C2) Challenges in unfolding principle algorithms: While deep unfolding typically enhances conventional iterative algorithms, not all can be effectively unfolded. This limitation arises from the core principle of deep unfolding, which structures the iterations of an iterative algorithm as layers within a DNN. Complex operations like matrix inversion, matrix decomposition, or the use of convex optimization tools such as CVX or YALMIP (with solvers like MOSEK, SeDuMi, etc.) are difficult to implement in a DNN framework. In large-scale systems such as massive MIMO, even matrix multiplications result in high complexity for unfolding models.
Another challenge is parameter generalization, particularly with respect to the number of layers in the unfolding model. In iterative algorithms with nested loops (discussed in C1), both inner and outer loops run until convergence, but deep unfolding models require the number of layers to be predetermined before deployment. This is typically done through empirical tuning, which may not provide optimal values for all channel realizations. As a result, convergence is not guaranteed across different scenarios, potentially leading to performance degradation. Proper tuning is essential to achieve a favorable performance-complexity tradeoff in deep unfolding models.
(C3) Challenge in handling design constraints and objectives: One of the biggest challenges in developing deep unfolding models lies in managing the design constraints formulated in the problem. These constraints require the DNN layers’ outputs to meet specific requirements, such as total power, QoS, or other element-wise constraints. Simple constraints, like power or constant modulus, can often be enforced through normalization [118]. For instance, the $tanh$ activation function can be employed to impose constant-modulus constraints on analog beamforming coefficients or sensing waveforms, similar to black-box DL architectures. However, QoS constraints are significantly more complex. While they resemble activation functions, none of the existing activation functions can adequately manage such intricate requirements. One approach is to incorporate these constraints into the design objective, but this method often results in a problem that is not equivalent to the original constrained one, making the solution potentially infeasible. Properly modeling such constraints within a deep unfolding DNN remains highly challenging, hindering the development of more advanced deep unfolding models.
Not only do complex constraints pose challenges for developing deep unfolding models, but intricate objective functions also present difficulties. Generally, deep unfolding models are trained based on the objective function. For instance, unfolding architectures like DetNet [49], FSNet [152], and ScNet [288] for symbol detection are trained to minimize the Euclidean distance between the model’s output (symbol vector) and the true transmitted symbol vector. When labels are unavailable, as in beamforming problems, unfolding networks use the objective function as the loss function for minimization or its negative for maximization problems. Typical deep unfolding techniques [84, 75] require calculating the gradient of the objective function. However, this becomes challenging when the objective function is a complex, high-order, or fractional function of the design variables, especially when those variables are vectors or matrices. Therefore, more dynamic methodologies are needed for developing unfolding models that can handle such intricate functions efficiently.
(C4) Need for black-box models in sophisticated mappings: While deep unfolding is powerful, certain inference tasks still require black-box DNN models, especially when there is no clear relationship between the input data, output solution, and performance metrics. For instance, in multi-modal sensing-assisted beam tracking, tasks may involve complex mappings from environmental data—such as sensory information from sensors and cameras—to the optimal beam [289, 290]. In such cases, there is no explicit relationship between environmental images and the optimal beam, and standard communication metrics like throughput or SINR cannot be directly expressed as functions of these images. These problems cannot be formulated as optimization tasks solvable with principle algorithms, making it difficult to develop unfolding models that map environmental information to beam patterns or blockage states.
The above three challenges C1C4 also outline the scenarios wherein deep unfolding models cannot be built or does not offer much gain in performance-complexity tradeoff. These include when there is lack of suitable conventional iterative algorithms (C1, C2, C4) or when the design problems are highly complicated with constraints and objective function (C3). It can be seen that deep unfolding is not a game changer that will completely replace conventional principle algorithms or black-box ML models. However, the applicability of deep unfolding can be enhanced by overcoming those challenges, as will be discussed next.
XII-B Development of Deep Unfolding Techniques
XII-B1 Coexistence of Deep Unfolding Models, Conventional Solvers, and Black-Box Models
Deep unfolding doesn’t need to fully replace conventional algorithms or black-box ML models. Instead, these approaches can coexist, leveraging their combined strengths. In hybrid designs, deep unfolding can (i) reduce the number of sub-problems, (ii) lower the complexity of solving each sub-problem, and (iii) enhance the learning process in black-box models. Benefits (i) and (ii) correspond to challenges C1 and C2, while (iii) relates to C4.
Benefit (i) arises because deep unfolding can rely on techniques like PGA/PGD, which avoids transforming the original problem into convex sub-problems. Benefit (ii) comes from the fact that unfolded models typically exhibit lower complexity compared to methods like SCA, and their runtime is usually shorter than that of solvers like CVX. For example, in the AO framework (discussed in C1), the most challenging sub-problem often dominates the overall complexity. By applying an efficient deep unfolding model to this sub-problem, the overall efficiency improves, while simpler sub-problems can be handled by conventional solvers like CVX. Benefit (iii) is frequently seen when deep unfolding is used to preprocess input data for black-box models, as in multimodal sensing applications. The combination of deep unfolding structures with conventional solvers and black-box models leads to simpler yet more efficient designs.
XII-B2 Deep Unfolding for Multi-Objective Learning
Most existing deep unfolding models are developed for single-objective designs, such as maximizing throughput, SINR, or minimizing BER. However, in many scenarios, system designs involve multiple objectives. Traditional approaches often formulate these as constrained problems, such as maximizing energy efficiency under a QoS constraint (e.g., throughput or SINR). In ISAC designs, weighted sums of communication and sensing performance metrics are used when the goal is to balance both functions without prioritizing one over the other. The complexity grows in multi-static ISAC systems, where distributed sensing receivers have different performance metrics depending on their functions, such as SCNR, mutual information, Cramér-Rao lower bound, or probability of detection.
These designs are often difficult to solve using traditional constrained formulations, as the problem can easily become infeasible if constraints are not properly tuned. However, by incorporating all design metrics into the objective function with appropriate weights, to form a multi-objective design, the feasibility of the problem can be significantly improved [291, 84]. This multiobjective formulation also makes it easier to develop unfolding models, optimizing multiple objectives simultaneously, and improving performance trade-offs. However, developing unfolding models for multi-objective designs presents several challenges. First, as the number of terms in the objective function increases, training the unfolding DNN becomes more difficult. This is because the loss function is typically set to mirror the objective function (or its negative in maximization problems). Furthermore, balancing the changes of these objectives with respect to the design variables across layers is crucial [84, 292], necessitating careful design of the unfolding models.
XII-B3 Distributed Deep Unfolding Architectures
While deep unfolding models provide advantages in terms of complexity due to faster convergence compared to conventional iterative algorithms, they can still face high complexity issues due to challenge C2. This limits their applicability on platforms with constrained computational resources, such as edge devices or distributed access points (APs). Future wireless networks will advance beyond simple communications, evolving into large-scale, distributed systems that integrate advanced sensing, communications, and computing capabilities. However, most existing AI models, both data-driven and model-driven, were designed for centralized scenarios, where they are deployed at a central node with substantial computing power. These models are not well-suited for distributed scenarios, highlighting the need for the development of appropriate distributed unfolding architectures.
One potential solution is to develop distributed deep unfolding architectures. In this approach, rather than executing all the unfolded layers at every distributed node or edge device, the layers can be distributed across multiple platforms for parallel execution, thereby reducing latency. For example, to design $L$ beamformers for $L$ APs for their downlink transmission in a distributed system, a CPU with high computing power and hardware resources can handle the majority of the layers, generating high-quality beamformers. These beamformers are then sent to the APs, serving as initial solutions and inputs for smaller sub-unfolding models at the distributed APs. This ensures that reliable beamformers can be generated at the APs with only a few layers, minimizing complexity and runtime to meet the computational constraints of the APs. Conversely, in the uplink, a distributed unfolding detection model can be deployed at the APs to generate initial detected symbols, which are then forwarded to the CPU for further processing and final decision-making by the large unfolding detector. By such an unfolding framework, highly reliable symbols are generated with low complexity at distributed APs.
XII-B4 Deep Unfolding to Enhance Black-box Solution
As its name suggests, a black-box ML model operates with a black-box nature, making it unexplainable and resistant to re-structuring beyond tuning parameters like the number of layers or nodes. The performance and efficiency of such models are heavily dependent on the quality of input data. Deep unfolding can contribute by performing pre-processing steps to improve the input data quality. One example is in multi-modal sensing-assisted communication tasks, where data from distributed cameras, LiDAR, and radar sensors are used to enhance communication performance, such as in beam tracking and blockage prediction. However, significant challenges arise in collecting, transferring, and processing large sensory data. Specifically, high-resolution camera images generate considerable traffic overhead and latency, impacting real-time communication processing.
Deep unfolding models can be particularly useful in pre-processing such images. In image processing, deep unfolding has been successfully applied to tasks such as denoising, super-resolution, and image reconstruction, allowing for high-quality outcomes with fewer iterations than conventional algorithms. For example, unfolded algorithms based on the ISTA for sparse coding have been shown to significantly accelerate image reconstruction while maintaining high-quality results [293, 294, 295]. This makes deep unfolding a valuable tool for handling large-scale sensory data in real-time communication systems.
XIII Conclusions
The integration of traditional algorithmic principles with DL techniques has led to the development of deep unfolding, a method that overcomes the limitations associated with both conventional algorithms and standalone DL models in wireless communication. This article offers a comprehensive survey of deep unfolding techniques, examining their application across various facets of the physical layer in wireless communication systems. Key applications, such as signal detection, channel estimation, precoder design, sensing, communication decoding for error-correcting codes, power allocation, and security, are explored, with an emphasis on technologies anticipated for beyond-5G networks. The survey highlights the strengths of deep unfolding in addressing complex challenges, particularly in achieving low-latency, high-accuracy performance in dynamic environments. We conclude by discussing the key challenges in developing deep unfolding models, including issues related to scalability, computational efficiency, and robustness across diverse scenarios. Finally, we propose potential advancements to enhance the adaptability and resilience of deep unfolding, with the goal of solidifying its role as a robust and versatile technique for next-generation wireless communications.

### v3

@@
 Abstract
-The application of machine learning in wireless communications has been extensively explored, with deep unfolding emerging as a powerful model-based technique.
-Deep unfolding enhances interpretability by transforming complex iterative algorithms into structured layers of deep neural networks (DNNs).
-This approach seamlessly integrates domain knowledge with deep learning (DL), leveraging the strengths of both methods to simplify complex signal processing tasks in communication systems.
-To provide a solid foundation, we first present a brief overview of DL and deep unfolding.
-We then explore the applications of deep unfolding in key areas, including signal detection, channel estimation, beamforming design, decoding for error-correcting codes, sensing and communication, power allocation, and security.
+The massive surge in device connectivity demands higher data rates, increased capacity with low latency and high throughput.
+Hence, to provide ultra-reliable, low-latency communication with ubiquitous connectivity for Internet-of-Things (IoT) devices, next-generation wireless communication leverages the incorporation of machine learning tools.
+However, standard data-driven models often need large datasets and lack interpretability.
+To overcome this, model-driven deep learning (DL) approaches combine domain knowledge with learning to improve accuracy and efficiency.
+Deep unfolding is a model-driven method that turns iterative algorithms into deep neural network (DNN) layers.
+It keeps the structure of traditional algorithms while allowing end-to-end learning.
+This makes deep unfolding both interpretable and effective for solving complex signal processing problems in wireless systems.
+We first present a brief overview of the general architecture of deep unfolding to provide a solid foundation.
+We also provide an example to outline the steps involved in unfolding a conventional iterative algorithm.
+We then explore the application of deep unfolding in key areas, including signal detection, channel estimation, beamforming design, decoding for error-correcting codes, integrated sensing and communication, power allocation, and physical layer security.
 Each section focuses on a specific task, highlighting its significance in emerging 6G technologies and reviewing recent advancements in deep unfolding-based solutions.
@@
 Index Terms:
-Deep learning (DL), deep unfolding, model-driven machine learning, wireless communication signal processing, ISAC, IRS, THz, mmWave.
+Deep learning, Deep unfolding, Model-driven machine learning, Wireless communication signal processing, Integrated sensing and communication (ISAC), Intelligent reflecting surfaces (IRS), Terahertz, IoT.
 I Introduction
@@
 Given the inherent limitations of available resources, problem formulation must prioritize flexibility, adaptability, and efficient utilization.
-Significant advancements in physical layer technologies have enhanced wireless communication systems, including massive multiple-input multiple-output (MIMO), intelligent reflecting surfaces (IRS), millimeter-wave (mmWave) communication, physical layer security, and non-orthogonal multiple access (NOMA).
-Additionally, orthogonal time-frequency space (OTFS) modulation is emerging as a promising alternative to orthogonal frequency-division multiplexing (OFDM), further improving spectral efficiency and resilience in next-generation networks.
-In the communication system, traditional model-based analytical algorithms often employ iterative optimization techniques to achieve accurate solutions.
-Increasing the number of iterations improves performance as the algorithm converges closer to the optimal solution.
-Deriving an optimal solution in the form of a closed-form expression can be highly challenging in specific arbitrary communication scenarios.
-Hence, we resort to a sub-optimal solution in the traditional model-based method.
-Moreover, practical implementation on hardware imposes significant constraints on computational resources, latency, and power consumption, which limit the feasibility of deploying highly iterative methods.
-The classical iterative approach often relies on mathematically tractable models and algorithms for specific scenarios.
-Hence, adapting to dynamic and heterogeneous conditions, such as rapidly changing channels, interference, and diverse network configurations, can be challenging.
-Machine learning techniques offer a transformative approach by leveraging data-driven learning to extract patterns, adapt to new conditions, and approximate solutions efficiently.
-These methods can generalize across varying scenarios, learn optimal strategies from experience, and enable the design of intelligent communication systems that operate effectively in complex, real-world environments.
-These methods can potentially learn task-specific patterns and approximate solutions with significantly reduced computational complexity.
-Machine learning models can provide efficient and scalable alternatives to traditional iterative algorithms, enabling better performance under hardware limitations.
-This paradigm shift has opened new avenues for addressing complex problems in communication systems, bridging the gap between theoretical advancements and practical deployment.
-Deep Learning (DL), a machine learning subclass maps the relationship between input and output without explicitly modeling it in the architecture [4].
-It involves the use of deep neural networks (DNN) with multiple hidden layers (hence “deep”) to analyze and interpret data.
-Numerous architectures have been proposed such as multilayer perceptron (MLP), convolutional NN (CNN), deep reinforcement learning (DRL), generative adversarial networks (GANs), etc.
+Significant advancements have taken place in wireless communication due to various exciting physical layer technologies, including massive multiple-input multiple-output (MIMO), Internet-of-Things (IoT), integrated sensing and communication (ISAC), intelligent reflecting surfaces (IRS), millimeter-wave (mmWave) communication, physical layer security, and next-generation multiple access such as nonorthogonal multiple access (NOMA).
+Additionally, orthogonal time-frequency space (OTFS) modulation is emerging as a promising alternative to orthogonal frequency-division multiplexing (OFDM), further improving spectral efficiency (SE) and resilience in next-generation networks.
+Deep learning (DL), a machine learning subclass, maps the relationship between input and output without explicitly modeling it in the architecture [4].
+It involves the use of deep neural network (DNN) with multiple hidden layers (hence “deep”) to analyze and interpret data.
+Numerous architectures have been proposed such as multilayer perceptron (MLP), convolutional neural network (CNN), deep reinforcement learning (DRL), generative adversarial networks (GANs), etc.
 DNNs have been applied to various components of wireless communication, including detection, channel estimation, beamforming designs, and power allocation [5, 6, 7, 8, 9].
-DL architectures exhibit several inherent limitations.
-These include complex training procedures and a black-box nature.
+However, DL architectures exhibit several inherent limitations.
+These include complex training procedures and a black-box nature of algorithms.
 The training process for these models requires extensive datasets, considerable time, and intricate fine-tuning.
 The complexity of implementing DNN architectures often constrains their applicability in real-world systems.
-Therefore, DL models are highly sensitive to the specific configurations of wireless systems, such as the number of antennas, users, and operating frequencies, requiring retraining or reimplementation whenever these parameters change.
-These challenges are not typically encountered in conventional optimization techniques, underscoring the potential advantages of using traditional algorithms [10, 11].
+Therefore, DL models are highly sensitive to the specific configurations of wireless systems, such as the number of antennas, users, and operating frequencies, which require retraining or reimplementation whenever these parameters change.
+These challenges are not typically encountered in conventional optimization techniques, underscoring the potential advantages of using classical algorithms [10, 11].
 Communication problems are inherently influenced by dynamic physical constraints that evolve in real-time.
@@
 Different communication scenarios impose varying demands on bandwidth, power consumption, and latency, which fluctuate with user density, data rates, and service types.
-Interpreting and addressing these constraints using conventional DNNs presents significant challenges.
-Traditional data-driven neural network architectures were developed primarily to process language and vision data, rendering them suboptimal for wireless communication tasks that involve channels, signals, and outputs governed by complex constraints.
+Interpreting and addressing these constraints using DNNs presents significant challenges.
+Data-driven neural network (NN) architectures were developed primarily to process language and vision data, rendering them suboptimal for wireless communication tasks that involve channels, signals, and outputs governed by complex constraints.
 Although certain constraints can be handled through the application of specific activation functions, for example, employing the $tanh$ function to enforce a constant modulus constraint in analog beamforming, standard activation functions generally lack the versatility to manage the intricate constraints characteristic of wireless communication systems.
-To mitigate these challenges, model-driven DL approaches facilitate the development of problem-specific network architectures and customized activation functions, optimized for the unique signal processing requirements of wireless communication applications [12].
+To mitigate these challenges, model-driven DL approaches facilitate the development of problem-specific network architectures and customized activation functions, optimized for the unique signal processing requirements of wireless communication [12].
 Compared to the conventional fully-connected DL architectures, model-driven DL have fewer trainable parameters, implying less training time and computational complexity.
 Deep unfolding is an approach that falls into the category of model-driven DL [13].
-The main motive behind developing a hybrid model (model-driven DL) for decision mapping, is to utilize the benefits of both optimization and data-driven DL.
-Model-driven DL is categorized into three frameworks as given in [13] : (1) learned optimizers [14, 15] (2) deep unfolding [16] (3) DNN-aided optimizers [17].
-Here, we will focus on the second category i.e., deep unfolding.
-A deep unfolded network can be trained in an easier and faster fashion with smaller datasets than conventional DNNs.
+There are several types of model-driven DL approaches, including deep unfolding [14],[15], physics informed NN [16, 17], learned optimizers [18, 19], etc.
+A deep unfolded network can be trained in an easier and faster fashion with smaller datasets than DNNs.
 This is because deep unfolding incorporates domain knowledge from iterative algorithms into layers of DNNs, offering improved performance and interpretability [20].
-The main idea in solving communication signal processing tasks lies in optimizing key hyperparameters of conventional algorithms based on data training.
-This includes elements such as step sizes in gradient-based methods, initial solutions, searching directions, criteria for early termination, and other relevant parameters.
-Deep unfolding uses suitable loss functions to train learnable parameters that are interpretable.
-Each iteration of the iterative algorithm is incorporated as a hidden layer of the DNN.
-These layers are then stacked to form a network, mimicking the optimization flow of algorithms.
-Furthermore, in many cases, the trainable parameters are independent of the system configurations, implying that the trained unfolding model can work for various scenarios.
+TABLE I: Acronyms and corresponding full names.
+Abbreviation Description
+AO Alternate optimization
+AP Access point
+AMP Approximate message passing
+ADC Analog-to-digital converters
+ADMM Alternating direction method of multipliers
+BP Belief propagation
+BS Base station
+BER Bit error rate
+CD Coordinate descent
+CNN Convolutional neural network
+CSI Channel state information
+CGD Conjugate gradient descent
+DL Deep learning
+DNN Deep neural network
+DRL Deep reinforcement learning
+EM Expectation maximization
+EP Expectation propagation
+FDD Frequency division duplexing
+GAN Generative adversarial network
+GNN Graph neural network
+HBF Hybrid beamforming
+IRS Intelligent reflecting surface
+IoT Internet-of-things
+ISI Inter-symbol interference
+ISAC Integrated sensing and communication
+ISTA Iterative shrinkage-thresholding algorithm
+LR Low-resolution
+LS Least squares
+LLR Log-likelihood ratio
+LDPC Low-density parity check
+LMMSE Linear minimum mean square error
+MF Matched filter
+ML Maximum likelihood
+MO Manifold optimization
+MU Multi-user
+MAP Maximum a posteriori
+MPD Message passing detector
+MSE Mean square error
+MMSE Minimum mean square error
+MIMO Multiple input multiple output
+MISO Multiple input single output
+MSER Minimum symbol error rate
+mMTC Massive machine type communication
+mmWave millimeter-wave
+NN Neural network
+NOMA Non-orthogonal multiple access
+NMSE Normalized mean square error
+OAMP Orthogonal approximate message passing
+OFDM Orthogonal frequency division multiplexing
+OTFS Orthogonal time frequency space
+PGA Projected gradient ascent
+PGD Projected gradient descent
+QAM Quadrature amplitude modulation
+QoS Quality of service
+QPSK Quadrature phase shift keying
+RF Radio frequency
+RNN Recurrent neural network
+RSMA Rate splitting multiple access
+SD Sphere decoder
+SE Spectral efficiency
+SER Symbol error rate
+SCA Successive convex approximation
+SGD Stochastic gradient descent
+SIC Successive interference cancellation
+SNR Signal-to-noise ratio
+SSCA Stochastic successive convex approximation
+SINR Signal-to-interference and noise ratio
+THz Terahertz
+TPGD Trainable projected gradient descent
+WMMSE Weighted minimum mean square error
+ZF Zero forcing
 I-A Related Works
-The growing application of deep unfolding in communication systems highlights the need for a comprehensive survey in this area.
+The growing adoption of deep unfolding in communication systems has led to several surveys covering its diverse applications.
 Wang et al.
-in [19] presented a review of DL and deep unfolding approaches for addressing physical layer signal processing challenges, including MIMO detection, channel estimation, and channel decoding.
-The authors of [20] provided an overview of deep unfolding applications in transceiver design, MIMO detection, and CSI estimation and feedback.
-Similarly, in [21], the authors highlighted the role of DL in communication, with a focus on DNNs, GANs, RNNs, and end-to-end learning.
-This work also explored deep unfolding for Bayesian optimal estimation, using MIMO detection and channel estimation as case studies.
-The authors in [22] focused primarily on MIMO detection while exploring its use in multiuser MIMO precoding and belief propagation-based channel decoding.
-A survey on model-based, DL, and hybrid approaches in the context of wireless communication networks is provided in [23].
-In [16], Monga et al.
-conducted a survey of deep unfolding algorithms specifically implemented for applications in signal and image processing.
-The authors in [24] reviewed the implementation of deep unfolding in signal processing tasks.
-A brief review on deep unfolding in MIMO detection is presented in [17].
-The authors in [13] have presented a general framework for optimization that leverages the benefit of deep unfolding with signal detection as one of the case studies.
-In [12], a chapter is dedicated to model-based machine learning for communication.
-The deep unfolding networks have been discussed for signal detection, channel estimation, and channel coding.
-Machine learning-based optimization framework is discussed in [25] with an emphasis on deep unfolding in specific design paradigms such as signal detection.
-A comparative study between purely data-driven and deep unfolding is conducted with complete emphasis on hybrid beamforming optimization in [26].
-An overview of existing surveys on deep unfolding is shown in TABLE I.
+[21] reviewed DL and deep unfolding methods for physical layer tasks such as MIMO detection, channel estimation, and decoding.
+The authors of [22] discussed its use in transceiver design and channel state information (CSI) estimation.
+In [23], deep learning methods including DNNs, GANs, and recurrent NN (RNN) were reviewed, with deep unfolding applied to Bayesian estimation.
+The authors of [15] focused on MIMO detection, multiuser precoding, and belief propagation.
+A broader comparison of model-based, DL, and hybrid methods is found in [24].
+Monga et al.
+[14] focused on algorithm unrolling for signal/image processing.
+Additional works [25, 13] explored signal processing and optimization.
+Eldar et al.
+[12] and Shi et al.
+[26] addressed DL-based communication frameworks.
+In [27], the authors highlight the advantages of deep unfolding for physical receiver design, with particular emphasis on its applications in signal detection and channel estimation.
+A comparison between data-driven and unfolding-based beamforming appears in [28] and Table II.
 I-B Contributions
-Existing surveys lack coverage of recent studies addressing critical 6G technologies, such as mmwave spectrum, new air interface that includes massive MIMO, IRS, next generation multiple access, and are often limited to specific application areas within the physical layer.
+Existing surveys lack coverage of recent studies addressing critical 6G technologies, such as mmWave spectrum, new air interface that includes massive MIMO, IRS, next-generation multiple access, and are often limited to specific application areas.
 Consequently, a comprehensive survey is needed to encompass the broad applications of deep unfolding techniques across various aspects of the physical layer in wireless communication.
 This paper provides an in-depth review of deep unfolding methods applied to the physical layer of wireless communication.
-The contributions are summarized below.
+The contributions are summarized below:
   1. 1.
-    We provide a comprehensive overview of deep unfolding algorithms and their applications in physical layer wireless communication.
-Part A of the paper introduces the fundamentals of DL and deep unfolding to establish a foundation, outlining their general structures.
-We describe the DL architectures, including MLP, CNN, RNN, DRL, GNN, and GAN.
-Additionally, we detail a systematic, step-by-step approach for developing deep unfolding algorithms.
-Finally, we highlight the advantages and motivations for adopting hybrid approaches to tackle optimization tasks effectively.
+    This paper presents a systematic, step-by-step approach for developing deep unfolding algorithms, with a detailed tutorial on the alternating direction method of multipliers (ADMM)-based MIMO detection.
+The tutorial offers insights into the ADMM unfolding process, identifies trainable parameters, and demonstrates the design of the resulting unfolded network.
   2. 2.
-    We survey the application of deep unfolding in the areas of signal detection, channel estimation, precoder design, sensing and communication, decoding, power allocation, and physical layer security in Part B.
-Within each section, we present these areas with the key 6G technologies (like IRS, mmWave, etc.) used to aid readers to design frameworks that adapt to future technologies.
-Additionally, this survey explores the application of deep unfolding in integrated sensing and communication (ISAC).Also, we discuss the joint techniques in the physical layer in Part C.
-The different areas are joint detection and channel estimation, joint activity detection and channel estimation, joint channel estimation and precoder design, and joint detection and decoding.
+    We survey the application of deep unfolding across various domains, including signal detection, channel estimation, precoder design, ISAC, decoding, power allocation, and physical layer security.
+Each section highlights the integration of deep unfolding with key 6G technologies such as IRS, mmWave, and IoT, providing readers with guidance on designing frameworks adaptable to future wireless systems.
   3. 3.
-    We outline a set of challenges in developing deep unfolding models and propose future directions that could guide advancements in deep unfolding techniques to enhance their performance.
-TABLE I: Existing Survey Papers on Deep Unfolding for Communication.
+    In addition, each section concludes with a comparative analysis table that evaluates prominent unfolded networks across key metrics such as computational complexity, generalization capability, and convergence behavior.
+These tables highlight the practical trade-offs and deployment boundaries associated with each architecture.
+Furthermore, we discuss the key takeaways to provide readers with deeper insights into the practical applicability and design considerations of deep unfolding in physical layer communication.
+  4. 4.
+    This paper concludes the survey by outlining a set of challenges in developing deep unfolding models and propose future directions that could guide advancements in deep unfolding techniques to enhance their performance.
+TABLE II: Existing survey papers on deep unfolding for communication.
 SNo. Reference Year Detection Channel Estimation Precoder Design Sensing and Communication Decoding in Error Correcting Codes Power Allocation Security
@@
 13 Artificial Intelligence-Empowered Hybrid Multiple-Input/Multiple-Output Beamforming: Learning to Optimize for High-Throughput Scalable MIMO [28] 2024 $\checkmark$
+14 Model and Data Dual-Driven: A Promising Receiver Design for 6G Wireless Communications [27] 2025 $\checkmark$ $\checkmark$
 15 This Survey $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
 This survey paper is organized as follows:
-  • Part A: We discuss the general structure of DNN and deep unfolding models.
-Section II covers the architecture of DNN and outlines the steps in designing a DNN.
-Section III presents the deep unfolding framework, which combines classical iterative algorithms with DNNs.
-  • Part B: The applications of deep unfolding in different physical layer signal processing tasks are thoroughly reviewed.
-Specifically, Section IV explains signal detection with perfect knowledge of channel gains in the receiver, followed by channel estimation in Section V.
-Section VI covers the precoder design.
-Section VII reviews different deep unfolding approaches for ISAC.
-The decoding for error correction codes is discussed in Section VIII, followed by power allocation in Section IX and security aspects in Section X.
-  • Part C : Section XI surveys the joint techniques in the physical layer.
-In this section, we discuss the application of deep unfolding in different joint techniques such as joint detection and channel estimation, joint activity detection and channel estimation, joint channel estimation and precoder design, and joint detection and decoding.
-  • Part D : Section XII discusses the challenges in adopting the deep unfolding approach in wireless communication and the future scopes of research. Section XIII concludes the paper.
-TABLE II: Acronyms and Corresponding Full Names.
-Abbreviation Description Abbreviation Description
-AMP Approximate message passing ADMM Alternating direction method of multipliers
-BP Belief propagation BS Base station
-BER Bit error rate CIR Channel impulse response
-CNN Convolutional neural network CSI Channel state information
-DL Deep learning DNN Deep neural network
-DRL Deep reinforcement learning EM Expectation maximization
-EP Expectation propagation FDD Frequency division duplexing
-FFT Fast Fourier transform FDMA Frequency division multiple access
-GAN Generative adversarial network GNN Graph neural network
-HetNets Heterogeneous networks IRS Intelligent reflecting surface
-IoT Internet of things ICI Inter-carrier interference
-ISI Inter-symbol interference ISAC Integrated sensing and communication
-ISTA Iterative shrinkage-thresholding algorithm JCEHB Joint channel estimation hybrid beamforming
-JCDD Joint channel estimation detection and decoding LLR Log-likelihood ratio
-LDPC Low-density parity check LPBN Long term passive beamforming
-LMMSE Linear minimum mean square error LR Low resolution
-LS Least squares LMS Least mean squares
-LSTM Long short term memory MAP Maximum a posteriori
-MDP Markov decision process MF Matched filtering
-MO Manifold optimization MM Majorization minimization
-MS Mobile station MMSE Minimum mean square error
-MIMO Multiple input multiple output ML Maximum likelihood
-MPD Message passing detector MSE Mean square error
-MSER Minimum symbol error rate mMTC Massive machine type communication
-NN Neural network NOMA Non-orthogonal multiple access
-OAMP Orthogonal approximate message passing OFDM Orthogonal frequency division multiplexing
-OMP Orthogonal matching pursuit OTFS Orthogonal time frequency space
-PGA Projected gradient ascent PGD Projected gradient descent
-RF Radio frequency RNN Recurrent neural network
-RL Reinforcement learning RSMA Rate splitting multiple access
-SSCA Stochastic successive convex approximation SABN Short term active beamforming network
-SBRNN Sliding bidirectional recurrent neural network SCDMA Sparse code division multiple access
-SE Spectral efficiency SGD Stochastic gradient descent
-SIC Successive interfernce cancellation SISO Single-input single-output
-SIMO Single-input multiple-output SINR Signal-to-interference and noise ratio
-SNR Signal-to-noise ratio TAS Transmit antenna selection
-TDD Time division duplexing THz Tera Hertz
-UAV Unmanned aerial vehicle URLLC Ultra reliable low latency communication
-WMMSE Weighted minimum mean square error ZF Zero forcing
+  • Part A: Section II presents the deep unfolding framework, which combines classical iterative algorithms with DNNs.
+The understanding of deep unfolding is exemplified via ADMM based MIMO detection.
+A step-by-step approach is presented for developing the unfolding of ADMM specifically for MIMO detection.
+  • Part B: This part explores deep unfolding applications in physical layer tasks: signal detection (Section III), channel estimation (Section IV), precoder design (Section V), ISAC (Section VI), decoding for error correcting code (Section VII), power allocation (Section VIII), and security (Section IX).
+  • Part C : Section X discusses the challenges in adopting the deep unfolding approach in wireless communication and the future scope of research. Section XI concludes the paper.
 I-D Notation and Acronyms
 […]
-II Deep Neural Network: General Structure
-Artificial intelligence originated with the famous Turing test in the 1950s, marking a breakthrough in machine thinking capability [28].
-Over the decades, around the 1990s, machine learning formed a subclass under artificial intelligence.
-Machine learning requires manual feature extraction of structured data.
-Most machine learning algorithms can be categorized into supervised, unsupervised, and reinforcement learning [4].
-Figure 2: Feedforward NN (MLP) architecture.
-Supervised learning aims to learn how to map inputs to outputs to predict labels for new data.
-In unsupervised learning, the dataset includes many features that are not labeled.
-The algorithm tries to identify hidden patterns within the data without the guidance of labeled examples.
-The third category, reinforcement learning, consists of an agent that learns to make decisions based on the environment.
-Since there is no predefined dataset, the agent interacts with the environment and learns from the feedback it receives for every action [4],[29].
-The output of a machine learning algorithm can be of two types: regression and classification.
-In the 2010s, DL emerged as a crucial technology in designing intelligent networks with unstructured data [30].
-Researchers have adopted DL techniques to solve complex problems.
-It has been widely used in computer vision, natural language processing, and image processing.
-This success has inspired researchers to use DL in the physical layer of wireless communication, especially in various signal-processing tasks.
-An NN consists of interconnected layers of neurons that work together to learn patterns from the given data.
-Different types of artificial NN architectures have been developed for various learning tasks, such as feed-forward NNs, CNN, recurrent NN (RNN), graph NN (GNN), etc.
+II Deep Unfolding: General Architecture and Fundamentals
+Deep unfolding, or algorithm unrolling [14], bridges model-based iterative algorithms and data-driven DNNs [10], [22], [31], [32], [33], [34], [35], [36].
+It maps each iteration of an iterative algorithm to a DNN layer, enabling end-to-end learning while retaining interpretability.
+In deep unfolding, the learnable parameters include model-specific variables and regularization coefficients, optimized layer by layer.
 Fig.
-2 shows the feedforward NN architecture with $K$ ($K=2$ here) hidden layers and $N_{k}$ neurons in layer $k$, for all $k=0,\dots,K$.
-The architecture has an input layer, hidden layers, and an output layer.
-We now discuss the steps in a feedforward NN:
-  1. 1.
-    The network begins with an input vector
-    ${\bf{s}}=[s_{1},s_{2},\dots,s_{n}]$, at the input layer where $n$ is the number of input features.
-  2. 2.
-    For each neuron in a layer, the weighted sum of inputs is computed:
-    $z^{(k)}_{j}=\sum_{i=1}^{n}w_{ji}^{(k)}s_{j}+b_{j}^{(k)},$
-    where $w_{ji}^{(k)}$ is the weight connecting neuron $i$ to neuron $j$ in the $k^{\text{th}}$ layer and $b_{j}^{(k)}$ is the bias for neuron $j$ at the $k^{\text{th}}$ layer.
-  3. 3.
-    A non-linear activation function $\sigma(\cdot)$ is applied:
-    $a_{j}^{(k)}=\sigma(z^{(k)}_{j}),$
-    where $a_{j}^{(k)}$ is the activation output of the neuron $j$ in the layer $k$. Common activation functions include ReLU, sigmoid, tanh, softmax etc. [4].
-  4. 4.
-    At the output layer, after applying the weighted sum and activation process, the output is given by:
-    $\hat{\bf{y}}=\sigma(\mathbf{W}^{K}\mathbf{s}^{K}+\mathbf{b}^{K})$
-    where ${\bf{W}}$ is the weight matrix at the output layer and $\bf{b}$ is the bias vector at the output layer.
-The activation function, ($\sigma(\cdot)$) for the output often depends on the task, whether it is a classification or regression task.
-  5. 5.
-    The loss is computed to measure the network’s performance:
-    $\mathcal{L}(\hat{\bf{y}},\bf{y})$
-    where $\hat{\bf{y}}$ is the predicted output, and $\bf{y}$ is the true output. The common loss functions are mean squared error (MSE), binary cross entropy, categorical cross entropy, etc.
-To train the NN, backpropagation is used by adjusting the parameters of the network, i.e., the weight and biases to minimize the loss.
-The parameters are updated using an optimization algorithm like stochastic gradient descent (SGD).
-On the other hand, CNN is designed to process grid-like data by leveraging spatial structure.
-It uses layers of convolutional filters to extract features like edges followed by pooling layers to reduce spatial dimensions and retain key information [31].
-The extracted features are passed through fully connected layers.
-One of the early applications of CNN was modulation classification based on signal constellation diagrams; the authors have used the CNN-based DL method [32].
-They have developed and analyzed methods to represent modulated signals in data formats with grid-like topologies to conform with the CNN architecture.
-RNN is another type of NN designed explicitly for processing sequential data.
-In RNN, the output from the previous step is fed as an input to the current step, incorporating sequential dependencies between the nodes.
-A recurrent neuron performs this function by maintaining a hidden state.
-Long short-term memory (LSTM) is one version of RNN [4].
-The LSTM network was adopted into the NOMA framework in [33] to learn the channel characteristics of the NOMA system.
-DRL integrates RL with DL to develop intelligent agents capable of making sequential decisions in complex environments.
-In DRL, an agent interacts with its environment to maximize cumulative rewards, learning optimal policies through exploration and exploitation.
-DNNs are employed to approximate value functions or policies, allowing DRL to efficiently handle high-dimensional state and action spaces.
-Key techniques such as Deep Q-Networks (DQN), policy gradient methods, and actor-critic frameworks enable agents to improve their performance iteratively through trial and error, guided by reward signals [4, 34].
-In the domain of communications and networking, DRL has emerged as a powerful tool for addressing challenges such as resource allocation, spectrum management, and network optimization.
-Modern network architectures, including the Internet of Things (IoT), heterogeneous networks (HetNets), and unmanned aerial vehicle (UAV) networks, exhibit decentralized, dynamic, and autonomous characteristics, making them well-suited for reinforcement learning-based solutions [35].
-The concept of using autoencoders to design end-to-end communication systems has been widely studied [36].
-End-to-end learning trains a single model to map raw input directly to the desired output.
-It avoids intermediate steps or manual feature engineering.
-Autoencoders are neural networks that optimize both the transmitter and receiver together in this framework.
-In communication systems, the goal is to design the transmitter and receiver jointly for a given channel model.
-This helps achieve better performance, such as faster convergence and lower bit error rates [37].
-GNNs are specialized networks designed for graph-structured data, where nodes are connected by edges [38].
-Unlike traditional neural networks, GNNs work on non-Euclidean data and capture relationships between graph nodes.
-They learn rich representations of nodes, edges, or the whole graph by aggregating and transforming features from neighboring nodes.
-This ability to use graph structure helps GNNs perform better in tasks that rely on relational data.
-In 2020, Goodfellow et al.
-proposed a new AI algorithm for generative modeling [39].
-GANs use two neural networks: a generator and a discriminator.
-The generator creates realistic outputs, while the discriminator identifies if they are real or fake.
-By competing and learning together, GANs can produce high-quality outputs that are often indistinguishable from real data.
-For further reading on application of deep learning in physical layer of communication, we invite readers to refer [40, 41, 42, 27].
-III Deep Unfolding: General Structure
-Deep unfolding, also known as algorithm unrolling [16], is a typical model-driven DL method [10],[20],[43, 44, 45, 46, 47, 48].
-It has been frequently considered for various tasks such as signal detection, design of beamforming matrices, and many more optimization tasks in wireless communication.
-Owing to reduced training time and dataset, deep unfolding is one promising alternate approach to purely model-driven or data-driven.
-Figure 3: Schematics of data-driven, classical and deep-unfolding approaches.
-Deep unfolding integrates model-based approaches and domain knowledge into the design of DNN architectures.
-The training is carried out in an end-to-end manner.
-In deep unfolding, an iterative algorithm is mapped to DNN layers.
-Then, DL techniques such as SGD are used to learn the parameters.
-Fig.
-3 shows a schematic diagram comparing the classical approach, data-driven approach, and the deep unfolding approach.
-Furthermore, Fig.
-4 depicts the deep unfolding approach in greater details.
-Figure 4: General framework of deep unfolding.
-For any iterative algorithm, the update equation for each iteration is given as:
-${\bf{s}}^{(t+1)}=h({\bf{s}}^{t};{\bf{\Psi}}^{t})$
-where ${\bf{s}}^{t}\in{\mathbb{R}}^{n}$ denotes the variable vector to be updated in each iteration in the range $t=1,\dots,T$.
-The iterative function for the iterative algorithm is represented by $h(\cdot;\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
-The trainable parameters, ${\bf{\Psi}}^{t}\in\mathbb{R}^{n}$ include the model parameters and the step size for the regularization coefficients of the algorithm [25].
-Each iteration function $h(\cdot;\cdot)$ is mapped to a single network layer, and a finite number of layers are stacked together as shown in Fig.
-4.
-To learn the trainable parameters ${\bf{\Psi}}^{t}$ in an end-to-end approach, the following optimization is carried out:
-$\text{minimize}_{{\bf{\Psi}}_{T}}\mathcal{L}({\bf{s}}^{T}({\bf{\Psi}}_{T}))$
-where $\mathcal{L}(\cdot)$ denotes the loss function for training, ${\bf{\Psi}}_{T}=\{\Psi^{t}\}_{t=0}^{T}$ is the set of trainable parameters of the total $T$-layer network and ${\bf{s}}^{T}(\cdot)$ is the output of the unfolded network.
+2 illustrates the process of unfolding an iterative algorithm.
+In general, for any iterative algorithm, the update equation for each iteration is given as:
+${\bf{s}}^{l+1}=h({\bf{s}}^{l};{\boldsymbol{\theta}}^{l})$ (1)
+where $\left\{{\bf{s}}^{l}\right\}_{l=1}^{L}\in{\mathbb{R}}^{n}$ denotes the sequence of variable vector updated across $L$ layers of the algorithm.
+At each layer, the update is performed by executing $h(\cdot;\cdot)$.
+Repeated application of this update rule over $L$ layers, yields the unfolded network architecture, $h(\cdot;{\boldsymbol{\theta}}^{l})_{l=1}^{L}$.
+The trainable parameters, $\left\{\boldsymbol{\theta}^{l}\right\}_{l=1}^{L}$ (shown in red fonts in Fig.
+2) includes the model parameters and the step-size for the regularization coefficients of the algorithm [26].
+Figure 2: General framework of training for deep unfolding. The learnable parameters are shown in red color.
+To learn the trainable parameters $\left\{\boldsymbol{\theta}\right\}_{l=1}^{L}$ in an end-to-end approach, the loss $\ell({\bf{s}}^{L},\bf{s})$ is computed, where ${\bf{s}}^{L}$ is the output of the network and $\bf{s}$ is the ground truth target.
+The key idea behind end-to-end learning is that all parameters across $L$ layers are simultaneously adjusted through backpropagation, where the loss is minimized using gradient-based optimization.
 Feeding the data forward through the unfolded network with trainable parameters in the testing stage is similar to executing the parameter-optimized iterative algorithm for a finite number of iterations.
@@
 Within these fixed layers, the model is trained to optimize the trainable parameters, enabling the problem to be addressed using a structured algorithm with parameters tailored for optimal performance.
-The procedure to solve an optimization problem using a deep unfolding approach is outlined below.
-  1. 1.
-    Identify a principle algorithm that serves as the foundation for deep unfolding.
-  2. 2.
-    Select trainable parameters or specific blocks in the principle algorithm. Replace these with data-trained operations optimized through deep unfolding.
-  3. 3.
-    Define the loss function and represent the algorithm as a DNN.
-Each layer corresponds to an iteration of the original algorithm.
-Train the network using supervised or unsupervised methods.
-Since the number of iterations is predetermined, the network’s layers match these iterations.
-The number of layers can also be fine-tuned to balance performance and computational complexity.
-TABLE III: Key technologies envisioned for 6G using deep unfolding.
-Key technology Task Reference
-Massive MIMO Detection [49],[50], [51], [52], [53], [44], [54], [55], [56], [57], [58], [59], [60], [61]
-Channel estimation [62], [63], [64], [65], [66], [67]
-Precoder Design [68], [69], [70], [71], [72], [73], [74], [75], [76], [77], [78], [79], [80], [81]
-ISAC [82], [83], [84]
-mmWave Channel Estimation [85]
-Precoder design [86], [87], [88]
-ISAC [89]
-Modern Multicarrier Modulation OFDM Detection [90],[91], [92]
-OTFS Detection [46]
-OFDM Channel Estimation [93], [94], [95], [45], [96], [97], [98], [99], [47] [100]
-OTFS Channel Estimation [101]
-OFDM-ISAC [102], [103]
-OTFS-ISAC [104]
-Next Generation Multiple Access Detection [105], [106]
-Precoder design [107]
-Orbital Angular Momentum Detection [108]
-IRS Channel Estimation [109], [110], [111], [112]
-Precoder design [113], [114], [115], [116]
-ISAC [117]
-THz Precoder design [118], [119], [120]
-Principle algorithms are typically iterative and based on optimization techniques.
-However, they are often complex, computationally demanding, and slow to converge.
-Deep unfolding improves these algorithms by leveraging data training.
-The key contributions of deep unfolding are:
-  1. 1.
-    Optimizing hyperparameters to enhance convergence, leading to better objective values or faster processing [26, 76, 84, 11].
-  2. 2.
-    Simplifying complex tasks, such as matrix inversion or decomposition, with lower computational complexity [121, 70, 122].
-The first application is common in unfolded projected gradient ascent/descent (PGA/PGD) methods, while the second is widely used in unfolded weighted minimum mean square error (WMMSE) methods.
-Another approach involves learning the output at each layer, which corresponds to the iterative solution of the principal algorithm.
-This technique is widely applied in detection networks [123, 118, 49], following the PGA/PGD framework.
-The main advantage of deep unfolding is its ability to incorporate domain knowledge into the network architecture.
-This improves performance while reducing complexity by using fewer parameters and a fixed number of layers.
-As a result, deep unfolding requires a smaller training dataset compared to generic DNNs.
-Next, we provide a comprehensive survey of deep unfolding algorithms designed for various signal processing tasks in wireless communication.
-Additionally, TABLE III lists key 6G technologies relevant to deep unfolding-based signal processing.
-IV Signal Detection
+Example 1.
+To develop an understanding on deep unfolding, we take ADMM as an example.
+ADMM decomposes a complex optimization problem into simpler subproblems, which are solved iteratively.
+The algorithm combines the principles of dual ascent and the method of multipliers.
+Each iteration involves solving the subproblems for the primal variables in an alternating manner, followed by updating the dual variables (Lagrange multipliers) to enforce consistency with the original problem’s constraints [37].
+This iterative process ensures that the subproblem solutions converge to the global solution of the main optimization problem.
+ADMM in its unfolded version has been widely used in MIMO signal detection [25], [38] and decoding in error correction codes [39, 40].
+Here, we show the unfolding of ADMM for MIMO detection as an illustrative example.
+Figure 3: Deep unfolding training framework for ADMM-based MIMO detection. The learnable parameters are shown in red color.
+Consider a MIMO system with $M$ transmit antennas and $N$ receive antennas. The received signal vector at the receiver can be given by
+${\bf{\bar{y}}}={\bf{\bar{H}\bar{x}}}+{\bf{\bar{n}}}$
+where ${\bf{\bar{H}}}\in\mathbb{C}^{N\times M}$ represents the channel matrix, ${\bf{\bar{x}}}\in{\cal{\bar{X}}}^{M}$ is the transmitted symbol vector, and ${\bf{\bar{n}}}\in\mathbb{C}^{N}$ denotes additive white Gaussian noise (AWGN) vector at the receiver.
+Here, $\cal{\bar{X}}$ represents the complex constellation set.
+In this example, we consider the QPSK modulation scheme, and thus, $\left({{\cal{\bar{X}}}=\left\{{1+i,-1+i,-1-i,1-i}\right\}}\right)$.
+The maximum likelihood (ML) detection rule is given as
+$\min_{\bf{\bar{x}}\in{\cal{\bar{X}}}^{M}}\left\lVert{\bf{\bar{y}}}-{\bf{\bar{H}\bar{x}}}\right\rVert_{2}^{2}$ (2)
+Denote the following real-valued vectors and matrix:
+$\mathbf{y}\triangleq\begin{bmatrix}\mathfrak{R}(\bar{\mathbf{y}})\\ \mathfrak{I}(\bar{\mathbf{y}})\end{bmatrix},\quad\mathbf{x}\triangleq\begin{bmatrix}\mathfrak{R}(\bar{\mathbf{x}})\\ \mathfrak{I}(\bar{\mathbf{x}})\end{bmatrix},\quad\mathbf{n}\triangleq\begin{bmatrix}\mathfrak{R}(\bar{\mathbf{n}})\\ \mathfrak{I}(\bar{\mathbf{n}})\end{bmatrix},$
+[…]
+where $\mathfrak{R}(\cdot)$ and $\mathfrak{I}(\cdot)$ denote the real and imaginary part of a complex vector or matrix, respectively.
+With this transformation, each element of $\mathbf{x}$ is drawn from the discrete space {-1, 1} .
+We then relax this space to the continuous domain ${\cal{X}}=[-1,1]$.
+Consequently, (2) can be relaxed to the following problem in the real domain: $\min_{{\bf{x}}\in{{\cal{X}}}^{2M}}\left\lVert{\bf{y}}-{\bf{Hx}}\right\rVert_{2}^{2}.$ We further introduce an auxiliary variable $\mathbf{z}$ to reformulate this problem as:
+$\underset{\begin{subarray}{c}{\bf{x}}\in{{\cal{X}}}^{2M}\end{subarray}}{\min}$ $\quad\left\lVert{\bf{y}}-{\bf{Hx}}\right\rVert_{2}^{2}$
+subject to $\quad\mathbf{x}=\mathbf{z},\mathbf{z}\in{{\cal{X}}}^{2M}.$
+The augmented Lagrangian is given as:
+$\mathcal{L}_{\rho}(\mathbf{x},\mathbf{z},\boldsymbol{\lambda})=\left\lVert\mathbf{y}-\mathbf{Hx}\right\rVert_{2}^{2}+\boldsymbol{\lambda}^{T}(\mathbf{x}-\mathbf{z})\\ +\frac{\rho}{2}\left\|\mathbf{x}-\mathbf{z}\right\|_{2}^{2}$ (4)
+where $\boldsymbol{\lambda}\in\mathbb{R}^{2M}$ is the dual variable (or Lagrange multiplier) and $\rho>0$ is the penalty parameter.
+The iterative steps of the ADMM involve alternating minimization of the augmented Lagrangian with respect to the primal variables $\bf x$ and $\bf z$, followed by an update of the dual variable, $\boldsymbol{\lambda}$ [41, 42].
+The closed-form solutions for $\bf x$ and $\bf z$ can be obtained as:
+${\mathbf{x}}^{l+1}$ $=\arg\min_{\mathbf{x}}\mathcal{L}_{\rho}(\mathbf{x},\mathbf{z}^{l},\boldsymbol{\lambda}^{l})$
+$=\arg\min_{\bf{x}}\left\lVert\mathbf{y}-\mathbf{Hx}\right\rVert_{2}^{2}+{(\boldsymbol{\lambda}^{l})}^{T}({\mathbf{x}}-{\mathbf{z}}^{l})+\frac{\rho}{2}\left\lVert{\mathbf{x}}-{\mathbf{z}}^{l}\right\rVert_{2}^{2}$
+$=\left({\mathbf{H}}^{T}\mathbf{H}+\rho\mathbf{I}\right)^{-1}\left({\mathbf{H}}^{T}\mathbf{y}+\rho{\mathbf{z}}^{l}-{\boldsymbol{\lambda}}^{l}\right),$ (5)
+${\mathbf{z}}^{l+1}$ $=\arg\min_{\mathbf{z}}\mathcal{L}_{\rho}(\mathbf{x}^{l+1},\mathbf{z},{\boldsymbol{\lambda}}^{l})$
+$=\arg\min_{\bf{z}}-{(\boldsymbol{\lambda}}^{l})^{T}{\mathbf{z}}+\frac{\rho}{2}\left\lVert{\mathbf{x}}^{l+1}-{\mathbf{z}}\right\rVert_{2}^{2}$
+$=\arg\min_{\bf{z}}\frac{\rho}{2}\left\lVert{\mathbf{z}}-({\mathbf{x}}^{l+1}+\frac{{\boldsymbol{\lambda}}^{l}}{\rho})\right\rVert_{2}^{2}$
+$=\Pi_{[-1,1]}\left(\mathbf{x}^{l+1}+\frac{\boldsymbol{\lambda}^{l}}{\rho}\right),\quad\quad{\text{and}}$ (6)
+${\boldsymbol{\lambda}}^{l+1}$ $={\boldsymbol{\lambda}}^{l}+\rho\left({\mathbf{x}}^{l+1}-{\mathbf{z}}^{l+1}\right)\quad\forall l$ (7)
+where $\Pi_{[-1,1]}({\boldsymbol{a}})$ denotes an element-wise projection of $\boldsymbol{a}$ onto the set $[-1,1]$. More specifically, if ${\boldsymbol{b}}=\Pi_{[-1,1]}(\boldsymbol{a})$, we have
+$b_{i}=\begin{cases}+1,&\text{if }a_{i}>1\\ -1,&\text{if }a_{i}<1\\ a_{i},&\text{otherwise}.\end{cases}$
+The detailed procedure for the unfolded ADMM approach is provided in Algorithm 1.
+In deep unfolding, to ensure gradient flow during backpropagation in the training stage, all operations must be differentiable.
+To enable differentiable approximation, a soft projection is used that allows gradient to flow while keeping outputs near the desired constellation points.
+Consequently, (6) becomes:
+${\mathbf{z}}^{l+1}$ $=\Pi_{\text{soft}}\left(\mathbf{x}^{l+1}+\frac{\boldsymbol{\lambda}^{l}}{\rho}\right).$ (8)
+Furthermore, the layer-wise unfolding of ADMM for MIMO detection is illustrated in Fig.
+3.
+The received symbol vector $\mathbf{y}$ and the channel matrix $\mathbf{H}$ are treated as fixed inputs and are provided to every layer of the network.
+Each layer of the unfolded network corresponds to one iteration of the ADMM procedure, specifically the update steps in equations (5), (7), and (8).
+Corresponding to each of these equations, a specific $\rho$ is considered to enable learning.
+These parameters form the trainable set ${\boldsymbol{\theta}}=\left\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{2}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{3}^{(l)}}\right\}_{l=1}^{L}$, where each layer $l$ has its own set of penalty coefficients.
+In the final layer $L$, the network produces the estimate ${\mathbf{z}}^{L}=\hat{\bf{z}}$.
+For $n=1,\dots,N$, data samples, the output $\hat{\bf{z}}_{n}({\bf{y}}_{n},{\bf{H}};\rho_{1}^{(l)},\rho_{2}^{(l)},\rho_{3}^{(l)})$ is retrieved from the network as the predicted transmitted symbol for ${\bf{y}}$.
+The network training loss function formed between the predicted output $\hat{\bf z}$, and ground truth transmitted symbol ${\bf z}_{n}$ is given as:
+$\ell(\rho_{1}^{(l)},\rho_{2}^{(l)},\rho_{3}^{(l)})=\frac{1}{N}\sum_{n=1}^{N}\left\lVert\hat{\bf{z}}_{n}({\bf{y}}_{n},{\bf{H}};\rho_{1}^{(l)},\rho_{2}^{(l)},\rho_{3}^{(l)})-{\bf z}_{n}\right\rVert_{2}^{2}.$ (9)
+The network is trained by minimizing the loss with gradient-based learning techniques such as stochastic gradient descent (SGD) to learn $\rho_{1}^{(l)}$, $\rho_{2}^{(l)}$ and $\rho_{3}^{(l)}$.
+Algorithm 1 Unfolded ADMM for MIMO detection.
+1:  Input: Initial values $\mathbf{z}^{(0)}$, ${\boldsymbol{\lambda}}^{(0)}$
+2:  Given: $\mathbf{y}$ and $\mathbf{H}$
+3:  Learnable parameters $\boldsymbol{\theta}=\left\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{2}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{3}^{(l)}}\right\}_{l=1}^{L}$
+4:  for $l=1$ to $L$ do
+5:     $\mathbf{x}^{l+1}=\left(\mathbf{H}^{T}\mathbf{H}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}^{(l)}}\mathbf{I}\right)^{-1}\left(\mathbf{H}^{T}\mathbf{y}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}^{(l)}}\mathbf{z}^{l-1}-{\boldsymbol{\lambda}}^{l}\right)$
+6:     $\mathbf{z}^{l+1}=\Pi_{\text{soft}}\left(\mathbf{x}^{l+1}+\frac{{\boldsymbol{\lambda}}^{l-1}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\rho_{2}^{(l)}}}}\right)$
+7:     ${\boldsymbol{\lambda}}^{l+1}={\boldsymbol{\lambda}}^{l}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{3}^{(l)}}\left({\mathbf{x}}^{l+1}-{\mathbf{z}}^{l+1}\right)$
+8:  end for
+9:  Output: Final estimate $\hat{\bf{z}}={\bf{z}}^{L}$
+[…]
+Figure 4: BER performance comparison for 16$\times$16 MIMO system. For soft projection, $tanh$ is used.
+Fig.
+4 presents the bit error rate (BER) performance comparison of the unfolded ADMM detector with zero forcing (ZF), minimum mean squared error (MMSE), ADMM and sphere decoder (SD), under QPSK modulation.
+We consider a $16\times 16$ MIMO system with Rayleigh fading channels.
+The results demonstrate that the unfolded ADMM detector outperforms ZF, MMSE, and standard ADMM detectors, except for SD.
+The unfolded ADMM detector performs better due to the use of deep unfolding, where learnable parameters are adaptively trained.
+This example has been included to provide the readers with an illustrative understanding of the deep unfolding process.
+The comparative analysis highlights the advantage of the unfolded method over classical approaches.
+Additional performance improvements can be achieved by optimizing or introducing trainable parameters in the unfolded network.
+In the following sections, we provide a comprehensive survey of deep unfolding algorithms for a variety of critical wireless communication tasks.
+III Signal Detection
 This section discusses the application of deep unfolding techniques for signal detection. We segregate these signal detection techniques into various key technologies, as described in the following.
-IV-A Massive MIMO
-MIMO transmission uses multiple antennas at both the transmitter and receiver.
-It improves spectral and energy efficiency by utilizing a larger spatial domain.
-Adding more antennas (tens or hundreds) introduces the concept of massive MIMO [3].
-However, detecting multiple symbols becomes more complex due to large-scale inter-symbol interference.
-Designing a low-complexity, high-performance detector for massive MIMO remains a challenge.
-For a MIMO communication system with $n_{\text{t}}$ transmit and $n_{\text{r}}$ receive antennas, the received signal can be given as:
-$\mathbf{y}=\mathbf{Hx}+\mathbf{n}$ (1)
-where $\mathbf{x}\in\mathbb{C}^{n_{\text{t}}}$ is the transmitted baseband signal vector, $\mathbf{H}\in\mathbb{C}^{n_{\text{r}}\times n_{\text{t}}}$ is the channel matrix and $\mathbf{y}\in\mathbb{C}^{n_{\text{r}}}$ is the received baseband signal vector.
-The detector extracts the desired transmitted information $\mathbf{x}$ from the received signal vector $\mathbf{y}$ with given $\mathbf{H}$ and the Gaussian noise vector $\mathbf{n}$.
-The maximum a posteriori (MAP) and maximum likelihood (ML) estimations are Bayesian-based detection methods.
-These are the most commonly used classical detection techniques.
-The MAP rule applies when the transmitted signals have unequal a priori probabilities.
-The ML rule is used when these probabilities are equal [124].
-ML minimizes the average error probability, providing an optimal solution.
-The ML detection rule is given as:
-$\mathbf{\hat{x}_{\text{ML}}}=\text{arg}\min_{\mathbf{x}\in\mathcal{A}^{n_{t}}}||\mathbf{y}-\mathbf{Hx}||^{2}.$ (2)
-This rule minimizes the Euclidean distance between $\bf{y}$ and $\bf{Hx}$.
-The minimization is performed over all possible combinations of the transmitted symbol vector $\bf{x}$ from the modulation alphabet $\mathcal{A}$.
-Since ML performs an exhaustive search and examines all possible signals, the computational complexity increases exponentially as $n_{\text{t}}$ increases.
+III-A Massive MIMO
+Massive MIMO uses numerous antennas at both the transmitter and receiver, improving spectral and energy efficiency, but making symbol detection more complex due to inter-symbol interference (ISI) [3].
+Designing low-complexity, high-performance detectors is still challenging.
 MIMO detection can be linear or non-linear.
-Common linear detectors include matched filter (MF), zero-forcing (ZF), and minimum mean squared error (MMSE) [125].
-These detectors require matrix inversion.
-Lattice reduction techniques can improve the performance of ZF and MMSE.
-Non-linear detectors, such as those using successive interference cancellation (SIC), perform interference estimation and removal in multiple stages [3], [126].
-SIC and parallel interference cancellation are well-known techniques for this.
-In SIC, a detector like MF, ZF, or MMSE is used to detect the transmitted symbol in a data stream, leading to MF-SIC, ZF-SIC, or MMSE-SIC schemes.
-The data streams are arranged based on the received SNR.
-After detecting the symbol and estimating the channel coefficients, the interference caused by the detected symbol is estimated.
-This interference is then removed from the received signal, which is used to detect the next data stream in terms of received SNR.
-This process repeats until all symbols are detected.
-Sphere decoding is another non-linear detector [127].
-It is an approximate ML detection method, where detection happens in a subspace of the modulation alphabet.
-The method searches within a bounded hypersphere around the received lattice point, using a radius for iterative symbol searching.
-The logarithmic likelihood ratio (LLR) is computed based on the symbol candidate found.
-This search region is smaller than the exhaustive ML search [128, 129].
-The channel matrix for a MIMO fading channel can be modeled as a graph.
-The MIMO detection problem is framed as an inference task to calculate the posterior probability of system variables.
-Low-complexity algorithms are designed to compute these probabilities efficiently.
-The graph model combines probability and graph theory, showing the relationships between different nodes.
-Common graphical models include Bayesian belief networks, Markov random fields, and factor graphs [3].
-Message-passing algorithms, like belief propagation (BP), are crucial for representing the MIMO system’s graph.
-Probability messages are exchanged between nodes representing transmitted symbols and received signals along the graph edges.
-The system is often modeled as a factor graph, where performance depends on the graph structure and algorithm convergence.
-Iterative algorithms, like message-passing detectors (MPD) with BP, approximate message passing (AMP), and channel hardening-exploiting message passing, provide solutions with lower complexity.
-Expectation propagation (EP) extends BP to handle larger constellations and higher-dimensional MIMO systems.
-EP approximates the posterior distribution of transmitted symbols iteratively, enabling efficient detection in polynomial time [130, 131].
-The conventional detectors (linear and non-linear) involve computationally intensive matrix inversions or factorizations (like singular value decomposition or QR decomposition).
-Although these detectors yield nearly optimal solutions, they come at the cost of high complexity and high latency.
-Hence, with the aid of DL techniques, a deep unfolding approach has been adopted in literature to achieve low complexity and low latency.
-PGD is an iterative optimization algorithm.
-Each step includes a gradient descent update to minimize a cost function, followed by a projection step to enforce constraints like sparsity or feasibility.
-In MIMO detection, PGD minimizes the error between the predicted and actual received signals while projecting the solution onto a feasible set, such as valid constellation points [132].
-To achieve ML optimization, the PGD algorithm is unfolded into the DetNet architecture for MIMO detection [49, 133].
-The gradient step size is learned across layers.
-The authors claimed that this approach reduces computational complexity compared to the sphere decoder and is robust to ill-conditioned channels.
-By leveraging the benefits of graphical models, He et al.
-proposed model-driven DL networks, OAMP-Net [10] and OAMP-Net2 [11], for MIMO detection with imperfect channel state information (CSI).
+Linear detectors, like matched filter (MF), ZF, and MMSE, involve matrix inversion and can be enhanced with lattice reduction [43].
+Non-linear detectors, such as successive interference cancellation (SIC) and parallel interference cancellation (PIC), remove interference in stages [3], [44].
+SIC schemes (e.g., MF-SIC, ZF-SIC, MMSE-SIC) use linear detectors to detect symbols, estimate interference, and remove it before detecting the next symbol.
+SD, a non-linear method, approximates ML detection by searching a bounded hypersphere around the received lattice point, offering a more efficient search than exhaustive methods [45], [46], [47].
+The projected gradient descent (PGD) method is a popular optimization-driven approach for MIMO detection.
+In each iteration, PGD performs gradient descent to minimize detection error, then projects the solution onto a feasible, valid constellation set [48].
+One of the first papers to use deep unfolding in detection is the DetNet architecture [49].
+To achieve ML optimization, the PGD algorithm is unfolded where the gradient’s step size is learned across layers.
+The authors in [49] claimed that the computational complexity is reduced compared to the SD.
+DetNet nearly matches SD’s performance (within 0.5 order of BER at 18 dB).
+The computational complexity is linear ($\mathcal{O}(L.K^{2})$), where $L$ and $K$ are the number of layers and constellation size, respectively.
+On the other hand, the SD exhibits ($\mathcal{O}(2^{K})$) for worst-case complexity.
+However, the proposed network is not flexible to changes in the size of the constellation or the number of users.
+To address MIMO detection under imperfect CSI, He et al.
+introduced model-driven DL networks, orthogonal approximate message passing (OAMP) network, OAMP-Net [10], and its enhanced variant OAMP-Net2 [11].
 OAMP-Net2 outperforms OAMP-Net by adapting to varying channel conditions.
-The orthogonal AMP (OAMP) algorithm is unfolded, and the LMMSE channel estimator is used to design the receiver.
-Learnable parameters, such as the step size for updating the prior mean and variance in the MMSE estimator, are trained using SGD.
+The OAMP algorithm is unfolded, and the linear MMSE (LMMSE) channel estimator is used to design the receiver in [11].
+Learnable parameters, such as the step-size for updating the prior mean and variance in the MMSE estimator, are trained using SGD.
 These parameters in OAMP-Net2 help compensate for channel estimation errors, improving performance compared to DetNet [49].
-In [134], the authors introduced MMNet, a deep unfolded detector designed for a wide range of realistic channel matrices, including ill-conditioned and non-i.i.d.
-scenarios.
-MMNet is based on an iterative soft-thresholding algorithm, alternating between a linear detector and a non-linear denoising step.
-It balances architectural flexibility and computational complexity, outperforming DetNet [49] and OAMP-Net2 [11].
-Overloaded massive MIMO systems, where the number of transmit antennas exceeds the number of receive antennas, require detectors with low complexity and good performance.
-In [61], the trainable projected gradient (TPG) algorithm was introduced for these systems.
-TPG uses a DL-aided iterative PGD method with trainable parameters, such as step size, optimized via backpropagation and SGD.
-TPG achieves lower computational cost and fewer trainable parameters compared to DetNet [49], OAMP-Net2 [11], and others.
-Additionally, a trainable detector based on the Hubbard–Stratonovich transformation is discussed in [135].
-Takabe et al.
-avoided matrix inversion in this approach.
-Compared to the TPG detector [61] and OAMP-Net2 [10], the proposed network achieves faster training and lower execution cost.
-In [136], the concrete MAP method is introduced, based on a continuous relaxation of the probability mass function for discrete random variables, as derived in [137].
-This approach solves the MAP detection problem using continuous optimization instead of exhaustive search, approximating the MAP solution for any probabilistic nonlinear model.
-By unfolding the gradient descent algorithm, the detector provides soft outputs (detection probabilities), enabling better decoding compared to DetNet [49], OAMP-Net2 [10], and MMNet [134].
+As the number of transmit antennas exceeds the number of receive antennas, low complexity detectors are required for such overloaded massive MIMO systems.
+In [50], the trainable PGD (TPGD) algorithm was introduced for such systems.
+TPGD uses a DL-aided iterative PGD method with trainable parameters, such as step-size, which are optimized via backpropagation and SGD.
+TPGD achieves lower computational cost with fewer trainable parameters compared to DetNet [49], OAMP-Net2 [10].
 Yun et al.
@@
 Momentum helps guide parameter updates in the right direction and reduces oscillations during PGD optimization [52].
-This modification improves convergence speed and overall performance, outperforming DetNet [49] and OAMP-Net2 [10, 11], while maintaining generality.
-In wireless communication, physical layer signal processing involves continuous-time analog waveforms, which must be quantized for hardware implementation.
-Low-resolution one-bit analog-to-digital converters are commonly used in massive MIMO systems due to their high sampling rates, low cost, and low energy consumption [140].
-To recover high-dimensional signals from noisy, low-resolution measurements, Khobahi et al.
-[141] developed a low-resolution detection network.
-This network extracts signals from noisy one-bit measurements without requiring prior knowledge of the channel matrix.
-It is based on unfolding the PGD method for symbol detection using the ML estimator, eliminating the need for CSI.
-The network outperforms traditional ML estimators [142] and DL networks like DeepSIC [143], even with small datasets of about 500 samples.
-In [59], a generalized PGD algorithm was proposed, eliminating the need for separate loss functions for projection and gradient steps.
-Instead of alternating after each step, projection is applied after several gradient steps, achieving a better balance.
-By unfolding this approach, the authors created a self-correcting auto detector that enhances performance [59].
-Homotopy optimization transforms the original problem into a different one using a homotopy map.
-This method solves a series of simpler problems step by step, gradually progressing to the harder ones [144].
-In MIMO detection, non-convex algorithms often struggle with the ML detection problem due to bad local minima [50].
-By using a homotopy map to transition from easier to harder problems, detection performance is significantly improved.
-The study in [50] focuses on ML MIMO detection with one-bit quantized observations and binary symbol constellations.
-Using homotopy optimization, the authors apply a non-convex continuous reformulation of binary optimization [145] to address the detection problem.
-Their algorithm is designed to run efficiently and adapt to different channels.
-Building on this work, the authors introduced a proximal gradient-based homotopy method in [53].
-This approach connects the structure of DetNet [49] with homotopy optimization for improved MIMO detection.
-GNNs, as structured DNNs, excel in adaptability and flexibility.
-In MIMO detection, the graph structure of wireless networks makes GNNs particularly effective.
-Tan et al.
-[55] unfold MPD iterations to update messages at symbol nodes using prior and posterior LLR.
-Learnable parameters, such as prior probabilities and correction factors (damping, rescaling, offset), are optimized via mini-batch SGD to minimize the loss function.
-A greedy search determines the number of layers to balance performance and complexity.
-In [60], a low-complexity AMP-GNN model is proposed for massive MIMO detection.
-It outperforms OAMP-Net2 [10] by avoiding matrix inversion.
-For single-carrier FDMA systems, common in 5G to reduce peak-to-average power ratios, message-passing algorithms are computationally demanding.
-To address this, the unfolded network in [146] simplifies generalized AMP iterations [147], maintaining high performance even with channel mismatches.
-To enhance EP-based detection, [51] optimizes parameter selection using a modified moment-matching scheme, showing superior performance in high modulation cases compared to OAMP-Net2.
-In [52], approximate EP from [148] is unfolded, achieving better detection performance and robust convergence.
-Liao et al.
-[44] incorporated interference cancellation into DNN layers using iterative embedding.
-Inspired by [149], their sequential detector trains residual error coefficients with MSE loss, reducing complexity by avoiding matrix inversion and relying on matrix addition and multiplication.
-Compared to DetNet [49, 133], which requires 90 layers for convergence, their model achieves convergence with just 8 layers, offering significant computational simplicity.
-Massive MIMO systems improve spectral efficiency (SE), coverage, and range but demand higher computational power.
-LcgNet, a model-driven DL architecture for massive MIMO detection, unfolds conjugate gradient descent iterations and learns step sizes through offline training [54].
-Its quantized version, QLcgNet, minimizes memory usage with negligible performance degradation.
-LcgNet outperforms DetNet in terms of symbol error rate, computational complexity, and memory cost, making it a robust solution for massive MIMO systems.
-The quantized version, QLcgNetV, reduces memory requirements while maintaining high detection performance.
+This modification improves convergence speed and overall performance, outperforming DetNet [49] and OAMP-Net2 [11].
+Furthermore, a trainable detector based on the Hubbard–Stratonovich transformation (to avoid matrix inversion) is discussed in [53].
+Compared to the TPGD detector [50] and OAMP-Net2 [11], the network in [53] achieves faster training and lower execution cost.
+In [54], the authors introduced the MIMO detection network (MMNet), an unfolded detector designed for ill-conditioned channels and non i.i.d.
+scenarios.
+MMNet is based on an iterative shrinkage-thresholding algorithm (ISTA), alternating between a linear detector and a nonlinear denoising step.
+The architectural flexibility and computational complexity are balanced, outperforming DetNet [49] and OAMP-Net [10].
+The MMNet outperforms the classical LMMSE detector with a performance gain of approximately $4-6$ dB at a symbol error rate (SER) of $10^{-3}$ for different channel conditions.
+In [55], the CONCRETE MAP method is proposed, based on a CONtinuous relaxation of the probability mass function for disCRETE random variables, as derived in [56].
+This approach solves the maximum a posteriori (MAP) detection problem using continuous optimization instead of exhaustive search, approximating the MAP solution for any probabilistic nonlinear model.
+By unfolding the gradient descent algorithm, the detector is able to generate soft outputs, making it suitable for integration with soft-input decoders.
+The network in [55] outperforms DetNet [49], OAMP-Net2 [11], and MMNet [54], in terms of BER under correlated channels.
+In [57], conjugate gradient descent (CGD) iterations are unfolded to design the network, LcgNet.
+The authors also proposed a quantized version, QLcgNet, which minimizes memory usage with negligible performance degradation.
+LcgNet outperforms DetNet [49] in terms of SER, computational complexity, and memory cost, making it a robust solution for massive MIMO systems.
 To enhance LcgNet, Olutayo et al.
 [58] incorporated a preconditioner during training, designing a preconditioned LcgNet.
-Preconditioning improves the condition number of a matrix [151], optimizing the receiver filter’s spectrum and minimizing eigenvalue spread.
-This enhancement significantly boosts the convergence speed compared to the original LcgNet [54].
-Additionally, the dynamic conjugate gradient network utilizes self-supervised learning with forward error correction codes, adapting effectively to non-stationary, time-varying channels.
-Sphere decoding achieves ML optimality but suffers from high computational complexity in massive MIMO systems.
-To address this, Nguyen et al.
-[123] proposed a fast DL-aided sphere decoding algorithm inspired by [152].
-The DNN generates reliable initial candidates for sphere decoding, prioritizing candidates closer to the DNN’s output for testing.
+Preconditioning improves the condition number of a matrix [59], optimizing the receiver filter’s spectrum.
+This enhancement significantly increases the convergence speed compared to the original LcgNet in [57].
+Additionally, the dynamic conjugate gradient network utilizes self-supervised learning with forward error correction codes, effectively adapting to non-stationary, time-varying channels.
+Physical layer signal processing involves continuous-time analog waveforms, which must be quantized for hardware implementation.
+Low-resolution (LR) one-bit analog-to-digital converters (ADC) are commonly used at radio frequency (RF) ends due to their high sampling rates, low cost, and low energy consumption [60].
+To recover high-dimensional signals from noisy measurements, Khobahi et al.
+[61] developed an LR detection network (LoRDNet).
+This network extracts signals from noisy one-bit measurements with unknown CSI.
+It is based on unfolding the PGD method for symbol detection using the ML estimator.
+The channel effects are embedded in the training data, so LoRDNet implicitly learns a representation of the channel behavior, eliminating the need for CSI.
+The network outperforms classical ML estimators [62] and DL networks like DeepSIC [63], even with small datasets of about 500 samples.
+In [64], the authors proposed a generalized PGD algorithm.
+The generalized PGD algorithm executes multiple gradient descent steps before applying the projection, allowing for a better balance between convergence speed and detection performance.
+The authors introduced the self-correcting auto detector (SAD) by unfolding the generalized PGD algorithm.
+The architecture has two modules: the self-correcting module that executes multiple gradient descent steps to refine the detection estimate, and a denoising autoencoder, a learnable projection module.
+The numerical results demonstrate that SAD [64] achieves a better trade-off between detection accuracy and complexity compared to DetNet [49].
+Although DetNet requires $15$ projection steps, SAD only performs $4$, significantly reducing detection latency and computational cost.
+The channel matrix for a MIMO fading channel can be modeled as a graph.
+Graph models, such as Bayesian belief networks, Markov random fields, and factor graphs, combine the posterior probability of transmitted data variables and graph theory, establishing a relationship between different nodes for proper detection [3].
+The performance of these methods largely depends on the underlying graph structure and convergence properties.
+Iterative algorithms, like message-passing detectors (MPD) with belief propagation (BP), approximate message passing (AMP), and channel hardening-exploiting message passing (CHEMP), further reduce computational complexity.
+Expectation propagation (EP), as an extension of BP, enables efficient detection in higher-dimensional MIMO systems and large constellations.
+EP approximates the posterior distribution of transmitted symbols iteratively, enabling efficient detection in polynomial time [65, 66].
+Tan et al.
+[67] unfolded MPD iterations to update messages at symbol nodes using prior and posterior log-likelihood ratio (LLR).
+Learnable parameters, such as prior probabilities and correction factors (damping, scaling factor, LLR rescaling), are optimized via mini-batch SGD.
+The authors unfolded the existing classical MPDs: BP and CHEMP, where the correction factors are learned instead of heuristic tuning.
+A greedy search determines the number of layers.
+The unfolded networks adapt better to correlated channels and perform consistently in higher-order modulation than classical MPDs.
+In [68], a low-complexity AMP-graph NN (GNN) model (AMP-GNN) is proposed, which outperforms OAMP-Net2 [11] by avoiding matrix inversion.
+Due to the high computational complexity, message-passing algorithms are generally unsuitable for single-carrier frequency division multiple access (FDMA) systems, where a low peak-to-average power ratio is a key design objective for efficient transmission.
+To address this, the unfolded network in [69] simplifies generalized AMP iterations as discussed in [70].
+The proposed network simplified generalized AMP iterations by approximating per-subcarrier channel statistics through statistical averaging.
+The unfolded network [69] maintains high performance even with channel mismatches and achieves $0.83$ dB at BER$=10^{-3}$ gain over classical generalized AMP [70].
+To enhance EP-based detection, the authors in [71] proposed a network, a modified EP MIMO detector (MEPD).
+In the network, parameter selection is optimized using a modified moment-matching scheme and shows superior performance in high modulation cases compared to OAMP-Net2 [11].
+In [72], an approximate EP inspired by [73] is unfolded, achieving better detection performance and robust convergence.
+To address high computational complexity of SD, Nguyen et al.
+[74] proposed a fast DL-aided SD algorithm inspired by [75].
+The DNN generates reliable initial candidates for SD, prioritizing candidates closer to the DNN’s output for testing.
 This approach accelerates sphere shrinkage, reducing computational cost.
 The sparse network architecture eliminates matrix inversion, relying only on element-wise matrix multiplications.
-The proposed method performs better in highly correlated channels than DetNet [49, 133], OAMP-Net2 [10, 11], MMNet [134], and LcgNet [54].
+The proposed method performs better in highly correlated channels than DetNet [49], OAMP-Net2 [11], MMNet [54], and LcgNet [57].
 A MIMO detector based on annealed Langevin dynamics is introduced in [76].
@@
 To mitigate the computational costs of matrix inversion in classical detectors like ZF and MMSE, Berra et al.
-[58] proposed unfolding accelerated Chebyshev successive over-relaxation and over-relaxation algorithms.
+[77] proposed unfolding accelerated Chebyshev successive over-relaxation algorithm.
 This network requires fewer trainable parameters and features a fast, stable training scheme.
-The unfolded detector achieves better performance compared to DetNet [133], OAMP-Net2 [10], TPG-detector [61], MEPD [51], EP-Net [46], and other methods like [53, 44].
+The unfolded detector achieves better performance compared to DetNet [49], OAMP-Net2 [11], TPGD-detector [50], MEPD [71], EP-Net [34], and other methods like [32].
+TABLE III: Summary of typical deep unfolding models for signal detection and their advantages and disadvantages.
+Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
+DetNet [49] First deep unfolding model for MIMO signal detection based on unrolling PGD method.
+In a $60\times 30$ MIMO system, DetNet achieves $3$ dB SNR gain over ZF for BER $=10^{-3}$, performs comparable to SDR and AMP, and approaches SD.
+Advantages: Explainable (follows PGD optimization flow), performs well when $N_{r}\gg N_{t}$.
+Disadvantages: Complex architecture (high computational cost), poor generalization under correlated channels.
+LcgNet [57] Unfolded CGD for massive MIMO detection.
+Achieves $1.52$ dB SNR gain over LMMSE in $32\times 128$ MIMO at BER $=10^{-4}$.
+Outperforms DetNet with lower complexity and faster training.
+Advantages: Fast training (practical for low-end CPUs).
+Disadvantages: Trained on Rayleigh fading, suffers significant BER degradation when tested on a spatially correlated channel.
+OAMP-Net2 [11] Lightweight unfolding of OAMP with soft input/output.
+Outperforms DetNet and AMP in 4$\times$4 and 8$\times$8 MIMO system, robust under imperfect CSI.
+Advantages: Fast training (4 parameters).
+Supports turbo equalization (soft I/O).
+Disadvantages: Underperforms in 3GPP channels (assumes unitarily-invariant matrix).
+MMNet [54] Online-trained unfolded ISTA for 3GPP channels.
+Achieves $8$ dB SNR gain over MMSE for SER $=10^{-3}$ (and $2.5$ dB over OAMP-Net2 [11] under imperfect CSI).
+Advantages: Robust to correlated channels, high-order QAM, and estimation errors.
+Disadvantages: Online training adds latency (unsuitable for fast-varying channels).
+ADMM-Net [38] Inexact ADMM-based unfolding with soft $tanh$ demodulation (supports $16/64$ QAM).
+Achieves SER within $1.2$ dB of SD in $30\times 60$ MIMO, outperforms DetNet with lower complexity.
+Advantages: Robust to channel correlation and imperfect CSI.
+Disadvantages: More layers (5205205–205 – 20) improve accuracy but increase computation, hence a tradeoff between performance and runtime.
+LoRDNet [61] Unfolding PGD for 1-bit ADC receivers without CSI.
+Outperforms classical ML detectors with just 500 samples.
+Advantages: Works with low-resolution ADCs; data-efficient.
+Disadvantages: Periodic retraining required for block fading channels (reduced SE due to pilot overhead) .
 To avoid the computational burden incurred due to matrix inversion, the authors in [78] have proposed a sparse refinement architecture for symbol detection in uplink massive MIMO.
-The authors have used deep unfolding for various crucial components in matrix inversion, significantly reducing the detection complexity.
-For detection, the alternating direction method of multipliers (ADMM) algorithm is utilized in [153].
-ADMM decomposes a complex optimization problem into simpler subproblems, which are solved iteratively.
-The algorithm combines the principles of dual ascent and the method of multipliers.
-Each iteration involves solving the subproblems for the primal variables in an alternating manner, followed by updating the dual variables (Lagrange multipliers) to enforce consistency with the original problem’s constraints [154].
-This iterative process ensures that the subproblem solutions converge to the global solution of the main optimization problem.
+The authors incorporated deep unfolding for components in matrix inversion, to reduce the detection complexity.
+The sparse weight matrix associated with the Jacobian iteration is trained across the layers of the network.
+Liao et al.
+[32] incorporated interference cancellation into DNN layers.
+Inspired by [79], their sequential detector trains residual error coefficients with mean square error (MSE) loss, reducing complexity by avoiding matrix inversion and relying on matrix addition and multiplication.
+Compared to DetNet [49], which requires $90$ layers for convergence, their model achieves convergence with just $8$ layers, offering significant computational simplicity.
 The authors in [38] developed an unfolded ADMM-based network for MIMO detection.
-They employed an inexact ADMM update [154, 156] to reduce complexity while achieving near-optimal performance.
-Compared to DetNet [49, 133], the proposed network requires fewer trainable parameters, resulting in improved performance in terms of symbol error.
-IV-B Modern Multicarrier Modulation
-An OFDM-based MIMO detector faces many challenges.
-The increased number of antennas at the transceiver exhibits enhanced spatial multiplexing but at the cost of increased inter-symbol interference (ISI).
-In standard OFDM multicarrier modulation, the cyclic prefix is inserted between OFDM blocks to mitigate ISI.
-Zhou et al.
-in [90], have unfolded a conjugate gradient method based OAMP detector for the OFDM-MIMO system.
-To address the issues faced in detecting coarsely quantized or low-resolution signals in OFDM-based MIMO systems, Shao et al.
-[91] have unfolded the expectation maximization (EM) algorithm.
-By unfolding the likelihood ascent search algorithm, Ullah et al.
-[92] are able to jointly detect and estimate the soft output in MIMO OFDM systems.
-Li et al.
-[46] proposed a deep unfolding network for OTFS detection by converting the iterative EP algorithm into structured network layers.
-The core idea is to approximate the conditional probability distributions $p(\bf{y}|\bf{x},\bf{H})$ and $p(\bf{x})$ where $p(\cdot)$ represents the probability, and $\bf{y},\bf{x},\bf{H}$ denote the received signal vector, transmitted signal vector, and channel matrix, respectively.
-The trainable parameters in this network include the step size, damping factor, and offset values for the mean ($m$) and variance ($v$) of the probability distribution.
-During offline training, the number of layers in the unfolding network is selected to balance performance and computational complexity.
-To enhance convergence and prevent the network from being trapped in local optima, the Adam optimizer with warm restarts is employed.
-This technique adaptively adjusts the step sizes, improving both training stability and performance.
-In [101], an unfolded OAMP-based detector for OTFS is introduced to handle channel estimation errors.
-This detector is trained using imperfect CSI and optimizes three trainable parameters: step sizes for updating the linear estimate, nonlinear estimate, and noise variance.
-IV-C Next Generation Multiple Access
+They employed an inexact ADMM update as in [37, 80] to reduce complexity while achieving near-optimal performance.
+Compared to DetNet [49], the proposed network [38] requires fewer trainable parameters, resulting in improved performance in terms of symbol error.
+Since, in [38], perfect CSI was assumed for MIMO detection, the authors in [81], proposed an ADMM-based unfolded algorithm with channel estimation errors.
+A set of penalty and relaxation parameters are learned layer wise.
+Moreover, the complexity of the proposed network is reduced as compared to OAMP-Net2 [11].
+III-B Modern Multicarrier Modulation
 A cellular network must simultaneously support a growing number of active subscribers within a limited time-frequency spectrum [82].
@@
 Multi-carrier modulation techniques like OFDM provide an effective solution to mitigate ISI by dividing a wideband signal into a set of orthogonal narrowband subcarriers [83].
-This approach not only minimizes interference but also enhances robustness in multipath environments.
+MIMO-OFDM detectors face many challenges due to the complexity introduced by high dimensional signal space.
+A low complexity receiver for a cyclic prefix-free MIMO-OFDM system is explored in [84].
+The receiver unfolds the OAMP algorithm and scalable parameters are introduced for optimizing the channel estimate matrix.
+The network was tested in over the air experiments and it performed reliably with low synchronization errors under imperfect CSI.
+Similarly, detection of coarsely quantized or LR signals in one-bit OFDM-based MIMO systems is a challenge.
+Hence, to address this challenge Shao et al.
+[85] have proposed an unfolded expectation maximization (EM) algorithm.
+One-bit ADCs severely quantize received signals, discarding amplitude information, and hence, ML detection becomes intractable due to nonlinear quantization effects.
+Hence, EM iteratively estimates unquantized symbols by computing its conditional expectation and optimizes the parameters in the maximization step.
+At high SNR, unfolded EM [85] performs better than ZF for one bit MIMO-OFDM detection.
+Furthermore, by unfolding the likelihood ascent search algorithm, Ullah et al.
+[86] developed an unfolded network to jointly detect and estimate the soft output in MIMO-OFDM systems.
+The likelihood ascent search algorithm is a neighborhood search-based detector for the MIMO system.
+The algorithm begins with an initial symbol estimate from linear detectors (ZF or MMSE) and iteratively refines the solution by flipping bits in the detected symbols to minimize the likelihood cost function.
+Near-optimal detection is achieved with the unfolded network, $1.2$ dB gap with 16-QAM as compared to SD at BER of $10^{-5}$ at polynomial complexity.
+Zhang et al.
+developed an unfolding-based framework for MIMO-OFDM detectors in [87].
+The EP is unfolded with layer-specific damping factors that are trained to adjust effectively to varying channel conditions.
+The proposed network achieved near-optimal detection across a wide range of channels and noise levels and demonstrated robustness in over-the-air experiment tests.
+Underwater acoustic communication is widely used for long-range transmissions, reaching up to tens of kilometers [88, 89, 90].
+Sound waves are used to transmit data through water since radio and light waves attenuate rapidly underwater.
+Channel estimation in underwater acoustic communication is particularly challenging due to doubly-selective channels that exhibit both severe time variation (from Doppler effects) and frequency selectivity (from long multipath delays).
+Hence, for OFDM systems in complex doubly-selective underwater acoustic channels, Zhao et al.
+in [35] proposed an unfolded equalizer (UDNet).
+The unfolded algorithm is based on unfolding the classical MMSE equalizer.
+In terms of performance, UDNet achieved better generalization and maintained lower computational complexity than DetNet [49].
+Li et al.
+[34] proposed an unfolded EP network (EP-Net) for OTFS detection.
+The trainable parameters include the step-size, damping factor, and offset values for the mean and variance of the posterior conditional probability distribution.
+To enhance convergence and prevent the network from being trapped in local optima, the Adam optimizer with warm restarts is employed.
+In [91], an unfolded OAMP-based detector for OTFS is designed to handle channel estimation errors.
+This detector is trained using imperfect CSI and optimizes three trainable parameters: step-sizes for updating the linear estimate, nonlinear estimate, and noise variance.
+Numerical analysis demonstrates that at SNR of $30$ dB, the proposed detector achieves a BER of $2.4\times 10^{-4}$ while EP-Net [34] achieves BER of $1.2\times 10^{-3}$.
+III-C Next Generation Multiple Access
 Multiple access techniques allow multiple users to share the same communication channel and resources, such as time, frequency, or code, to transmit data efficiently.
-Traditional multiple access schemes achieve this by dividing signaling dimensions orthogonally across time, frequency, or code [159].
+Usually multiple access schemes achieve this by dividing signaling dimensions orthogonally across time, frequency, or code [92].
 This orthogonality enables low-complexity multi-user detection at the receiver.
@@
 NOMA is categorized into power-domain NOMA and code-domain NOMA [93].
-In power-domain NOMA, the transmitter superimposes information symbols into a single waveform using superposition coding.
-At the receiver, SIC is employed to iteratively decode signals until the desired signal is recovered.
-A deep unfolded detector, named the learned preconditioned conjugate gradient descent network with SIC, is proposed for downlink MIMO-NOMA systems in [105].
-To achieve a robust model with low dataset requirements, the algorithm incorporates a data transfer learning strategy.
-This approach adapts the model efficiently to variations in environmental factors, such as channel models, modulation schemes, or power allocation, ensuring consistent performance across diverse scenarios.
-In sparse spread CDMA, each user is assigned spreading codes with only a few non-zero elements.
-This sparsity minimizes overlap and enables multiple users to transmit simultaneously on the same resources with minimal interference.
-In sparse CDMA, the spreading matrix contains sparse non-zero elements, as initially studied in [161].
-Each user’s transmitted symbols are encoded using unique sparse signature sequences, which facilitate decoding at the BS [162].
+Power domain NOMA multiplexes users by assigning different power levels on the same resource, such as time, frequency, or code, and relies on SIC at the receiver.
+Whereas, code domain NOMA multiplexes users by assigning different non-orthogonal spreading codes on the same resource block.
+A deep unfolded detector, named the learned preconditioned CGD network with SIC, is proposed for downlink MIMO-NOMA systems in [94].
+The algorithm incorporates a data transfer learning strategy to achieve a robust model with low data set requirements.
+This approach efficiently adapts the model to variations in environmental factors, such as channel models, modulation schemes, or power allocation, ensuring consistent performance in diverse scenarios.
+Tanner-graph-based transmission schemes for NOMA are discussed in [95].
+To enhance detection efficiency, a message-passing detector is unfolded, allowing the detector to improve as more information about transmitted data is learned during training epochs.
+In sparse code division multiple access (CDMA), each user transmits using sparse signature codes to reduce interference and enable efficient multiuser access [96, 97].
 Takabe et al.
-proposed a complex sparse trainable projected gradient detector (C-STPG) for SCDMA networks in [106].
-The detector builds upon the PGD framework, delivering significantly improved detection performance compared to conventional LMMSE detectors.
-Simulation results demonstrate reduced computational complexity during execution and minimal training costs due to the limited number of trainable parameters.
-The authors also introduced gradual sparsification—a joint learning technique that designs sparse signature sequences using deep unfolding.
-This technique optimizes the weights of the signature matrix while reducing error during training through MSE loss.
-The trainable parameters include the step size for gradient descent ($\gamma$) and a softness parameter ($\alpha$) to adjust the strength of the soft projection.
-IV-D Orbital Angular Momentum
-Orbital angular momentum (OAM) MIMO multiplexing leverages electromagnetic waves with different orthogonal OAM modes [163].
-These modes, theoretically infinite in number, form an orthogonal basis that can be used to multiplex distinct signals.
-A comprehensive review of OAM wave generation and detection is presented in [153].
-In the context of OAM MIMO systems, Kamiya et al.
-[108] developed a signal detection algorithm specifically designed for systems employing uniform circular array antennas.
-These systems often face inter-mode interference, particularly when the transmitter and receiver antennas are misaligned, which degrades both MMSE capacity and detection performance.
-To address this issue, the ADMM algorithm is unfolded with parallel interference cancellation.
-This approach significantly enhances detection performance, even in scenarios where alignment imperfections introduce severe inter-mode interference.
-In addition to detection, managing the peak-to-average power ratio (PAPR) presents a significant challenge in OAM multiplexing systems.
-To address this, Kamiya et al.
-[164] proposed a deliberate amplitude clipping method for OAM systems employing uniform circular array antennas.
-This approach involves applying amplitude clipping to digital baseband signals before OAM beamforming and pulse-shaping filtering.
-By doing so, it effectively reduces peak power, minimizes out-of-band radiation, and preserves the orthogonality of OAM modes.
-To counteract the bit error rate (BER) degradation introduced by clipping, the authors developed an iterative distortion recovery algorithm.
-Drawing inspiration from clipping noise cancellation techniques used in OFDM systems, the algorithm is implemented using deep unfolding.
-This implementation allows for trainable parameters that adaptively mitigate clipping-induced distortions, leading to enhanced detection performance and improved robustness in practical OAM systems.
-V Channel Estimation
-Estimating the channel in massive MIMO systems presents significant challenges, primarily because of the large number of antennas at the transceiver.
-The complexity of estimation grows with the scale of the system, and noise further exacerbates the issue, resulting in imperfect channel estimates and rendering estimation errors unavoidable [3], [165], [166], [167].
-Pilot-based channel estimation techniques are commonly used to address these challenges.
-These techniques involve transmitting known pilot symbols, followed by processing at the receiver to estimate the channel’s characteristics and compensate for its effects.
-A block diagram illustrating the sequence of operations in a pilot-based channel estimation technique is provided in Fig.
-5.
-It typically includes the generation and transmission of pilot symbols, reception at the receiver, estimation of the channel matrix using algorithms such as least squares (LS) or MMSE, and application of the estimated channel to process the received signals.
-Figure 5: Block diagram for pilot-based channel estimation.
+proposed the complex sparse TPGD (C-STPG) detector in [98], achieving superior detection over LMMSE with low complexity and minimal training due to few trainable parameters.
+They also introduced gradual sparsification, a deep unfolding method that jointly learns sparse signature matrices by minimizing MSE, tuning step size and a softness parameter.
+III-D Discussions and Takeaways
+Deep unfolding has proven to be a compelling solution for signal detection in complex and high-dimensional problems, such as those in massive MIMO, low-resolution ADC, and time-varying systems.
+By transforming iterative algorithms such as gradient descent variants, ISTA, ADMM, and AMP into structured networks, unfolding approaches bridge interpretability and performance.
+Table III summarizes the novelties, performance gains, advantages and disadvantages of the most popular deep unfolding models for signal detection.
+While pioneering unfolding architectures such as DetNet [49] and LcgNet [57] have demonstrated the advantages of data-driven signal detection with explainable DNN structures, they exhibit limited generalization in challenging scenarios, such as correlated channels.
+More recent unfolding-based approaches, including MMNet [54] and ADMM-Net [38], address these limitations by offering robustness to both channel correlation and imperfect CSI.
+However, methods like MMNet [54] and LoRDNet [61] introduce additional latency and training complexity due to their reliance on online training and periodic retraining, making them less suitable for fast-varying channel environments.
+The research community has actively tested deep unfolded detectors in real-world scenarios.
+Over-the-air test platforms [87, 84, 35] have been employed to validate the performance of unfolded detector networks under practical conditions.
+These real-world experiments confirm the feasibility of model-driven deep unfolding networks that are trained offline and subsequently deployed on software-defined radios or software-based receivers for real-time inference.
+The results demonstrate that deep unfolding techniques in signal detection can achieve robust BER performance in practical indoor environments, outperforming both traditional and purely data-driven methods.
+IV Channel Estimation
+Channel estimation is essential for signal detection because it allows the receiver to compensate for distortions introduced by the transmission medium such as fading, delay, and phase shifts.
 In the following, we discuss deep unfolding aided channel estimation methods.
-V-A Massive MIMO
+IV-A Massive MIMO
 MMSE-based channel estimation is commonly used in MIMO systems to estimate the channel response in noisy environments.
-It minimizes the mean squared error between the actual channel and its estimate using the second-order statistics of the channel and noise.
-In massive MIMO systems with FDD, reducing CSI feedback overhead is challenging.
-A method in [62] combines deep learning and superposition coding to address this.
-Downlink CSI is encoded, superimposed on uplink user data, and sent to the base station.
-A multi-task neural network unfolds the MMSE criterion to reduce interference and reconstruct the CSI efficiently.
-The iterative shrinkage-thresholding algorithm (ISTA) is commonly used for sparse channel estimation, enabling efficient CSI feedback.
-Each ISTA iteration includes a linear operation followed by a nonlinear soft-thresholding step, similar to the ReLU activation function [16].
+In massive MIMO systems with frequency division duplexing (FDD), reducing CSI feedback overhead is challenging.
+A method in [99] combines DL and superposition coding to address this.
+Downlink CSI is encoded, superimposed on uplink user data, and sent to the BS.
+A multi-task NN unfolds the MMSE process to reduce interference and reconstruct the CSI efficiently.
+ISTA is commonly used for sparse channel estimation, enabling efficient CSI feedback.
+Each ISTA iteration includes a linear operation followed by a nonlinear soft-thresholding step, similar to the rectified linear unit (ReLU) activation function [20].
 In [100], Chen et al.
-proposed a trainable deep learning architecture for sparse recovery problems using an adaptive depth approach.
+proposed a trainable DL architecture for sparse recovery problems using an adaptive depth approach.
 Instead of a fixed number of layers, the method dynamically adjusts the depth based on the problem’s complexity.
 Highly sparse channels use fewer layers, while more complex channels require additional layers.
-This adaptive approach improves accuracy and lowers computational costs compared to traditional methods and fixed-depth networks.
+This adaptive approach improves accuracy and lowers computational costs compared to classical methods and fixed-depth networks.
 Results in [100] demonstrate that the method reduces reconstruction error and speeds up channel estimation, making it more efficient.
@@
 It employs a deep unfolding-based feedback network to iteratively refine the feedback process and enhance feature extraction.
-Similarly, the decoding network uses a deep unfolding structure inspired by the ISTA algorithm.
+Similarly, the decoding network uses a deep unfolding structure inspired by ISTA.
 This iterative approach improves CSI recovery accuracy, making the process more efficient and adaptable to changing conditions.
 In massive MIMO systems, where the channel matrix is often sparse, the majorization-minimization (MM) approach is effective for addressing sparse channel estimation [102].
-In [64], the authors address the critical challenge of downlink channel estimation in massive MIMO, essential for efficient communication.
-Traditional methods often rely on angular-domain channel sparsity to reduce pilot overhead but struggle in dynamic environments.
-To overcome this, the authors propose an online two-timescale joint uplink/downlink dictionary learning and channel estimation algorithm.
-This method leverages channel sparsity and angle reciprocity between uplink and downlink transmissions to adapt to dynamic scenarios and significantly reduce pilot overhead.
-The two-timescale strategy includes learning a long-term dictionary to capture spatial channel characteristics and frequently estimating short-term sparse channel vectors.
-The short-term channel estimation subproblems are solved using an unfolded MM algorithm, while the long-term dictionary optimization uses standard convex methods.
-Deep unfolding is employed to jointly optimize the dictionary learning and channel estimation process for both uplink and downlink.
-Simulations in [64] show that this approach reduces pilot overhead and outperforms existing methods in static and dynamic environments, especially in scenarios with mobile users.
+In [103], the authors address the challenge of downlink channel estimation in massive MIMO.
+A two-timescale joint uplink/downlink dictionary learning and channel estimation algorithm is proposed.
+The short-term subproblems focuses on estimating sparse channel vectors for each timeslot.
+The MM algorithm is unfolded, and the sparse channel vectors are treated as trainable parameters.
+In contrast, the long-term subproblem involves updating the dictionary to adapt to dynamic environments for which a constrained stochastic successive convex approximation (SSCA) framework is employed without deep unfolding.
+Simulations results in [103] demonstrate that the proposed approach reduces pilot overhead and outperforms existing methods in both static and dynamic environments, particularly in environments with mobile users.
 The matching pursuit algorithm is a greedy iterative method often applied in sparse channel scenarios like massive MIMO.
@@
 This network estimates the multiuser massive MIMO channel in an unsupervised manner, initializing with a dictionary of imperfect steering vectors.
-The system operates in time division duplex (TDD) mode to leverage channel reciprocity.
+The system operates in time division duplex mode to leverage channel reciprocity.
 The network dynamically adjusts the number of layers based on incoming data, ensuring efficient performance in realistic environments.
@@
 In [105], the authors observed that existing sparse channel estimation schemes for downlink CSI often deliver suboptimal reconstruction due to the use of random measurement matrices.
-To address this, they proposed deep unfolding-based autoencoders that optimize the measurement matrix by unfolding the basis pursuit algorithm.
+To address this, they proposed deep unfolding-based autoencoders, that optimize the measurement matrix by unfolding the basis pursuit algorithm.
 This customized design enhances reconstruction performance in sparse channel estimation tasks.
@@
 Hu et al.
-[67] introduced a deep unfolding framework called deep deterministic policy gradient (DDPG) for multi-user MIMO systems.
+[107] introduced a deep unfolding framework called deep deterministic policy gradient (DDPG) for MU-MIMO systems.
 DDPG, a variant of deep Q-networks, is designed to handle problems in continuous action spaces [108].
-Unlike traditional deep unfolding neural networks with fixed depth, the authors leveraged DDPG to dynamically adjust the network depth based on the complexity of the problem.
+Unlike conventional deep unfolding NNs with fixed depth, the authors leveraged DDPG to dynamically adjust the network depth based on the complexity of the problem.
 This adaptive design enhances the efficiency and performance of sparse channel recovery in varying scenarios.
 The channel estimation problem is formulated using an off-grid sparse Bayesian learning model.
-The sparse Bayesian learning algorithm is unfolded layer by layer, introducing trainable parameters such as grid points and off-grid variables.
-Simulation results show that the proposed method significantly outperforms fixed-depth models, achieving a 30% reduction in the number of layers while improving normalized mean square error performance.
-V-B mmWave Massive MIMO
-mmWave massive MIMO operates in the high-frequency range of 30 GHz to 300 GHz, allowing higher dimensionality and the arrangement of many antennas in a compact space due to compressed wavelengths.
+The sparse Bayesian learning algorithm is unfolded with trainable parameters grid points and off-grid variables.
+Simulation results show that the proposed method significantly outperforms fixed-depth models, achieving a 30% reduction in the number of layers while improving normalized MSE (NMSE) performance.
+IV-B mmWave Massive MIMO
+mmWave massive MIMO operates in the high-frequency range of 30 GHz to 300 GHz, offering abundant spectral resources to alleviate the bandwidth crunch at sub-6 GHz frequencies [82].
+However, mmWave signals suffer from severe propagation and penetration losses.
+To compensate for these losses, large-scale antenna arrays with tens to hundreds of elements are required to achieve high power gains.
+The short wavelength (1–10 mm) of mmWave signals allows compact integration of these arrays into small transceivers.
+These features introduce unique challenges in designing physical-layer transmission algorithms [109].
 The BS, equipped with numerous antennas, must acquire downlink CSI.
-However, the large amount of feedback required creates significant overhead, posing a challenge in mmWave massive MIMO systems.
-Compressed sensing techniques are widely used to reduce feedback overhead by exploiting channel sparsity [172].
+However, the requirement of large amount of feedback creates significant overhead, posing a challenge in channel estimation in mmWave massive MIMO systems.
+Compressive sensing techniques are widely used to reduce feedback overhead by exploiting channel sparsity [110].
 The limited scattering properties of mmWave propagation and the power-focusing capability of lens antenna arrays transform the spatial domain channel into a sparse beamspace representation.
@@
 However, performing beamspace channel estimation in massive MIMO systems with limited RF chains remains a challenging task.
-Compressive sensing aided deep unfolding is explored in [173], where the authors apply Onsager correction for AMP [174] and construct an unfolded model similar to [18].
-The network in [175] uses deep unfolding to learn beamspace channel estimates.
-The authors demonstrate that this algorithm performs excellently with large antenna arrays and limited RF chains.
+Compressive sensing aided deep unfolding is explored in [111], where the authors apply Onsager correction for AMP as explained in [112].
+The architecture of the unfolded model, learned AMP (LAMP) [111] is similar to unfolded ISTA in [20].
+A learned denoising-based AMP network (LDAMP) is proposed in [113] to learn beamspace channel estimates.
+The authors demonstrate that the iterative sparse signal recovery algorithm with denoising achieves robustness in channel estimation particularly in scenarios with large antenna arrays and limited RF chains.
 Yi et al.
-in [176] introduced an unfolded network called LampResNet (Learned AMP with deep residual learning).
+in [114] introduced an unfolded network called learned AMP with deep residual learning (LampResNet).
 In this model, the AMP algorithm is unfolded to estimate the channel, and deep residual learning is employed to mitigate channel noise.
-Normalized MSE analysis shows that the proposed networks outperform LAMP [173] and achieve performance similar to LDAMP [175].
+NMSE analysis shows that the proposed network in [114] outperforms LAMP [111] and achieves performance comparable to LDAMP [113] in terms of estimation accuracy.
 However, exploiting channel sparsity does not always yield optimal estimation performance.
-To address this, the authors in [85] investigate mmWave channel sparsity in the beamspace.
+To address this, the authors of [115] investigated the sparsity of the mmWave channel in the beamspace.
 They formulate the estimation problem as a sparse signal recovery task and improve convergence speed by unfolding the iterative PGD algorithm.
-In this approach, the gradient descent step size is treated as a trainable parameter.
-V-C Modern Multicarrier Modulation
-For frequency-selective channels, OFDM is a preferred multicarrier modulation technique.
-It spreads the data across multiple orthogonal subcarriers in the frequency domain, helping to mitigate the effects of frequency-selective fading.
-The channel is typically modeled on a per-subcarrier basis.
-To enable accurate CSI feedback by transforming the channel matrix into a sparse representation, Yiyun et al.
-[93] proposed a neural network consisting of compression and reconstruction modules for CSI compression and feedback.
-They designed the reconstruction module as an unfolded ISTA network and found that higher SNR improves reconstruction performance compared to the traditional BP method.
-Hu et al.
-[94] introduced an unfolded algorithm called the learnable optimization and regularization algorithm (LORA) to enhance CSI feedback speed and accuracy.
-In this approach, the conventional ISTA algorithm is unfolded, and the regularization term is learned through end-to-end training.
-They modeled the regularization term as a learnable, differentiable transform, further refined in [94].
-By incorporating this learnable regularization module, the model better captures CSI characteristics and adapts to different levels of channel estimation error.
-To improve CSI feedback accuracy and efficiency in OFDM FDD-based massive MIMO systems, while reducing complexity at the user equipment, Yangyang et al.
+In this approach, the step-size of the descent step is treated as a trainable parameter.
+In [116], an iterative trimmed-ridge regression algorithm is unfolded to improve beamspace channel estimation.
+The proposed algorithm involves parallel computation to achieve negligible latency.
+These algorithms were applied to both narrowband and wideband massive MIMO systems, to ensure versatility in channel estimation.
+IV-C Modern Multicarrier Modulation
+Yiyun et al.
+[117] proposed a CSI recovery network for FDD downlink massive MIMO-OFDM systems.
+ISTA is unfolded and the network consists of compression and reconstruction modules for CSI compression and feedback.
+Building on ISTA, Hu et al.
+[118] introduced an unfolded algorithm called the learnable optimization and regularization algorithm (LORA) to enhance the precision of CSI feedback.
+This approach also unfolds ISTA, and the regularization term is learned through end-to-end training.
+The model captures CSI characteristics and adapts to different levels of channel estimation error.
+Furthermore, to improve CSI feedback accuracy and efficiency, Yangyang et al.
 [119] proposed an unfolded ISTA network.
-Their study showed that the real and imaginary components of the channel differential term share similar correlations, simplifying network design.
-Unfolding ISTA significantly enhances the CSI encoder’s reconstruction performance.
+Their study showed that the real and imaginary components of the channel differential terms, i.e, the difference between CSIs at adjacent time slots, exhibit similar correlation patterns in both the angular and delay domains.
+This property enables a shared model to process both components, thereby reducing network complexity and enhancing the efficiency of the CSI encoder.
+Each of the networks proposed in [117, 118, 119] enhance CSI feedback performance for FDD massive MIMO-OFDM system.
 In [33], Guo et al.
 proposed a two-stage low-rank CSI feedback scheme for a low-rank OFDM mmWave channel.
-In massive MIMO systems operating in FDD mode, CSI is estimated by unfolding the fast ISTA algorithm.
-Fast ISTA is an improved version of the conventional ISTA algorithm.
-A set of thresholds is derived from the input features of the network, and to reduce the uncertainty in CSI reconstruction, the iterative steps are mapped to layers of a neural network, with $L$ stages corresponding to $L$ iterations of fast ISTA.
-In [96], the authors explore quantized compressed CSI at the user equipment before feedback.
-They propose a deep unfolding-based bit-level CSI feedback method to mitigate the effects of accumulated quantization errors.
-Building upon ISTA, the authors use tied learned ISTA to reduce the number of learnable parameters, offering a faster and more accurate solution to the signal reduction problem.
-In tied learned ISTA, the trainable parameters are shared across all layers [177].
+In massive MIMO systems operating in FDD mode, CSI is estimated by unfolding the fast ISTA.
+Fast ISTA is an improved version of the classical ISTA due to the introduction of momentum term that accelerates the convergence rate.
+The network, FISTA-Net, includes trainable step-size and thresholds to achieve accurate reconstruction performance.
+In [120], the authors explore quantized compressed CSI at the user equipment before feedback for massive MIMO-OFDM system.
+They propose a deep unfolding-based bit-level CSI feedback network, TiLISTA, to mitigate the effects of accumulated quantization errors.
+Building upon ISTA, they used tied learned ISTA as discussed in [121].
+Since the trainable parameters are shared across all layers, the number of learnable parameters are reduced, offering a faster and more accurate solution to the signal reduction problem.
+FISTA-Net [33] focuses on faster convergence while TiLISTA [120] targets robustness to quantization errors.
+In [122], an unfolded sparse Bayesian learning algorithm is derived to capture channel sparsity in a mmWave massive MIMO-OFDM system with hybrid beamforming (HBF).
+The proposed network achieved a lower NMSE of $0.033$ outperforming the classical sparse Bayesian learning algorithm which recorded an NMSE of $0.1203$.
 Jianqiao et al.
-[97] developed an unfolded inverse-free variational Bayesian learning method to reduce the high computational complexity of matrix inversion.
-The sparse channel recovery performance is enhanced by learning the a-priori parameters of the sparsely represented dictionary for channel estimation.
-In [98], based on approximate Bayesian inference and an extension of expectation consistent approximation, the authors proposed a generalized expectation consistent algorithm, an iterative message-passing approach for sparse channel estimation.
-In a lens-based beamspace massive MIMO system, this approach compensates for large channel attenuation in the mmWave band.
-The authors introduced a deep unfolded unsupervised learning network called the learned denoising-based generalized expectation consistent signal recovery network.
-This network unfolds the generalized expectation consistent signal recovery algorithm and formulates the wideband beamspace channel estimation problem as a compressed image recovery task.
-Stein’s unbiased risk estimator is used as the loss function.
-In [99], the authors leveraged angle-domain channel sparsity to reduce uplink pilot overhead during channel estimation.
-They designed an unfolded multiple-measurement-vectors learned AMP network by jointly training the phase shift network and channel estimator, integrating the channel sparsity for improved performance.
-Underwater acoustic communication is widely used for long-range transmissions, reaching up to tens of kilometers [178, 179, 180].
-The main challenges in underwater communication include frequency attenuation, time-varying multipath, and large Doppler effects.
-Zhao et al.
-[47] explored integrating the MMSE structure within the layers of a DNN to create an interpretable model.
-This approach aims to build an equalizer framework for improving underwater communication performance.
-Figure 6: Analog, digital and hybrid beamformer [181].
-Cognitive radio is a wireless technology that continuously monitors the spectrum to identify unused channels, or spectrum holes, allowing for more efficient spectrum usage.
-In [100], a wideband signal compressed spectrum reconstruction algorithm is proposed to detect spectrum holes across a large band using a sub-Nyquist sampling rate.
-The authors integrate ISTA into layers of a CNN, improving the performance in both the probability of detection and the probability of false alarm for spectrum hole detection.
-V-D IRS
+[123] proposed an unfolded inverse-free variational Bayesian learning framework to address the high computational complexity of matrix inversion in mmWave MIMO-OFDM systems.
+The sparse channel recovery performance is enhanced by learning the a priori parameters of the sparsely represented dictionary for channel estimation.
+In [124], based on approximate Bayesian inference and an extension of expectation-consistent approximation, the authors proposed a generalized expectation consistent algorithm, an iterative message-passing approach for sparse channel estimation.
+In a lens-based beamspace massive MIMO-OFDM system, this approach compensates for large channel attenuation in the mmWave band.
+The authors developed a deep unfolded unsupervised learning network called the learned denoising-based generalized expectation consistent signal recovery network.
+The generalized expectation consistent signal recovery algorithm as discussed in [125] is unfolded.
+Both the networks in [123] and [124] provide accurate channel recovery estimates for mmWave channel with few RF chains.
+In [126], the authors leveraged angle-domain channel sparsity to reduce uplink pilot overhead during channel estimation for mmWave MIMO-OFDM systems.
+They propose an unfolded multiple-measurement-vectors-based AMP network.
+The phase shift network and channel estimator are jointly trained, thus integrating the channel sparsity for improved performance.
+Gao et al.
+[127] proposed a network, ComNet, that unfolds an OFDM receiver into two subnetworks: channel estimation and signal detection.
+The LMMSE and ZF are unfolded for channel estimation and signal detection, respectively.
+The weights of the LMMSE and ZF matrices are trained layer-wise.
+The simulation results show that the BER performance is better compared to the classical LMMSE estimation.
+Furthermore, Wang et al.
+in [128] proposed an unfolded least squares (LS) estimation to enhance channel estimation for OTFS MIMO system under fractional Doppler condition.
+The network, CENet, is structured as parallel sub-networks, each responsible for estimating one row (delay tap) of the delay-time channel matrix.
+At BER= $3\times 10^{-3}$, CENet achieved $\sim 5$ dB SNR gain over LS estimation algorithm.
+TABLE IV: Summary of typical deep unfolding models for channel estimation and their advantages and disadvantages.
+Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
+LDAMP [113] Denoising-based unfolded AMP model for sparse-beamspace channel estimation in mmWave massive MIMO.
+Outperforms classical denoising AMP method with small number of RF chains at the receiver.
+Advantages: Simple architecture and fast convergence (within $5$ layers), require a few RF chains.
+Disadvantages: Fixed parameters post-training restrict the model’s adaptability, performance may degrade in varying scenarios.
+LampResNet [114] Combines unfolded AMP with deep residual learning to enhance robustness against noise.
+Achieves close performance to LDAMP [113] at high SNR regime ($25$ dB).
+Advantages: Resilient to noise, consistent NMSE improvement under SNR variations.
+Disadvantages: Rely on prior information of beamspace channel, not consistent with realistic channel models, and network may fail to converge.
+FISTA-Net [33] Fast ISTA-based deep unfolding for low-rank CSI feedback in mmWave OFDM systems.
+Outperforms the classical FISTA with lower complexity.
+Advantages: Fast convergence ($20$ layers).
+Disadvantages: Requires retraining for major environment variations.
+mpNet [104] Deep unfold model based on the matching pursuit method for multiuser massive MIMO systems.
+Outperforms classical LS estimator in terms of NMSE under varying SNR.
+Advantages: Avoids ground-truth channel data by using unsupervised online learning on noisy LS estimates.
+Disadvantages: Need prior knowledge of physical wave propagation model, as small uncertainty of system parameters lead to high estimation performance loss.
+LORA [118] The deep unfolding model based on ISTA introduces a learnable regularization module to adaptively capture the CSI structure.
+Improved CSI reconstruction accuracy by $>2$ dB NMSE over TiLISTA [117].
+Advantages: LORA exhibits strong denoising capability under channel estimation error.
+Disadvantages: The number of layers is critical: using fewer than four layers degrades reconstruction accuracy, while increasing the number beyond four leads to higher computational complexity with only marginal improvements in NMSE.
+IV-D IRS
 To mitigate the blockage effect and extend the connectivity range for mmWave communication, IRS is a key technology proposed to enhance both energy and SE.
 IRS consists of many reflecting elements that steer the propagation of electromagnetic waves by adjusting their reflection coefficients.
-However, to unlock the full potential of IRS, accurate channel estimation is required for both individual and cascaded links.
+However, to unlock the full potential of IRS, accurate channel estimation is required for both direct and cascaded links.
 This is a challenging task because IRS-based channel estimation involves estimating multiple channels simultaneously: the direct channels between the BS and each user, the channel between the BS and IRS elements, and the channel between IRS elements and users.
 In [129], Jiguang et al.
-proposed a method for estimating the rank-deficient cascaded channel in an IRS-assisted SIMO system.
+proposed a method for estimating the rank-deficient cascaded channel in an IRS-assisted single-input multiple-output system.
 They developed a deep unfolding network based on a gradient descent algorithm.
@@
 The proposed network addresses this challenge by compensating for the computational complexity.
-Parallel factor decomposition is an efficient technique for estimating multiple large channel matrices in massive MIMO communication systems.
-It allows a high-dimensional matrix to be decomposed into a linear combination of several rank-one matrices, making the estimation of low-rank matrices more computationally efficient [182].
-In [110], the parallel factor decomposition is unfolded to estimate multiple large channel matrices in the uplink of an IRS-empowered multi-user MISO communication system.
 Awais et al.
-[111] developed two networks based on the MAP estimation criterion.
+[130] developed two networks based on the MAP estimation criterion to address the challenges in IRS cascaded channel estimation.
 The first network is a deep denoising network designed for noise-dominated SNR regions, providing support at various noise levels.
 The second network is a deep unfolded half-quadratic splitting network that alternates between optimizing extrapolation and denoising for interference-dominated regions.
-As the SNR increases, the proposed network shows significant improvement in normalized MSE.
+As the SNR increases, the proposed network shows significant improvement in NMSE.
 Under severe channel variations, the network delivers a performance gain of 20% in the high SNR range.
@@
 [131] proposed a two-stage learned AMP network with row compression to address the joint estimation problem of direct and cascaded channels in IRS-aided mmWave systems.
-By utilizing the deep unfolding technique, the network combines the advantages of compressive sensing and deep learning.
+By utilizing the deep unfolding technique, the network combines the advantages of compressive sensing and DL.
 This approach enables reliable channel estimation with reduced training, which is crucial for efficiently reconfiguring wireless environments under IRS reflection control.
-VI Precoder Design
-This section reviews the applications of deep unfolding for precoder design used along with key 6G technologies.
-VI-A Massive MIMO
-In MIMO systems, beamforming enhances signal transmission and reception by directing signals toward specific directions using multiple antennas.
-This improves the received SNR [181, 183].
-There are three types of beamformers: analog, digital, and hybrid.
-Fig.
-6 illustrates the architectures for analog, digital, and hybrid beamforming [181].
-The hybrid beamformer system model is described in [184].
-A digital beamformer assigns one RF chain per antenna.
-However, in large-scale arrays, it has high power consumption and complex architecture.
-Analog beamformers use time delay units or phase shifters for antenna weighting but struggle with inter-user interference in multi-user scenarios.
-To overcome these issues, hybrid beamforming was introduced [185, 186].
-Beamforming design is an optimization problem where the objective is to maximize a network utility function, such as sum rate, mean rate, or harmonic rate.
-This optimization is subject to resource and hardware constraints, including power budget and constant modulus analog beamforming coefficients.
-Consider a base station (BS) with $n_{\text{t}}$ antennas serving $M$ single-antenna users [27]. The received signal at user $m$ is given by:
-$y_{m}=\boldsymbol{{h}}_{m}^{H}\sum_{m=1}^{M}\boldsymbol{{f}}_{m}x_{m}+n_{m}$ (3)
-where, $x_{m}$ is the transmitted signal with an average power of $\mathbb{E}\{|x_{m}|^{2}\}=1$, $\mathbf{h}_{m}\in\mathbb{C}^{n_{t}\times 1}$ is the channel vector, $n_{m}\sim\mathcal{CN}(0,\sigma^{2})$ represents the AWGN noise at user $m$, and $\boldsymbol{f}_{m}\in\mathbb{C}^{n_{t}\times 1}$ is the beamforming vector for user $m$.
-The signal-to-interference-plus-noise ratio (SINR) at user $m$ is given by:
-$\text{SINR}_{m}=\frac{\left|{\boldsymbol{{h}}_{m}^{H}\boldsymbol{{f}}_{m}}\right|^{2}}{\sum_{i=1,i\neq m}^{M}\left|{\boldsymbol{{h}}_{i}^{H}\boldsymbol{{f}}_{i}}\right|^{2}+\sigma^{2}}.$ (4)
-This expression quantifies the received signal quality by measuring the desired signal power relative to interference and noise.
-The hybrid beamformer design problem can be formulated as the maximization of the sum rate, given by:
-$(\mathcal{P}_{\text{HBF}})\quad\max_{\boldsymbol{F}^{\text{DB}},\boldsymbol{F}^{\text{AB}}}$ $\sum_{m=1}^{M}\log_{2}\left(1+\text{SINR}_{m}\right)$ (5)
-s.t. $\left\lVert\boldsymbol{F}^{\text{AB}}\boldsymbol{F}^{\text{DB}}\right\rVert_{F}^{2}\leq P,\quad\boldsymbol{F}^{\text{AB}}\in\mathcal{F},$
-$|\boldsymbol{F}^{\text{AB}}|^{2}=1.$
-where $\boldsymbol{f}_{m}=\mathbf{F}^{\text{AB}}\boldsymbol{f}_{m}^{\text{DB}}$, $\mathbf{F}^{\text{AB}}\in\mathbb{C}^{n_{t}\times N_{\text{RF}}}$ is the analog precoding matrix with $N_{\text{RF}}$ RF chains, belonging to a codebook $\mathcal{F}$.
-The digital precoding vector for the $m^{\text{th}}$ user is $\boldsymbol{f}_{m}^{\text{DB}}\in\mathbb{C}^{N_{\text{RF}}\times 1}$, and the digital precoding matrix $\boldsymbol{F}^{\text{DB}}$ consists of column vectors $\boldsymbol{f}_{m}^{\text{DB}}$.
-These problems are nonconvex and NP-hard, meaning they lack closed-form solutions.
-The unit-modulus constraints on the analog precoding matrix make precoder design even more challenging.
-To address this, the algorithm unrolling technique combines model-driven and data-driven approaches.
-It optimizes beamforming matrices by integrating the interpretability of model-based methods with the adaptability of data-driven techniques.
+IV-E IoT devices
+Joint activity detection and channel estimation are critical in grant-free random access for massive machine type communication (mMTC), where numerous IoT devices access the network sporadically without scheduling.
+To enhance detection under imperfect CSI, Qiang et al.
+[31] proposed AMP-Net, an unfolded AMP based network.
+Each AMP iteration alternates between MMSE-based denoising and residual updates to minimize the MSE.
+The network uses four trainable scalar parameters to estimate activity probability, channel variance, regulate the denoiser, and control the state evolution.
+AMP-Net achieved accurate detection of the active device by extracting the CSI from the estimated device state matrix with a $4$ dB gain over classical AMP at SNR$=15$ dB.
+The authors in [132] proposed a grant-free access protocol for mmWave cell-free massive MIMO using a hybrid structure based on learned vector AMP, reducing reliance on prior channel knowledge and improving convergence over classical AMP.
+In [133], Ma et al.
+introduced a one-phase non-coherent scheme for mMTC.
+The data bits are embedded in pilot sequences to enable joint activity and data detection without channel estimation.
+They used unfolded AMP to leverage pilot correlation.
+In [134], an unfolded AMP with backpropagation was applied to an asynchronous random access scheme with data length diversity, addressing activity detection, channel estimation, and data recovery in mMTC.
+The authors in [135] proposed an ADMM-based deep unfolding network for joint channel estimation and active user detection in massive IoT systems.
+The trainable parameters include an auxiliary variable for matrix inversion, a regularization term balancing sparsity and accuracy, and a shrinkage threshold controlling output sparsity per iteration.
+Similarly, in [136], the ADMM algorithm is unfolded to improve convergence rate and recovery accuracy in grant-free vehicular networks, particularly for BS with a limited number of antennas.
+In [137], two unfolded networks were introduced for sporadic user detection and channel estimation in mMTC: ADMM-Net and VAMP-ISTA-Net.
+ADMM-Net unfolds a linearized ADMM [80] to address constrained convex problems, using a single iteration of the proximal gradient method to avoid solving subproblems explicitly.
+VAMP-ISTA-Net unfolds a vector AMP enhanced with an ISTA-based denoiser.
+Both architectures required only $5$ layers, significantly reducing the iteration count from $300$ in classical algorithms and achieving faster convergence.
+Overall, the approaches in [135, 136, 137] tackle the multiple measurement vector sparse recovery problem using unfolded ADMM frameworks for improved reconstruction performance.
+Shi et al.
+[138] considered grant-free massive access in an IoT network.
+A large number of single-antenna IoT devices sporadically transmit non-orthogonal sequences to a multi-antenna BS.
+An unfolded ISTA with trainable weight matrices is designed.
+The network trains faster due to less number of trainable parameters.
+Zou et al.
+in [139] formulated the joint activity detection and channel estimation problem as a group-row-sparse matrix recovery problem.
+A proximal gradient-based method is unfolded and it achieves better convergence than the network in [138].
+IV-F Discussions and Takeaways
+Channel estimation in modern wireless systems, especially in massive MIMO and mmWave settings, poses distinct challenges due to high dimensionality, feedback overhead, and sparse propagation characteristics.
+Deep unfolding addresses these issues by embedding structural priors into learning pipelines, as demonstrated by models like LDAMP [113] and LampResNet [114].
+Table  IV presents notable unfolded networks for channel estimation with their performance gains, advantages, and disadvantages.
+Nonetheless, deep unfolding for channel estimation remains sensitive to model assumptions (e.g., sparsity or statistical independence).
+Performance degradation in dense or correlated environments, retraining overhead, and reliance on fixed unfolding depth are key limitations.
+Moreover, model-agnostic unfolding and joint learning of estimation and feedback (especially in IRS and wideband settings) will be crucial for making deep unfolding a mainstream tool in physical layer design.
+Unlike deep unfolding models for signal detection, the practical deployment of such models for channel estimation remains limited.
+Moreover, their performance has yet to be thoroughly validated through real-world over the air experiments.
+V Precoder Design
+This section reviews the applications of deep unfolding for precoder design used in key 6G technologies.
+V-A Massive MIMO
+Beamforming in MIMO systems enhances signal transmission by steering signals toward the intended receivers using multiple antennas, thereby improving received SNR [140, 141].
+Beamforming techniques can be broadly divided into three types: analog, digital, and hybrid [140].
+Digital beamforming uses a digital precoding matrix with one RF chain per antenna, leading to high power consumption and complexity in large arrays.
+Analog beamforming employs phase shifters but struggles with inter-user interference.
+HBF [142, 143] addresses these issues by combining digital and analog methods, reducing RF chain count while maintaining performance and efficiency which is ideal for mmWave and massive MIMO in 5G and beyond.
 Shlezinger et al.
-[26] discuss the architectural constraints and challenges of hybrid beamforming optimization using both conventional and data-driven methods.
-The classic iterative WMMSE algorithm is widely used for precoder design to maximize the system’s weighted sum rate.
-It assigns different weights to the MSE of each user based on priority or quality of service.
-By adjusting these weights iteratively, WMMSE balances user performance and improves overall efficiency.
-The algorithm provides closed-form, iterative update rules to solve a surrogate optimization problem with the same local optima as the sum-rate maximization problem [122].
-Maximizing the weighted sum rate while satisfying a power constraint is a nonconvex problem [187].
-The approach proposed in [187] exploits the separable structure of power constraints.
-To achieve a locally optimal beamforming strategy, the authors introduce a simple and efficient distributed inexact cyclic coordinate descent method.
-A deep unfolded version of the iterative WMMSE algorithm is proposed for precoder design in multi-user MIMO systems [69].
-The authors call this network architecture the iterative algorithm induced deep unfolding network (IAIDNN).
+[28] discuss HBF optimization using classical, data-driven, and deep unfolding methods, highlighting deep unfolding’s efficiency with limited RF chains.
+For precoder design, the weighted MMSE (WMMSE) algorithm [144] maximizes the weighted sum rate via iterative updates involving matrix inversions and eigendecompositions.
+Given the nonconvexity of the power-constrained problem [145], a distributed inexact cyclic coordinate descent (CD) method is proposed for efficient locally optimal solutions.
+A deep unfolded architecture, iterative algorithm induced deep unfolding network (IAIDNN) is proposed for precoder design in MU-MIMO systems in [146].
+The iterative WMMSE algorithm is unfolded.
 To improve efficiency, WMMSE is reformulated to approximate matrix inversions.
@@
 Using this rule, the gradients of trainable parameters are computed effectively.
-The WMMSE algorithm [122] involves matrix inversions, eigendecompositions, and bisection searches in its iterations.
-Since any MMSE solution requires matrix inversion, WMMSE also relies on it.
-To improve convergence while avoiding matrix inversion, a downlink beamforming method using an unfolded PGD-WMMSE algorithm for multi-user MISO systems is proposed in [68].
+To improve convergence while avoiding matrix inversion, a downlink beamforming method using an unfolded PGD-WMMSE algorithm for multi-user multiple-input single-output (MISO) system is proposed in [147].
 The authors replace matrix inversions with PGD and compute gradients in parallel.
@@
 This limitation is addressed in [148].
-An unfolded WMMSE algorithm for transceiver design in multicell MU-MIMO interference channels with local CSI is proposed in [70].
-The authors use graph convolutional networks (GCNs) to unfold WMMSE, enabling coordination in multicell MU-MIMO systems.
-Graph filters and convolutional networks approximate certain WMMSE variables using graph learners.
+An unfolded WMMSE algorithm for transceiver design in multicell multi-user MIMO (MU-MIMO) interference channels with local CSI is proposed in [148].
+The authors use graph convolutional networks (a variant of GNN) to unfold WMMSE, enabling coordination in multicell MU-MIMO systems.
+Graph filters and convolutional networks approximate WMMSE-related variables, such as weighted beamformer matrices, by leveraging graph learning techniques.
 However, the approach in [148] lacks flexibility in learning the exact transformation of WMMSE variables.
@@
 The WMMSE algorithm is unfolded to enable distributed and synchronized precoding across cooperative BSs.
-Numerical analysis compares this approach with the networks in [69, 68].
 The methods in [146, 147] rely on centralized optimization, leading to high complexity and backhaul overhead.
-In contrast, the proposed network in [188] reduces computational cost while maintaining efficiency.
-The gradient descent algorithm is a widely used method for finding a stationary point.
+In contrast, due to the distributed precoding, the proposed network in [150] reduces computational cost while maintaining efficiency.
+An unfolded WMMSE-based NN for massive MU-MIMO under imperfect CSI is studied in [151].
+The proposed network enhances performance as compared to [144], with low complexity by leveraging CSI statistics for near-optimal precoder solutions, thus accelerating convergence.
 Shi et al.
-[72] designed a hybrid analog-digital transceiver using iterative gradient descent to obtain minimum symbol error rate (MSER) in massive MIMO systems.
-They unfolded the iterative gradient descent algorithm into a DNN, where learnable parameters serve as step sizes for the gradients.
-By using a fixed number of DNN layers with optimized step sizes, the approach reduces the number of iterations and speeds up convergence compared to conventional methods.
-The input to this MSER-based algorithm includes a user’s transmit beamforming matrix, an analog transmit beamforming matrix, an analog receive beamforming matrix, and a digital receive beamforming matrix.
-The step sizes guide the gradient descent update for each matrix.
-The gradients for these parameters are updated using SGD.
+[152] designed a hybrid analog-digital transceiver using iterative gradient descent to obtain minimum SER (MSER) in massive MIMO systems.
+They unfolded the iterative gradient descent algorithm into a DNN, where step-sizes for the gradients are the learnable parameters.
 The authors claim that their proposed network matches the performance of the MSER-based iterative gradient descent algorithm while significantly reducing complexity.
+Gradient-based methods have been successfully applied to beamforming optimization with reduced complexity.
+However, these approaches can be further improved using advanced optimization frameworks like PGD or projected gradient ascent (PGA).
+A robust frequency-selective precoding algorithm that maximizes the expected weighted sum rate while considering per-antenna power constraints is developed for wideband MIMO-OFDM system in [153].
+This frequency-selective precoding algorithm is then unfolded into a deep network, with the aid of PGD.
+The trainable parameters include step sizes and projections in each PGD layer.
+The proposed network achieves near-optimal sum rate with lower computational complexity.
 Zhu et al.
 [154] proposed a beamforming learning architecture by unfolding a parallel gradient projection algorithm for MISO heterogeneous systems.
-Their design optimizes scenarios where network parameters, such as the number of antennas, BSs, or overall network size, differ from those used during training.
-Gradient-based methods have been successfully applied to beamforming optimization with reduced complexity.
-However, these approaches can be further improved using advanced optimization frameworks like PGA.
-A fast and robust precoding network for multi-user MIMO systems with imperfect CSI and per-antenna power constraints is proposed in [74].
-The objective function is formulated as a geometric program that minimizes the MSE between the transmitted and recovered symbol vectors at the user.
-This optimization leverages the duality between downlink and uplink MSE feasible regions.
-The authors map the iterative optimization process into a layer-wise deep unfolding framework, enabling fast and robust downlink precoding prediction.
-PGA is a variant of the gradient ascent algorithm used to maximize an objective function under constraints.
-It updates the solution in the gradient direction to increase the objective function and then projects it back onto the feasible set.
+The trainable set includes the step-size for the gradient ascent update and weighted sum rate coefficients used to construct the beamforming gradient direction.
+The proposed network generalizes well across different numbers of antennas, different numbers of BSs, and different network sizes especially without retraining.
+PGA is a variant of the gradient ascent algorithm.
+It updates the solution in the gradient direction to maximize the objective function and then projects it back onto the feasible set.
 Lavi et al.
@@
 To handle noisy CSI, they implemented the projected conceptual mirror prox algorithm [156, 157].
-The step sizes for updating the analog and digital precoder matrices are tuned using mini-batch SGD, a data-driven approach.
+The step-sizes for updating the analog and digital precoder matrices are tuned using mini-batch SGD, a data-driven approach.
 The solution is computed within a fixed number of iterations, defined as network layers, ensuring a balance between interpretability and efficiency.
@@
 proposed an efficient analog beamforming design to maximize the achievable rate of massive MIMO in-band full-duplex systems.
-The beamforming problem is non-convex due to constant modulus constraints on analog beamforming coefficients.
-To address this, they introduced an unfolded algorithm combining PGA with ZF projection onto convex sets, achieving faster convergence.
+The underlying beamforming problem is non-convex, primarily due to constant modulus constraints imposed by the use of analog phase shifters.
+To overcome this challenge, they developed an unfolded algorithm based on PGA combined with ZF projection onto convex sets.
+The convex feasible set arises due to the requirements of the ZF constraint and unit norm beamforming vectors.
+This approach leverages the structure of the zero-forcing conditions and the constant modulus constraints, achieving significantly faster convergence and improved performance compared to classical iterative methods.
+Besides PGA and PGD, CGD is also an iterative algorithm that falls under the gradient-based optimization approach.
+In [159], the CGD algorithm is unfolded for constant envelope precoding to reduce computational overhead.
+The Riemannian manifold optimization (MO)-based CGD is implemented across network layers.
+Simulations show that this approach effectively suppresses multi-user interference, improving overall performance.
+A fast and robust precoding network for MU-MIMO systems with imperfect CSI and per-antenna power constraints is proposed in [160].
+The objective function is a geometric program that minimizes the MSE between the transmitted and recovered symbol vectors at the user.
+The iterative optimization framework is derived from the MSE uplink-downlink duality-based algorithm, which is unfolded to design a robust MSE duality network (RMSED-Net).
+RMSED-Net is robust to channel estimation errors and achieves sum rate close to the upper bound of its benchmark.
+In [161], the interior point method is unfolded for precoding design.
+It leverages interference in multi-user MISO systems to minimize power while meeting signal-to-interference and noise ratio (SINR) constraints by training the proximal operators and Lagrangian multipliers.
+The authors show that the execution time is reduced by ${}50-70\%$ compared to the classical method.
+However, the network is limited to phase shift keying (PSK) modulation, requiring reformulation for higher order and QAM modulations.
+A constructive interference-based symbol-level precoding for multi-user MISO downlink transmission is discussed in [162].
+The proximal Jacobian ADMM is unfolded.
+The unfolded network reduces average transmit power and execution time by 98% compared to proximal Jacobian ADMM.
+The ADMM algorithm struggles with nonlinear discrete optimization due to biased alternating updates.
+To address this, iterative discrete estimation (IDE) is introduced in [163], ensuring unbiased updates and better convergence.
 Wang et al.
-[191] proposed a universal algorithm for downlink massive MU-MIMO systems with finite-alphabet precoding to minimize inter-user interference.
-The algorithm is an improved version of the original ADMM framework.
-The authors observed that the standard ADMM struggles with nonlinear discrete optimization due to biased alternating updates.
-To address this, they introduced iterative discrete estimation [191], ensuring unbiased updates and better convergence.
+in [163] proposed an algorithm for downlink massive MU-MIMO systems with finite-alphabet precoding to minimize inter-user interference.
 To reduce hardware costs and power consumption in multi-user massive MIMO, He et al.
-[77] developed an unfolded network using the IDE-based algorithm with finite-alphabet precoding.
-Additionally, constructive interference-based symbol-level precoding for multi-user MISO downlink transmission is discussed in [78].
-In [79], the conjugate gradient algorithm is unfolded for constant envelope precoding to reduce computational overhead.
-The Riemannian manifold optimization (MO)-based conjugate gradient descent is implemented across network layers.
-Simulations show that this approach effectively suppresses multi-user interference, improving overall performance.
-In [192], the authors use deep unfolding to optimize user scheduling, BS clustering, and beamforming in a dense C-RAN system.
-Reducing the number of BSs per user helps ease backhaul congestion.
-The goal is to maximize the weighted sum rate while minimizing BS clustering, all within power constraints.
-The PGD algorithm is unfolded with step sizes learned in an unsupervised manner.
-This approach avoids matrix inversion and bisection search, reducing iterations and computational complexity.
-In [80], the interior point method is unfolded in an unsupervised manner for precoding design.
-It leverages interference in multi-user MISO systems to minimize power while meeting SINR constraints.
-In [81], the authors address the challenges of mutual coupling in beamforming.
-They propose mitigating its effects through covariance matrix reconstruction by unfolding an augmented coprime array.
-VI-B mmWave
-mmWave wireless communication, operating in the 30–300 GHz band, offers abundant spectral resources to alleviate the bandwidth crunch at sub-6 GHz frequencies [157].
-However, mmWave signals suffer from severe propagation and penetration losses.
-To compensate for these losses, large-scale antenna arrays with tens to hundreds of elements are required to achieve high power gains.
-The short wavelength (1–10 mm) of mmWave signals allows compact integration of these arrays into small transceivers.
-These features introduce unique challenges in designing physical-layer transmission algorithms [193].
+[164] developed an unfolded network using the IDE-based algorithm in [163].
+The unfolded network with learnable step-size and damping factor achieves lower BER across a range of SNR.
+In [165] a learning-based channel semantic acquisition and beamforming solution is proposed for cell-free massive MIMO.
+The beamformer is constructed by unfolding successive over-relaxation-based linear beamforming algorithm which reduces estimation errors in HBF.
+The proposed approach reduces computational complexity and enhances robustness to imperfect CSI.
+TABLE V: Summary of typical deep unfolding models for precoder design and their advantages and disadvantages.
+Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
+Unfolded WMMSE [166] Pioneering deep unfolding model based on the WMMSE algorithm for downlink precoding in MU-MIMO systems.
+Achieves $10\%$ improvement in sum rate metric compared with the conventional WMMSE at SNR $=45$ dB.
+Advantages: The network has a few trainable parameters, leading to fast convergence.
+Disadvantages: Performance loss with ill-conditioned channel matrices because the matrix inversion in WMMSE is replaced by a trainable approximation.
+IAIDNN [146] Unfolded WMMSE model for MU-MIMO precoding.
+Converges after $2.94$ seconds compared to $291.01$ seconds required in classical WMMSE and reduces WMMSE computational cost by up to $30\%$.
+Advantages: IAIDNN generalizes to varying system sizes (numbers of users and antennas) with minor adjustments.
+Disadvantages: High training computational complexity due to the increased number of trainable parameters.
+PGA-Net [155] Unfolds PGA with trainable step sizes for hybrid precoding.
+Proposed network achieves better sum-rate than classical PGA with fewer iterations.
+Advantages: Fast convergence and generalize well with varying channel size, effective for wideband channels.
+Disadvantages: Lacks inter-user interference cancellation and involves high gradient computation complexity.
+ManNet [181] Deep unfolding models for HBF (with both fully- and sub-connected architectures) based on Riemannian MO method.
+Over $600$ times faster and $6$ times complexity reduction compared to MO-AltMin algorithm [182] ,has lower complexity than PGA-Net [155].
+Advantages: Sparsely connected network architecture, and thus, has small number of training parameters.
+Disadvantages: The dynamic RF chain-antenna connections has not been incorporated into the unfolding models.
+V-B mmWave MIMO
 mmWave MIMO systems rely on beamforming to mitigate multi-user interference while enhancing diversity and spectrum efficiency.
@@
 [166] proposed a joint design approach combining DL and model-driven techniques.
-DRL optimizes beam selection, while a deep unfolding neural network, based on iterative WMMSE, optimizes the digital precoding matrix.
+DRL optimizes beam selection, while a deep unfolding NN, based on iterative WMMSE, optimizes the digital precoding matrix.
 The goal is to maximize the system sum rate while satisfying beam selection constraints.
@@
 Backpropagation and SGD are used for training and parameter updates.
+The authors verified the proposed framework outperforms benchmarks in sum-rate, complexity, and robustness while maintaining interpretability and scalability.
 Lin et al.
-[87] studied a linear MMSE transceiver design for uplink massive MU-MIMO with one-bit ADCs.
+[167] studied a LMMSE transceiver design for uplink massive MU-MIMO with one-bit ADCs.
 To minimize the system MSE by optimizing the precoder at each user equipment, the PGD algorithm is unfolded.
-The authors focused on training only the step size in each PGD iteration, significantly reducing the training overhead.
-In [88], the authors combined deep generative models with unfolding techniques to achieve near-optimal hybrid beamforming while minimizing feedback and computational complexity.
-They represented the channel as a low-dimensional manifold using a GAN, enabling efficient optimization of digital and analog precoders.
-The hybrid beamforming problem is formulated as a bilevel optimization, where an iterative rule updates both precoders.
-The upper-level optimization refines the digital precoder, leveraging full knowledge of the lower-level task constraints.
-Simultaneously, the lower-level optimization adjusts the analog precoder based on the digital precoder’s decisions.
-This structured optimization approach ensures efficient beamforming design with reduced computational costs.
-VI-C Next Generation Multiple Access
-Rate-Splitting Multiple Access (RSMA) is a flexible multiple access technology that merges the strengths of space division multiple access (SDMA) and NOMA.
+The authors focused on training only the step-size in each PGD iteration, significantly reducing the training overhead.
+In [168], the authors combined deep generative models with unfolding techniques to achieve near-optimal HBF while minimizing feedback and computational complexity.
+The receiver estimates the channel using GAN, enabling efficient optimization of digital and analog precoders.
+The HBF problem is formulated as a bilevel optimization.
+The algorithm unfolds an iterative bi-level optimization problem into the layers of the CNN.
+There are two levels are upper and lower layer.
+In the upper layer digital precoder is optimized given a fixed analog precoder while in the lower layer the analog precoder is optimized given a fixed digital precoder.
+This alternating iterative steps are then unfolded.
+The trainable parameter set include the weights of the CNN that maps digital precoder to the analog precoder.
+The unfolded bilevel optimization algorithm achieved near optimal HBF with low feedback and computational cost.
+V-C Next Generation Multiple Access
+Rate-splitting multiple access (RSMA) is a flexible multiple access technology that merges the strengths of space division multiple access (SDMA) and NOMA [169].
 It uses linearly precoded rate-splitting at the transmitter and SIC at the receivers.
-RSMA balances interference management by partially decoding some interference while treating the rest as noise, bridging the gap between SDMA (which treats all interference as noise) and NOMA (which fully decodes interference) [194].
+RSMA balances interference management by partially decoding some interference while treating the rest as noise, bridging the gap between SDMA (which treats all interference as noise) and NOMA (which fully decodes interference).
 In [170], the authors applied fractional programming with hyperplane fixed-point iteration to design the RSMA beamformer.
-The algorithm is unfolded into a neural network, replacing its inner loop with a small-scale DNN to predict Lagrangian dual variables.
+The algorithm is unfolded into an NN, replacing its inner loop with a small-scale DNN to predict Lagrangian dual variables.
 Simulations show that the proposed approach achieves high performance while reducing computational complexity.
-VI-D IRS
-An IRS consists of a programmable array of passive elements that dynamically adjust the phase shifts of incoming signals.
-This enables signals to align in phase at the receiver, enhancing reception.
-IRS can mitigate interference, create alternative propagation paths to bypass obstacles, and extend signal coverage to previously unreachable areas.
-The phase shifts are optimized to reflect signals coherently toward the intended user while minimizing interference in other directions.
-Therefore, jointly optimizing IRS phase control and transmit beamforming enhances the performance of IRS-assisted systems [195].
+V-D IRS
 With the aid of IRS, Chen et al.
-[113] addressed the joint design of the hybrid beamforming matrix and the phase-shifting matrix of the IRS elements.
-They proposed the deep unfolded WMMSE MO algorithm (DU-WMMSE-MO) using a deep unfolding methodology.
-The algorithm showed improved performance in IRS phase-shift design, utilizing two loops to jointly update the step size of the beamforming and IRS matrices.
-Compared to the Armijo backtracking line search method used in conventional WMMSE-MO [196], DU-WMMSE-MO provides more flexibility but with a higher runtime.
-Training is performed in an unsupervised manner, with a loss function that aligns with the objective function.
-The trainable parameters include the weights for the precoder, phase configuration of the IRS elements and combiner, and biases for the precoder and phase-shift matrix of the IRS.
+[171] addressed the joint design of the HBF matrix and the phase-shifting matrix of the IRS elements.
+The authors proposed the deep unfolded WMMSE-MO network (DU-WMMSE-MO), for IRS-assisted mmWave MIMO-OFDM systems.
+The trainable parameters include the weights for the precoder, phase configuration of the IRS elements and the combiner, and biases for the precoder and phase-shift matrix of the IRS.
+The algorithm showed improved performance in IRS phase-shift design, utilizing two loops to jointly update the step-size of the beamforming and IRS matrices.
+Compared to the Armijo backtracking line search method used in classical WMMSE-MO [172], DU-WMMSE-MO [171] provides more flexibility but with a higher runtime.
 An IRS-assisted multiuser MIMO full-duplex system is explored in [173].
-The active beamforming matrices at the BS and uplink users, along with the passive beamforming matrix at the IRS, are jointly optimized to maximize the weighted sum-rate of the system.
-To address the high computational complexity of matrix inversion in the conventional stochastic successive convex approximation (SSCA)-based optimization algorithm, a deep-unfolding neural network is proposed.
-Fig.
-7 shows two such netwroks from [114] .
-The network consists of a long-term passive beamforming network (LPBN) and a short-term active beamforming network (SABN).
-The learnable parameters of the LPBN and SABN are the IRS phase shift or passive beamforming vector $\boldsymbol{\theta}$ and the active beamforming vector $\boldsymbol{\Psi}$, respectively.
-Figure 7: Two networks LPBN and SABN [114].
-In [115], Ming-Min et al.
+The active beamforming matrices at the BS and uplink users, along with the passive beamforming matrix at the IRS, are jointly optimized to maximize the weighted sum rate of the system.
+A deep unfolding approach is proposed to approximate the stochastic SSCA-based beamforming algorithm, reducing the complexity of matrix inversion while maintaining near-optimal performance.
+In [174], Min et al.
 introduced a two-timescale transmission approach to reduce the signal processing complexity and channel training overhead for the IRS-aided multiuser MISO system.
-To minimize the average transmit power at the AP while ensuring the individual quality of service constraints for the users, a primal-dual decomposition-based two-timescale hybrid beamforming method is proposed and unfolded.
+To minimize the average transmit power at the analog precoder while ensuring the individual quality of service (QoS) constraints for the users, a primal-dual decomposition-based two-timescale HBF method is proposed and unfolded.
 This approach allows for efficient beamforming optimization by separating the optimization into two timescales, improving both computational efficiency and performance.
-In [197], a low-complexity IRS-assisted downlink MISO system is designed using an unfolded block coordinate descent algorithm to address the joint optimization problem of transmit beamforming and IRS phase shifts.
-The multi-ratio fractional programming structure of the objective function is handled by decoupling the problem into two subproblems: one for optimizing the beamforming at the base station and another for optimizing the phase shifts of the IRS.
+In [175], a low-complexity IRS-assisted downlink MISO system is designed using an unfolded block CD algorithm to address the joint optimization problem of transmit beamforming and IRS phase shifts.
+The multi-ratio fractional programming structure of the objective function is handled by decoupling the problem into two subproblems: one for optimizing the beamforming at the BS and another for optimizing the phase shifts of the IRS.
 This is achieved through the use of Lagrange dual transformation, effectively simplifying the optimization process.
-The numerical analysis validated the achieved weighted sum rate is comparable to [198] but with reduced computational complexity.
+The numerical analysis validated that the achieved weighted sum rate is comparable to that of [176] but with reduced computational complexity.
 Furthermore, the authors in [177], have developed an unfolded network to address the joint beamforming task for an IRS-assisted MISO system.
-The WMMSE and power iteration algorithm are used as an alternating optimization strategy to derive the local optimal solution for the active and passive beamforming.
-The proposed network reduce the computational runtime to approximately 3% of the block coordinate descent algorithm in [198].
+The WMMSE and power iteration algorithm are used as an alternating optimization (AO) strategy to derive the local optimal solution for the active and passive beamforming.
+The proposed network reduce the computational runtime to approximately 3% of the block CD algorithm in [176].
 In [178], an IRS-assisted MISO covert symbiotic radio communication system is explored, and a deep unfolding algorithm based on gradient descent is proposed for beamforming design.
-The goal is to optimize the achievable transmission rate of the primary transmitter by refining both active and passive beamforming strategies while satisfying two key constraints: (1) ensuring covertness and (2) maintaining the SNR of the secondary transmission.
-The complex optimization problem, along with its constraints, is transformed into the Lagrangian dual domain.
-Each layer in the deep unfolded network consists of four modules that update: (1) the active beamforming, (2) the passive beamforming, and (3) two Lagrange multipliers corresponding to the two constraints.
-Gangyong et al.
-in [200] have designed the IRS phase shift matrix by maximizing the sum path gain of the channel.
-This maximization is achieved by unfolding the Riemannian gradient descent algorithm.
-The network takes channel matrices and the initial reflection coefficients as inputs and outputs the optimized IRS coefficients.
-By configuring the IRS element phase shifts, the system effectively directs signal propagation paths to maximize the sum-path-gain, which is a key metric for evaluating beamforming quality and learning optimal IRS phase shifts to enhance the overall beamforming pattern.
-VI-E THz
+The goal is to jointly optimize active and passive beamforming under covert communication constraints.
+The step-size for updating the active and passive beamformers, and Lagrange multipliers form the trainable set.
+In contrast, Gangyong et al.
+in [179] focused purely on passive beamforming and unfolded the Riemannian gradient descent algorithm.
+They have designed the IRS phase shift matrix by maximizing the sum path gain of the channel.
+In this network, the only trainable parameter is the step-size associated with each Riemannian gradient descent iteration.
+The network in [179] is simpler compared to that of [178], due to less number of parameters to train but is specific to optimizing the phase shift of the IRS.
+V-E Terahertz
 6G wireless networks are expected to achieve terabit-per-second (Tbps) data rates to support ultra-reliable low-latency communication.
-To enhance spectral and energy efficiency, THz-enabled massive MIMO has emerged as a promising solution, leveraging the abundant bandwidth at THz frequencies to significantly boost data rates.
-Additionally, hybrid beamforming offers a cost-effective approach by reducing the number of RF chains while maintaining high performance, making large-scale deployment of high-frequency systems more practical [201].
+To enhance spectral and energy efficiency, terahertz (THz)-enabled massive MIMO is a promising solution, leveraging the abundant bandwidth at THz frequencies to significantly boost data rates [180].
 Nguyen et al.
-[118] proposed a deep unfolding framework by unrolling the iterations of Riemannian manifold optimization (Riemannian MO), specifically the MO-AltMin algorithm [202].
-Instead of directly minimizing WMMSE and maximizing SE, the problem is transformed into an approximate matrix factorization problem, which is then solved using LS.
-This approach is computationally more efficient, as it requires fewer layers to compute the weights of the precoding matrix, significantly reducing the training time.
-Compared to the approach in [75], the proposed network achieves better performance while maintaining much lower time and computational complexity.
+[181] proposed a deep unfolding framework that unfolds the iterations of Riemannian MO (specifically the MO-AltMin algorithm) [182].
+Instead of directly solving the original SE maximization or WMMSE minimization problems, which are highly nonconvex and computationally intensive, they reformulated the HBF design as an approximate matrix factorization problem.
+This factorization leads to LS formulation for analog and digital beamforming enabling a more efficient solution.
+By leveraging this structure, their proposed ManNet-based approach [181] requires significantly fewer layers to estimate the analog precoder, resulting in faster training and lower computational complexity as compared to the approach in [155].
 Minghui et al.
-[119] proposed a DL-based RSMA scheme for IRS-aided THz multi-user MIMO systems, focusing on enhancing robustness against CSI imperfections to improve system performance.
-The approximate WMMSE algorithm for RSMA is unfolded in this approach to boost precoding performance while reducing computational complexity.
-The deep unfolding algorithm processes the estimated equivalent channel through a series of layers to optimize the digital active precoding.
-This strategy aims to mitigate performance loss caused by model mismatches in real-world systems.
-By deep unfolding the approximate WMMSE algorithm, the method achieves improved precoding performance and lower computational complexity compared to the original WMMSE approach.
-This iterative deep unfolding process refines the precoding strategy based on the estimated equivalent channel, optimizing performance in IRS-aided THz massive MIMO-OFDM systems.
-In [120], Adam et al.
-introduced a method to jointly optimize active beamforming, passive beamforming, and UAV trajectory to minimize power consumption while accounting for UAV jitters and imperfect hardware constraints.
-They designed a hybrid semi-unfolding DNN to address these constraints during the real-time CSI gain estimation for IRS and UAV links.
-The approach is termed “semi-unfolded” because it combines two types of networks: one based on deep unfolding and the other based on a multi-kernel convolutional LSTM (MK-CLSTM) NN.
-For active beamforming, the authors employed semidefinite programming to obtain a closed-form solution, which was then transformed into NN layers to replicate the iterative beamforming solution in real time.
-The Lagrangian functions with well-defined Karush-Kuhn-Tucker (KKT) conditions were derived to design the layers of the deep unfolded architecture.
-The UAV trajectory and beamforming vectors are the parameters updated within these layers.
-For passive beamforming, a MK-CLSTM-based NN was used.
-Training was done using mini-batch SGD.
-The authors claim that their model-driven approach achieved $99.24\%$ accuracy and is $26.21\%$ faster than the classical iterative methods, offering significant improvements in both efficiency and speed for UAV-based communication systems.
-VII Sensing and Communication
-The independent operation of sensing and communication as separate systems has been widely studied [203].
-However, with the deployment of wireless communication services in higher frequency bands, such as mmWave and THz, spectrum congestion has become a major issue.
-In the 6G era, independent sensing systems experience interference from communication signals, leading researchers to recognize the shared system architecture and signal processing techniques between sensing and communication [204], [205], [206].
-This realization has driven the development of ISAC, where both functionalities are merged into a unified system.
-ISAC can be categorized based on sensing methods into active sensing and passive sensing [207].
-  • Active sensing (Mono-static sensing): The transmitter and receiver are co-located, allowing the system to extract target parameters using both the transmitted signal and received echoes.
-However, this approach suffers from significant self-interference at the ISAC receiver due to the strong reflections from the transmitted signals.
-  • Passive sensing (Bi-static sensing): The sensing echoes are received at a different location than the transmitter, eliminating the need for in-band full-duplex operation at the receiver.
-This approach is more suitable for applications where reducing self-interference is a priority.
-ISAC provides a promising solution for optimizing spectral efficiency and enhancing system performance in future wireless networks.
-However, understanding the fundamental performance limits of ISAC systems is crucial for their practical implementation.
-The estimation accuracy of key sensing parameters, such as direction of arrival, Doppler frequency, velocity, and target position, was thoroughly investigated in [208].
-This study provided a detailed analysis of both theoretical and practical limits of ISAC systems.
-The authors emphasized that the performance limits of ISAC channels cannot be derived by simply merging the existing bounds from independent sensing and communication systems.
-Instead, a unified framework is necessary to characterize these limits effectively.
-To address this, the authors proposed deriving unified Cramér-Rao bounds for ISAC, which establish the fundamental lower bounds on sensing accuracy given a specific allocation of orthogonal resources between communication and sensing.
-These unified expressions provide insights into how ISAC systems can optimize their resource allocation strategies to achieve superior performance in joint sensing and communication tasks.
-VII-A MIMO ISAC
-The integration of deep unfolding in ISAC systems has shown promising advancements in optimizing waveform design, beamforming, and communication-sensing tradeoffs.
-Ziwei et al.
-[82] applied deep unfolding techniques to optimize waveforms in MIMO radar systems, specifically focusing on minimizing the range-integrated sidelobe level.
-This is crucial for reducing interference and improving target detection performance.
-Their deep unfolding network efficiently optimizes radar waveforms, enhancing overall radar performance.
-In [83], the challenge of designing constant-modulus waveforms for multiuser MIMO ISAC systems was explored.
-Since constant-modulus waveforms help avoid signal distortion in nonlinear power amplifiers, their design is crucial.
+[183] proposed an unfolded network to enhance RSMA precoding in IRS-aided THz MU-MIMO systems under CSI imperfections.
+They unfold an approximate WMMSE algorithm to optimize the digital active precoding layers based on the estimated equivalent channels.
+The trainable set includes the transformer weights associated with the private and common streams and a regularization parameter.
+Compared to the original approximate WMMSE algorithm, the proposed unfolding method achieves superior precoding performance with substantially lower computational complexity.
+V-F Discussions and Takeaways
+Precoding design in massive MIMO and HBF systems involves solving non-convex optimization problems under hardware constraints and channel uncertainty.
+Popular unfolded networks for precoding design are listed in Table V.
+For instance, IAIDNN [146] unfolds the WMMSE algorithm and replaces complex operations like matrix inversion with a learnable block.
+This results in a significant reduction in computational complexity while achieving performance close to traditional WMMSE solutions.
+Deep unfolding enables the incorporation of domain-specific constraints such as the constant modulus constraint in analog beamformers through tailored activation functions.
+This approach significantly reduces computational complexity by avoiding matrix inversion and accelerates convergence, making it well-suited for real-time hybrid precoding under practical hardware limitations.
+Despite their benefits, current unfolded precoding models often assume fixed system dimensions and ideal CSI, which limits generalization and applications in practical systems.
+Future work should focus on scalable, adaptive unfolding frameworks that can operate reliably in time-varying, multi-user, and wideband settings.
+Integration with channel estimation modules and robust designs under CSI uncertainty will be key for practical deployment in next-generation wireless networks.
+Moreover, applying such unfolded frameworks for precoders in real systems and analyzing the performance in real experiments will be very beneficial.
+VI Sensing and Communication
+The independent operation of sensing and communication faces challenges in the 6G era due to spectrum congestion in high-frequency bands like mmWave and THz [184].
+This has driven the development of ISAC systems, which unify both functions using shared architectures and processing techniques [185], [186], [187].
+ISAC includes active (mono-static) and passive (bi-static) sensing methods [188].
+While ISAC enhances SE and system performance, understanding its fundamental limits is crucial.
+Key sensing parameter estimations have been thoroughly analyzed in [189], revealing theoretical and practical boundaries.
+TABLE VI: Summary of typical deep unfolding models for ISAC and their advantages and disadvantages.
+Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
+Unfolded PGA network [191] Unfolds PGA to maximize the system sum rate while minimizing the sensing beampattern error.
+Achieves $33.5\%$ higher sum rate, $2.5$ dB lower radar beam pattern error compared with SCA, and $65\%$ reduction in runtime and complexity compared with the conventional PGA.
+Advantages: Ensure convergence in the challenging ISAC HBF designs, perform well in both single use and MU systems.
+Disadvantages: The convergence speed is relatively slow, high complexity to compute the gradients.
+ISAC-Net [194] Unfolds the denoising AMP method for communication and DFT-based sensing in narrowband OFDM ISAC.
+Achieve communication performance close to OAMP-Net2 [11] and better sensing performance than classical 2D-DFT algorithm.
+Advantages: Leverages shared resources to overcome challenges in signal processing without prior knowledge of transmitted signals, enhanced accuracy and security.
+Disadvantages: Beam squint effect is not considered leading to discrepancies extracting sensing parameters (angle of arrival or angle of departure) across various subcarriers in OFDM.
+ADMM-Net for OTFS ISAC [196] Unfolds ADMM to estimate delay-Doppler targets and demodulate OTFS symbols under high mobility.
+With $2$ targets, ADMM-Net achieve $4$ dB gain in NMSE compared with the classical ADMM and ISTA at SNR=$14$ dB.
+Advantages: Suitable for vehicular/high-mobility scenarios and robust to delay-Doppler spread.
+Disadvantages: Real time CSI estimation is not addressed.
+Double loop unfolded network [197] Unfolds gradient descent to jointly optimize IRS phase shifts in ISAC system.
+Minimizes worst case CRB for sensing targets.
+Advantages: Has low computational complexity and fast convergence (within $15$ layers).
+Disadvantages: Offline training ($1000$ channel realization, $20$ epochs) may not be feasible for frequently changing environments.
+VI-A MIMO ISAC
+The integration of deep unfolding in ISAC systems has shown promising advances in optimizing waveform design, beamforming, and communication-sensing trade-offs.
+In [190], the challenge of designing constant-modulus waveforms for MU-MIMO ISAC systems was explored.
+Since constant-modulus waveforms help avoid signal distortion in non-linear power amplifiers, their design is crucial.
 However, the non-convex nature of the constant-modulus constraint makes optimization challenging.
-To address this, the authors employed a deep unfolded network with a sparsely connected structure, trained in an unsupervised manner, ensuring efficient waveform design under these constraints.
-Hybrid beamforming in ISAC systems poses additional complexity due to the need to balance communication rates and sensing accuracy.
+To address this, the authors employed a deep unfolded network based on PGD algorithm, ensuring efficient waveform design under these constraints.
+HBF in MIMO ISAC systems poses additional complexity due to balancing communication rates and sensing accuracy.
 The authors in [191] tackled this challenge by unfolding the iterative PGA algorithm.
 Their deep unfolding-based method accelerated convergence and enhanced overall system performance, achieving an optimal balance between communication and sensing objectives.
-Further, joint optimization of radar-communication beamforming was addressed in [209].
+In addition, joint optimization of radar communication beamforming in MIMO system was addressed in [192].
 The authors incorporated symbol-level precoding with constructive interference constraints to improve energy efficiency by exploiting multi-user interference.
-They developed iterative algorithms to leverage problem separability, enabling parallel and recursive computation.
-Deep unfolding was utilized to reduce computational complexity, improve scalability, and support real-time implementation.
-VII-B mmWave MIMO ISAC
-mmWave MIMO is a key technology in 6G wireless communication, offering high transmission capacity and sub-meter user localization.
-Its large-scale antenna arrays enable precise angular resolution.
-Traditional localization methods estimate position-related parameters before reconstructing the user’s location.
+Since, the objective problem is computationally complex due to non-convexity and higher dimensional matrix inversions, a custom iterative algorithm is developed.
+The iterative algorithm is based on block CD and primal-dual optimization (augmented Lagrangian method).
+Therefore, the trainable parameter set for training the deep unfolded network include step-sizes to update the primal, dual and auxiliary variable, as well as a correction matrix to approximate matrix inversion.
+The proposed network reduces computational complexity due to small number of trainable parameters, improve scalability, and support real-time implementation.
+The large-scale antenna arrays in mmWave MIMO systems enable sub-meter user localization by providing high angular resolution.
+However, these methods often incur significant computational complexity, particularly in high dimensional mmWave MIMO settings.
 To improve this, Fan et al.
-proposed a fast direct localization method in [89].
-They use an unfolded ADMM solver with low complexity for direct localization.
-The trainable parameters include $\rho$ (penalty parameter), $w$ (weight coefficient for channel gain between the base station and user), and $\tau_{1}$, $\tau_{2}$ (two positive proximal parameters).
-VII-C Modern Multicarrier Modulation for ISAC
-5G and beyond 5G wireless communication systems operate at carrier frequencies in the microwave regime, where radar applications are dominant.
-This has led to the development of joint waveforms for sensing and communication to address limited spectral resources [210].
-OFDM-based radars in LTE and 5G NR mobile networks’ BS are studied for radar and sensing in [211], [212], [213].
-OTFS modulation for joint radar parameter estimation and communication is explored in [214], [215], [216].
-An ISAC signal processing problem was proposed in [102] to process pilot and data signals together.
-The optimization problem requires the simultaneous recovery of CSI, transmitted data symbols, and target information, making it non-convex.
-To solve this, the authors used an iterative optimization algorithm that alternates between passive sensing and communication detection.
-Jingcai et al.
-proposed a deep unfolded network for target parameter estimation in MIMO-OFDM-based ISAC systems [103].
-Their method leverages the sparsity of target parameters, using a compressed sensing approach.
-The problem is reformulated as a least absolute shrinkage and selection operator (LASSO) problem.
-The fast ISTA algorithm is then unfolded, with learnable parameters enabling high-precision solutions with low computational complexity.
+proposed, a fast and accurate direct localization method in [193].
+An ADMM solver is developed to solve the optimization problem with low complexity.
+The ADMM iterations are unfolded to learn the trainable parameters that include penalty and proximal parameters, and weight coefficient for channel gain between the BS and user.
+The unfolded network converges faster than classical ADMM and computational time is reduced.
+VI-B Modern Multicarrier Modulation for ISAC
+A joint passive sensing and communication demodulation for OFDM is addressed in [194].
+The ISAC-Net is designed by unfolding two algorithms: a denoising-based AMP algorithm for the communication module and the 2D discrete Fourier transform (DFT) for the sensing module.
+The step-size to update the channel and DFT matrix form the trainable set to balance the communication demodulation and sensing accuracy.
+Furthermore, Jingcai et al.
+proposed an unfolded network (FISTA-Net) for joint target parameter estimation and communication in MIMO-OFDM systems [195].
+The fast ISTA is unfolded, with learnable parameters such as step-size for weight matrix, enabling high-precision solutions with low computational complexity.
+The proposed network converges with $8$ layers as compared to $2000$ iteration in fast ISTA.
+Also, the network achieves an NMSE of $-20.12$ dB while the classical complex ISTA achieved $-16.8$ dB.
 The problem of estimating parameters of moving targets in high-speed scenarios using OTFS modulation for ISAC systems is explored in [196].
-The target parameter estimation is reformulated as a compressed sensing problem.
-For the deep unfolded network, ADMM iterations are unfolded into the network’s layers.
-VII-D IRS-Assisted ISAC
+The ADMM iterations are unfolded (ADMM-Net) to accelerate convergence and enhance sparse recovery of target position.
+Regarding parameter estimation, ADMM-Net [196] and FISTA-Net [195] reduce NMSE significantly.
+Also, ISAC-Net [194] and FISTA-Net [195] show faster convergence than iterative baselines.
+VI-C IRS-Assisted ISAC
 Liu et al.
-explored beamforming design for an IRS-assisted ISAC system with bounded angle uncertainty of sensing targets and statistical CSI uncertainty related to user equipment in [117].
-They formulated a bi-level optimization problem, unfolding the projection method and Riemannian MO approach.
-These methods handle the maximum transmit power and unit-modulus constraints.
-The bi-level problem consists of two parts: the lower-level problems focus on identifying the worst-case CSI and target sensing angles, while the upper-level problem addresses beamforming design and optimization of Lagrange multipliers based on the given CSI and target angles.
-A double-loop deep unfolding NN is applied, unfolding the iterative gradient descent into a multi-layer structure, with each layer containing an upper-level and a lower-level loop.
-Trainable step sizes and offset parameters are introduced as network parameters.
-VIII Decoding in Error Correcting Codes
-Error control codes are essential for reliable and efficient communication.
-During data transmission, bits may flip due to channel noise, causing errors.
-Error correction coding helps detect and correct these errors at the receiver.
-Claude Shannon revolutionized channel coding with his landmark paper [217].
-Mobile communication has evolved through several generations with significant technological advancements.
-In the first generation (1G), communication was primarily analog, leaving no need for channel coding.
-With the second generation (2G), convolutional and block codes were introduced to compress voice signals and convert them into digital form.
-Turbo codes were adopted as standard coding schemes in the third (3G) and fourth (4G) generations, offering near-Shannon limit performance with moderate decoding complexity [218].
-Low-density parity-check (LDPC) codes, first introduced in the 1960s, were later rediscovered and extensively studied [219], [220], [221].
-They are widely used in 4G, 5G, and beyond.
-LDPC codes are represented using a bipartite graph, placing them in the “codes on graphs” category, designed for capacity-approaching codes with iterative decoding algorithms.
-Since LDPC codes use sparse graphs, the number of edges grows linearly with block length, making them low-density and less complex [222].
-Polar codes, introduced by Erdal Arıkan in 2009, are a modern class of error-correcting codes designed to achieve the capacity of symmetric binary-input discrete memoryless channels [223].
-They are based on channel polarization and are known for their capacity-achieving property, meaning they can approach the Shannon limit as block length increases.
-BP is widely used to decode modern error correction codes [224].
-This decoder is constructed from the Tanner graph, introduced by Michael Tanner in his landmark paper [225].
-He developed the min-sum and max-sum algorithms, which are derived from the generic sum-product decoding [226].
-Polar codes use successive cancellation decoding, where received bits are decoded sequentially by leveraging frozen bits and previously decoded bits.
-More advanced techniques, such as successive list decoding assisted by cyclic redundancy check (CRC), further improve performance [227].
-In practical wireless communication, channels often deviate from mathematical models or are not precisely known at the decoder.
-In such cases, a DL-based decoder offers a robust and adaptive solution.
-Additionally, model-driven DL is well-suited for providing an interpretable analysis of iterative decoding algorithms.
+addressed a robust beamforming design for an IRS-assisted ISAC system under imperfect CSI and target direction estimates in [197].
+The objective is to minimize the sum of worst-case Cramér-Rao bounds (CRB) for all sensing targets.
+They formulated a bi-level optimization structure that consists of two levels: the lower-level problem finds the worst-case CSI and worst-case target sensing angles.
+The upper-level problem optimizes the active and passive beamforming matrices.
+A double-loop deep unfolding NN is designed to tackle the bi-level optimization problem.
+The CSI and angle uncertainties are updated via PGD in the lower-level loop.
+The beamforming variables are updated in the upper-level loop using PGD and Riemannian MO optimization.
+The proposed network effectively achieves near-optimal performance by handling statistical CSI errors and bounded target direction errors.
+VI-D Discussions and Takeaways
+Deep unfolding offers a structured and interpretable framework for tackling challenging optimization problems in ISAC systems.
+By unrolling classical algorithms like PGD and WMMSE into NN layers, unfolding enables the joint design of beamformers and waveforms that balance sensing accuracy and communication performance.
+These approaches reduce computational complexity and provide faster convergence compared to traditional iterative solutions.
+We summarize typical deep unfolding models for ISAC in Table VI.
+It is observed that deep unfolding exhibits advantages in constant modulus waveform design, optimization of dual-functional beamformers, and power allocation.
+Indeed, the deep unfolding technique can effectively handle non-convex constraints or objective functions like SINR, sum rate, and CRB, which are infeasible with black-box models.
+In [198], the authors have provided techniques such as gradient calculation for estimating position that can be employed to facilitate practical deployment.
+However, the application of deep unfolding in ISAC is still underexplored.
+Indeed, most existing unfolding solutions are developed to optimize the transmit waveform in monostatic ISAC scenarios.
+In sensing functions, the processing of the received echo signal plays an important role in translating the observation into a decision on target detection and localization.
+Well-designed deep unfolding models can perform these tasks with low complexity and high accuracy, which is important for real-time sensing applications.
+There is limited literature addressing the practical deployment of unfolded networks in ISAC systems.
+VII Decoding in Error Correcting Codes
+Mobile communication has evolved across generations.
+1G used analog transmission (without any coding), and 2G introduced digital voice coding with convolutional and block codes.
+Turbo codes, adopted in 3G and 4G, offer near-Shannon limit performance with moderate complexity [199].
+Low-density parity check (LDPC) codes, rediscovered in the 1960s, are now widely used in 4G and 5G [200], [201], [202].
+Polar codes, introduced in 2009, achieve the capacity of symmetric binary-input discrete memoryless channels [203], approaching the Shannon limit as the block length increases.
+They have been adopted in 5G new radio for both uplink and downlink control channels [204].
+BP operates on a bipartite graph known as Tanner graph, introduced by Michael Tanner [205] The min-sum and max-sum algorithms are derived from the generic sum-product decoding [206].
+Polar codes use successive cancellation decoding, where the received bits are decoded sequentially by leveraging frozen bits and previously decoded bits.
+More advanced techniques, such as successive list decoding assisted by cyclic redundancy check (CRC), further improve the performance of polar codes [207].
 Wei et al.
-[228] proposed an unfolded ADMM-based penalized algorithm for binary linear codes.
-Their approach introduces two networks that transform the ADMM penalty parameters $\{\mu,\alpha\}$ into iteration-dependent parameters, along with an adjustable penalty function.
-Each stage of the learnable ADMM decoding network (LADN) corresponds to one ADMM iteration, where iteration-dependent parameters enhance convergence and reduce the number of iterations.
-The key idea is that increasing the number of learnable parameters improves generalization capability.
-For the second network, a piecewise linear penalty function with adjustable slopes is designed to suppress pseudo codewords effectively, improving decoding performance beyond conventional $\ell_{1}$ and $\ell_{2}$ norm functions.
-Parameter updates are trained using backpropagation via the Adam optimizer.
-A customized loss function based on MSE ensures effectiveness at every iteration by considering outputs from all layers.
-The authors claim that their decoder outperforms the conventional ADMM decoder.
+in [39] proposed an unfolded ADMM-based penalized algorithm for binary linear codes to enhance the decoding performance.
+The learnable ADMM decoding network (LADN) trains the ADMM penalty parameters.
 Additionally, decoding irregular binary LDPC codes via unfolded ADMM is explored in [40].
-In [48], the authors propose the TurboNet architecture, which combines DL with the max-log MAP algorithm.
-While most existing model-driven decoding algorithms rely on BP and are iterative, the authors enhance the traditional MAP algorithm by unfolding it, with each iteration represented as a DNN decoding unit.
-As a result, the soft-input soft-output decoder, which traditionally uses the max-log MAP algorithm and extrinsic LLRs, is replaced by two subnets based on a neural log-MAP algorithm for each iteration.
-In [230], the authors tackle the decoding challenges of marker codes in insertion and deletion channels, with a focus on overcoming CSI uncertainty.
-To improve detection performance, they propose DL techniques, implementing both RNNs and a deep unfolded network.
-The forward-backward algorithm is unfolded into multiple layers within the network to enhance the overall decoding process.
+The LADN [39] outperforms the original ADMM-penalized decoder in block error rate for LDPC codes, while the network in [40] outperforms the original ADMM-penalized decoder in block error rate for irregular LDPC codes.
+In [36], the authors proposed the TurboNet architecture, to improve turbo decoding by unfolding the max-log MAP algorithm.
+Trainable parameters were inserted into the max-log-MAP decoding flow like weights for prior LLRs.
+Numerical analysis demonstarted that TurboNet outperforms classical turbo decoding based on the max-log-MAP algorithm, particularly at moderate-to-high SNRs.
+Furthermore, TurboNet has been validated through over the air experiments confirming its robustness to SNR mismatches.
+Moreover, the authors have extended the work by including joint detection and decoding in [208].
+A CGD-based OAMP detector (CG-OAMPNet) for the MIMO-OFDM system is proposed by unfolding EP.
+The robustness of CG-OAMPNet was further validated through over-the-air experiments, demonstrating practical resilience to environmental variations.
+In correlated channel scenario, CG-OAMPNet outperforms OAMP-Net2 [11] by performance gain of $\sim 2$ dB at BER$=10^{-3}$.
+The authors in [209] tackle the decoding challenges of marker codes in insertion and deletion channels without perfect CSI.
+They proposed a network called FBNet to improve detection performance; the forward-backward algorithm is unfolded for marker codes.
+The trainable weights are the CSI probabilities for insertion/deletion channels.
+FBNet performs quite close to the classical forward-backward algorithm but is robust to CSI mismatch.
+FBNet [209] is based on RNN structure and target synchronization errors (insertions/deletions) without accurate CSI, while TurboNet [36] is a deep feedforward DNN structure and targets substitution errors in turbo codes under AWGN channel.
 Gao et al.
-[231] have unfolded the factor graph of the BP decoder using a one-bit quantizer for polar codes.
+[210] have unfolded the polar BP algorithm to improve polar code decoding performance under one-bit quantization.
+The network with trainable weights for the factor graph edges outperforms standard BP with two bit quantizer by $0.2$ dB.
 In [211], a double-loop iterative decoding algorithm for LDPC codes is unfolded within a penalty dual decomposition framework.
-Additionally, by unfolding the iterative decoding process between check nodes and variable nodes into a feed-forward propagation network, a deep unfolded normalized min-sum LDPC decoder is derived in [233].
-Tanner-graph-based transmission schemes for massive multiple access, incorporating NOMA and grant-based random access, are discussed in [234].
-To enhance decoding efficiency, a message-passing decoder is unfolded for both transmission schemes, allowing the decoder to improve as more information about the transmitted data is learned during the training epochs.
-In [235], theoretical results for the unfolded BP decoder are presented.
-The authors define the generalization gap of a decoder as the difference between the empirical and expected bit-error rates.
+The network alternately updates primal and dual variables in a double loop structure.
+The penalty coefficients, penalty parameters and over relaxation factor form the trainable parameter set.
+The network has lower computational complexity than ADMM based decoders.
+Additionally, by unfolding the iterative decoding process between check nodes and variable nodes for LDPC codes, a deep unfolded normalized min-sum LDPC decoder is implemented in [212].
+The unfolded network outperforms the classical normalized min-sum LDPC decoder in BER and is robust across code lengths, rates and different channel conditions.
+The networks in [211, 212] focus on LDPC codes while the network in [210] consider polar codes.
+Moreover, the networks in [211, 212] use multi-SNR training for robustness.
+In [213], theoretical results for the unfolded BP decoder for Tanner codes are presented.
+The authors define the generalization gap of a decoder as the difference between the empirical and expected BER.
 They provide a set of theoretical results that bound this gap and examine its dependence on decoder complexity, which is influenced by code parameters (such as block length, message length, and variable/check node degrees), the number of decoding iterations, and the size of the training dataset.
-IX Power allocation
-In a multi-antenna wireless communication system with multiple users and limited resources, ensuring reliable transmission of massive amounts of data presents significant challenges.
-One of the primary constraints is the power budget, which must be allocated efficiently among data streams to maintain a high quality of service.
-The water-filling algorithm is a widely recognized technique for optimal power allocation across subchannels, aiming to maximize the sum capacity or sum rate.
-Since power allocation directly impacts channel capacity, it is formulated as an optimization problem, where maximizing the capacity rate leads to improved system performance.
-The power allocation problem can thus be formulated as an optimization problem as shown in (6) below:
-$\max_{\left\{P_{i}\right\}}$ $\frac{\sum_{i=1}^{K}w_{i}R_{i}(P_{i})}{\sum_{i=1}^{K}P_{i}}$ (6)
-s.t. $R_{i}(P_{i})\geq R_{\min,i},\quad\forall i,$
-$\sum_{i=1}^{K}P_{i}\leq P_{\text{total}},$
-$P_{i}\geq 0,\quad\forall i.$
-where $K$ is the number of users, $w_{i}$ is the weight associated to user $i$ decided by user priority or fairness, $R_{i}(P_{i})$ is the rate achieved by user $i$ given the power $P_{i}$ allocated to them, $R_{\text{mini},i}$ is the minimum required rate for fairness and $P_{\text{total}}$ is the sum of all users’ power values.
+Traditionally, channel estimation, signal detection, and channel decoding in a MIMO receiver are performed sequentially and independently.
+This disjoint approach leads to suboptimal performance.
+To address this, Sun et al.
+[214] propose an unfolded ADMM-based receiver for joint detection and decoding.
+The detection and LDPC decoding components are jointly unfolded from ADMM iterations.
+Simulations show that the proposed receiver outperforms classical turbo receivers.
+In [215], Zhang et al.
+proposed a joint signal detection and channel decoding for MIMO systems.
+Signal detection is achieved by unfolding the EP algorithm.
+The channel decoding is based on an unfolded turbo code decoding network, TurboNet, as proposed in [36].
+The damping factors in EP and decoder parameters in turbo decoding are learned in the training stage.
+The networks showed fast adaptation and strong BER gains in over-the-air experiments.
+TABLE VII: Summary of typical deep unfolding models for decoding and their advantages and disadvantages.
+Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
+LADN [39] Introduces learnable ADMM-based decoder with trainable penalties for binary LDPC decoding.
+Achieves a block error rate reduction of 10$\sim$ 10$\times$ 10 × at SNR $=2$ dB over the baseline ADMM decoder.
+Advantages: Has small number of learning parameters and fast convergence.
+Disadvantages: Limited by code length.
+Inherent codeword structure not exploited leading to high computational complexity.
+TurboNet [36] Unfolds max-log MAP decoding steps for turbo codes with learnable weights.
+Outperforms the conventional max-log MAP scheme by $0.25-0.5$ dB BER.
+Advantages: Compatible with 3GPP turbo codes and validated through practical experiment.
+Disadvantages: Training requires fixed interleaver and careful SNR selection.
+Quantized Polar BP [210] Learns BP decoding under 1-bit quantization for polar codes over an AWGN channel.
+Outperforms standard BP with one-bit quantization by $1.2-2$ dB (block error rate).
+Advantages: Hardware-friendly and applicable to IoT systems with coarse ADCs.
+Disadvantages: Less scalable.
+Require large number of trainable weights assigned to each edge of factor graph for decoding.
+ADMM-Net [40] Extends LADN [39] to irregular LDPC codes.
+Unfolded ADMM, learn different penalty weights for each variable node in decoding.
+Achieve better BER performance than BP and ADMM penalized decoders for irregular LDPC codes.
+Advantages: Suitable for irregular code graphs.
+Disadvantages: The initialization strategy for the trainable parameters assumes prior knowledge of the code structure, which makes it less generalizable.
+FBNet [209] Introduce a RNN-based unfolding model based on forward-backward decoding for Marker codes over insertion-deletion channels.
+Outperforms the classical forward-backward algorithm under imperfect CSI.
+Advantages: Robust to synchronization errors with flexible sequence handling.
+Disadvantages: Degraded performance under burst insertion-deletion channels.
+VII-A Discussions and Takeaways
+In Table VII, we summarize the most popular unfolded network for the decoding of channel codes.
+Deep unfolding spans their applications in modern error-correcting codes, such as LDPC, turbo, and polar codes.
+The soft probabilistic estimates are progressively refined through learnable decoding layers in deep unfolding while enforcing parity constraints, enabling near-optimal hard decisions without the need for manual hyperparameter tuning.
+Hence, it merges model based optimization, for structured soft-output refinement and codeword validity with data-driven learning, for adaptive penalty tuning, outperforming traditional decoders in accuracy and efficiency.
+Despite these advantages, there are several limitations in existing deep unfolding architectures for decoding and channel coding.
+For example, the unfolded decoders in [39] is tied to specific code structures, requiring retraining when parameters such as code length or interleaving patterns change.
+Some methods demand significant training effort or hyperparameter tuning to maintain generalization.
+There are several works [215, 36, 208] that developed a prototyping system to experimentally validate the effectiveness and practical feasibility of unfolded networks in real-world channel conditions.
+The experimental results indicated that the unfolded network demonstrated strong robustness to environmental variations and operate effectively without the need for frequent retraining.
+Future research should aim to create more adaptable unfolding architectures that can handle a variety of code structures, channel models, and SNR conditions.
+Joint optimization with modulation and detection, as well as hardware-aware training, may further improve the viability of deep unfolding for end-to-end error correction in next-generation wireless communication systems.
+VIII Power Allocation
+Power allocation in wireless communication refers to the process of distributing available transmit power across different communication channels or users to optimize system performance.
+It aims to maximize throughput, minimize interference, and ensure fairness, while considering factors like channel conditions, user requirements, and energy efficiency for improved overall network efficiency.
 Several classical approaches, such as the Dinkelbach algorithm, Lagrange dual decomposition, interference pricing, successive convex approximation (SCA), WMMSE, and matching theory, have been employed for power allocation [144, 216, 217, 218, 219, 220].
@@
 The deep unfolding approach mitigates this issue by reducing the number of iterations required to achieve a near-optimal solution, thereby enhancing computational efficiency and energy savings.
-Wireless networks expand in terms of spatial coverage, but traditional DNN architectures, such as multi-layer perceptrons and CNNs, struggle to capture the inherent topology of these networks.
-To address this limitation, Chowdhury et al.
-[241] propose leveraging GNNs to parameterize the power allocation function.
-In their approach, the WMMSE algorithm is unfolded into layers, resulting in a distributed, efficient, and scalable solution.
-The unfolded WMMSE method integrates domain-specific elements with trainable components, which are parameterized by GNNs to exploit the natural graph structure of wireless communication networks.
-The objective is to map the channel matrix and approximate optimal power allocation by training the GNN using a block coordinate descent algorithm.
-In [242], Li et al.
-unroll the SCA algorithm into NN layers for energy-efficient power allocation in wireless interference networks.
-The proposed unfolded algorithm is highly adaptable, as it generalizes across varying network sizes and channel distributions while also enabling transfer learning between different path loss models.
-Given that the system operates in a multi-user wireless interference network, it can be naturally modeled as a graph neural network GNN.
-The power allocation policy is formulated based on the CSI matrix ${\bf{H}}$ and the total power constraint $P_{\text{total}}$, within a layered architecture incorporating trainable weights.
+Chowdhury et al.
+in [221] unfolded WMMSE into GNN based architecture to learn a power allocation policy from the channel matrix.
+The key feature of their architecture is that graph topology is leveraged with GNN for efficient rate utility maximization.
+The unfolded WMMSE method integrates domain-specific elements with trainable components, which are parameterized by generated adaptive weigh of the GNN.
 Hu et al.
@@
 The power allocation framework is built upon this attentive graph representation, as described in [223, 224].
-Meanwhile, the deep unfolded WMMSE maps the iterative steps of the WMMSE algorithm into neural network layers within an unsupervised learning paradigm.
-This approach effectively combines deep unfolding with graph representation learning, significantly reducing the search space for the objective function.
-As a result, the model achieves improved interpretability and accelerated training while eliminating the need for supervised data.
-X Physical Layer Security
-Physical layer security is a critical research area aimed at ensuring secure and reliable transmission between the transmitter and receiver while preventing eavesdropping and impersonation attacks.
-Due to the broadcast nature of wireless communication, adversarial users pose a significant threat.
-These adversaries generally fall into two categories: eavesdroppers (or middleman attackers), who attempt to extract information from an ongoing transmission without detection, and jammers, who deliberately degrade the signal quality to prevent the receiver from successfully decoding the intended message [246].
-Developing robust security mechanisms at the physical layer is essential to mitigate these threats and enhance communication reliability.
-The fundamental network model for analyzing security and confidentiality is a three-terminal system, as illustrated in Fig.
-8.
-In this setup, the legitimate transmitter (Alice), the intended receiver (Bob), and the unauthorized eavesdropper (Eve) are equipped with $N_{\text{t}}$, $N_{\text{r}}$ and $N_{\text{e}}$ antennas, respectively.
-As introduced by Wyner in his seminal work on wiretap channels [247], this model captures the imperfections introduced by the communication channel.
-Alice transmits an information signal, $\bf{x}$, a random vector intended for Bob, over the main channel, which is modeled as a discrete memoryless channel.
-However, due to the broadcast nature of wireless communication, the signal received by Bob, $\bf{y}$, is also partially intercepted by Eve through an additional wiretap channel, resulting in her received signal, $\bf{z}$, as shown in Fig.
-8.
-Figure 8: General wiretap channel [246].
-Traditionally, secret key-based encryption has been the standard approach to ensuring message confidentiality, where the encryption key is shared exclusively between the transmitter and the intended receiver.
-As defined in [248], a secrecy system can be abstractly described as a set of transformations that map a set of possible messages to a set of possible cryptograms.
-Each transformation corresponds to an encryption process that utilizes a unique, secure key to encipher the message, ensuring that unauthorized entities cannot decipher the transmitted information.
-The primary objective of a secure communication system is to maximize the secrecy rate, ensuring that the transmission rate to the intended user remains high while minimizing information leakage to the eavesdropper.
-The secrecy capacity $C_{s}$ is defined as the maximum rate $R_{s}$ at which secure communication can be achieved, guaranteeing that the eavesdropper receives negligible information about the transmitted message.
-This capacity represents the fundamental limit of secure data transmission, balancing efficiency and confidentiality in the presence of adversarial interception.
-The secrecy capacity can be expressed as:
-$C_{s}=\max_{P_{X}}\left\{I\left({\bf{X};\bf{Y}}\right)-I\left({\bf{X};\bf{Z}}\right)\right\}$
-where, $I(\bf{X};\bf{Y})$ is the mutual information between the transmitter and the intended user, and $I(\bf{X};\bf{Z})$ represents the information leakage to the eavesdropper.
-The transmitter can maximize the secrecy capacity by carefully choosing the probability distribution $P_{X}$ for the transmitted signal $\bf{x}$ to maximize the difference $I(\bf{X};\bf{Y})$-$I(\bf{X};\bf{Z})$ i.e, the rate $R_{s}$.
+The model [222] achieves improved interpretability and accelerated training over [221] while eliminating the need for supervised data.
+In [225], Li et al.
+unroll the SCA algorithm into NN layers for energy-efficient power allocation in wireless interference networks.
+The proposed unfolded algorithm is highly adaptable, as it generalizes across varying network sizes and channel distributions.
+The power allocation policy is formulated based on the CSI matrix and the total power constraint, within a layered architecture incorporating trainable weights.
+Each of the networks discussed in [221, 222] and [225] leverages the GNN structure in the deep unfolding approach and achieve near optimal performance of their benchmark iterative algorithm i.e, WMMSE and SCA respectively.
+IX Physical Layer Security
+Physical layer security aims to ensure secure and reliable communication while preventing eavesdropping and impersonation attacks.
+Wireless communication’s broadcast nature exposes it to adversaries like eavesdroppers, who intercept transmissions, and jammers, who degrade signal quality.
+The primary objective is to maximize the secrecy rate, ensuring high transmission rates for the intended user while minimizing information leakage to the eavesdropper.
+The secrecy capacity represents the maximum rate at which secure communication can be achieved, balancing efficiency and confidentiality.
+It is defined as the difference between mutual information between transmit and receive signals and information leakage to the eavesdropper [226].
 Jamming attacks pose a serious threat to the reliable operation of wireless communication systems [226, 227].
 A jamming attacker deliberately transmits interference or noise on the same frequencies as legitimate users.
-This disrupts communication channels, reduces signal quality, and can block data transmission entirely.
-The interference makes it harder for users to decode messages, leading to connection losses or forced retransmissions.
-As a result, network efficiency decreases, and latency increases.
 Marti et al.
-in [251] explored solutions to combat smart jamming in massive multi-user MIMO systems.
-In these systems, jamming attacks can selectively interfere at specific times to avoid detection.
-The authors proposed an optimization-based approach for joint jammer estimation, channel estimation, and data detection, using the BS’s multi-antenna configuration.
+in [228] explored solutions to combat smart jamming in massive multiuser MIMO systems.
+The authors proposed an optimization-based approach for joint jammer estimation, channel estimation, and data detection, using the BS multiantenna configuration.
 This method does not require prior knowledge of the jammer’s behavior, making communication more resilient even during sporadic or bursty jamming.
-Using a deep unfolding approach, the authors developed two iterative algorithms.
-The first is mitigation estimation and detection (MAED), which uses forward-backward splitting for estimation.
-The second is soft-output MAED (SO-MAED), an enhanced version that applies deep unfolding to optimize parameters.
+The authors have unfolded an iterative algorithm, called as mitigation estimation and detection (MAED), that uses forward-backward splitting optimization to jointly estimate the jammer subspace, user equipment’s channels, and user equipment’s data within a coherence block.
+The soft-output MAED (SO-MAED), is the deep unfolded network.
+The trainable set include error precision (to represent noise in symbol estimation), gradient descent step-sizes, scaling factor and momentum weights.
 SO-MAED improves detection accuracy, especially in high-interference environments, and does not require knowledge of the attack type.
-Error precision and gradient step sizes are key parameters for improving learning stability.
-XI Joint Techniques
-This section reviews different joint techniques that are aided by deep unfolding method.
-XI-A Joint Detection and Channel Estimation
-An iterative method for estimating joint channel parameters and selecting symbols using the EM algorithm is presented in [252].
-This method performs symbol detection in the presence of Gaussian noise and intersymbol interference in a dispersive channel.
-The authors model the missing data using a Markov chain, and channel parameter estimation follows the ML criterion.
-The Markovian properties of the channel state sequence allow for calculating the required likelihoods using the forward-backward algorithm.
-In [253], the authors explore a V-BLAST-type combination of OFDM with MIMO to improve SE and multi-user downlink throughput.
-They propose a joint data detection and channel estimation algorithm for MIMO-OFDM that integrates the QR decomposition M-algorithm and Kalman filter.
-The M-algorithm, combined with QR decomposition, reduces the computational complexity of tree-based search methods [254].
-As mentioned in [6], the data-driven approach uses a fully connected network, where each sub-block of the OFDM transceiver is designed.
-Gao et al.
-[255] proposed a deep unfolding approach to address the issue of DNN performance being dependent on large training datasets.
-In contrast to the approach in [6], where a DL architecture completely replaces the conventional modules, the ComNet architecture shown in Fig.
-9 incorporates both channel estimation and detection modules.
-Both modules use conventional communication solutions as inputs to the network.
-The DL network then refines these coarse inputs.
-For example, the LS channel estimate, $\bf{\hat{h}_{\text{LS}}}$, is refined using pilot symbols ${\bf{x}}_{p}$ and corresponding received symbols ${\bf{y}}_{p}$ as inputs to the LS sub-network within the channel estimation module.
-The channel estimate, $\bf{\hat{h}_{\text{LS}}}$, is passed from the LS sub-network to the LS Refine Net sub-network to generate an accurate channel estimate, $\bf{\hat{h}}$.
-A single-layer DNN is used to design the LS Refine Net.
-The refined channel estimate, $\bf{\hat{h}}$, from the LS Refine Net is then passed as input to the ZF detector sub-network in the signal detection module.
-The ZF detector sub-network is also fed with the received data vector, $\bf{y_{\text{D}}}$, and the LS channel estimate, $\bf{\hat{h}_{\text{LS}}}$.
-This is followed by the ZF Refine Net.
-The refined channel estimate, $\bf{\hat{h}}$, the received data vector, $\bf{y_{\text{D}}}$, and the ZF estimate, $\bf{\hat{x}_{\text{ZF}}}$, from the ZF detector sub-network are passed to the ZF Refine Net to estimate the transmit binary data, $\hat{b}$.
-The short dotted line from the LS sub-network to the ZF detector sub-network in Fig.
-9 provides a robust candidate for the recovered binary symbols, $\hat{b}^{{}^{\prime}}$, at deQAM.
-Both the channel estimation and detection modules are trained using an $\ell_{2}$ loss between the estimated binary data, $\hat{b}$, and the raw transmit binary symbol, $b$.
-Simulations have shown that this design outperforms both LMMSE and FC-DNN [6].
-Figure 9: ComNet architecture [255].
-An artificial intelligence-aided receiver for a cyclic prefix-free OFDM system with joint detection and channel estimation is explored in [256].
-Zhang et al.
-[257] unfolded the OAMP algorithm for generalized frequency division multiplexing (GFDM) systems with low-resolution analog-to-digital converters.
-In [258], a general EM algorithm is derived and unfolded to address the impact of channel estimation errors on signal detection.
-The expectation step estimates the CSI, while the maximization step detects the transmitted signal using the estimated CSI.
-A modified, trainable projected gradient detector is used in the maximization step.
-In [259], an ADMM-based unfolded algorithm is proposed for robust MIMO detection, incorporating channel estimation errors.
-To reduce error propagation due to channel aging, a robust data-aided Kalman filter-based channel tracking method is developed.
-The pre-estimated data symbols are used to improve channel estimation accuracy, enhancing robust detection performance.
-Recently, the energy harvesting technique has gained significant attention as a promising feature for green communication networks.
-Ambient backscatter communication is an energy-efficient technology that allows devices to transmit data by reflecting existing ambient signals, such as those from Wi-Fi routers or cellular networks, instead of generating their own radio waves.
-A backscatter device modulates the signal over the ambient radio frequency without an active radio frequency component.
-However, the backscatter link is weaker than the direct link, which limits its coverage and data rate [260].
-In [261], Wang et al.
-explored an IRS-based ambient backscatter communication system.
-The proposed network consists of two parts.
-The first subnetwork calculates the IRS phase shift matrix for binary symbols.
-The phase shift matrix weights, or phase angle vectors, are updated using a DNN training tool.
-These updated phase angle vectors are then used to compute the received signal, which is passed to the second subnetwork, the detection module.
-The EM algorithm is unfolded for symbol detection.
-XI-B Joint Activity Detection and Channel Estimation
-In [262], Ke et al.
-addressed massive access in massive MIMO systems by introducing an adaptive scheme for active user detection and channel estimation using compressive sensing techniques.
-This scheme takes advantage of the sparse traffic patterns of many user devices and the sparse nature of massive MIMO channels in the virtual angular domain.
-The access latency is reduced due to the use of non-orthogonal pseudo-random pilots for uplink communication.
-The problem of detecting active users and estimating their channels is formulated as a generalized multiple measurement vector compressive sensing problem.
-A generalized multiple measurement vector AMP (GMMV-AMP) algorithm is designed to leverage the structured sparsity in the uplink channel matrix.
-Additionally, a Turbo-GMMV-AMP algorithm is developed to exploit channel sparsity in both the spatial and angular domains.
-In [263], an EP-based joint active user detection and channel estimation technique for mMTC networks is proposed.
-The EP algorithm iteratively minimizes the Kullback-Leibler divergence between the target and approximate distributions via moment matching, efficiently finding a tractable close approximation of the target distribution.
-Grant-free random access is used for active device detection at the BS to address the latency issue in massive access.
-The active device detection is formulated as a sparse signal recovery problem.
-To improve the performance of massive device detection, Qiang et al.
-[43] proposed a deep unfolded network called AMP-Net, as shown in Fig.
-10.
-The basic principle of AMP-Net is based on the AMP algorithm.
-The input variables passed through the $T$ layers include the state evolution parameter, $\tau^{2}$, the device state matrix $\bf{X}$, the received signal $\bf{Y}$ at the BS, and the residual matrix $\bf{R}$.
-The pilot vectors are denoted by $\bf{a_{\text{n}}}$.
-Each layer contains an MMSE estimator, $\eta_{t}(\cdot)$.
-The set of trainable parameters consists of four scalar values.
-The parameters $\lambda_{n}^{t}$ and $\gamma_{n}^{t}$ estimate the activity probability and channel variance, respectively.
-The parameter $\theta_{n}^{t}$ regulates the denoiser, while $\mu_{n}^{t}$ controls the state evolution parameter, helping facilitate an accurate prediction of the detected target.
-Figure 10: The structure of AMP-Net [43].
-Traditional orthogonal multiple access fails to achieve massive connectivity for mMTC systems.
-Random and sporadic data traffic means that only a small fraction of users access the BS during uplink transmission.
-The joint channel estimation and active user detection problem for uplink mMTC is investigated in [264].
-The authors developed a pilot transmission scheme for nonorthogonal identification.
-Due to the sporadic activity of users, where only a small fraction are active at any time, traditional orthogonal pilots are inefficient.
-Instead, nonorthogonal pilots are used, and the joint estimation and detection problem is formulated as a compressive sensing problem.
-The work considers single and multiple antenna configurations at the base station, modeling them as single measurement vector and multiple measurement vector problems, respectively.
-The authors designed a deep unfolding model based on the ADMM algorithm.
-For ADMM, the trainable parameters include: (1) an auxiliary variable to compute matrix inversions, (2) a regularization parameter that controls the penalty term in the ADMM optimization, influencing the balance between sparsity and accuracy in the recovered signal, and (3) a threshold parameter in the shrinkage function that determines the sparsity of the output in each iteration.
-In [265], two deep unfolded neural network architectures are proposed for sporadic user detection and channel estimation in mMTC.
-The first architecture unfolds a linearized ADMM, applied to a constrained convex problem.
-Linearized ADMM addresses computationally intractable block coordinate sub-problems by executing a single iteration of the proximal gradient method for each subproblem instead of solving them explicitly [156].
-The second architecture unfolds a vector AMP with a denoiser based on ISTA.
-This unfolded network reduces complexity and improves accuracy compared to the original algorithm.
-In [266], a deep unfolded algorithm is proposed for the joint design of compression modules, quantization modules at the APs, and the decompression and detection modules at the CPU for cell-free massive MIMO systems with capacity-limited fronthauls.
-The objective is to address activity detection in an end-to-end learning framework.
-Ma et al.
-[267] investigated joint device activity and data detection in mMTC.
-They proposed a one-phase non-coherent scheme, where data bits are embedded in the pilot sequences.
-The BS simultaneously detects active devices and their embedded data bits without explicit channel estimation.
-The AMP algorithm is unfolded to enhance detection performance by effectively exploiting the pilot activity correlation.
-In [268], a contention-based asynchronous random access scheme with data length diversity is considered for mMTC.
-The paper unfolds AMP with backpropagation to address the composite problem of activity detection, channel estimation, and data recovery.
-Zou et al.
-[269] formulated the joint activity detection and channel estimation problem as a group-row-sparse matrix recovery problem.
-A proximal gradient-based unfolding NN is developed to parameterize the algorithmic iterations and solve this problem.
-In [270], compressive sensing recovery algorithms are used to address the joint active device detection and channel estimation problem for uplink transmission in grant-free based vehicular networks.
-The ADMM algorithm is unfolded to improve convergence rate and recovery accuracy.
-Shi et al.
-[271] considered grant-free massive access in an IoT network.
-A large number of single-antenna IoT devices sporadically transmit non-orthogonal sequences to a multi-antenna BS.
-An unfolded DNN framework is developed using the multidimensional shrinkage thresholding operator inherited from the ISTA structure.
-This network simultaneously achieves low computational complexity and high robustness for solving the joint activity detection and channel estimation problem.
-Cell-free technology for massive MIMO enhances mmWave signal coverage [272].
-Using deep unfolding techniques, the authors in [273] developed a grant-free access protocol based on mmWave cell-free massive MIMO.
-The proposed hybrid structure leverages learned vector AMP to reduce the need for prior channel information and achieve faster convergence.
-XI-C Joint Channel Estimation and Precoder Design
-A codebook design for channel estimation in mmWave communications with a hybrid precoding structure is studied in [274].
-To evaluate the performance of a codeword, the authors propose a metric called generalized detection probability (GDP), which considers both the beam pattern’s flatness and the maximum transmission power.
-Yuan et al.
-[275] introduced a two-stage wideband hybrid beamforming scheme for THz wireless systems.
-The model features a multi-antenna BS communicating with a multi-antenna user over frequency-selective fading.
-Additionally, a probabilistic robust hybrid beamforming scheme is designed to mitigate channel estimation errors.
-(a) Long-term
-(b) Short-term
-Figure 11: JCEHB architecture [276].
-Instead of modeling and optimizing hybrid beamforming and channel estimation separately, Kang et al.
-[276] proposed the joint channel estimation hybrid beamforming (JCEHB) algorithm.
-The sum rate is maximized for TDD massive MIMO system in a mixed timescale framework.
-The algorithm’s architecture is shown in Fig.
-11.
-Observe from Fig.
-LABEL:sub@fig:fig1, the deep-unfolding NN architecture consists of two parts : channel estimation and hybrid beamforming.
-For channel estimation, the pilot training module is designed by unfolding the recursive least squares (RLS) algorithm.
-The SSCA algorithm is unfolded to replace the high complexity computation operations in hybrid beamforming.
-The long-term analog beamforming matrices, are optimized offline with CSI statistics that is constant over long time slots or a superframe.
-After fixing the analog beamformers, the deep-unfolding NN for channel estimation is trained at each time slot to optimize the digital beamformers, as shown in Fig.
-LABEL:sub@fig:fig2.
-In contrast to long-term analog beamforming, short-term digital beamforming matrices are optimized every time slot using real-time low-dimensional effective CSI.
-Analog beamforming matrices are optimized offline using CSI statistics, while digital beamforming matrices are updated within each time slot based on estimated low-dimensional CSI.
-The authors claimed that the hybrid approach in a mixed (long-term and short term) timescale framework reduce signaling overhead as compared to separate design of NN in a single timescale.
-An unfolded WMMSE-based NN for massive multi-user MIMO under imperfect CSI is studied in [277].
-The proposed network enhances performance with low complexity by leveraging CSI statistics for near-optimal precoder solutions, accelerating convergence.
-In [278], an unfolded sparse Bayesian learning algorithm is derived to capture channel sparsity in a mmWave massive MIMO system with hybrid beamforming.
-Similarly, [279] proposes a learning-based channel semantic acquisition and beamforming solution for cell-free massive MIMO.
-A multi-layer perceptron estimates the channel, while a successive over-relaxation-based linear beamforming scheme reduces estimation errors in hybrid beamforming.
-Wideband precoding under imperfect CSI is discussed in [280].
-A robust frequency-selective precoding algorithm maximizes the expected weighted sum rate while considering per-antenna power constraints.
-This algorithm is unfolded to reduce computational complexity.
-To improve beamspace channel estimation, [281] introduces a deep unfolding algorithm using the top $\ell_{K,2}$ norm.
-This norm, a variation of the traditional $\ell_{2}$ norm, focuses on the largest elements in the sparse beamspace channel vector [282].
-It helps determine sparsity levels, reducing noise and computational complexity.
-The iterative trimmed-ridge regression algorithm is unfolded to allow parallel computation.
-These algorithms are applicable to both narrowband and wideband massive MIMO systems, ensuring versatility in channel estimation.
-XI-D Joint Detection and Decoding
-An adaptive Bayesian receiver using sequential Monte Carlo for signal detection and decoding in fading channels with known channel statistics is proposed in [283].
-The transmitted signal is treated as missing data, and multiple signal samples are sequentially estimated based on observed signals.
-These estimated sequences approximate the Bayesian estimate of both transmitted signals and channel states.
-Adaptive receiver algorithms are developed for both uncoded and convolutionally coded systems.
-El Gamal et al.
-[284] present an iterative receiver for joint detection and decoding of CDMA signals.
-The multiuser detector follows the MMSE criterion.
-A modified symbol-aided demodulation technique extracts soft information from previous decoding iterations to improve channel estimates, enhancing joint detection and decoding efficiency.
-Traditionally, channel estimation, signal detection, and channel decoding in a MIMO receiver are performed sequentially and independently.
-This disjoint approach leads to suboptimal performance.
-To address this, Sun et al.
-[285] propose receivers that integrate these functions, reducing error propagation and processing delay.
-Their design targets MIMO ultra-reliable low-latency communication (URLLC) systems.
-The model uses short channel codes and a limited number of pilot symbols for channel estimation.
-The authors employ the ADMM algorithm to solve large-scale non-convex optimization problems, deriving closed-form solutions for each iteration.
-Two NNs, JCDDNet-G for Gaussian channels and JCDDNet-S for sparse mmWave channels, are built by unfolding the ADMM algorithms as shown in Fig.
-12.
-Simulations show that the proposed trainable JCDD receivers significantly outperform traditional turbo receivers.
-Figure 12: JCDDNet-S/G architecture [285].
-In [286], Zhang et al.
-propose a receiver framework based on a deep unfolded NN structure.
-The joint signal detection and channel decoding modules are modeled as multi-layer deep feed-forward networks.
-Signal detection is achieved by unfolding the EP algorithm, while channel decoding is based on unfolded turbo decoding.
-Building on this work, Zhang et al.
-develop an unfolding-based framework for MIMO detectors in [287].
-This framework adaptively determines trainable parameters, allowing the detector to adjust effectively to varying channel conditions.
-XII Challenges in Developing Deep Unfolding Models and Future Research Scopes
+There is limited literature on the application of deep unfolding techniques in physical layer security.
+One possible reason is that physical layer security strategies are often dynamic and must adapt to adversarial attacks.
+In contrast, deep unfolding methods typically rely on a structured iterative algorithm which are seldom available in scenarios such as jamming attacks, where the system behavior must adjust rapidly to unpredictable threats.
+X Challenges in Developing Deep Unfolding Models and Future Research Scopes
 The extensive review of recently emerging techniques and applications of deep unfolding have demonstrated its potential in addressing complex signal processing problems with reduced complexity and runtime.
@@
 In this section, we outline the main challenges in developing deep unfolding models, explore the reasons behind these challenges, and propose potential improvements to make deep unfolding a more powerful and widely applicable technique in wireless communications.
-XII-A Limitations of Deep Unfolding Techniques
+X-A Limitations of Deep Unfolding Techniques
 Although deep unfolding has found applications in various design problems, it has limitations.
 Most of these limitations arise from the fact that deep unfolding is not an independent design methodology; it must be built upon existing conventional optimization frameworks.
-Consequently, developing deep unfolding solutions faces the following challenges:
-(C1) Contrasts between conventional optimization and deep unfolding: There is a notable contradiction between the development methodologies of conventional optimizers and deep unfolding models.
-Most conventional solutions for transceiver design and signal processing in wireless communications are based on convex optimization techniques.
-However, many wireless design problems are inherently non-convex and highly coupled due to factors such as fractional SINR, constant-modulus constraints (e.g., for analog beamforming), and mixed-integer variables (e.g., in symbol detection or antenna selection).
-These challenges are typically addressed using a two-layer approach.
-In the first layer, design variables are decoupled through alternating optimization (AO) or block coordinate ascent/descent, breaking the original problem into smaller sub-problems, each focusing on a subset of variables that are solved alternately.
-In the second layer, techniques such as SCA or the ADMM are applied to solve each sub-problem, often introducing additional variables.
-Although this two-layer approach transforms the original non-convex problem into a series of convex ones that can be solved using convex optimization tools like CVX, it significantly increases the number of variables and sub-problems and leads to nested loops in the solution process.
-These loops, corresponding to the two layers, must iterate until convergence, adding considerable complexity.
-In contrast, deep unfolding models typically approach the problem as a unified whole, solving all variables in parallel, facilitating the development of unfolding architectures.
-This fundamental difference makes it challenging to develop suitable algorithms for unfolding.
-Additionally, the increased complexity in variables and sub-problems inherent in conventional methods further complicates the design of deep unfolding frameworks, directly contributing to their overall complexity.
-(C2) Challenges in unfolding principle algorithms: While deep unfolding typically enhances conventional iterative algorithms, not all can be effectively unfolded.
+Consequently, it faces the following challenges in the development and real-world applications:
+(C1) Methodological challenges: Deep unfolding methods are inspired by the iterative procedures of conventional optimization algorithms, where each iteration is typically interpreted as a layer in an NN.
+In traditional wireless transmitter-receiver design and signal processing, optimization problems are often non-convex and highly coupled due to factors such as fractional SINR, constant-modulus constraints (e.g., for analog beamforming), or mixed-integer variables (e.g., in symbol detection or antenna selection).
+They are typically solved using a two-layer approach.
+In the first layer, AO or block coordinate ascent/descent is used to decouple variables, breaking the original problem into smaller, more manageable sub-problems.
+In the second layer, convex optimization techniques such as SCA or the ADMM are employed to solve each sub-problem.
+While this hierarchical structure enables the use of convex solvers (e.g., CVX), it often introduces auxiliary variables and results in nested loops that must converge at each layer, significantly increasing algorithmic complexity.
+Deep unfolding models, on the other hand, aim to transform these iterative procedures into end-to-end trainable architectures.
+Instead of solving sub-problems sequentially, the unfolding framework typically approximates the entire optimization process in a more unified and parallelizable manner.
+This shift in design philosophy, from explicitly solving decoupled sub-problems to learning a unified mapping, poses non-trivial challenges, especially when attempting to translate complex multi-stage optimizers with auxiliary variables into a compact, learnable structure.
+Therefore, while deep unfolding builds upon conventional iterative methods, the methodological differences in structure and implementation represent a major challenge in developing effective unfolding models.
+(C2) Unfolding principle algorithms: While deep unfolding typically enhances conventional iterative algorithms, not all can be effectively unfolded.
 This limitation arises from the core principle of deep unfolding, which structures the iterations of an iterative algorithm as layers within a DNN.
-Complex operations like matrix inversion, matrix decomposition, or the use of convex optimization tools such as CVX or YALMIP (with solvers like MOSEK, SeDuMi, etc.) are difficult to implement in a DNN framework.
+Complex operations like matrix decomposition or the use of convex optimization tools such as CVX or YALMIP (with solvers like MOSEK, SeDuMi, etc.) are difficult to implement in a DNN framework.
 In large-scale systems such as massive MIMO, even matrix multiplications result in high complexity for unfolding models.
@@
 Proper tuning is essential to achieve a favorable performance-complexity tradeoff in deep unfolding models.
-(C3) Challenge in handling design constraints and objectives: One of the biggest challenges in developing deep unfolding models lies in managing the design constraints formulated in the problem.
+(C3) Handling design constraints and objectives: One of the biggest challenges in developing deep unfolding models lies in managing the design constraints formulated in the problem.
 These constraints require the DNN layers’ outputs to meet specific requirements, such as total power, QoS, or other element-wise constraints.
@@
 While they resemble activation functions, none of the existing activation functions can adequately manage such intricate requirements.
-One approach is to incorporate these constraints into the design objective, but this method often results in a problem that is not equivalent to the original constrained one, making the solution potentially infeasible.
-Properly modeling such constraints within a deep unfolding DNN remains highly challenging, hindering the development of more advanced deep unfolding models.
-Not only do complex constraints pose challenges for developing deep unfolding models, but intricate objective functions also present difficulties.
-Generally, deep unfolding models are trained based on the objective function.
+One approach is to incorporate these constraints into the design objective.
+However, this technique often results in an nonequivalent problem with a more complicated objective function [191].
+Not only do complex constraints pose challenges for developing deep unfolding models, but intricate objective functions also present difficulties in training the models.
+This is because their loss functions are generally formulated based on the objective function.
 For instance, unfolding architectures like DetNet [49], FSNet [75], and ScNet [229] for symbol detection are trained to minimize the Euclidean distance between the model’s output (symbol vector) and the true transmitted symbol vector.
-When labels are unavailable, as in beamforming problems, unfolding networks use the objective function as the loss function for minimization or its negative for maximization problems.
+In particular, when labels are unavailable, as in beamforming problems, unfolding networks use the objective function as the loss function for minimization or its negative for maximization problems.
 Typical deep unfolding techniques [191, 155] require calculating the gradient of the objective function.
 However, this becomes challenging when the objective function is a complex, high-order, or fractional function of the design variables, especially when those variables are vectors or matrices.
-Therefore, more dynamic methodologies are needed for developing unfolding models that can handle such intricate functions efficiently.
-(C4) Need for black-box models in sophisticated mappings: While deep unfolding is powerful, certain inference tasks still require black-box DNN models, especially when there is no clear relationship between the input data, output solution, and performance metrics.
-For instance, in multi-modal sensing-assisted beam tracking, tasks may involve complex mappings from environmental data—such as sensory information from sensors and cameras—to the optimal beam [289, 290].
+(C4) Sophisticated mappings: While deep unfolding is powerful, certain inference tasks still require black-box DNN models, especially when there is no clear relationship between the input data, the output solution, and performance metrics.
+For instance, in multi-modal sensing-assisted beam tracking, tasks may involve complex mappings from environmental data, such as sensory information from sensors and cameras, to the optimal beam [230, 231].
 In such cases, there is no explicit relationship between environmental images and the optimal beam, and standard communication metrics like throughput or SINR cannot be directly expressed as functions of these images.
 These problems cannot be formulated as optimization tasks solvable with principle algorithms, making it difficult to develop unfolding models that map environmental information to beam patterns or blockage states.
-The above three challenges C1C4 also outline the scenarios wherein deep unfolding models cannot be built or does not offer much gain in performance-complexity tradeoff.
-These include when there is lack of suitable conventional iterative algorithms (C1, C2, C4) or when the design problems are highly complicated with constraints and objective function (C3).
+(C5) Real-world applications: Deep unfolding architectures, while promising in mimicking iterative algorithms through trainable layers, often entail considerable depth and parameter complexity [191].
+This imposes severe challenges for real-time inference under strict latency and throughput constraints, especially in systems with multiple and large dimensions like massive MIMO-OFDM.
+From a hardware perspective, deep unfolding designs are not naturally aligned with the requirements for area-efficient and parallelizable architectures needed in scalable OFDM deployments.
+For example, the non-uniform and sequential computations common in unfolding detectors can result in inefficient silicon utilization and high power consumption, limiting the ability to instantiate multiple parallel detection cores across subcarriers [232].
+Furthermore, in real-world scenarios, the data distribution is highly dynamic because of user mobility, varying propagation environments, and hardware impairments.
+These variations often violate the assumptions underlying many unfolding-based transmitter and receiver designs, which are typically developed under idealized conditions.
+However, this assumption breaks down in practical deployments, where channels often exhibit strong line-of-sight components and spatial correlation.
+As a result, detectors relying on such assumptions can suffer significant performance degradation.
+To maintain robustness in these conditions, deep unfolding models must be redesigned or trained using data that captures the complexity and variability of realistic, deployment-specific channels.
+The above five challenges (C1C5) also outline the scenarios wherein deep unfolding models cannot be built or does not offer much gain in performance-complexity tradeoff.
+These include when there is lack of suitable conventional iterative algorithms (C1, C2, C4), when the design problems are highly complicated with constraints and objective function (C3), or when the exist imperfection in real-world propagation environments and hardware (C5).
 It can be seen that deep unfolding is not a game changer that will completely replace conventional principle algorithms or black-box ML models.
 However, the applicability of deep unfolding can be enhanced by overcoming those challenges, as will be discussed next.
-XII-B Development of Deep Unfolding Techniques
-XII-B1 Coexistence of Deep Unfolding Models, Conventional Solvers, and Black-Box Models
+X-B Development of Deep Unfolding Techniques
+X-B1 Coexistence of Deep Unfolding Models, Conventional Solvers, and Black-Box Models
 Deep unfolding doesn’t need to fully replace conventional algorithms or black-box ML models.
@@
 The combination of deep unfolding structures with conventional solvers and black-box models leads to simpler yet more efficient designs.
-XII-B2 Deep Unfolding for Multi-Objective Learning
+X-B2 Deep Unfolding for Multi-Objective Learning
 Most existing deep unfolding models are developed for single-objective designs, such as maximizing throughput, SINR, or minimizing BER.
@@
 In ISAC designs, weighted sums of communication and sensing performance metrics are used when the goal is to balance both functions without prioritizing one over the other.
-The complexity grows in multi-static ISAC systems, where distributed sensing receivers have different performance metrics depending on their functions, such as SCNR, mutual information, Cramér-Rao lower bound, or probability of detection.
+The complexity grows in multi-static ISAC systems, where distributed sensing receivers have different performance metrics depending on their functions, such as signal clutter noise ratio (SCNR), mutual information, CRB, or probability of detection.
 These designs are often difficult to solve using traditional constrained formulations, as the problem can easily become infeasible if constraints are not properly tuned.
@@
 Furthermore, balancing the changes of these objectives with respect to the design variables across layers is crucial [191, 234], necessitating careful design of the unfolding models.
-XII-B3 Distributed Deep Unfolding Architectures
+X-B3 Distributed Deep Unfolding Architectures
 While deep unfolding models provide advantages in terms of complexity due to faster convergence compared to conventional iterative algorithms, they can still face high complexity issues due to challenge C2.
 This limits their applicability on platforms with constrained computational resources, such as edge devices or distributed access points (APs).
-Future wireless networks will advance beyond simple communications, evolving into large-scale, distributed systems that integrate advanced sensing, communications, and computing capabilities.
+Future wireless networks will advance beyond simple communications, evolving into large-scale, distributed systems such as cell-free massive MIMO that integrate advanced sensing, communications, and computing capabilities.
 However, most existing AI models, both data-driven and model-driven, were designed for centralized scenarios, where they are deployed at a central node with substantial computing power.
-These models are not well-suited for distributed scenarios, highlighting the need for the development of appropriate distributed unfolding architectures.
+These models are not well-suited for distributed scenarios, highlighting the need for appropriate distributed unfolding architectures.
 One potential solution is to develop distributed deep unfolding architectures.
@@
 By such an unfolding framework, highly reliable symbols are generated with low complexity at distributed APs.
-XII-B4 Deep Unfolding to Enhance Black-box Solution
+X-B4 Deep Unfolding to Enhance Black-box Solution
 As its name suggests, a black-box ML model operates with a black-box nature, making it unexplainable and resistant to re-structuring beyond tuning parameters like the number of layers or nodes.
@@
 Specifically, high-resolution camera images generate considerable traffic overhead and latency, impacting real-time communication processing.
-Deep unfolding models can be particularly useful in pre-processing such images.
+Deep unfolding models can be particularly useful in pre-processing such high resolution images.
 In image processing, deep unfolding has been successfully applied to tasks such as denoising, super-resolution, and image reconstruction, allowing for high-quality outcomes with fewer iterations than conventional algorithms.
-For example, unfolded algorithms based on the ISTA for sparse coding have been shown to significantly accelerate image reconstruction while maintaining high-quality results [293, 294, 295].
+For example, unfolded algorithms based on ISTA for sparse coding have been shown to significantly accelerate image reconstruction while maintaining high-quality results [235, 236, 237].
 This makes deep unfolding a valuable tool for handling large-scale sensory data in real-time communication systems.
-XIII Conclusions
-The integration of traditional algorithmic principles with DL techniques has led to the development of deep unfolding, a method that overcomes the limitations associated with both conventional algorithms and standalone DL models in wireless communication.
+XI Conclusions
+The integration of classical algorithmic principles with DL techniques has led to the development of deep unfolding, a method that overcomes the limitations associated with both classical algorithms and standalone DL models in wireless communication.
 This article offers a comprehensive survey of deep unfolding techniques, examining their application across various facets of the physical layer in wireless communication systems.
-Key applications, such as signal detection, channel estimation, precoder design, sensing, communication decoding for error-correcting codes, power allocation, and security, are explored, with an emphasis on technologies anticipated for beyond-5G networks.
+Key applications, such as signal detection, channel estimation, precoder design, ISAC, decoding for error-correcting codes, power allocation, and physical layer security, are explored, with an emphasis on technologies anticipated for beyond-5G networks.
+The comparisons of the unfolded networks have also been given, which reflect the performance gains over the benchmark iterative algorithms.
+Also, we have analyzed the advantages and limitations of the design of various unfolding models.
+Moreover, we have explored the practical feasibility of deep unfolding networks across key application areas of wireless communications.
 The survey highlights the strengths of deep unfolding in addressing complex challenges, particularly in achieving low-latency, high-accuracy performance in dynamic environments.
 We conclude by discussing the key challenges in developing deep unfolding models, including issues related to scalability, computational efficiency, and robustness across diverse scenarios.
-Finally, we propose potential advancements to enhance the adaptability and resilience of deep unfolding, with the goal of solidifying its role as a robust and versatile technique for next-generation wireless communications.
+Finally, this paper presents potential advancements to enhance the adaptability and resilience of deep unfolding, with the goal of solidifying its role as a robust and versatile technique for next-generation wireless communications.

### v4

@@
 Finally, we discuss the challenges associated with developing deep unfolding techniques and propose potential improvements to enhance their applicability across diverse wireless communication scenarios.
-Index Terms:
-Deep learning, Deep unfolding, Model-driven machine learning, Wireless communication signal processing, Integrated sensing and communication (ISAC), Intelligent reflecting surfaces (IRS), Terahertz, IoT.
 I Introduction
@@
 Data-driven neural network (NN) architectures were developed primarily to process language and vision data, rendering them suboptimal for wireless communication tasks that involve channels, signals, and outputs governed by complex constraints.
-Although certain constraints can be handled through the application of specific activation functions, for example, employing the $tanh$ function to enforce a constant modulus constraint in analog beamforming, standard activation functions generally lack the versatility to manage the intricate constraints characteristic of wireless communication systems.
+Although certain constraints can be handled through the application of specific activation functions, for example, employing the tanhtanh function to enforce a constant modulus constraint in analog beamforming, standard activation functions generally lack the versatility to manage the intricate constraints characteristic of wireless communication systems.
 To mitigate these challenges, model-driven DL approaches facilitate the development of problem-specific network architectures and customized activation functions, optimized for the unique signal processing requirements of wireless communication [222].
@@
 I-A Related Works
-The growing adoption of deep unfolding in communication systems has led to several surveys covering its diverse applications.
+The growing application of deep unfolding in communication systems highlights the need for a comprehensive survey in this area.
+There are limited survey articles that comprehensively discuss deep unfolding applications in the context of the physical layer of communication.
+Here, we present a few related survey articles on deep unfolding that explore some particular task such as detection, channel estimation, or precoder design.
 Wang et al.
-[21] reviewed DL and deep unfolding methods for physical layer tasks such as MIMO detection, channel estimation, and decoding.
-The authors of [22] discussed its use in transceiver design and channel state information (CSI) estimation.
-In [23], deep learning methods including DNNs, GANs, and recurrent NN (RNN) were reviewed, with deep unfolding applied to Bayesian estimation.
-The authors of [15] focused on MIMO detection, multiuser precoding, and belief propagation.
-A broader comparison of model-based, DL, and hybrid methods is found in [24].
-Monga et al.
-[14] focused on algorithm unrolling for signal/image processing.
-Additional works [25, 13] explored signal processing and optimization.
-Eldar et al.
-[12] and Shi et al.
-[26] addressed DL-based communication frameworks.
-In [27], the authors highlight the advantages of deep unfolding for physical receiver design, with particular emphasis on its applications in signal detection and channel estimation.
-A comparison between data-driven and unfolding-based beamforming appears in [28] and Table II.
+in [199] presented a review of DL and deep unfolding approaches that address physical layer signal processing challenges, including MIMO detection, channel estimation, and channel decoding.
+The authors of [69] provided an overview of deep unfolding applications in transceiver design, MIMO detection, and channel state information (CSI) estimation and feedback.
+Similarly, in [161], the authors highlighted the role of DL in communication, with a focus on DNNs, GANs, recurrent neural network (RNN), and end-to-end learning.
+They also explored deep unfolding for Bayesian optimal estimation, using MIMO detection and channel estimation as case studies.
+The authors in [14] focused primarily on MIMO detection while exploring its use in multiuser MIMO precoding and belief propagation-based channel decoding.
+A survey on model-based, DL, and hybrid approaches in the context of wireless communication networks is provided in [227].
+In [146], Monga et al.
+conducted a survey of deep unfolding algorithms specifically implemented for applications in signal and image processing.
+The authors in  [92] reviewed the implementation of deep unfolding in signal processing tasks.
+A brief review on deep unfolding in MIMO detection is presented in [177].
+The authors in [174] have presented a general framework for optimization that leverages the benefit of deep unfolding with signal detection as one of the case studies.
+In [222], a chapter is dedicated to model-based machine learning for communication.
+The deep unfolding networks have been discussed for signal detection, channel estimation, and channel coding.
+Machine learning-based optimization framework is discussed in [173] with an emphasis on deep unfolding in specific design paradigms such as signal detection.
+Quoc et al.
+have presented a survey on application of deep unfolding in signal detection, channel estimation and beamforming in [158].
+The authors in [126] explored the application of learning-based optimization techniques to enhance the capacity of communication networks.
+The survey reviews key domains such as encoding-decoding, equalization, power control, and beamforming, emphasizing the impact of optimization through learning on network performance.
+A comparative study between purely data-driven and deep unfolding is conducted with complete emphasis on HBF optimization in [176] The significant rise in application of deep unfolding has led to development of theoretical framework to validate and support practical experiments.
+There are several works [165, 33, 34] which primarily focus on the convergence rate of the unfolded iterative shrinkage-thresholding algorithm (ISTA) and its variants such as learned ISTA (LISTA) [33] and analytic LISTA (ALISTA) [34].
+The authors of [35] discuss the motivation and design methodology of deep unfolding, using LISTA as a representative case study.
+They have studied the design of learned parameter, techniques to design the architecture of the network and parameterize the resulting unfolded network.
+Certain open mathematical problems pertaining to deep unfolding are also explored in [35].
 I-B Contributions
-Existing surveys lack coverage of recent studies addressing critical 6G technologies, such as mmWave spectrum, new air interface that includes massive MIMO, IRS, next-generation multiple access, and are often limited to specific application areas.
-Consequently, a comprehensive survey is needed to encompass the broad applications of deep unfolding techniques across various aspects of the physical layer in wireless communication.
+Recently, there has been a surge of application of deep unfolding in various areas of physical layer of wireless communication.
+The existing survey contributions such as [14, 199, 69], all published before 2020, focus on some particular areas of the physical layer of wireless communication, such as detection, channel estimation, precoder design, or channel coding.
+In [176], the authors have provided a survey on application of deep unfolding only in the area of hybrid beamforming design.
+However, the realm of deep unfolding has recently expanded into emerging areas such as sensing and communication, which have not been addressed in any of the previous works.
+Most of the previous literature discuss the general approach of deep unfolding without dedicated analysis concerning communication perspective.
+Only a few works such as [146, 173] have given a tutorial description but with a biased emphasis towards signal processing tasks.
+Therefore, we provide a detailed tutorial describing the deep unfolding in communication perspective and discuss convergence proofs that impact algorithmic stability.
+In addition, in this work, the application of deep unfolding is explored in the context of critical 6G technologies that are not addressed in existing survey articles such as [14, 199, 69].
+Therefore, there is a lack of coverage of recent studies addressing critical 6G technologies, such as mmWave spectrum, new air interface that includes massive MIMO, IRS, next-generation multiple access, and are often limited to specific application areas.
+Consequently, a comprehensive survey is needed to encompass the broad applications of deep unfolding techniques across various aspects of the physical layer in wireless communication which has not been addressed in existing survey articles [14, 199, 69, 176].
 This paper provides an in-depth review of deep unfolding methods applied to the physical layer of wireless communication.
 The contributions are summarized below:
+Figure 1: Typical block diagram of a modern wireless communication system. The section number covering a particular component is shown as well.
   1. 1.
     This paper presents a systematic, step-by-step approach for developing deep unfolding algorithms, with a detailed tutorial on the alternating direction method of multipliers (ADMM)-based MIMO detection.
-The tutorial offers insights into the ADMM unfolding process, identifies trainable parameters, and demonstrates the design of the resulting unfolded network.
+The tutorial offers insights into the ADMM unfolding process, identification of the trainable parameters, and demonstrates the design of the resulting unfolded network.
+Furthermore, we discuss the theoretical foundation of the convergence properties and guidelines for selecting hyperparameters that impact algorithmic stability of deep unfolding.
+We have highlighted that the convergence properties can be studied under three categories: convergence order, generalization error and normalized mean squared error (NMSE).
   2. 2.
-    We survey the application of deep unfolding across various domains, including signal detection, channel estimation, precoder design, ISAC, decoding, power allocation, and physical layer security.
-Each section highlights the integration of deep unfolding with key 6G technologies such as IRS, mmWave, and IoT, providing readers with guidance on designing frameworks adaptable to future wireless systems.
+    Fig.
+1 depicts the principal components of the physical layer in contemporary wireless communication systems.
+Considering this block diagram, we undertake a systematic survey of each component in details.
+In particular, we investigate the role of deep unfolding for signal detection, channel estimation, precoder design, integrated sensing and communication (ISAC), decoding, power allocation, and physical-layer security.
+For each domain, we highlight the integration of deep unfolding with key enabling technologies for sixth-generation (6G) networks, such as intelligent reflecting surfaces (IRS), millimeter-wave (mmWave) communications, and the Internet of Things (IoT).
+The survey aims to provide a comprehensive perspective on the design of adaptable and forward-looking frameworks for next-generation wireless systems.
   3. 3.
-    In addition, each section concludes with a comparative analysis table that evaluates prominent unfolded networks across key metrics such as computational complexity, generalization capability, and convergence behavior.
+    In addition, each section concludes with a cross method comparison table that evaluates prominent unfolded networks across key metrics: computational complexity (floating point operations (FLOPs) and big $\mathcal{O}$ complexity), memory footprint (number of trainable parameters) and convergence speed.
 These tables highlight the practical trade-offs and deployment boundaries associated with each architecture.
@@
     This paper concludes the survey by outlining a set of challenges in developing deep unfolding models and propose future directions that could guide advancements in deep unfolding techniques to enhance their performance.
-TABLE II: Existing survey papers on deep unfolding for communication.
-SNo. Reference Year Detection Channel Estimation Precoder Design Sensing and Communication Decoding in Error Correcting Codes Power Allocation Security
-1 Deep Learning for Wireless Physical Layer: Opportunities and Challenges [21] 2017 $\checkmark$ $\checkmark$ $\checkmark$
-2 Model-Driven Deep Learning for Physical Layer Communications [22] 2019 $\checkmark$ $\checkmark$
-3 Deep Learning in Physical Layer Communications [23] 2019 $\checkmark$ $\checkmark$
-4 Deep Unfolding for Communications Systems: A Survey and Some New Directions [15] 2019 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-5 Wireless Networks Design in the Era of Deep Learning: Model-Based, AI-Based, or Both? [24] 2019 $\checkmark$ $\checkmark$
-6 Algorithm Unrolling: Interpretable, Efficient Deep Learning for Signal and Image Processing [14] 2021 $\checkmark$ $\checkmark$
-7 Intelligent Radio Signal Processing: A Survey [29] 2021 $\checkmark$ $\checkmark$ $\checkmark$
-8 Redefining Wireless Communication for 6G: Signal Processing Meets Deep Learning With Deep Unfolding [25] 2021 $\checkmark$ $\checkmark$ $\checkmark$
-9 Model-Based Deep Learning: Key Approaches and Design Guidelines [30] 2021 $\checkmark$
-10 Model-Based Deep Learning: On the Intersection of Deep Learning and Optimization [13] 2022 $\checkmark$
-11 Machine Learning and Wireless Communications [12] 2022 $\checkmark$ $\checkmark$ $\checkmark$
-12 Machine Learning for Large-Scale Optimization in 6G Wireless Networks [26] 2023 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-13 Artificial Intelligence-Empowered Hybrid Multiple-Input/Multiple-Output Beamforming: Learning to Optimize for High-Throughput Scalable MIMO [28] 2024 $\checkmark$
-14 Model and Data Dual-Driven: A Promising Receiver Design for 6G Wireless Communications [27] 2025 $\checkmark$ $\checkmark$
-15 This Survey $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+We point out that challenges in designing a deep unfolding model mainly emerge from scenarios where these models does not offer significant gain in performance-complexity tradeoff.
+However, we also provide potential solutions that deep unfolding models can benefit from to overcome those challenges.
+  5. 5.
+    We also provide open source representative works with its github links for different design problems such as detection, channel estimation, ISAC, among others.
+We believe that this will enable readers to reproduce and develop new unfolding methods.
 I-C Paper Outline
-The skeleton of this paper is depicted in Fig.  1.
+The skeleton of this paper is depicted in Fig.  2.
 Figure 2: Outline of the paper.
@@
 A step-by-step approach is presented for developing the unfolding of ADMM specifically for MIMO detection.
+Also a brief understanding on how trainable parameters impact the convergence order of the deep unfolded network is discussed.
   • Part B: This part explores deep unfolding applications in physical layer tasks: signal detection (Section III), channel estimation (Section IV), precoder design (Section V), ISAC (Section VI), decoding for error correcting code (Section VII), power allocation (Section VIII), and security (Section IX).
-  • Part C : Section X discusses the challenges in adopting the deep unfolding approach in wireless communication and the future scope of research. Section XI concludes the paper.
+  • Part C : Section X discusses the limitations of existing deep unfolding approaches and the challenges associated with their adoption in wireless communications.
+We further outline potential directions and solutions to address these limitations, highlighting promising avenues for future research.
+Section XI concludes the paper.
 I-D Notation and Acronyms
-[…]
+Lower case, bold-face lower case, and bold-face upper case letters denote scalars, vectors, and matrices respectively.
+$(\cdot)^{T}$ and $(\cdot)^{H}$ denote transpose and Hermitian transpose, respectively.
+$\|\cdot\|$ denotes the Euclidean norm of a vector.
+$\mathbb{R}^{n}$ and $\mathbb{C}^{n}$ denote $n$-dimensional real and complex vector spaces, respectively.
+$\mathcal{N}(\mu,\Sigma)$ denotes the multivariate Gaussian distribution with mean vector $\mu$ and covariance matrix $\Sigma$, while $\mathcal{CN}(\mu,\Sigma)$ denotes the complex multi-variate Gaussian distribution.
+Table I lists the main acronyms used in this paper.
 II Deep Unfolding: General Architecture and Fundamentals
@@
 The trainable parameters, $\left\{\boldsymbol{\theta}^{l}\right\}_{l=1}^{L}$ (shown in red fonts in Fig.
-2) includes the model parameters and the step-size for the regularization coefficients of the algorithm [26].
+3) includes the model parameters and the step-size for the regularization coefficients of the algorithm [173].
 Figure 3: General framework of training for deep unfolding. The learnable parameters are shown in red color.
@@
 $\mathbf{y}\triangleq\begin{bmatrix}\mathfrak{R}(\bar{\mathbf{y}})\\ \mathfrak{I}(\bar{\mathbf{y}})\end{bmatrix},\quad\mathbf{x}\triangleq\begin{bmatrix}\mathfrak{R}(\bar{\mathbf{x}})\\ \mathfrak{I}(\bar{\mathbf{x}})\end{bmatrix},\quad\mathbf{n}\triangleq\begin{bmatrix}\mathfrak{R}(\bar{\mathbf{n}})\\ \mathfrak{I}(\bar{\mathbf{n}})\end{bmatrix}$,
-[…]
+$\mathbf{H}\triangleq\begin{bmatrix}\mathfrak{R}(\bar{\mathbf{H}})&-\mathfrak{I}(\bar{\mathbf{H}})\\ \mathfrak{I}(\bar{\mathbf{H}})&\mathfrak{R}(\bar{\mathbf{H}})\end{bmatrix}$,
 where $\mathfrak{R}(\cdot)$ and $\mathfrak{I}(\cdot)$ denote the real and imaginary part of a complex vector or matrix, respectively.
@@
 Consequently, (2) can be relaxed to the following problem in the real domain: $\min_{{\bf{x}}\in{{\cal{X}}}^{2M}}\left\lVert{\bf{y}}-{\bf{Hx}}\right\rVert_{2}^{2}.$ We further introduce an auxiliary variable $\mathbf{z}$ to reformulate this problem as:
-$\underset{\begin{subarray}{c}{\bf{x}}\in{{\cal{X}}}^{2M}\end{subarray}}{\min}$ $\quad\left\lVert{\bf{y}}-{\bf{Hx}}\right\rVert_{2}^{2}$
-subject to $\quad\mathbf{x}=\mathbf{z},\mathbf{z}\in{{\cal{X}}}^{2M}.$
+$\begin{IEEEeqnarraybox}[]{[}][c]{rCl}\underset{{\bf{x}}\in{{\cal{X}}}^{2M}}{\min}&&\quad\quad\left\lVert{\bf{y}}-{\bf{Hx}}\right\rVert_{2}^{2}\\ \text{subject to}&&\quad\quad\mathbf{x}=\mathbf{z},\quad\mathbf{z}\in{{\cal{X}}}^{2M}.\end{IEEEeqnarraybox}$ (3)
 The augmented Lagrangian is given as:
@@
 ${\boldsymbol{\lambda}}^{l+1}$ $={\boldsymbol{\lambda}}^{l}+\rho\left({\mathbf{x}}^{l+1}-{\mathbf{z}}^{l+1}\right)\quad\forall l$ (7)
-where $\Pi_{[-1,1]}({\boldsymbol{a}})$ denotes an element-wise projection of $\boldsymbol{a}$ onto the set $[-1,1]$. More specifically, if ${\boldsymbol{b}}=\Pi_{[-1,1]}(\boldsymbol{a})$, we have
+where $\Pi_{[-1,1]}({\boldsymbol{a}})$ denotes an element-wise projection of $\boldsymbol{a}$ onto the set [1,1][-1,1]. More specifically, if ${\boldsymbol{b}}=\Pi_{[-1,1]}(\boldsymbol{a})$, we have
 $b_{i}=\begin{cases}+1,&\text{if }a_{i}>1\\ -1,&\text{if }a_{i}<1\\ a_{i},&\text{otherwise}.\end{cases}$
@@
 Furthermore, the layer-wise unfolding of ADMM for MIMO detection is illustrated in Fig.
-3.
+4.
 The received symbol vector $\mathbf{y}$ and the channel matrix $\mathbf{H}$ are treated as fixed inputs and are provided to every layer of the network.
@@
 These parameters form the trainable set ${\boldsymbol{\theta}}=\left\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{2}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{3}^{(l)}}\right\}_{l=1}^{L}$, where each layer $l$ has its own set of penalty coefficients.
-In the final layer $L$, the network produces the estimate ${\mathbf{z}}^{L}=\hat{\bf{z}}$.
+In the final layer $L$, the network produces the estimate $^{\mathbf{z}}^{L}=\hat{\bf{z}}$.
 For $n=1,\dots,N$, data samples, the output $\hat{\bf{z}}_{n}({\bf{y}}_{n},{\bf{H}};\rho_{1}^{(l)},\rho_{2}^{(l)},\rho_{3}^{(l)})$ is retrieved from the network as the predicted transmitted symbol for $\bf{y}$}.
@@
 9:Output: Final estimate $\hat{\bf{z}}={\bf{z}}^{L}$
+TABLE II: Simulation Parameters.
+Parameter Name Value
+Number of transmit antennas 16
+Number of receive antennas 16
+Constellation 4 QAM
+Propagation model EVA [49]
+Training data size 10000
+Testing data size 10000
+Epochs 20
+Batch size 300
+Number of Layers 5
+$\Pi_{\text{soft}}$ tanhtanh
 […]
-Figure 4: BER performance comparison for 16$\times$16 MIMO system. For soft projection, $tanh$ is used.
+Figure 5: BER performance comparison for $\times16$ MIMO system in EVA model [49]. For soft projection, tanhtanh is used.
 Fig.
 5 presents the bit error rate (BER) performance comparison of the unfolded ADMM detector with zero forcing (ZF), minimum mean squared error (MMSE), ADMM and sphere decoder (SD), under QPSK modulation.
-We consider a $16\times 16$ MIMO system with Rayleigh fading channels.
-The results demonstrate that the unfolded ADMM detector outperforms ZF, MMSE, and standard ADMM detectors, except for SD.
-The unfolded ADMM detector performs better due to the use of deep unfolding, where learnable parameters are adaptively trained.
+For the unfolded ADMM, the penalty parameters and step sizes are considered as learnable parameters.
+We consider a $16\times 16$ MIMO system in extended vehicular (EVA) propagation model.
+Table II summarizes the key simulation parameters.
+The power delay profile of EVA propagation model [49] is given by [0,1.5,1.4,3.6,0.6,9.1,7.0,12.0,16.90,-1.5,-1.4,-3.6,-0.6,-9.1,-7.0,-12.0,-16.9] dB against the excess tap delays [$0,30,150,310,370,710,1090,1730,2510$] ns.
+The results demonstrate that the unfolded ADMM detector outperforms ZF, MMSE, and classical ADMM detectors.
+The unfolded ADMM detector performs better due to the use of deep unfolding, where learnable parameters: penalty parameters and step sizes, are adaptively trained.
+The SD achieves near ML performance and is used often as a benchmark scheme for validating the performance of various detectors.
 This example has been included to provide the readers with an illustrative understanding of the deep unfolding process.
@@
 Additional performance improvements can be achieved by optimizing or introducing trainable parameters in the unfolded network.
+The introduction of trainable parameters make it challenging to derive theoretical insights for unfolded algorithms as compared to classical optimization algorithms.
+For any optimization method, it is crucial to perform convergence analysis.
+In the literature, the convergence studies of various methods can be classified into three broad categories:
+  1. 1.
+    Convergence order: In this category, the convergence properties of unfolded networks are explored to provide a comparison with classical methods.
+Under idealized choices of trainable parameters, different unfolded architectures are studied that yield an accurate solution with a faster convergence order.
+It is well known that the classical ISTA exhibits sublinear convergence [16].
+The work in [165] discusses how learned ISTA (LISTA) or unfolded ISTA produces a sequence of trainable parameters per layer exhibiting a linear convergence order.
+As discussed in [33], in the unfolded ISTA, due to the transformation of trainable parameters into learned weights, a linear convergence order is achieved.
+The authors in [35] point out that in deep unfolding, the most essential step in achieving better empirical performance and acceleration is to design the learnable parameters and analyze how the learnable parameters influence each other.
+For instance, in [33], the parameters related to dictionary based matrices in sparse signal recovery, are not shared across the LISTA network’s layers.
+The learnable parameters are instead trained in each layer, similar to untying them across the layers.
+The increased complexity in parameterization enables the network to learn complex mapping and exhibit better acceleration in converging towards the solution.
+In [34], the authors have proposed analytic LISTA (ALISTA), in which the weight matrices are precomputed through an optimization procedure, thereby eliminating the need for learning them from data.
+Only a small number of scalar parameters: the layer-wise step size, threshold, and momentum factor are learned.
+Building upon this framework, the subsequent HyperLISTA model dynamically adapts these parameters in response to the reconstruction error, retraining them for each instance to match the current problem state.
+This instance adaptive strategy yields a superlinear convergence order, a property that cannot be achieved when a uniform set of optimal parameters is learned across the entire dataset.
+Although theoretical studies have been attempted on the dependency of parameters on convergence order and stability of the algorithm for ISTA, there is a significant scope to explore theoretical studies on how other classical algorithms such as alternating direction of method of multipliers (ADMM), approximate message passing (AMP), etc., perform compared to their respective unfolded versions.
+For instance, channel estimation for integrated sensing and communication (ISAC) in intelligent transportation system is proposed in [215].
+An unfolded network for channel estimation based on ISTA is designed with trainable parameters, that includes the parameters related to dictionary based matrices in each ISTA iteration and the learnable threshold parameter used in the soft-thresholding operator.
+In this work, the authors theoretically analyze the proposed unfolded network and establish it achieves a linear convergence order.
+  2. 2.
+    Generalization error: A fundamental question for any learned model is its reconstruction ability.
+Especially, the degree to which it can accurately recover a parameter from incomplete or corrupted data.
+Generalization error or bounds give a probabilistic estimate on the difference between the true error (evaluated over the underlying data distribution) and the empirical error computed from the available training samples for a given hypothesis function.
+Thus, such bounds predict how well a learned model performs on unseen data and assess whether the training process has led to appropriate generalization [18].
+The underlying machine learning algorithm of a learned model can be viewed as stochastic transformation which maps training data to hypotheses.
+The algorithm stability largely depends on the generalization capability of learning algorithm, and thus is of theoretical interest [23].
+As described in [18], based on the training sequence $\mathcal{S}$ and given hypothesis space $\mathcal{H}$, a learning algorithm yields a function $h_{S}\in\mathcal{H}$.
+The function $h_{S}$ is defined as $h_{S}:\mathcal{X}\rightarrow\mathcal{Y}$ for a training sequence from a distribution $\mathcal{D}$ on $\mathcal{X}\times\mathcal{Y}$.
+Let us consider a training sequence $\mathcal{S}=\{({\mathbf{x}}_{i},{\mathbf{y}}_{i})\}_{i=1,\dots,m}$ with $m$ i.i.d.
+samples drawn from the distribution $\mathcal{D}$.
+Therefore, the empirical loss $\hat{\mathcal{L}}(h)$ of a hypothesis $h$ is the reconstruction error on the training sequence $\mathcal{S}$, as given below:
+    $\hat{\mathcal{L}}(h)=\frac{1}{m}\sum_{i=1}^{m}\ell(h,{\mathbf{x}}_{i},{\mathbf{y}}_{i}).$ (10)
+    The loss function $\ell$ measures the reconstruction error.
+Mean squared error (MSE) is the most commonly used loss function.
+The true loss $\mathcal{L}(h)$ defined as the risk of a hypothesis $h$ is computed by evaluating expectation, $\mathbb{E}[\cdot]$ over a distribution $\mathcal{D}$ as shown below:
+    $\mathcal{L}(h)=\mathbb{E}_{\mathbf{x},\mathbf{y}\sim\mathcal{D}}[\ell(h,\mathbf{x},\mathbf{y})].$ (11)
+    The generalization error of the hypothesis $h_{\mathcal{S}}$ is based on the training samples $\mathcal{S}$.
+It is defined as the difference between the empirical loss and the true loss as shown below:
+    $\text{GE}(h_{\mathcal{S}})=|\hat{\mathcal{L}}(h_{\mathcal{S}})-\mathcal{L}(h_{\mathcal{S}})|.$ (12)
+    In [23], the authors have identified several definitions to quantify the stability of a learning algorithm such as: hypothesis stability, point-wise hypothesis stability, error stability, and uniform stability.
+    The stability of a learning algorithm is quantified by the bound on the generalization error which must be tight.
+The trainable parameters in deep unfolding are trained via DL tools so it is imperative that algorithmic stability will be impacted by the choice of these parameters.
+Thus, the notion of algorithmic stability can offer a theoretical bridge between optimization and learning in deep unfolding.
+A proper parameterization is fundamental to the empirical performance and acceleration achieved during training of the network.
+Hence, to maintain convergence and algorithmic stability with robust generalization, the learnable parameters must be carefully designed.
+Consequently, exploring stability aware parameterization and training strategies presents a promising direction for advancing the theoretical foundation of deep unfolding method.
+    Moreover, in [18], the authors have discussed generalization error that relies on Rademacher complexity analysis.
+The result presented in [18], suggested that deep unfolding models may exhibit better generalization capabilities than purely data driven neural networks, in line with empirical results.
+For example, the generalization error bound for LISTA depends on the number of layers only logarithmically, whereas generalization error bound for purely data driven neural network scale exponentially in the number of layers.
+There have been works [178, 104] that discuss generalization error and Rademacher complexity, involving different learnable parameters and a variety of network architectures.
+  3. 3.
+    NMSE: The convergence rates of different approaches: classical, purely data driven and deep unfolding are examined through the NMSE with respect to the number of layers or iterations required to achieve the optimal solution.
+Some existing works that use NMSE as a performance metric are as follows, for channel estimation [85, 243, 203], detection [204], precoder design [209], ISAC [94], joint activity detection and channel estimation [172, 250, 139].
+Generally, NMSE evaluates the performance of any approach by evaluating the MSE between ground truth, $\mathbf{x}$ and estimated value, $\hat{\mathbf{x}}$ followed by normalization as given below:
+[…]
+    NMSE is used as a unified performance metric to compare classical iterative algorithms, purely data-driven models, and deep unfolding methods.
+It provides a direct measure of convergence speed in terms of the number of layers or iterations required to reach a target accuracy.
+The NMSE-vs-layers plots reported in [172, 85, 243, 250, 203, 209, 139, 204, 94] visually illustrate the convergence behavior of unfolded networks relative to benchmark algorithms.
+These results consistently show that NMSE decreases as the number of layers or iterations increases.
+Classical iterative methods exhibit slow NMSE reduction because they typically require many iterations to converge.
+In contrast, deep unfolding achieves the same NMSE performance with significantly fewer layers, outperforming both purely data-driven models and classical approaches [85, 172].
+    While NMSE provides a direct measure of reconstruction accuracy with respect to the number of layers, it does not capture characteristics such as the convergence order of the algorithm or its susceptibility to overfitting.
+These aspects are instead characterized by convergence order analysis and generalization error, which together describe the theoretical convergence behavior and algorithmic stability of deep unfolding methods.
+However, NMSE can be complemented with these analyses to obtain an understanding of the efficiency and the reliability of unfolded architectures.
+The existing theoretical studies have examined the convergence order [165, 33, 34] and generalization error [18, 23, 178] of unfolded networks.
+However, the authors in [167] have provided optimization guarantees showing that unfolded networks can achieve near-zero training loss as the number of learning epochs increases.
+Their analysis focuses on how the training loss decreases with more epochs while keeping the number of layers fixed.
+In their numerical results, they show that purely data-driven networks require significantly more parameters to achieve the same near-zero training loss compared to LISTA and ADMM-based compressive sensing networks.
+This highlights the parameter efficiency and training advantages of unfolded architectures.
 In the following sections, we provide a comprehensive survey of deep unfolding algorithms for a variety of critical wireless communication tasks.
 III Signal Detection
-This section discusses the application of deep unfolding techniques for signal detection. We segregate these signal detection techniques into various key technologies, as described in the following.
+This section discusses the applications of deep unfolding for signal detection. We segregate these signal detection techniques into various key technologies, as described in the following.
 III-A Massive MIMO
@@
 On the other hand, the SD exhibits ($\mathcal{O}(2^{K})$) for worst-case complexity.
+The FLOPs count for QPSK, 8PSK and 16 QAM constellation with varying transmitted symbol size, show that DetNet has lower FLOP count than those for semidefinite relaxation (SDR) based MIMO detection but has higher FLOP count than the AMP detector.
 However, the proposed network is not flexible to changes in the size of the constellation or the number of users.
@@
 OAMP-Net2 outperforms OAMP-Net by adapting to varying channel conditions.
+However, the robustness of the OAMP-Net2 under hardware impairment has not been verified in [73].
 The OAMP algorithm is unfolded, and the linear MMSE (LMMSE) channel estimator is used to design the receiver in [73].
@@
 These parameters in OAMP-Net2 help compensate for channel estimation errors, improving performance compared to DetNet [163].
-As the number of transmit antennas exceeds the number of receive antennas, low complexity detectors are required for such overloaded massive MIMO systems.
-In [50], the trainable PGD (TPGD) algorithm was introduced for such systems.
-TPGD uses a DL-aided iterative PGD method with trainable parameters, such as step-size, which are optimized via backpropagation and SGD.
-TPGD achieves lower computational cost with fewer trainable parameters compared to DetNet [49], OAMP-Net2 [10].
-Yun et al.
-in [51] enhanced the unfolded PGD design by adding momentum components to the iterative hidden layers.
+Furthermore, for overloaded massive MIMO systems, low-complexity detectors are essential.
+The trainable PGD (TPGD) algorithm [183] unfolds the PGD method with trainable parameters (step size) optimized via backpropagation and SGD.
+To further improve PGD with an increase in convergence speed, Yun et al.
+in [226] added momentum components to the iterative hidden layers.
 Momentum helps guide parameter updates in the right direction and reduces oscillations during PGD optimization [159].
-This modification improves convergence speed and overall performance, outperforming DetNet [49] and OAMP-Net2 [11].
-Furthermore, a trainable detector based on the Hubbard–Stratonovich transformation (to avoid matrix inversion) is discussed in [53].
-Compared to the TPGD detector [50] and OAMP-Net2 [11], the network in [53] achieves faster training and lower execution cost.
-In [54], the authors introduced the MIMO detection network (MMNet), an unfolded detector designed for ill-conditioned channels and non i.i.d.
-scenarios.
-MMNet is based on an iterative shrinkage-thresholding algorithm (ISTA), alternating between a linear detector and a nonlinear denoising step.
-The architectural flexibility and computational complexity are balanced, outperforming DetNet [49] and OAMP-Net [10].
-The MMNet outperforms the classical LMMSE detector with a performance gain of approximately $4-6$ dB at a symbol error rate (SER) of $10^{-3}$ for different channel conditions.
-In [55], the CONCRETE MAP method is proposed, based on a CONtinuous relaxation of the probability mass function for disCRETE random variables, as derived in [56].
-This approach solves the maximum a posteriori (MAP) detection problem using continuous optimization instead of exhaustive search, approximating the MAP solution for any probabilistic nonlinear model.
-By unfolding the gradient descent algorithm, the detector is able to generate soft outputs, making it suitable for integration with soft-input decoders.
-The network in [55] outperforms DetNet [49], OAMP-Net2 [11], and MMNet [54], in terms of BER under correlated channels.
+Both networks in [183, 226] achieve lower complexity and outperform DetNet [163] and OAMP-Net2 [71].
+A detector based on the Hubbard–Stratonovich transformation [182] avoids matrix inversion, offering faster training and lower execution cost compared to TPGD [183] and OAMP-Net2 [73].
+MMNet [101] unfolds ISTA, alternating linear detection and nonlinear denoising, designed for ill-conditioned and non i.i.d.
+channels.
+MMNet balances complexity and flexibility, outperforming DetNet [163], OAMP-Net [71].
+The MMNet outperforms the classical LMMSE detector with a performance gain of approximately 464-6 dB at a symbol error rate (SER) of $10^{-3}$ for different channel conditions.
+The CONCRETE MAP method in [17] solves the maximum a posteriori (MAP) detection problem via CONtinuous relaxation of disCRETE random variables, approximating MAP solutions using gradient descent unfolding.
+It generates soft outputs for integration with soft-input decoders and outperforms DetNet [163], OAMP-Net2 [73], and MMNet [101] in BER under correlated channels.
 In [204], conjugate gradient descent (CGD) iterations are unfolded to design the network, LcgNet.
 The authors also proposed a quantized version, QLcgNet, which minimizes memory usage with negligible performance degradation.
-LcgNet outperforms DetNet [49] in terms of SER, computational complexity, and memory cost, making it a robust solution for massive MIMO systems.
-To enhance LcgNet, Olutayo et al.
-[58] incorporated a preconditioner during training, designing a preconditioned LcgNet.
+LcgNet surpasses DetNet [163] in SER, complexity, and memory cost, making it suitable for massive MIMO system.
+Olutayo et al.
+[154] improved LcgNet by adding a preconditioner during training, enhancing convergence speed by optimizing the receiver filter’s spectrum.
 Preconditioning improves the condition number of a matrix [30], optimizing the receiver filter’s spectrum.
 This enhancement significantly increases the convergence speed compared to the original LcgNet in [204].
-Additionally, the dynamic conjugate gradient network utilizes self-supervised learning with forward error correction codes, effectively adapting to non-stationary, time-varying channels.
-Physical layer signal processing involves continuous-time analog waveforms, which must be quantized for hardware implementation.
-Low-resolution (LR) one-bit analog-to-digital converters (ADC) are commonly used at radio frequency (RF) ends due to their high sampling rates, low cost, and low energy consumption [60].
-To recover high-dimensional signals from noisy measurements, Khobahi et al.
-[61] developed an LR detection network (LoRDNet).
-This network extracts signals from noisy one-bit measurements with unknown CSI.
-It is based on unfolding the PGD method for symbol detection using the ML estimator.
-The channel effects are embedded in the training data, so LoRDNet implicitly learns a representation of the channel behavior, eliminating the need for CSI.
-The network outperforms classical ML estimators [62] and DL networks like DeepSIC [63], even with small datasets of about 500 samples.
-In [64], the authors proposed a generalized PGD algorithm.
-The generalized PGD algorithm executes multiple gradient descent steps before applying the projection, allowing for a better balance between convergence speed and detection performance.
-The authors introduced the self-correcting auto detector (SAD) by unfolding the generalized PGD algorithm.
-The architecture has two modules: the self-correcting module that executes multiple gradient descent steps to refine the detection estimate, and a denoising autoencoder, a learnable projection module.
-The numerical results demonstrate that SAD [64] achieves a better trade-off between detection accuracy and complexity compared to DetNet [49].
-Although DetNet requires $15$ projection steps, SAD only performs $4$, significantly reducing detection latency and computational cost.
-The channel matrix for a MIMO fading channel can be modeled as a graph.
-Graph models, such as Bayesian belief networks, Markov random fields, and factor graphs, combine the posterior probability of transmitted data variables and graph theory, establishing a relationship between different nodes for proper detection [3].
-The performance of these methods largely depends on the underlying graph structure and convergence properties.
-Iterative algorithms, like message-passing detectors (MPD) with belief propagation (BP), approximate message passing (AMP), and channel hardening-exploiting message passing (CHEMP), further reduce computational complexity.
-Expectation propagation (EP), as an extension of BP, enables efficient detection in higher-dimensional MIMO systems and large constellations.
-EP approximates the posterior distribution of transmitted symbols iteratively, enabling efficient detection in polynomial time [65, 66].
-Tan et al.
-[67] unfolded MPD iterations to update messages at symbol nodes using prior and posterior log-likelihood ratio (LLR).
-Learnable parameters, such as prior probabilities and correction factors (damping, scaling factor, LLR rescaling), are optimized via mini-batch SGD.
-The authors unfolded the existing classical MPDs: BP and CHEMP, where the correction factors are learned instead of heuristic tuning.
-A greedy search determines the number of layers.
-The unfolded networks adapt better to correlated channels and perform consistently in higher-order modulation than classical MPDs.
-In [68], a low-complexity AMP-graph NN (GNN) model (AMP-GNN) is proposed, which outperforms OAMP-Net2 [11] by avoiding matrix inversion.
-Due to the high computational complexity, message-passing algorithms are generally unsuitable for single-carrier frequency division multiple access (FDMA) systems, where a low peak-to-average power ratio is a key design objective for efficient transmission.
-To address this, the unfolded network in [69] simplifies generalized AMP iterations as discussed in [70].
-The proposed network simplified generalized AMP iterations by approximating per-subcarrier channel statistics through statistical averaging.
-The unfolded network [69] maintains high performance even with channel mismatches and achieves $0.83$ dB at BER$=10^{-3}$ gain over classical generalized AMP [70].
-To enhance EP-based detection, the authors in [71] proposed a network, a modified EP MIMO detector (MEPD).
-In the network, parameter selection is optimized using a modified moment-matching scheme and shows superior performance in high modulation cases compared to OAMP-Net2 [11].
-In [72], an approximate EP inspired by [73] is unfolded, achieving better detection performance and robust convergence.
+A dynamic conjugate gradient network further adapts to time-varying channels using self-supervised learning with forward error correction codes.
+Physical layer signal processing handles continuous-time analog waveforms, which must be quantized for hardware implementation.
+Low-resolution (LR) one-bit analog-to-digital converters (ADC) are widely used at radio frequency (RF) ends for high sampling rates, low cost, and low power [7].
+Khobahi Khobahi et al.
+[102] developed an LR detection network (LoRDNet), which unfolds the PGD method for symbol detection using the ML estimator, learning channel behavior implicitly from training data without requiring CSI.
+LoRDNet outperforms classical ML [39] and DL networks like DeepSIC [175], even with small datasets ($\sim 500$ samples).
+A generalized PGD algorithm was proposed in [77], where multiple gradient steps are performed before projection to balance convergence and performance.
+The self-correcting auto detector (SAD) unfolds the PGD algorithm, by combining a self-correcting gradient module and a learnable denoising autoencoder projection.
+SAD [77] achieves better detection-complexity trade-off than DetNet [163], using only $4$ projection steps versus DetNet’s $15$ projection steps, significantly reducing latency and computational cost.
+[…]
 To address high computational complexity of SD, Nguyen et al.
@@
 The unfolded detector achieves better performance compared to DetNet [163], OAMP-Net2 [73], TPGD-detector [183], MEPD [28], EP-Net [110], and other methods like [114].
-TABLE III: Summary of typical deep unfolding models for signal detection and their advantages and disadvantages.
-Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
-DetNet [49] First deep unfolding model for MIMO signal detection based on unrolling PGD method.
-In a $60\times 30$ MIMO system, DetNet achieves $3$ dB SNR gain over ZF for BER $=10^{-3}$, performs comparable to SDR and AMP, and approaches SD.
-Advantages: Explainable (follows PGD optimization flow), performs well when $N_{r}\gg N_{t}$.
-Disadvantages: Complex architecture (high computational cost), poor generalization under correlated channels.
-LcgNet [57] Unfolded CGD for massive MIMO detection.
-Achieves $1.52$ dB SNR gain over LMMSE in $32\times 128$ MIMO at BER $=10^{-4}$.
-Outperforms DetNet with lower complexity and faster training.
-Advantages: Fast training (practical for low-end CPUs).
-Disadvantages: Trained on Rayleigh fading, suffers significant BER degradation when tested on a spatially correlated channel.
-OAMP-Net2 [11] Lightweight unfolding of OAMP with soft input/output.
-Outperforms DetNet and AMP in 4$\times$4 and 8$\times$8 MIMO system, robust under imperfect CSI.
-Advantages: Fast training (4 parameters).
-Supports turbo equalization (soft I/O).
-Disadvantages: Underperforms in 3GPP channels (assumes unitarily-invariant matrix).
-MMNet [54] Online-trained unfolded ISTA for 3GPP channels.
-Achieves $8$ dB SNR gain over MMSE for SER $=10^{-3}$ (and $2.5$ dB over OAMP-Net2 [11] under imperfect CSI).
-Advantages: Robust to correlated channels, high-order QAM, and estimation errors.
-Disadvantages: Online training adds latency (unsuitable for fast-varying channels).
-ADMM-Net [38] Inexact ADMM-based unfolding with soft $tanh$ demodulation (supports $16/64$ QAM).
-Achieves SER within $1.2$ dB of SD in $30\times 60$ MIMO, outperforms DetNet with lower complexity.
-Advantages: Robust to channel correlation and imperfect CSI.
-Disadvantages: More layers (5205205–205 – 20) improve accuracy but increase computation, hence a tradeoff between performance and runtime.
-LoRDNet [61] Unfolding PGD for 1-bit ADC receivers without CSI.
-Outperforms classical ML detectors with just 500 samples.
-Advantages: Works with low-resolution ADCs; data-efficient.
-Disadvantages: Periodic retraining required for block fading channels (reduced SE due to pilot overhead) .
+TABLE III: Comparison of Classical vs Unfolded methods for Signal Detection.
+$T$: Layers, $I$: Iterations, $M$: Symbol size, $N_{t}$: No. of transmit antenna, $N_{r}$: No. of receive antenna.
+Unfolded Computational Complexity Number of Parameters (Unfolded) Convergence speed (Number of layers)
+Model Classical Unfolded Simulation Setup Classical Unfolded
+DetNet [163] Sphere Decoder: $\mathcal{O}(M^{N_{t}})$ $\mathcal{O}(TN_{t}^{2})$ $(6N_{r}N_{t}+2N_{r}+N_{r})T$ QPSK, 8-PSK, 16-QAM, $0.55$-Toeplitz channel, $N_{t}=\{15,20,30\}$, $N_{r}=\{20,25,30\}$.
+$\sim 1000$ iterations $30$ Layers
+LcgNet [204] CGD: $\mathcal{O}(I(8N_{t}^{2}+14N_{t}+8))$ $\mathcal{O}(T(4N_{t}^{2}+6N_{t}))$ $(6N_{t}+4N_{t}^{2}+4T)$ MIMO $N_{r}=32$, $N_{t}=64$, Rayleigh fading channel.
+$\sim 40$ iterations 77 Layers
+OAMP-Net2 [73] OAMP: $\mathcal{O}(N_{t}^{3})$ $\mathcal{O}(TN_{t}^{3})$ $4T$ MIMO Rayleigh fading channel, QPSK and 16-QAM, $N_{t}=N_{r}=4$ and $N_{t}=N_{r}=8$. $1000$ iterations 44 Layers
+MMNet [101] MMSE: $\mathcal{O}(N_{r}^{3})$ $\mathcal{O}(N_{r}^{3})$ $2N_{t}(N_{r}+1)$ Massive MIMO, QAM modulation, 3GPP 3D channel, $N_{t}=\{16,32\}$, $N_{r}=64$. $50$ iterations $10$ Layers
+ADMM-Net [103] ADMM: $\mathcal{O}(M^{3})$ $\mathcal{O}(M^{2})$ $5T$ MIMO Rayleigh fading channel, BPSK, 4-QAM and 16-QAM, $N_{t}=N_{r}=50$. $30$ iterations <20<20 Layers
 To avoid the computational burden incurred due to matrix inversion, the authors in [45] have proposed a sparse refinement architecture for symbol detection in uplink massive MIMO.
@@
 [114] incorporated interference cancellation into DNN layers.
-Inspired by [79], their sequential detector trains residual error coefficients with mean square error (MSE) loss, reducing complexity by avoiding matrix inversion and relying on matrix addition and multiplication.
+Inspired by [137], their sequential detector trains residual error coefficients with MSE loss, reducing complexity by avoiding matrix inversion and relying on matrix addition and multiplication.
 Compared to DetNet [163], which requires $90$ layers for convergence, their model achieves convergence with just $8$ layers, offering significant computational simplicity.
@@
 Moreover, the complexity of the proposed network is reduced as compared to OAMP-Net2 [73].
+The authors in [9] have proposed an unfolded coherent detector for optical communication.
+They have developed a MIMO filter architecture to compensate for hardware impairments such as transmitter and receiver in-phase/quadrature skew.
+The learnable parameters include the filter coefficients of the MIMO filter and are trained using the method in [183].
+The results showed that the unfolded network performed well in the presence of chromatic dispersion, polarization rotation, and frequency offset accumulated over the fiber channel.
 III-B Modern Multicarrier Modulation
 A cellular network must simultaneously support a growing number of active subscribers within a limited time-frequency spectrum [96].
-The current trend in wireless communication is to increase the signal bandwidth to achieve higher transmission rates.
-However, increasing bandwidth results in shorter symbol periods due to the inverse relationship between bandwidth and symbol period.
-Consequently, in a multi-path fading channel, the reduced symbol period exacerbates ISI.
-Multi-carrier modulation techniques like OFDM provide an effective solution to mitigate ISI by dividing a wideband signal into a set of orthogonal narrowband subcarriers [83].
-MIMO-OFDM detectors face many challenges due to the complexity introduced by high dimensional signal space.
-A low complexity receiver for a cyclic prefix-free MIMO-OFDM system is explored in [84].
-The receiver unfolds the OAMP algorithm and scalable parameters are introduced for optimizing the channel estimate matrix.
+By increasing bandwidth the symbol period is reduced which worsens the ISI in multi-path fading channels.
+OFDM mitigates ISI by splitting wideband signals into orthogonal narrowband subcarriers [208].
+MIMO-OFDM detectors faces challenges due to high-dimensional signal space.
+An unfolded OAMP-based low-complexity receiver for cyclic prefix-free MIMO-OFDM is explored in [234].
 The network was tested in over the air experiments and it performed reliably with low synchronization errors under imperfect CSI.
 Similarly, detection of coarsely quantized or LR signals in one-bit OFDM-based MIMO systems is a challenge.
-Hence, to address this challenge Shao et al.
-[85] have proposed an unfolded expectation maximization (EM) algorithm.
 One-bit ADCs severely quantize received signals, discarding amplitude information, and hence, ML detection becomes intractable due to nonlinear quantization effects.
-Hence, EM iteratively estimates unquantized symbols by computing its conditional expectation and optimizes the parameters in the maximization step.
-At high SNR, unfolded EM [85] performs better than ZF for one bit MIMO-OFDM detection.
-Furthermore, by unfolding the likelihood ascent search algorithm, Ullah et al.
-[86] developed an unfolded network to jointly detect and estimate the soft output in MIMO-OFDM systems.
+Shao et al.
+[168] proposed an unfolded expectation maximization (EM) algorithm that iteratively estimates unquantized symbols, outperforming ZF detection at high SNR.
+Furthermore, Ullah et al.
+[191] unfolded a likelihood ascent search algorithm for joint detection and soft-output estimation.
 The likelihood ascent search algorithm is a neighborhood search-based detector for the MIMO system.
@@
 Zhang et al.
-developed an unfolding-based framework for MIMO-OFDM detectors in [87].
-The EP is unfolded with layer-specific damping factors that are trained to adjust effectively to varying channel conditions.
-The proposed network achieved near-optimal detection across a wide range of channels and noise levels and demonstrated robustness in over-the-air experiment tests.
-Underwater acoustic communication is widely used for long-range transmissions, reaching up to tens of kilometers [88, 89, 90].
-Sound waves are used to transmit data through water since radio and light waves attenuate rapidly underwater.
-Channel estimation in underwater acoustic communication is particularly challenging due to doubly-selective channels that exhibit both severe time variation (from Doppler effects) and frequency selectivity (from long multipath delays).
-Hence, for OFDM systems in complex doubly-selective underwater acoustic channels, Zhao et al.
-in [35] proposed an unfolded equalizer (UDNet).
-The unfolded algorithm is based on unfolding the classical MMSE equalizer.
+[235] unfolded expectation propagation (EP) with trainable layer-specific damping factors, providing robust detection under varying channels and noise, validated in over-the-air tests.
+In underwater acoustic communication, where radio or light waves fail, Zhao et al.
+in [238] proposed UDNet, an unfolded MMSE equalizer that generalizes well with low complexity in doubly-selective channels.
 In terms of performance, UDNet achieved better generalization and maintained lower computational complexity than DetNet [163].
 Li et al.
-[34] proposed an unfolded EP network (EP-Net) for OTFS detection.
-The trainable parameters include the step-size, damping factor, and offset values for the mean and variance of the posterior conditional probability distribution.
-To enhance convergence and prevent the network from being trapped in local optima, the Adam optimizer with warm restarts is employed.
+[110] proposed EP-Net for OTFS detection, training step-size, damping, and posterior distribution parameters to improve detection performance.
 In [225], an unfolded OAMP-based detector for OTFS is designed to handle channel estimation errors.
@@
 By transforming iterative algorithms such as gradient descent variants, ISTA, ADMM, and AMP into structured networks, unfolding approaches bridge interpretability and performance.
-Table III summarizes the novelties, performance gains, advantages and disadvantages of the most popular deep unfolding models for signal detection.
+In Table III, we summarize the computational complexity, number of trainable parameters and convergence speed of the most popular deep unfolding models against those of classical methods for signal detection.
 While pioneering unfolding architectures such as DetNet [163] and LcgNet [204] have demonstrated the advantages of data-driven signal detection with explainable DNN structures, they exhibit limited generalization in challenging scenarios, such as correlated channels.
@@
 The results demonstrate that deep unfolding techniques in signal detection can achieve robust BER performance in practical indoor environments, outperforming both traditional and purely data-driven methods.
+Although the aforementioned unfolded detectors achieved improved performance with reduced complexity under both perfect and imperfect CSI, their robustness over practical hardware and synchronization impairments have not been investigated.
+The performance analysis in this direction for different unfolded detectors are worth investigating in future works.
 IV Channel Estimation
@@
 This iterative approach improves CSI recovery accuracy, making the process more efficient and adaptable to changing conditions.
-In massive MIMO systems, where the channel matrix is often sparse, the majorization-minimization (MM) approach is effective for addressing sparse channel estimation [102].
-In [103], the authors address the challenge of downlink channel estimation in massive MIMO.
-A two-timescale joint uplink/downlink dictionary learning and channel estimation algorithm is proposed.
-The short-term subproblems focuses on estimating sparse channel vectors for each timeslot.
-The MM algorithm is unfolded, and the sparse channel vectors are treated as trainable parameters.
-In contrast, the long-term subproblem involves updating the dictionary to adapt to dynamic environments for which a constrained stochastic successive convex approximation (SSCA) framework is employed without deep unfolding.
-Simulations results in [103] demonstrate that the proposed approach reduces pilot overhead and outperforms existing methods in both static and dynamic environments, particularly in environments with mobile users.
-The matching pursuit algorithm is a greedy iterative method often applied in sparse channel scenarios like massive MIMO.
-It identifies channel components (basis functions) by matching them to the residual, which is the difference between the observed signal and the signal generated by the current channel estimate.
+In massive MIMO, where channels exhibit sparsity, majorization-minimization (MM) is effective for sparse channel estimation [122].
+In [242], the authors propose a two-timescale joint uplink/downlink dictionary learning and channel estimation framework.
+The short-term subproblem estimates sparse channel vectors by unfolding the MM algorithm and treating these vectors as trainable parameters.
+The long-term subproblem updates the dictionary using a constrained stochastic successive convex approximation (SSCA) method without unfolding.
+Simulation results show reduced pilot overhead and superior performance in both static and dynamic scenarios, especially with mobile users.
+The matching pursuit algorithm is a greedy method used for sparse channel estimation, iteratively identifying channel components by correlating the residual with dictionary elements.
 Yassine et al.
-[104] introduced a deep unfolded algorithm, called mpNet, based on the classical matching pursuit algorithm.
-This network estimates the multiuser massive MIMO channel in an unsupervised manner, initializing with a dictionary of imperfect steering vectors.
-The system operates in time division duplex mode to leverage channel reciprocity.
-The network dynamically adjusts the number of layers based on incoming data, ensuring efficient performance in realistic environments.
+[219] proposed mpNet, a deep unfolded matching pursuit based network for unsupervised multiuser massive MIMO channel estimation, initialized with a dictionary of imperfect steering vectors.
+The network dynamically adjusts its number of layers based on the received data, providing efficient estimation in realistic settings.
+Building upon the unfolded network: mpNet [219], Chatelier et al.
+[27] have considered hardware impairments in the shifting of the subcarrier frequency to achieve SISO-OFDM channel estimation.
 The basis pursuit algorithm is a widely used method for sparse signal recovery, particularly in compressed sensing.
@@
 The sparse Bayesian learning algorithm is unfolded with trainable parameters grid points and off-grid variables.
-Simulation results show that the proposed method significantly outperforms fixed-depth models, achieving a 30% reduction in the number of layers while improving normalized MSE (NMSE) performance.
+Simulation results show that the proposed method significantly outperforms fixed-depth models, achieving a 30% reduction in the number of layers while improving NMSE performance.
 IV-B mmWave Massive MIMO
@@
 In this approach, the step-size of the descent step is treated as a trainable parameter.
-In [116], an iterative trimmed-ridge regression algorithm is unfolded to improve beamspace channel estimation.
-The proposed algorithm involves parallel computation to achieve negligible latency.
-These algorithms were applied to both narrowband and wideband massive MIMO systems, to ensure versatility in channel estimation.
+In [210], an unfolded iterative trimmed ridge regression algorithm is proposed for beamspace channel estimation.
+The method supports parallel computation for near zero latency and is applicable to both narrowband and wideband massive MIMO systems, ensuring versatile channel estimation performance.
 IV-C Modern Multicarrier Modulation
 Yiyun et al.
-[117] proposed a CSI recovery network for FDD downlink massive MIMO-OFDM systems.
-ISTA is unfolded and the network consists of compression and reconstruction modules for CSI compression and feedback.
-Building on ISTA, Hu et al.
-[118] introduced an unfolded algorithm called the learnable optimization and regularization algorithm (LORA) to enhance the precision of CSI feedback.
-This approach also unfolds ISTA, and the regularization term is learned through end-to-end training.
-The model captures CSI characteristics and adapts to different levels of channel estimation error.
-Furthermore, to improve CSI feedback accuracy and efficiency, Yangyang et al.
-[119] proposed an unfolded ISTA network.
-Their study showed that the real and imaginary components of the channel differential terms, i.e, the difference between CSIs at adjacent time slots, exhibit similar correlation patterns in both the angular and delay domains.
-This property enables a shared model to process both components, thereby reducing network complexity and enhancing the efficiency of the CSI encoder.
+[202] proposed a CSI recovery network for FDD downlink massive MIMO-OFDM by unfolding ISTA, with compression and reconstruction modules for CSI feedback.
+Hu et al.
+[86] improved it with learnable optimization and regularization algorithm (LORA), by learning the regularization term via end-to-end training to adapt to different channel estimation errors.
+In [237], Yangyang et al.
+exploited the correlation between real and imaginary parts of CSI differential terms, enabling a shared model to reduce complexity and improve encoder efficiency.
 Each of the networks proposed in [202, 86, 237] enhance CSI feedback performance for FDD massive MIMO-OFDM system.
-In [33], Guo et al.
-proposed a two-stage low-rank CSI feedback scheme for a low-rank OFDM mmWave channel.
-In massive MIMO systems operating in FDD mode, CSI is estimated by unfolding the fast ISTA.
-Fast ISTA is an improved version of the classical ISTA due to the introduction of momentum term that accelerates the convergence rate.
-The network, FISTA-Net, includes trainable step-size and thresholds to achieve accurate reconstruction performance.
-In [120], the authors explore quantized compressed CSI at the user equipment before feedback for massive MIMO-OFDM system.
-They propose a deep unfolding-based bit-level CSI feedback network, TiLISTA, to mitigate the effects of accumulated quantization errors.
+Guo et al.
+[63] proposed FISTA-Net, unfolding fast ISTA with momentum, trainable step-sizes, and thresholds for faster convergence and accurate low-rank OFDM mmWave CSI feedback.
+For quantized CSI feedback, the authors in [25] introduced TiLISTA, a deep unfolding bit-level feedback network using tied learned ISTA, reducing learnable parameters and enhancing robustness to quantization errors.
 Building upon ISTA, they used tied learned ISTA as discussed in [124].
@@
 FISTA-Net [63] focuses on faster convergence while TiLISTA [25] targets robustness to quantization errors.
-In [122], an unfolded sparse Bayesian learning algorithm is derived to capture channel sparsity in a mmWave massive MIMO-OFDM system with hybrid beamforming (HBF).
-The proposed network achieved a lower NMSE of $0.033$ outperforming the classical sparse Bayesian learning algorithm which recorded an NMSE of $0.1203$.
+In [54], an unfolded sparse Bayesian learning method is proposed for mmWave massive MIMO-OFDM with hybrid beamforming, effectively capturing channel sparsity and reducing NMSE from $0.1203$ to $0.033$.
 Jianqiao et al.
-[123] proposed an unfolded inverse-free variational Bayesian learning framework to address the high computational complexity of matrix inversion in mmWave MIMO-OFDM systems.
-The sparse channel recovery performance is enhanced by learning the a priori parameters of the sparsely represented dictionary for channel estimation.
-In [124], based on approximate Bayesian inference and an extension of expectation-consistent approximation, the authors proposed a generalized expectation consistent algorithm, an iterative message-passing approach for sparse channel estimation.
-In a lens-based beamspace massive MIMO-OFDM system, this approach compensates for large channel attenuation in the mmWave band.
-The authors developed a deep unfolded unsupervised learning network called the learned denoising-based generalized expectation consistent signal recovery network.
-The generalized expectation consistent signal recovery algorithm as discussed in [125] is unfolded.
-Both the networks in [123] and [124] provide accurate channel recovery estimates for mmWave channel with few RF chains.
-In [126], the authors leveraged angle-domain channel sparsity to reduce uplink pilot overhead during channel estimation for mmWave MIMO-OFDM systems.
-They propose an unfolded multiple-measurement-vectors-based AMP network.
-The phase shift network and channel estimator are jointly trained, thus integrating the channel sparsity for improved performance.
+[29] introduced an unfolded inverse-free variational Bayesian learning framework to avoid costly matrix inversions, improving sparse channel recovery by learning the prior parameters of the dictionary.
+In [70], a generalized expectation consistent algorithm is unfolded for sparse channel estimation in lens based beamspace massive MIMO-OFDM, mitigating severe mmWave attenuation.
+In [29] and [70] accurate channel recovery with few RF chains is achieved.
+Additionally, in [133] the author proposed an unfolded multiple measurement vector AMP network that exploits angle-domain sparsity to reduce uplink pilot overhead, jointly training the phase-shift network and channel estimator for improved estimation performance.
 Gao et al.
@@
 At BER= $3\times 10^{-3}$, CENet achieved $\sim 5$ dB SNR gain over LS estimation algorithm.
-TABLE IV: Summary of typical deep unfolding models for channel estimation and their advantages and disadvantages.
-Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
-LDAMP [113] Denoising-based unfolded AMP model for sparse-beamspace channel estimation in mmWave massive MIMO.
-Outperforms classical denoising AMP method with small number of RF chains at the receiver.
-Advantages: Simple architecture and fast convergence (within $5$ layers), require a few RF chains.
-Disadvantages: Fixed parameters post-training restrict the model’s adaptability, performance may degrade in varying scenarios.
-LampResNet [114] Combines unfolded AMP with deep residual learning to enhance robustness against noise.
-Achieves close performance to LDAMP [113] at high SNR regime ($25$ dB).
-Advantages: Resilient to noise, consistent NMSE improvement under SNR variations.
-Disadvantages: Rely on prior information of beamspace channel, not consistent with realistic channel models, and network may fail to converge.
-FISTA-Net [33] Fast ISTA-based deep unfolding for low-rank CSI feedback in mmWave OFDM systems.
-Outperforms the classical FISTA with lower complexity.
-Advantages: Fast convergence ($20$ layers).
-Disadvantages: Requires retraining for major environment variations.
-mpNet [104] Deep unfold model based on the matching pursuit method for multiuser massive MIMO systems.
-Outperforms classical LS estimator in terms of NMSE under varying SNR.
-Advantages: Avoids ground-truth channel data by using unsupervised online learning on noisy LS estimates.
-Disadvantages: Need prior knowledge of physical wave propagation model, as small uncertainty of system parameters lead to high estimation performance loss.
-LORA [118] The deep unfolding model based on ISTA introduces a learnable regularization module to adaptively capture the CSI structure.
-Improved CSI reconstruction accuracy by $>2$ dB NMSE over TiLISTA [117].
-Advantages: LORA exhibits strong denoising capability under channel estimation error.
-Disadvantages: The number of layers is critical: using fewer than four layers degrades reconstruction accuracy, while increasing the number beyond four leads to higher computational complexity with only marginal improvements in NMSE.
 IV-D IRS
@@
 Tsai et al.
-[131] proposed a two-stage learned AMP network with row compression to address the joint estimation problem of direct and cascaded channels in IRS-aided mmWave systems.
-By utilizing the deep unfolding technique, the network combines the advantages of compressive sensing and DL.
-This approach enables reliable channel estimation with reduced training, which is crucial for efficiently reconfiguring wireless environments under IRS reflection control.
+[190] proposed a two-stage learned AMP network with row compression for joint estimation of direct and cascaded channels in IRS-aided mmWave systems.
+Using deep unfolding, the method integrates compressive sensing with learning to achieve reliable channel estimation with reduced training overhead, facilitating efficient IRS reconfiguration.
 IV-E IoT devices
@@
 The network uses four trainable scalar parameters to estimate activity probability, channel variance, regulate the denoiser, and control the state evolution.
-AMP-Net achieved accurate detection of the active device by extracting the CSI from the estimated device state matrix with a $4$ dB gain over classical AMP at SNR$=15$ dB.
+AMP-Net achieved accurate detection of the active device by extracting the CSI from the estimated device state matrix with a $4$ dB gain over classical AMP at SNR=15=15 dB.
 The authors in [111] proposed a grant-free access protocol for mmWave cell-free massive MIMO using a hybrid structure based on learned vector AMP, reducing reliance on prior channel knowledge and improving convergence over classical AMP.
@@
 A proximal gradient-based method is unfolded and it achieves better convergence than the network in [172].
+TABLE IV: Comparison of Classical vs Unfolded methods for Channel Estimation.
+$T$: Layer, $N_{f}$: No.
+of subcarriers, $N$: No.
+of devices, $M$: No.
+of BS antenna, $m$: Size of transmitted signal vector, $n$: Size of measurement vector, $L$: Pilot length, $G$: No.
+of discrete angular domain grid of channel clusters.
+Unfolded Computational Complexity Number of Parameters (Unfolded) Convergence speed (Number of layers)
+Model Classical Unfolded Simulation Setup Classical Unfolded
+LISTA-Net [32] ISTA: $\mathcal{O}(nm)$ $\mathcal{O}(Tnm)$ $n\times n+2T$ MIMO, $n=250$, $m=500$ 2020 iterations $14$ Layers
+ADMM-Net [44] ADMM: $\mathcal{O}(N^{2}M)$ $\mathcal{O}(LNM)$ $3T$ Massive MIMO, $L=90$, $M=4$, $N=300$ $40$ iterations 2020 Layers
+ComNet [56] LMMSE: $1.6$ mega FLOPs $0.37$ mega FLOPs $1.2$ MBytes OFDM, $N_{f}=64$, $64$-QAM, WINNER II channel model. $200$ iterations 55 Layers
+SBL-CE-Net [54] Sparse Bayesian learning: $0.6115$ tera FLOPs $0.0189$ tera FLOPs $TG^{2}N_{f}$ MIMO, $N_{f}=8$, $N_{t}=N_{r}=12$, $G=48$ $100$ iterations 33 Layers
 IV-F Discussions and Takeaways
@@
 Deep unfolding addresses these issues by embedding structural priors into learning pipelines, as demonstrated by models like LDAMP [72] and LampResNet [206].
-Table  IV presents notable unfolded networks for channel estimation with their performance gains, advantages, and disadvantages.
+Table IV summarizes the computational complexity, number of trainable parameters, and convergence speed (in terms of layers required for convergence) of key unfolded algorithms for channel estimation.
+Additionally, studies reporting complexity in FLOPs are included to highlight the implementation cost.
 Nonetheless, deep unfolding for channel estimation remains sensitive to model assumptions (e.g., sparsity or statistical independence).
@@
 Additionally, it achieves a high weighted sum rate in fully loaded scenarios, where the number of users matches the number of transmit antennas at the BS.
-In [146], learnable parameters are used to approximate matrix multiplication and inversion in WMMSE update rules.
-However, the network fails to exploit the inherent graph structure of wireless networks.
-This limitation is addressed in [148].
-An unfolded WMMSE algorithm for transceiver design in multicell multi-user MIMO (MU-MIMO) interference channels with local CSI is proposed in [148].
-The authors use graph convolutional networks (a variant of GNN) to unfold WMMSE, enabling coordination in multicell MU-MIMO systems.
-Graph filters and convolutional networks approximate WMMSE-related variables, such as weighted beamformer matrices, by leveraging graph learning techniques.
-However, the approach in [148] lacks flexibility in learning the exact transformation of WMMSE variables.
-To overcome this, [149] introduces a hybrid unfolded WMMSE algorithm for transmit beamforming in MU-MIMO ad-hoc networks.
-This method leverages GNNs to learn functional transformations of WMMSE parameters.
-Compared to the classical WMMSE algorithm, the proposed network reduces the number of layers, leading to faster convergence of the optimization problem.
-In [150], the joint optimization of precoders for multicell MU-MIMO systems is studied.
-The WMMSE algorithm is unfolded to enable distributed and synchronized precoding across cooperative BSs.
-The methods in [146, 147] rely on centralized optimization, leading to high complexity and backhaul overhead.
-In contrast, due to the distributed precoding, the proposed network in [150] reduces computational cost while maintaining efficiency.
-An unfolded WMMSE-based NN for massive MU-MIMO under imperfect CSI is studied in [151].
-The proposed network enhances performance as compared to [144], with low complexity by leveraging CSI statistics for near-optimal precoder solutions, thus accelerating convergence.
+In [83], learnable parameters approximate matrix operations in WMMSE, but the method does not exploit the graph structure of wireless networks.
+To address this, in [166] the authors unfolded WMMSE using graph convolutional networks for multicell MU-MIMO, enabling coordination with only local CSI, though with limited flexibility in learning exact transformations.
+In [41], the authors improved this by proposing a hybrid unfolded WMMSE that uses GNNs to learn functional mappings, reducing the number of layers and accelerating convergence.
+Distributed precoding across cooperative BSs was achieved in [221], lowering complexity and backhaul load compared to centralized approaches [83, 156].
+Finally, in [195] unfolded WMMSE is applied to massive MU-MIMO with imperfect CSI by leveraging CSI statistics, achieving faster convergence and near-optimal precoding relative to the classical WMMSE algorithm [170].
 Shi et al.
@@
 The authors claim that their proposed network matches the performance of the MSER-based iterative gradient descent algorithm while significantly reducing complexity.
-Gradient-based methods have been successfully applied to beamforming optimization with reduced complexity.
-However, these approaches can be further improved using advanced optimization frameworks like PGD or projected gradient ascent (PGA).
-A robust frequency-selective precoding algorithm that maximizes the expected weighted sum rate while considering per-antenna power constraints is developed for wideband MIMO-OFDM system in [153].
-This frequency-selective precoding algorithm is then unfolded into a deep network, with the aid of PGD.
-The trainable parameters include step sizes and projections in each PGD layer.
-The proposed network achieves near-optimal sum rate with lower computational complexity.
+Gradient based methods are widely used for beamforming optimization due to their reduced complexity, and further gains have been achieved using advanced frameworks such as PGD and projected gradient ascent (PGA).
+In [214], a frequency selective precoding algorithm for wideband MIMO-OFDM was unfolded using PGD with trainable step sizes and projection parameters, achieving near-optimal sum rate at low complexity.
 Zhu et al.
-[154] proposed a beamforming learning architecture by unfolding a parallel gradient projection algorithm for MISO heterogeneous systems.
-The trainable set includes the step-size for the gradient ascent update and weighted sum rate coefficients used to construct the beamforming gradient direction.
-The proposed network generalizes well across different numbers of antennas, different numbers of BSs, and different network sizes especially without retraining.
-PGA is a variant of the gradient ascent algorithm.
-It updates the solution in the gradient direction to maximize the objective function and then projects it back onto the feasible set.
+[248] unfolded a parallel gradient projection method for MISO heterogeneous networks, where trainable step sizes and weighted sum-rate coefficients enable generalization across antenna and network configurations.
+PGA itself performs gradient ascent followed by projection onto the feasible set.
 Lavi et al.
-[155] unfolded the PGA algorithm to optimize hybrid precoding for multiuser downlink MIMO, assuming perfect CSI availability.
-To handle noisy CSI, they implemented the projected conceptual mirror prox algorithm [156, 157].
-The step-sizes for updating the analog and digital precoder matrices are tuned using mini-batch SGD, a data-driven approach.
-The solution is computed within a fixed number of iterations, defined as network layers, ensuring a balance between interpretability and efficiency.
-In [158], Bilbao et al.
-proposed an efficient analog beamforming design to maximize the achievable rate of massive MIMO in-band full-duplex systems.
-The underlying beamforming problem is non-convex, primarily due to constant modulus constraints imposed by the use of analog phase shifters.
-To overcome this challenge, they developed an unfolded algorithm based on PGA combined with ZF projection onto convex sets.
-The convex feasible set arises due to the requirements of the ZF constraint and unit norm beamforming vectors.
-This approach leverages the structure of the zero-forcing conditions and the constant modulus constraints, achieving significantly faster convergence and improved performance compared to classical iterative methods.
-Besides PGA and PGD, CGD is also an iterative algorithm that falls under the gradient-based optimization approach.
-In [159], the CGD algorithm is unfolded for constant envelope precoding to reduce computational overhead.
-The Riemannian manifold optimization (MO)-based CGD is implemented across network layers.
-Simulations show that this approach effectively suppresses multi-user interference, improving overall performance.
-A fast and robust precoding network for MU-MIMO systems with imperfect CSI and per-antenna power constraints is proposed in [160].
-The objective function is a geometric program that minimizes the MSE between the transmitted and recovered symbol vectors at the user.
-The iterative optimization framework is derived from the MSE uplink-downlink duality-based algorithm, which is unfolded to design a robust MSE duality network (RMSED-Net).
-RMSED-Net is robust to channel estimation errors and achieves sum rate close to the upper bound of its benchmark.
-In [161], the interior point method is unfolded for precoding design.
-It leverages interference in multi-user MISO systems to minimize power while meeting signal-to-interference and noise ratio (SINR) constraints by training the proximal operators and Lagrangian multipliers.
-The authors show that the execution time is reduced by ${}50-70\%$ compared to the classical method.
-However, the network is limited to phase shift keying (PSK) modulation, requiring reformulation for higher order and QAM modulations.
-A constructive interference-based symbol-level precoding for multi-user MISO downlink transmission is discussed in [162].
-The proximal Jacobian ADMM is unfolded.
-The unfolded network reduces average transmit power and execution time by 98% compared to proximal Jacobian ADMM.
-The ADMM algorithm struggles with nonlinear discrete optimization due to biased alternating updates.
-To address this, iterative discrete estimation (IDE) is introduced in [163], ensuring unbiased updates and better convergence.
-Wang et al.
-in [163] proposed an algorithm for downlink massive MU-MIMO systems with finite-alphabet precoding to minimize inter-user interference.
-To reduce hardware costs and power consumption in multi-user massive MIMO, He et al.
-[164] developed an unfolded network using the IDE-based algorithm in [163].
-The unfolded network with learnable step-size and damping factor achieves lower BER across a range of SNR.
-In [165] a learning-based channel semantic acquisition and beamforming solution is proposed for cell-free massive MIMO.
-The beamformer is constructed by unfolding successive over-relaxation-based linear beamforming algorithm which reduces estimation errors in HBF.
-The proposed approach reduces computational complexity and enhances robustness to imperfect CSI.
-TABLE V: Summary of typical deep unfolding models for precoder design and their advantages and disadvantages.
-Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
-Unfolded WMMSE [166] Pioneering deep unfolding model based on the WMMSE algorithm for downlink precoding in MU-MIMO systems.
-Achieves $10\%$ improvement in sum rate metric compared with the conventional WMMSE at SNR $=45$ dB.
-Advantages: The network has a few trainable parameters, leading to fast convergence.
-Disadvantages: Performance loss with ill-conditioned channel matrices because the matrix inversion in WMMSE is replaced by a trainable approximation.
-IAIDNN [146] Unfolded WMMSE model for MU-MIMO precoding.
-Converges after $2.94$ seconds compared to $291.01$ seconds required in classical WMMSE and reduces WMMSE computational cost by up to $30\%$.
-Advantages: IAIDNN generalizes to varying system sizes (numbers of users and antennas) with minor adjustments.
-Disadvantages: High training computational complexity due to the increased number of trainable parameters.
-PGA-Net [155] Unfolds PGA with trainable step sizes for hybrid precoding.
-Proposed network achieves better sum-rate than classical PGA with fewer iterations.
-Advantages: Fast convergence and generalize well with varying channel size, effective for wideband channels.
-Disadvantages: Lacks inter-user interference cancellation and involves high gradient computation complexity.
-ManNet [181] Deep unfolding models for HBF (with both fully- and sub-connected architectures) based on Riemannian MO method.
-Over $600$ times faster and $6$ times complexity reduction compared to MO-AltMin algorithm [182] ,has lower complexity than PGA-Net [155].
-Advantages: Sparsely connected network architecture, and thus, has small number of training parameters.
-Disadvantages: The dynamic RF chain-antenna connections has not been incorporated into the unfolding models.
+[107] applied unfolded PGA to hybrid precoding, with trainable step sizes for analog and digital updates; for noisy CSI, they used a projected mirror-prox scheme.
+Bilbao et al.
+[21] unfolded PGA with ZF projection for analog beamforming under constant-modulus constraints in full-duplex massive MIMO, exploiting convex ZF-feasible sets for faster convergence and improved performance.
+They have considered the CSI error to account for hardware impairment and non linearity in the channel.
+Beyond PGD and PGA, CGD has also been unfolded in [79].
+The authors incorporated CGD with Riemannian manifold optimization for constant-envelope precoding, reducing complexity while effectively mitigating multi-user interference.
+In [213], a fast and robust precoding network is proposed for MU-MIMO with imperfect CSI and per-antenna power constraints.
+The MSE minimizing geometric program and its uplink–downlink duality-based iterative solution are unfolded to create RMSED-Net, which offers robustness to channel estimation errors and achieves sum-rate performance close to the benchmark upper bound.
+In [144], an unfolded interior-point method is used for symbol-level precoding in multi-user MISO, where trainable proximal operators and Lagrange multipliers enable power minimization under SINR constraints with a 50–70% reduction in runtime.
+Its main limitation is applicability only to PSK modulation.
+In [216], a constructive-interference-based precoding scheme unfolds the proximal Jacobian ADMM, yielding 98% lower transmit power and significantly reduced execution time.
+Since ADMM suffers from biased updates in discrete optimization, iterative discrete estimation (IDE) for unbiased convergence in finite-alphabet precoding is introduced in [194].
+He et al.
+[75] unfolded this IDE-based algorithm for massive MU-MIMO, with learnable step sizes and damping achieving improved BER across SNRs.
+Finally, in [57] the authors proposed a learning-based channel semantic acquisition and beamforming framework for cell-free massive MIMO, where a successive over-relaxation-based beamforming algorithm is unfolded to reduce estimation error, lower complexity, and enhance robustness under imperfect CSI.
 V-B mmWave MIMO
@@
 Lin et al.
-[167] studied a LMMSE transceiver design for uplink massive MU-MIMO with one-bit ADCs.
-To minimize the system MSE by optimizing the precoder at each user equipment, the PGD algorithm is unfolded.
-The authors focused on training only the step-size in each PGD iteration, significantly reducing the training overhead.
-In [168], the authors combined deep generative models with unfolding techniques to achieve near-optimal HBF while minimizing feedback and computational complexity.
-The receiver estimates the channel using GAN, enabling efficient optimization of digital and analog precoders.
-The HBF problem is formulated as a bilevel optimization.
-The algorithm unfolds an iterative bi-level optimization problem into the layers of the CNN.
-There are two levels are upper and lower layer.
-In the upper layer digital precoder is optimized given a fixed analog precoder while in the lower layer the analog precoder is optimized given a fixed digital precoder.
-This alternating iterative steps are then unfolded.
-The trainable parameter set include the weights of the CNN that maps digital precoder to the analog precoder.
-The unfolded bilevel optimization algorithm achieved near optimal HBF with low feedback and computational cost.
+[116] unfolded the PGD algorithm for LMMSE transceiver design in uplink massive MU-MIMO with one-bit ADCs, training only the step-size per iteration to minimize system MSE and reduce training overhead.
+In [15], deep generative models are combined with unfolding to solve HBF as a bilevel optimization problem.
+The upper layer optimizes the digital precoder given a fixed analog precoder, and the lower layer optimizes the analog precoder given a fixed digital precoder.
+Trainable CNN weights map between the precoders.
+This unfolded bilevel approach achieves near-optimal HBF with low feedback and computational complexity.
 V-C Next Generation Multiple Access
@@
 V-D IRS
-With the aid of IRS, Chen et al.
-[171] addressed the joint design of the HBF matrix and the phase-shifting matrix of the IRS elements.
-The authors proposed the deep unfolded WMMSE-MO network (DU-WMMSE-MO), for IRS-assisted mmWave MIMO-OFDM systems.
-The trainable parameters include the weights for the precoder, phase configuration of the IRS elements and the combiner, and biases for the precoder and phase-shift matrix of the IRS.
-The algorithm showed improved performance in IRS phase-shift design, utilizing two loops to jointly update the step-size of the beamforming and IRS matrices.
-Compared to the Armijo backtracking line search method used in classical WMMSE-MO [172], DU-WMMSE-MO [171] provides more flexibility but with a higher runtime.
-An IRS-assisted multiuser MIMO full-duplex system is explored in [173].
-The active beamforming matrices at the BS and uplink users, along with the passive beamforming matrix at the IRS, are jointly optimized to maximize the weighted sum rate of the system.
-A deep unfolding approach is proposed to approximate the stochastic SSCA-based beamforming algorithm, reducing the complexity of matrix inversion while maintaining near-optimal performance.
-In [174], Min et al.
-introduced a two-timescale transmission approach to reduce the signal processing complexity and channel training overhead for the IRS-aided multiuser MISO system.
-To minimize the average transmit power at the analog precoder while ensuring the individual quality of service (QoS) constraints for the users, a primal-dual decomposition-based two-timescale HBF method is proposed and unfolded.
-This approach allows for efficient beamforming optimization by separating the optimization into two timescales, improving both computational efficiency and performance.
-In [175], a low-complexity IRS-assisted downlink MISO system is designed using an unfolded block CD algorithm to address the joint optimization problem of transmit beamforming and IRS phase shifts.
-The multi-ratio fractional programming structure of the objective function is handled by decoupling the problem into two subproblems: one for optimizing the beamforming at the BS and another for optimizing the phase shifts of the IRS.
-This is achieved through the use of Lagrange dual transformation, effectively simplifying the optimization process.
-The numerical analysis validated that the achieved weighted sum rate is comparable to that of [176] but with reduced computational complexity.
-Furthermore, the authors in [177], have developed an unfolded network to address the joint beamforming task for an IRS-assisted MISO system.
-The WMMSE and power iteration algorithm are used as an alternating optimization (AO) strategy to derive the local optimal solution for the active and passive beamforming.
-The proposed network reduce the computational runtime to approximately 3% of the block CD algorithm in [176].
-In [178], an IRS-assisted MISO covert symbiotic radio communication system is explored, and a deep unfolding algorithm based on gradient descent is proposed for beamforming design.
-The goal is to jointly optimize active and passive beamforming under covert communication constraints.
-The step-size for updating the active and passive beamformers, and Lagrange multipliers form the trainable set.
+Chen et al.
+[31] proposed the deep unfolded WMMSE-MO (DU-WMMSE-MO) network for IRS-assisted mmWave MIMO-OFDM systems, jointly designing the HBF and IRS phase-shifting matrices.
+Trainable parameters include precoder and IRS phase weights and biases.
+The method improves IRS phase-shift design by jointly updating step-sizes, offering more flexibility than classical WMMSE-MO with Armijo backtracking line search method [240] but with higher runtime.
+In [128], the authors studied IRS-assisted multiuser MIMO full-duplex systems, using deep unfolding to approximate the stochastic SSCA beamforming algorithm, reducing matrix inversion complexity while achieving near-optimal weighted sum rate performance.
+Min et al.
+[239] proposed a two-timescale unfolded primal-dual HBF method for IRS-aided multiuser MISO systems to reduce signal processing complexity and channel training overhead.
+By separating optimization into two timescales, the approach efficiently minimizes transmit power while meeting individual quality of service (QoS) constraints, improving computational efficiency and performance.
+In [212], a low-complexity IRS-assisted downlink MISO system is designed using an unfolded block CD algorithm, decoupling the joint optimization of transmit beamforming and IRS phase shifts into two subproblems via Lagrange dual transformation.
+The numerical analysis validated that the achieved weighted sum rate is comparable to that of classical approach in [62] but with reduced computational complexity.
+In [97], an unfolded network combining WMMSE and power iteration was proposed for joint beamforming, reducing computational runtime to about 3% of the classical approach in [62].
+In [78], a deep unfolding algorithm based on gradient descent is proposed for IRS-assisted MISO covert symbiotic radio systems, jointly optimizing active and passive beamforming under covert constraints, with trainable step-sizes and Lagrange multipliers.
 In contrast, Gangyong et al.
-in [179] focused purely on passive beamforming and unfolded the Riemannian gradient descent algorithm.
-They have designed the IRS phase shift matrix by maximizing the sum path gain of the channel.
-In this network, the only trainable parameter is the step-size associated with each Riemannian gradient descent iteration.
+in [247] focused only on passive beamforming, using unfolded Riemannian gradient descent to optimize the IRS phase shift matrix by maximizing the channel path gain.
 The network in [247] is simpler compared to that of [78], due to less number of parameters to train but is specific to optimizing the phase shift of the IRS.
@@
 Compared to the original approximate WMMSE algorithm, the proposed unfolding method achieves superior precoding performance with substantially lower computational complexity.
+In THz communication systems, ultra-large antenna arrays will be employed to form narrow beams and mitigate severe path loss.
+This objective naturally leads to the combination of massive MIMO and wideband THz, which introduces significant scalability challenges.
+In particular, deep unfolding algorithms for precoding must efficiently scale with the large number of antennas and multiple users, whereas existing designs have primarily focused on point-to-point massive MIMO [152] or single-antenna users [209].
+Moreover, the frequency selectivity of THz channels makes the design and optimization of beamformers (both analog and digital) more challenging, as they must operate across a large number of subcarriers in wideband systems.
+The computational complexity of deep unfolding generally grows linearly with the number of subcarriers [152, 209].
+This increase in complexity raises concerns about scalability in practical deployments.
+Moreover, THz systems face hardware constraints such as phase noise, low-resolution ADCs, and synchronization errors, which remain largely unexplored in current deep unfolding designs and present opportunities for further optimization.
+TABLE V: Comparison of Classical vs Unfolded methods for Precoder design.
+$T$: Layers, $I$: Outer iterations, $J$: Inner iterations, $I_{\text{net}}$: Iterations of network, $N_{RF}$: No.
+of RF chains, $N_{t}$: No.
+of transmit antenna, $N_{r}$: No.
+of receive antenna, $K$: No.
+of users, $M$: Symbol size, $N_{f}$: No.
+of subcarriers.
+Unfolded Computational Complexity Number of Parameters (Unfolded) Convergence speed (Number of layers)
+Model Classical Unfolded Simulation Setup Classical Unfolded
+IAIDNN [83] WMMSE:  $\mathcal{O}(N_{RF}^{3})$ $\mathcal{O}(T(K^{2}N_{t}N_{r}^{2}+K^{2}N_{t}^{2}N_{r}+KN_{t}^{2.37}+KN_{r}^{3})$ $TK(3N_{r}^{2}+3M^{2}+MN_{r})+(T-1)K(3N_{t}^{2}+MN_{t})$ Massive MIMO, $N_{t}=128$, $N_{r}=2$ and $K=30$ $100$ iterations 77 Layers
+PGA-Net [107] PGA:  $\mathcal{O}(IN_{f}N_{t}(KN_{RF}+N_{t}(K+N_{RF})))$ $\mathcal{O}(TN_{f}N_{t}(KN_{RF}+N_{t}(K+N_{RF})))$ $T(N_{f}+1)$ Single cell downlink MIMO with Rayleigh fading ($N_{t}=12$, $N_{r}=1$, $K=6$ $N_{f}=8$, $N_{RF}=10$) $100$ iterations 55 Layers
+ManNet [152] MO-AltMin algorithm:  $I\mathcal{O}(N_{t}N_{f}N_{RF}^{2}+J(3N_{t}N_{RF}+2N_{f}(N_{RF}^{2}+N_{RF})M))$ $(I_{\text{net}}-1)\mathcal{O}(N_{t}N_{f}N_{RF}^{2})+\mathcal{O}(N_{t}N_{f}N_{RF})+I_{\text{net}}\mathcal{O}(2N_{f}N_{RF}^{2}M+T(3N_{t}N_{RF}+2N_{f}N_{RF}M)$ $4TI_{\text{net}}N_{t}$ THz massive MIMO, $N_{t}=\{16,32,64,128\}$, $N_{r}=N_{RF}=M=2$ and $N_{f}=128$ $500$ iterations $30$ Layers
 V-F Discussions and Takeaways
 Precoding design in massive MIMO and HBF systems involves solving non-convex optimization problems under hardware constraints and channel uncertainty.
-Popular unfolded networks for precoding design are listed in Table V.
-For instance, IAIDNN [146] unfolds the WMMSE algorithm and replaces complex operations like matrix inversion with a learnable block.
-This results in a significant reduction in computational complexity while achieving performance close to traditional WMMSE solutions.
+Table V lists some popular unfolded networks for precoding design and their performance metrics.
 Deep unfolding enables the incorporation of domain-specific constraints such as the constant modulus constraint in analog beamformers through tailored activation functions.
@@
 Integration with channel estimation modules and robust designs under CSI uncertainty will be key for practical deployment in next-generation wireless networks.
-Moreover, applying such unfolded frameworks for precoders in real systems and analyzing the performance in real experiments will be very beneficial.
+Moreover, applying such unfolded frameworks in real systems and analyzing the performance in real experiments will be very beneficial.
+On the other hand, more emphasis should be placed on ensuring the scalability of deep unfolded precoding algorithms, especially for massive MIMO wideband systems.
+This requires not only unrolling conventional iterative algorithms, but also rethinking the unfolding architecture and optimization so that the models scale effectively with the array size, bandwidth, and number of users.
+Possible solutions include designing unfolding networks with sparse inter-layer connections [152] and employing adaptive pruning or approximated computations to reduce redundant operations across layers [143, 10].
+For distributed systems, scalability can also be achieved through hierarchical unfolding architectures that offload the main computational tasks to a central unit, as further discussed in Section X-B3.
 VI Sensing and Communication
@@
 Key sensing parameter estimations have been thoroughly analyzed in [118], revealing theoretical and practical boundaries.
-TABLE VI: Summary of typical deep unfolding models for ISAC and their advantages and disadvantages.
-Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
-Unfolded PGA network [191] Unfolds PGA to maximize the system sum rate while minimizing the sensing beampattern error.
-Achieves $33.5\%$ higher sum rate, $2.5$ dB lower radar beam pattern error compared with SCA, and $65\%$ reduction in runtime and complexity compared with the conventional PGA.
-Advantages: Ensure convergence in the challenging ISAC HBF designs, perform well in both single use and MU systems.
-Disadvantages: The convergence speed is relatively slow, high complexity to compute the gradients.
-ISAC-Net [194] Unfolds the denoising AMP method for communication and DFT-based sensing in narrowband OFDM ISAC.
-Achieve communication performance close to OAMP-Net2 [11] and better sensing performance than classical 2D-DFT algorithm.
-Advantages: Leverages shared resources to overcome challenges in signal processing without prior knowledge of transmitted signals, enhanced accuracy and security.
-Disadvantages: Beam squint effect is not considered leading to discrepancies extracting sensing parameters (angle of arrival or angle of departure) across various subcarriers in OFDM.
-ADMM-Net for OTFS ISAC [196] Unfolds ADMM to estimate delay-Doppler targets and demodulate OTFS symbols under high mobility.
-With $2$ targets, ADMM-Net achieve $4$ dB gain in NMSE compared with the classical ADMM and ISTA at SNR=$14$ dB.
-Advantages: Suitable for vehicular/high-mobility scenarios and robust to delay-Doppler spread.
-Disadvantages: Real time CSI estimation is not addressed.
-Double loop unfolded network [197] Unfolds gradient descent to jointly optimize IRS phase shifts in ISAC system.
-Minimizes worst case CRB for sensing targets.
-Advantages: Has low computational complexity and fast convergence (within $15$ layers).
-Disadvantages: Offline training ($1000$ channel realization, $20$ epochs) may not be feasible for frequently changing environments.
 VI-A MIMO ISAC
@@
 To address this, the authors employed a deep unfolded network based on PGD algorithm, ensuring efficient waveform design under these constraints.
-HBF in MIMO ISAC systems poses additional complexity due to balancing communication rates and sensing accuracy.
-The authors in [191] tackled this challenge by unfolding the iterative PGA algorithm.
-Their deep unfolding-based method accelerated convergence and enhanced overall system performance, achieving an optimal balance between communication and sensing objectives.
-In addition, joint optimization of radar communication beamforming in MIMO system was addressed in [192].
-The authors incorporated symbol-level precoding with constructive interference constraints to improve energy efficiency by exploiting multi-user interference.
-Since, the objective problem is computationally complex due to non-convexity and higher dimensional matrix inversions, a custom iterative algorithm is developed.
-The iterative algorithm is based on block CD and primal-dual optimization (augmented Lagrangian method).
-Therefore, the trainable parameter set for training the deep unfolded network include step-sizes to update the primal, dual and auxiliary variable, as well as a correction matrix to approximate matrix inversion.
-The proposed network reduces computational complexity due to small number of trainable parameters, improve scalability, and support real-time implementation.
-The large-scale antenna arrays in mmWave MIMO systems enable sub-meter user localization by providing high angular resolution.
-However, these methods often incur significant computational complexity, particularly in high dimensional mmWave MIMO settings.
-To improve this, Fan et al.
-proposed, a fast and accurate direct localization method in [193].
-An ADMM solver is developed to solve the optimization problem with low complexity.
-The ADMM iterations are unfolded to learn the trainable parameters that include penalty and proximal parameters, and weight coefficient for channel gain between the BS and user.
-The unfolded network converges faster than classical ADMM and computational time is reduced.
+HBF in MIMO ISAC systems introduces complexity by balancing communication rates and sensing accuracy.
+The authors in [153] addressed this by unfolding the iterative PGA algorithm, accelerating convergence and improving system performance.
+Similarly, the authors in [231] tackled joint radar-communication beamforming using symbol-level precoding with constructive interference constraints to enhance energy efficiency.
+To handle the non-convex and high-dimensional optimization, a custom iterative algorithm based on block CD and primal-dual optimization was proposed.
+The deep unfolded network trains step-sizes and a correction matrix for matrix inversion, reducing complexity, improving scalability, and enabling real-time implementation.
+In mmWave MIMO systems, large antenna arrays allow sub-meter localization but with high computational cost.
+Fan et al.
+[50] proposed a fast direct localization method using an unfolded ADMM solver with trainable penalty, proximal parameters, and channel gain weights.
+Also in [232], the authors have unfolded ADMM to design a transceiver for ISAC in a cluttered environment.
+The learnable parameters consist of the per layer step sizes in the network and the ADMM penalty coefficients.
+Both these approaches [50, 232] achieve faster convergence and lower computational time than classical ADMM.
 VI-B Modern Multicarrier Modulation for ISAC
@@
 The proposed network converges with $8$ layers as compared to $2000$ iteration in fast ISTA.
-Also, the network achieves an NMSE of $-20.12$ dB while the classical complex ISTA achieved $-16.8$ dB.
+Also, the network achieves an NMSE of 20.12-20.12 dB while the classical complex ISTA achieved 16.8-16.8 dB.
 The problem of estimating parameters of moving targets in high-speed scenarios using OTFS modulation for ISAC systems is explored in [117].
@@
 The proposed network effectively achieves near-optimal performance by handling statistical CSI errors and bounded target direction errors.
+TABLE VI: Comparison of Classical vs Unfolded methods for ISAC.
+II: Inner Layer, $J$: Outer Layer, $N_{RF}$: No.
+of RF chains, $N_{t}$: No.
+of transmit antenna, $N_{r}$: No.
+of receive antenna, $N_{b}$: No.
+of BS antenna, $G$: No.
+of uniform grid location, $B$: No.
+of BS, $L_{b}$: No.
+of uniform grid of angle, $M$: No.
+of OFDM symbols, $N$: No.
+of subcarriers, $P$: No.
+of multipaths.
+Unfolded Computational Complexity Number of Parameters (Unfolded) Convergence speed (Number of layers)
+Model Classical Unfolded Simulation Setup Classical Unfolded
+Unfolded PGA [153] PGA: $IJ\mathcal{O}(4N_{\text{RF}}N_{t}^{2}+3N_{t})$ $\mathcal{O}(IJN_{t}^{2}N_{r})$ $2IJ$ Massive MIMO, $N_{t}=64$, $N_{r}=$, $N_{\text{RF}}=4$. $\sim 500$ iterations $40$ Layers
+Unfolded ADMM for localization [50] Direct Source Localization [58]: $\mathcal{O}((GB+\sum_{b=1}^{B}L_{b})^{3.5}\sum_{b=1}^{B}N_{b})$ $\mathcal{O}(((BG)^{2}+(\sum_{b=1}^{B}L_{b})^{2})\sum_{b=1}^{B}N_{b})$ $3J+B$ 5G 3GPP Urban Macro MIMO mmWave channel, $B=4$, $N_{b}=50$, $G=900$, $L_{b}=100$.
+$30$ iterations 55 Layers
+ISAC-Net [94] 2D FFT and cyclic cross correlation [207]: $\mathcal{O}(J(M^{2}+N^{2}))$ $\mathcal{O}(TP(N_{t}N_{r}+MN+N\log(N)+M\log(M)+2N_{t}N_{r}(MN)^{2}))$ $5J$ MIMO OFDM ($N_{t}=N_{r}=8$), 16-QAM modulation, $M=256$, $N=1024$.
+$500$ iterations 55 Layers
 VI-D Discussions and Takeaways
@@
 These approaches reduce computational complexity and provide faster convergence compared to traditional iterative solutions.
-We summarize typical deep unfolding models for ISAC in Table VI.
+We summarize typical deep unfolding models for ISAC in Table VI and present a comparison in terms of computational metrics.
 It is observed that deep unfolding exhibits advantages in constant modulus waveform design, optimization of dual-functional beamformers, and power allocation.
@@
 The LADN [205] outperforms the original ADMM-penalized decoder in block error rate for LDPC codes, while the network in [64] outperforms the original ADMM-penalized decoder in block error rate for irregular LDPC codes.
-In [36], the authors proposed the TurboNet architecture, to improve turbo decoding by unfolding the max-log MAP algorithm.
-Trainable parameters were inserted into the max-log-MAP decoding flow like weights for prior LLRs.
-Numerical analysis demonstarted that TurboNet outperforms classical turbo decoding based on the max-log-MAP algorithm, particularly at moderate-to-high SNRs.
-Furthermore, TurboNet has been validated through over the air experiments confirming its robustness to SNR mismatches.
-Moreover, the authors have extended the work by including joint detection and decoding in [208].
-A CGD-based OAMP detector (CG-OAMPNet) for the MIMO-OFDM system is proposed by unfolding EP.
-The robustness of CG-OAMPNet was further validated through over-the-air experiments, demonstrating practical resilience to environmental variations.
-In correlated channel scenario, CG-OAMPNet outperforms OAMP-Net2 [11] by performance gain of $\sim 2$ dB at BER$=10^{-3}$.
-The authors in [209] tackle the decoding challenges of marker codes in insertion and deletion channels without perfect CSI.
-They proposed a network called FBNet to improve detection performance; the forward-backward algorithm is unfolded for marker codes.
-The trainable weights are the CSI probabilities for insertion/deletion channels.
-FBNet performs quite close to the classical forward-backward algorithm but is robust to CSI mismatch.
-FBNet [209] is based on RNN structure and target synchronization errors (insertions/deletions) without accurate CSI, while TurboNet [36] is a deep feedforward DNN structure and targets substitution errors in turbo codes under AWGN channel.
+[…]
 Gao et al.
@@
 The networks showed fast adaptation and strong BER gains in over-the-air experiments.
-TABLE VII: Summary of typical deep unfolding models for decoding and their advantages and disadvantages.
-Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
-LADN [39] Introduces learnable ADMM-based decoder with trainable penalties for binary LDPC decoding.
-Achieves a block error rate reduction of 10$\sim$ 10$\times$ 10 × at SNR $=2$ dB over the baseline ADMM decoder.
-Advantages: Has small number of learning parameters and fast convergence.
-Disadvantages: Limited by code length.
-Inherent codeword structure not exploited leading to high computational complexity.
-TurboNet [36] Unfolds max-log MAP decoding steps for turbo codes with learnable weights.
-Outperforms the conventional max-log MAP scheme by $0.25-0.5$ dB BER.
-Advantages: Compatible with 3GPP turbo codes and validated through practical experiment.
-Disadvantages: Training requires fixed interleaver and careful SNR selection.
-Quantized Polar BP [210] Learns BP decoding under 1-bit quantization for polar codes over an AWGN channel.
-Outperforms standard BP with one-bit quantization by $1.2-2$ dB (block error rate).
-Advantages: Hardware-friendly and applicable to IoT systems with coarse ADCs.
-Disadvantages: Less scalable.
-Require large number of trainable weights assigned to each edge of factor graph for decoding.
-ADMM-Net [40] Extends LADN [39] to irregular LDPC codes.
-Unfolded ADMM, learn different penalty weights for each variable node in decoding.
-Achieve better BER performance than BP and ADMM penalized decoders for irregular LDPC codes.
-Advantages: Suitable for irregular code graphs.
-Disadvantages: The initialization strategy for the trainable parameters assumes prior knowledge of the code structure, which makes it less generalizable.
-FBNet [209] Introduce a RNN-based unfolding model based on forward-backward decoding for Marker codes over insertion-deletion channels.
-Outperforms the classical forward-backward algorithm under imperfect CSI.
-Advantages: Robust to synchronization errors with flexible sequence handling.
-Disadvantages: Degraded performance under burst insertion-deletion channels.
+TABLE VII: Comparison of Classical vs Unfolded methods for Decoding in Error Correcting Codes.
+$T$: Layers, $I$: Iterations, $n$: No.
+of variable nodes, $\Gamma_{a}$: No.
+of auxiliary variables, $m$: No.
+of check nodes, $d_{\text{max}}$: Maximum degree of check nodes, $N_{t}$: No.
+of transmit antenna, $N_{r}$: No.
+of receive antenna, $t$: Time slot.
+Unfolded Computational Complexity Number of Parameters (Unfolded) Convergence speed (Number of layers)
+Model Classical Unfolded Simulation Setup Classical Unfolded
+ADMM-DL Network [64] ML decoder: $\mathcal{O}(2^{n})$ $\mathcal{O}(n+\Gamma_{a})$ $T(n+\Gamma_{a}+1)$ Irregular LDPC code $[n=2048,m=512]$ 3GPP protocol and $[n=576,m=288]$ WiMAX protocol, BPSK modulation.
+$24$ iterations 33 Layers
+PDD-Net [129] Standard BP decoder: $\mathcal{O}(md_{\text{max}}^{2})$ $\mathcal{O}(n+m(d_{\text{max}}-3))$ $3T$ LDPC code $[n=256,m=128]$, BPSK modulation. $2\times 10^{4}$ iterations $25$ Layers
+JCED Net [180] BP decoder: $\mathcal{O}(I(4md_{\text{max}}+md_{\text{max}}^{2}))$ $\mathcal{O}(T(8N{r}N_{t}^{2}+10N_{r}N_{t}+16N_{t}^{2}t+8N_{r}N_{t}t+(8+2d_{\text{max}})m2^{d_{\text{max}}-1}))$ $4T$ mmWave MIMO ($N_{r}=8$, $N_{t}=4$), regular LDPC code [$n=288$, $m=144$].
+$100$ iterations 202520-25 Layers
 VII-A Discussions and Takeaways
-In Table VII, we summarize the most popular unfolded network for the decoding of channel codes.
+Table VII compares various unfolded networks for decoding in error correcting codes based on computational complexity, trainable parameter count, and convergence speed (in number of layers).
 Deep unfolding spans their applications in modern error-correcting codes, such as LDPC, turbo, and polar codes.
@@
 X-A Limitations of Deep Unfolding Techniques
-Although deep unfolding has found applications in various design problems, it has limitations.
-Most of these limitations arise from the fact that deep unfolding is not an independent design methodology; it must be built upon existing conventional optimization frameworks.
-Consequently, it faces the following challenges in the development and real-world applications:
-(C1) Methodological challenges: Deep unfolding methods are inspired by the iterative procedures of conventional optimization algorithms, where each iteration is typically interpreted as a layer in an NN.
-In traditional wireless transmitter-receiver design and signal processing, optimization problems are often non-convex and highly coupled due to factors such as fractional SINR, constant-modulus constraints (e.g., for analog beamforming), or mixed-integer variables (e.g., in symbol detection or antenna selection).
-They are typically solved using a two-layer approach.
-In the first layer, AO or block coordinate ascent/descent is used to decouple variables, breaking the original problem into smaller, more manageable sub-problems.
-In the second layer, convex optimization techniques such as SCA or the ADMM are employed to solve each sub-problem.
-While this hierarchical structure enables the use of convex solvers (e.g., CVX), it often introduces auxiliary variables and results in nested loops that must converge at each layer, significantly increasing algorithmic complexity.
-Deep unfolding models, on the other hand, aim to transform these iterative procedures into end-to-end trainable architectures.
-Instead of solving sub-problems sequentially, the unfolding framework typically approximates the entire optimization process in a more unified and parallelizable manner.
-This shift in design philosophy, from explicitly solving decoupled sub-problems to learning a unified mapping, poses non-trivial challenges, especially when attempting to translate complex multi-stage optimizers with auxiliary variables into a compact, learnable structure.
-Therefore, while deep unfolding builds upon conventional iterative methods, the methodological differences in structure and implementation represent a major challenge in developing effective unfolding models.
-(C2) Unfolding principle algorithms: While deep unfolding typically enhances conventional iterative algorithms, not all can be effectively unfolded.
-This limitation arises from the core principle of deep unfolding, which structures the iterations of an iterative algorithm as layers within a DNN.
-Complex operations like matrix decomposition or the use of convex optimization tools such as CVX or YALMIP (with solvers like MOSEK, SeDuMi, etc.) are difficult to implement in a DNN framework.
-In large-scale systems such as massive MIMO, even matrix multiplications result in high complexity for unfolding models.
-Another challenge is parameter generalization, particularly with respect to the number of layers in the unfolding model.
-In iterative algorithms with nested loops (discussed in C1), both inner and outer loops run until convergence, but deep unfolding models require the number of layers to be predetermined before deployment.
-This is typically done through empirical tuning, which may not provide optimal values for all channel realizations.
-As a result, convergence is not guaranteed across different scenarios, potentially leading to performance degradation.
-Proper tuning is essential to achieve a favorable performance-complexity tradeoff in deep unfolding models.
-(C3) Handling design constraints and objectives: One of the biggest challenges in developing deep unfolding models lies in managing the design constraints formulated in the problem.
-These constraints require the DNN layers’ outputs to meet specific requirements, such as total power, QoS, or other element-wise constraints.
-Simple constraints, like power or constant modulus, can often be enforced through normalization [181].
-For instance, the $tanh$ activation function can be employed to impose constant-modulus constraints on analog beamforming coefficients or sensing waveforms, similar to black-box DL architectures.
-However, QoS constraints are significantly more complex.
-While they resemble activation functions, none of the existing activation functions can adequately manage such intricate requirements.
-One approach is to incorporate these constraints into the design objective.
-However, this technique often results in an nonequivalent problem with a more complicated objective function [191].
-Not only do complex constraints pose challenges for developing deep unfolding models, but intricate objective functions also present difficulties in training the models.
-This is because their loss functions are generally formulated based on the objective function.
-For instance, unfolding architectures like DetNet [49], FSNet [75], and ScNet [229] for symbol detection are trained to minimize the Euclidean distance between the model’s output (symbol vector) and the true transmitted symbol vector.
-In particular, when labels are unavailable, as in beamforming problems, unfolding networks use the objective function as the loss function for minimization or its negative for maximization problems.
-Typical deep unfolding techniques [191, 155] require calculating the gradient of the objective function.
-However, this becomes challenging when the objective function is a complex, high-order, or fractional function of the design variables, especially when those variables are vectors or matrices.
-(C4) Sophisticated mappings: While deep unfolding is powerful, certain inference tasks still require black-box DNN models, especially when there is no clear relationship between the input data, the output solution, and performance metrics.
-For instance, in multi-modal sensing-assisted beam tracking, tasks may involve complex mappings from environmental data, such as sensory information from sensors and cameras, to the optimal beam [230, 231].
-In such cases, there is no explicit relationship between environmental images and the optimal beam, and standard communication metrics like throughput or SINR cannot be directly expressed as functions of these images.
-These problems cannot be formulated as optimization tasks solvable with principle algorithms, making it difficult to develop unfolding models that map environmental information to beam patterns or blockage states.
-(C5) Real-world applications: Deep unfolding architectures, while promising in mimicking iterative algorithms through trainable layers, often entail considerable depth and parameter complexity [191].
-This imposes severe challenges for real-time inference under strict latency and throughput constraints, especially in systems with multiple and large dimensions like massive MIMO-OFDM.
-From a hardware perspective, deep unfolding designs are not naturally aligned with the requirements for area-efficient and parallelizable architectures needed in scalable OFDM deployments.
-For example, the non-uniform and sequential computations common in unfolding detectors can result in inefficient silicon utilization and high power consumption, limiting the ability to instantiate multiple parallel detection cores across subcarriers [232].
-Furthermore, in real-world scenarios, the data distribution is highly dynamic because of user mobility, varying propagation environments, and hardware impairments.
-These variations often violate the assumptions underlying many unfolding-based transmitter and receiver designs, which are typically developed under idealized conditions.
-However, this assumption breaks down in practical deployments, where channels often exhibit strong line-of-sight components and spatial correlation.
-As a result, detectors relying on such assumptions can suffer significant performance degradation.
-To maintain robustness in these conditions, deep unfolding models must be redesigned or trained using data that captures the complexity and variability of realistic, deployment-specific channels.
-The above five challenges (C1C5) also outline the scenarios wherein deep unfolding models cannot be built or does not offer much gain in performance-complexity tradeoff.
-These include when there is lack of suitable conventional iterative algorithms (C1, C2, C4), when the design problems are highly complicated with constraints and objective function (C3), or when the exist imperfection in real-world propagation environments and hardware (C5).
-It can be seen that deep unfolding is not a game changer that will completely replace conventional principle algorithms or black-box ML models.
-However, the applicability of deep unfolding can be enhanced by overcoming those challenges, as will be discussed next.
+Although deep unfolding has found applications in various design problems, it also has important limitations.
+Most of these stem from the fact that deep unfolding is not an independent design methodology; it must be built upon existing conventional optimization frameworks.
+Consequently, it faces the following challenges in both development and real-world applications:
+(C1) Methodological challenges: Deep unfolding methods are inspired by the iterative procedures of conventional optimization algorithms, where each iteration is typically interpreted as a layer in a neural network.
+In traditional transmitter-receiver design and signal processing, optimization problems are often non-convex and highly coupled due to factors such as fractional SINR, constant-modulus constraints, or mixed-integer variables.
+These are typically solved by a two-layer approach: in the first layer, alternating optimization or block coordinate ascent/descent is used to decouple variables; in the second layer, convex optimization techniques such as SCA or ADMM are applied to the resulting sub-problems.
+While this makes the problems solvable using convex solvers (e.g., CVX), it introduces auxiliary variables and nested loops that must converge at each stage, increasing algorithmic complexity.
+Deep unfolding, by contrast, attempts to approximate such iterative processes in a unified, trainable architecture, making it challenging to translate multi-stage solvers into compact unfolding models.
+(C2) Unfolding principle algorithms: Deep unfolding typically enhances conventional iterative algorithms, but not all can be effectively unfolded.
+Complex operations such as matrix decompositions or convex solver calls (e.g., CVX, YALMIP with MOSEK or SeDuMi) are difficult to represent in differentiable NN layers.
+Furthermore, large-scale systems such as massive MIMO incur high complexity even for basic operations like matrix multiplications.
+Another issue is the fixed number of layers: iterative algorithms run until convergence, whereas unfolding models require the number of layers to be fixed a priori, often through empirical tuning.
+This can lead to suboptimal performance or lack of guaranteed convergence across different scenarios.
+(C3) Handling design constraints and objectives: Another major challenge is enforcing problem-specific constraints and objectives in the unfolding framework.
+Simple constraints (e.g., power or constant modulus) can often be imposed by normalization or suitable activation functions, but more complex QoS constraints are difficult to embed.
+Similarly, intricate objective functions result in complicated loss functions, especially when gradients are hard to compute for high-order or fractional functions.
+This complicates training and may limit the range of problems amenable to unfolding.
+(C4) Sophisticated mappings: While deep unfolding is powerful for problems with well-defined iterative solvers, many inference tasks lack such formulations.
+For example, in multi-modal sensing-assisted beam tracking, there is no explicit relationship between environmental sensory data (e.g., images) and the optimal beam.
+Since these problems cannot be expressed as tractable optimization tasks, unfolding methods are not suitable, and black-box learning models remain necessary.
+(C5) Real-world applications: Deep unfolding architectures, while promising in theory, often require considerable network depth and parameter complexity.
+This creates challenges for real-time inference under stringent latency and throughput constraints, especially in high-dimensional systems such as massive MIMO-OFDM.
+Hardware-wise, unfolding designs are not naturally aligned with area-efficient or parallelizable architectures, leading to inefficient silicon utilization and high power consumption.
+Moreover, practical deployments face dynamic channel conditions, mobility, and hardware impairments that often violate the assumptions made in unfolding-based designs.
+As a result, models trained under idealized conditions may suffer significant performance degradation in practice.
+The above five challenges (C1C5) illustrate scenarios where deep unfolding is either not applicable or offers limited performance-complexity benefits.
+These include the absence of suitable iterative algorithms (C1, C2, C4), the presence of highly complicated constraints or objectives (C3), and practical imperfections in real-world propagation environments and hardware (C5).
+It is clear that deep unfolding is not a universal replacement for conventional algorithms or black-box machine learning models, but rather a complementary tool most effective when problem structures are well matched to unfolding principles.
+At the same time, several strategies can mitigate the negative impacts of these limitations in practice.
+Robustness can be improved by training on realistic datasets that incorporate mobility, propagation variability, and hardware impairments.
+Hybrid designs, in which unfolding is applied only to specific modules (e.g., replacing matrix inversion or accelerating convergence) while other parts rely on conventional or data-driven methods, can achieve better tradeoffs between interpretability and performance.
+Finally, hardware-aware unfolding architectures tailored for parallelism and energy efficiency can help overcome deployment challenges in large-scale systems.
+Moreover, the applicability of deep unfolding can be enhanced by overcoming those challenges, as will be discussed next.
 X-B Development of Deep Unfolding Techniques
 X-B1 Coexistence of Deep Unfolding Models, Conventional Solvers, and Black-Box Models
-Deep unfolding doesn’t need to fully replace conventional algorithms or black-box ML models.
+Deep unfolding doesn’t need to fully replace conventional algorithms or black-box machine learning models.
 Instead, these approaches can coexist, leveraging their combined strengths.
@@
 Furthermore, balancing the changes of these objectives with respect to the design variables across layers is crucial [153, 132], necessitating careful design of the unfolding models.
-X-B3 Distributed Deep Unfolding Architectures
-While deep unfolding models provide advantages in terms of complexity due to faster convergence compared to conventional iterative algorithms, they can still face high complexity issues due to challenge C2.
-This limits their applicability on platforms with constrained computational resources, such as edge devices or distributed access points (APs).
-Future wireless networks will advance beyond simple communications, evolving into large-scale, distributed systems such as cell-free massive MIMO that integrate advanced sensing, communications, and computing capabilities.
-However, most existing AI models, both data-driven and model-driven, were designed for centralized scenarios, where they are deployed at a central node with substantial computing power.
-These models are not well-suited for distributed scenarios, highlighting the need for appropriate distributed unfolding architectures.
-One potential solution is to develop distributed deep unfolding architectures.
-In this approach, rather than executing all the unfolded layers at every distributed node or edge device, the layers can be distributed across multiple platforms for parallel execution, thereby reducing latency.
-For example, to design $L$ beamformers for $L$ APs for their downlink transmission in a distributed system, a CPU with high computing power and hardware resources can handle the majority of the layers, generating high-quality beamformers.
-These beamformers are then sent to the APs, serving as initial solutions and inputs for smaller sub-unfolding models at the distributed APs.
-This ensures that reliable beamformers can be generated at the APs with only a few layers, minimizing complexity and runtime to meet the computational constraints of the APs.
-Conversely, in the uplink, a distributed unfolding detection model can be deployed at the APs to generate initial detected symbols, which are then forwarded to the CPU for further processing and final decision-making by the large unfolding detector.
-By such an unfolding framework, highly reliable symbols are generated with low complexity at distributed APs.
+X-B3 Scalable Distributed Deep Unfolding Architectures
+While deep unfolding models provide advantages in terms of complexity due to faster convergence compared to conventional iterative algorithms, they can still face scalability issues due to challenge C2, especially when deployed on platforms with constrained computational resources such as edge devices or distributed access points (APs).
+Future wireless networks will evolve into large-scale distributed systems, such as cell-free massive MIMO, that integrate advanced sensing, communications, and computing capabilities.
+However, most existing AI models, both data-driven and model-driven, are designed for centralized scenarios where they are executed at a powerful central node, making them poorly suited for distributed deployments.
+This highlights the need for scalable distributed unfolding architectures.
+One potential solution is to develop distributed deep unfolding frameworks.
+Instead of executing all the unfolded layers at every AP or edge device, the layers can be partitioned across multiple platforms for parallel execution, thereby reducing latency and computational burden.
+For example, to design $L$ precoders for $L$ APs, a CPU with strong computational resources can process the majority of the layers and generate high-quality beamformers.
+These beamformers are then sent to the APs as initial solutions, which are refined locally by lightweight sub-unfolding models requiring only a few layers.
+This ensures reliable beamformers at the APs while maintaining scalability by reducing their computational load.
+Conversely, in the uplink, a distributed unfolding model can be deployed at the APs to generate initial solutions, which are then forwarded to the CPU for final refinement by a larger unfolding model.
+Such a hierarchical CPU-AP cooperation framework directly addresses scalability challenges in distributed deployments by lowering the per-node computational cost, reducing fronthaul/backhaul signaling overhead, and enabling heterogeneous APs to operate with resource-aware lightweight sub-models.
+In practice, this can be further enhanced by tailoring the unfolding depth at each AP to its available hardware resources and by adopting split-learning style execution, where unfolding layers are adaptively divided between CPU and APs.
+In this way, reliable symbol detection and beamforming can be achieved with scalable complexity across distributed networks.
 X-B4 Deep Unfolding to Enhance Black-box Solution
-As its name suggests, a black-box ML model operates with a black-box nature, making it unexplainable and resistant to re-structuring beyond tuning parameters like the number of layers or nodes.
+As its name suggests, a black-box machine learning model operates with a black-box nature, making it unexplainable and resistant to re-structuring beyond tuning parameters like the number of layers or nodes.
 The performance and efficiency of such models are heavily dependent on the quality of input data.
@@
 This makes deep unfolding a valuable tool for handling large-scale sensory data in real-time communication systems.
+X-B5 Evaluation Protocols for Deep Unfolding
+To facilitate the development of deep unfolding techniques, it is necessary to establish standardized evaluation protocols.
+A natural and broadly applicable criterion is the convergence profile, i.e., how quickly an unfolding model achieves satisfactory performance within a limited number of layers [152, 153, 132, 107, 156, 157, 149].
+This metric reflects both computational complexity (since fewer layers correspond to fewer iterations and lower runtime) and model performance.
+Convergence profiles also allow direct comparison with conventional iterative algorithms, highlighting the performance complexity trade-offs of unfolding designs.
+Another important aspect is the methodology for employing different learnable hyperparameters.
+Most unfolding techniques learn scalar parameters such as step sizes [152, 153, 132, 107, 156, 157, 149], while others consider richer structures such as trainable matrices [83].
+Although the specific parameters vary, they share the common effect of shaping convergence speed and stability, which again makes the convergence profile a valid and unifying evaluation criterion.
+By contrast, training convergence criteria (e.g., number of epochs or early stopping) are less critical, as unfolding models typically converge quickly [152], and this detail is often omitted in the literature.
+Furthermore, reproducible open-source implementations of deep unfolding solutions with clear and consistent settings would not only enable fair comparisons but also facilitate the development of new unfolding models.
+In Table VIII, we present several available implementations for representative tasks.
+TABLE VIII: Open-source implementations of deep unfolding techniques. The GitHub links are embedded directly in model names as hyperlinks.
+Design Problem Representative Works
+Theoretical convergence of LISTA LISTA [33]
+Detection DetNet [163], FSNet [151], MMNet [101], LoRD-Net [102], CG-OAMP-Net [246]
+Channel Estimation SBL-Net [54]
+Precoding Unfolded WMMSE [156]
+Hybrid Beamforming ManNet, Sub-ManNet [152], Unfolded PGA [107], Modular PGA [108]
+ISAC Unfolded PGA-ISAC [153]
+Decoding in Error Correcting Codes JCED-Net [180], NMS-Net [198]
 XI Conclusions